### v5

# Torsion Vanishing for some Shimura Varieties

Abstract.
We generalize the torsion vanishing results of [7, 8, 37, 54]. Our results apply to the cohomology of general Shimura varieties $(\mathbf{G},X)$ of PEL type AC, localized at a suitable maximal ideal $\mathfrak{m}$ in the spherical Hecke algebra at primes $p$ such that $\mathbf{G}_{\mathbb{Q}_{p}}$ is a group for which we know the Fargues-Scholze local Langlands correspondence is the semi-simplification of a suitably nice local Langlands correspondence, as shown in [16, 27, 31, 47]. This is accomplished by combining Koshikawa’s technique [37], the theory of geometric Eisenstein series over the Fargues-Fontaine curve [24], the work of Santos [54] describing the structure of the fibers of the minimally and toroidally compactified Hodge-Tate period morphism for general PEL type Shimura varieties of type AC, and ideas developed by Zhang [68] on comparing Hecke correspondences on the moduli stack of $G$-bundles with the cohomology of Shimura varieties. In the process, we also establish a description of the generic part of the cohomology that bears resemblance to the work of Xiao-Zhu [67]. Moreover, we also construct a filtration on the compactly supported cohomology that differs from Mantovan’s filtration in the case that the Shimura variety is non-compact,allowing us to circumvent some of the circumlocutions taken in [8, 37]. Our method showcases a very general strategy for proving such torsion vanishing results, and should bear even more fruit once the inputs are generalized. Motivated by this, we formulate an even more general torsion vanishing conjecture (Conjecture 6.6).
1. Introduction
1.0.1. The Main Result
Let $\mathbf{G}$ be a connected reductive group over $\mathbb{Q}$ admitting a Shimura datum $(\mathbf{G},X)$ which we fix from now on. Fix a prime number $p>0$ and let $G:=\mathbf{G}_{\mathbb{Q}_{p}}$ be the base-change to $\mathbb{Q}_{p}$. We will assume that $G$ is unramified so that there exists a hyperspecial subgroup $K^{\mathrm{hs}}_{p}\subset G(\mathbb{Q}_{p})$ and a Borel $B\subset G$ surjecting onto a maximal torus $T\subset G$ which we now fix. We consider the open compact subgroup $K:=K^{p}K_{p}^{\mathrm{hs}}\subset\mathbf{G}(\mathbb{A}_{f})$, where $K^{p}\subset\mathbf{G}(\mathbb{A}_{f}^{p})$ denotes a sufficiently small level away from $p$. Let $\mathrm{Sh}(\mathbf{G},X)_{K}$ denote the corresponding Shimura variety defined over the reflex field $E$. Given a prime $p\neq\ell$, we will be interested in understanding the $\ell$-torsion cohomology groups
$R\Gamma_{c}(\mathrm{Sh}(\mathbf{G},X)_{K,\overline{E}},\overline{\mathbb{F}}_{\ell})$
and
$R\Gamma(\mathrm{Sh}(\mathbf{G},X)_{K,\overline{E}},\overline{\mathbb{F}}_{\ell}).$
In particular, since the level at $p$ is hyperspecial, the unramified Hecke algebra
[…]
will act on these complexes via the right action. Given a maximal ideal $\mathfrak{m}\subset H_{K_{p}^{\mathrm{hs}}}$, we can localize both of these cohomology groups at $\mathfrak{m}$. We will be interested in describing this localization. To do this, we recall that, given such a maximal ideal $\mathfrak{m}\subset H_{K_{p}^{\mathrm{hs}}}$, this defines an unramified $L$-parameter
$\phi_{\mathfrak{m}}:W_{\mathbb{Q}_{p}}\rightarrow\phantom{}^{L}G(\overline{\mathbb{F}}_{\ell})$
up to $\hat{G}-$conjugacy specified by a semisimple element $\phi_{\mathfrak{m}}(\mathrm{Frob}_{\mathbb{Q}_{p}})$. In particular, the $L$-parameter $\phi_{\mathfrak{m}}$ up to conjugacy is induced from a parameter $\phi_{\mathfrak{m}}^{T}:W_{\mathbb{Q}_{p}}\rightarrow\phantom{}^{L}T(\overline{\mathbb{F}}_{\ell})\subset\phantom{}^{L}G(\overline{\mathbb{F}}_{\ell})$ factoring through the $L$-group of the maximal torus. Now, recall that the irreducible representations of $\phantom{}{}^{L}T$ correspond to the $\Gamma$-orbits $\mathbb{X}_{*}(T_{\overline{\mathbb{Q}}_{p}})/\Gamma$ of geometric cocharacters of $T$. We have the following definition.
Definition 1.1.
[24, Definition 1.2] Given a toral $L$-parameter $\phi_{T}:W_{\mathbb{Q}_{p}}\rightarrow\phantom{}^{L}T(\overline{\mathbb{F}}_{\ell})$, we say that $\phi_{T}$ is generic if, for all $\alpha\in\mathbb{X}_{*}(T_{\overline{\mathbb{Q}}_{p}})/\Gamma$ corresponding to a $\Gamma$-orbit of coroots, we have that the complex $R\Gamma(W_{\mathbb{Q}_{p}},\alpha\circ\phi_{T})$ is trivial. Similarly, we say that $\mathfrak{m}$ is generic if the $L$-parameter $\phi_{\mathfrak{m}}^{T}$ is a generic toral parameter, where we note that this only depends on $\phi_{\mathfrak{m}}$ up to $\hat{G}-$conjugacy.
If $G=\mathrm{GL}_{n}$ then this coincides with the notion of decomposed generic considered in [7, Definition I.9]. We set $d=\mathrm{dim}(\mathrm{Sh}(\mathbf{G},X)_{K})$. Motivated by [7, Theorem 1.1] and [8, Theorem 1.1], we make the following conjecture.
Conjecture 1.2.
Let $(\mathbf{G},X)$ be a Shimura datum such that $G=\mathbf{G}_{\mathbb{Q}_{p}}$ is unramified and $K=K_{p}K^{p}$ is a sufficiently small level with $K_{p}=K_{p}^{\mathrm{hs}}$ hyperspecial. If $\mathfrak{m}\subset H_{K_{p}^{\mathrm{hs}}}$ is a generic maximal ideal then the cohomology of $R\Gamma(\mathrm{Sh}(\mathbf{G},X)_{K,\overline{E}},\overline{\mathbb{F}}_{\ell})_{\mathfrak{m}}$ (resp. $R\Gamma_{c}(\mathrm{Sh}(\mathbf{G},X)_{K,\overline{E}},\overline{\mathbb{F}}_{\ell})_{\mathfrak{m}}$) is concentrated in degrees $d\leq i\leq 2d$ (resp. $0\leq i\leq d$).
We first recall the motivating situation of Caraiani-Scholze [7, 8]. Let $F/\mathbb{Q}$ be a CM field, and let $(B,*,V,\langle\cdot,\cdot\rangle)$ be a PEL datum with $B$ a central simple $F$-algebra and $V$ a non-zero finite type left $B$-module. Let $(\mathbf{G},X)$ denote the Shimura datum attached to it, where $\mathbf{G}$ is the connected reductive group over $\mathbb{Q}$ defined by the automorphisms of $V$ preserving the choice of $B$-linear pairing $\langle\cdot,\cdot\rangle$ up to similitude. We have the following result.
Theorem 1.3.
[7, 8, 37, 54] Assume that $(\mathbf{G},X)$ is a PEL type Shimura datum of type $A$. If the prime $p$ splits completely in $F$ then Conjecture 1.2 is true.
Remark 1.4.
Koshikawa proved this under the assumption that $B=F$ and $V=F^{2n}$, and the global unitary group $\mathbf{G}$ is quasi-split, as well as in the case when $p$ is split in $F$ and the Shimura variety is compact. These additional assumptions were removed in the PhD thesis of Santos [54].
Remark 1.5.
Caraiani-Scholze actually proved a slightly different result. More precisely, let $S$ be a set of finite places not containing $p$ such that $\mathbf{G}$ is unramified and $K^{p}$ is hyperspecial away from $S$. Consider a maximal ideal $\mathfrak{m}\subset\mathbb{T}^{S}$ in the prime-to-$S$ Hecke algebra. This defines a global Galois representation $\rho_{\mathfrak{m}}$, and we can consider the restriction $\rho_{\mathfrak{m}}|_{\mathrm{Gal}(\bar{E}_{v}/E_{v})}$ which is unramified. If the associated toral $L$-parameter is generic as in Definition 1.1, then Caraiani-Scholze show (under some technical global assumptions) that the localization at $\mathfrak{m}^{p}\subset\mathbb{T}^{S\cup\{p\}}$ is concentrated in the same degrees as Conjecture 1.2.
Remark 1.6.
In the case of Harris-Taylor Shimura varieties, there is also work of Boyer [4], which describes the localization of the torsion cohomology at non-generic maximal ideals.
Remark 1.7.
We believe that Conjecture 1.2 is true under the weaker hypothesis that $H^{2}(W_{\mathbb{Q}_{p}},\alpha\circ\phi_{T})$ is trivial for all $\Gamma$-orbits of coroots $\alpha$, as is shown in [8, 54, 37] in their particular case. However, the theory of geometric Eisenstein series which we will invoke in this paper becomes more complicated in this case (See the discussion around [24, Conjecture 1.24]), and so a proof of this Theorem using our methods would require more deeply understanding geometric Eisenstein series when this assumption is dropped (cf. Remark 6.8). In certain cases however the theory of geometric Eisenstein series becomes unnecessary, and in this case an argument using compatibility (See Assumption 4.5) of Fargues-Scholze with some form of local Langlands is sufficient, as in the style of the arguments of [37, 54] (See Remark 4.28 for details).
Caraiani-Scholze [7, 8] proved their results under some technical restrictions of a global nature, which Koshikawa [37] was able to remove by using compatibility of the Fargues-Scholze local Langlands correspondence with the semi-simplification of the Harris-Taylor correspondence for $\mathrm{GL}_{n}$. In the process, Koshikawa exhibited a much more flexible method for proving Theorem 1.3. The goal of the current paper is to expand the scope of Koshikawa’s technique, motivated by work of the first author on geometric Eisenstein series in the Fargues-Fontaine setting [24]. We then carry the strategy out in some particular cases using work on local-global compatibility of the Fargues-Scholze local Langlands correspondence beyond the case of $\mathrm{GL}_{n}$, as studied in [24, 47]. Our main result is the following.
Theorem 1.8.
(Theorem 5.2) Suppose $(\mathbf{G},X)$ is a PEL datum of type AC satisfying Assumption 1.13 such that $\mathbf{G}_{\mathbb{Q}_{p}}$ is a product of groups in Table (1) with $p$ and $\ell$ satisfying the corresponding conditions. Let $\mathfrak{m}\subset H_{K_{p}^{\mathrm{hs}}}$ be a generic maximal ideal. Then, for a level $K=K^{p}K_{p}^{\mathrm{hs}}\subset\mathbf{G}(\mathbb{A}_{f})$, the cohomology of $R\Gamma(\mathrm{Sh}(\mathbf{G},X)_{K,\overline{E}},\overline{\mathbb{F}}_{\ell})_{\mathfrak{m}}$ (resp. $R\Gamma_{c}(\mathrm{Sh}(\mathbf{G},X)_{K,\overline{E}},\overline{\mathbb{F}}_{\ell})_{\mathfrak{m}}$) is concentrated in degrees $d\leq i\leq 2d$ (resp. $0\leq i\leq d$).
Remark 1.9.
This notably allows one to relax the assumption in [7, 8, 37, 54] that the prime $p$ splits in $F$ in the setting of Theorem 1.3, answering a question of Caraiani.
Here is the table summarizing our local constraints:
$G$ Constraint on $G$ $\ell$ $p$
$\mathrm{Res}_{L/\mathbb{Q}_{p}}(\mathrm{GL}_{n})$ $L/\mathbb{Q}_{p}$ unramified $(\ell,[L:\mathbb{Q}_{p}])=1$
$\mathrm{Res}_{L/\mathbb{Q}_{p}}(\mathrm{GSp}_{4})$ $L=\mathbb{Q}_{p}$
$L/\mathbb{Q}_{p}$ unramified $(\ell,[L:\mathbb{Q}_{p}]=1$ $p\neq 2$
$\mathrm{Res}_{L/\mathbb{Q}_{p}}(\mathrm{GU}_{2})$ $L/\mathbb{Q}_{p}$ unramified $(\ell,2[L:\mathbb{Q}_{p}])=1$
$G=\mathrm{U}_{n}(L/\mathbb{Q}_{p})$ $n$ odd $L$ unramified $\ell\neq 2$
$G=\mathrm{GU}_{n}(L/\mathbb{Q}_{p})$ $n$ odd $L$ unramified $\ell\neq 2$
$G(\mathrm{SL}_{2,L})$ $L/\mathbb{Q}_{p}$ unramified $(\ell,[L:\mathbb{Q}_{p}])=1$
$G(\mathrm{Sp}_{4,L})$ $L/\mathbb{Q}_{p}$ unramified, $L\neq\mathbb{Q}_{p}$ $(\ell,[L:\mathbb{Q}_{p}])=1$ $p\neq 2$
The groups $G(\mathrm{SL}_{2,L})$ and $G(\mathrm{Sp}_{4,L})$ are the similitude subgroup of $\mathrm{Res}_{L/\mathbb{Q}_{p}}(\mathrm{GL}_{2})$ (resp. $\mathrm{Res}_{L/\mathbb{Q}_{p}}(\mathrm{GSp}_{4})$), i.e. the subgroup of elements such that the similitude factor lies in $\mathbb{Q}_{p}$. We will recall the definition of these groups in §4.4.
We can also easily deduce a result for some abelian type Shimura varieties, such as Hilbert modular varieties, from the above result, which recovers results similar to that of Caraiani-Tamiozzo [9, Theorem B] (See Corollary 5.7).
Corollary 1.10.
(Corollary 5.6) Suppose $(\mathbf{G},X)$ is an abelian-type Shimura datum which has an associated PEL-type datum $(\mathbf{G}_{1},X_{1})$ of type AC satisfying Assumption 1.13 and such that $\mathbf{G}^{\mathrm{der}}\simeq\mathbf{G}_{1}^{\mathrm{der}}$, $G_{1}:=\mathbf{G}_{1,\mathbb{Q}_{p}}$ is unramified and $G:=\mathbf{G}_{\mathbb{Q}_{p}}$ is a product of groups as in Table (1) with $p$ and $\ell$ satisfying the corresponding conditions. We write $G:=\mathbf{G}_{\mathbb{Q}_{p}}$ and $G_{1}:=\mathbf{G}_{1,\mathbb{Q}_{p}}$. We assume that $\mathfrak{m}\subset H_{K_{p}^{\mathrm{hs}}}$ is a generic maximal ideal. Then, for a level $K=K^{p}K_{p}^{\mathrm{hs}}\subset\mathbf{G}(\mathbb{A}_{f})$, the cohomology $R\Gamma(\mathrm{Sh}(\mathbf{G},X)_{K,\overline{E}},\overline{\mathbb{F}}_{\ell})_{\mathfrak{m}}$ (resp. $R\Gamma_{c}(\mathrm{Sh}(\mathbf{G},X)_{K,\overline{E}},\overline{\mathbb{F}}_{\ell})_{\mathfrak{m}}$) is concentrated in degrees $d\leq i\leq 2d$ (resp. $0\leq i\leq d$).
We now explain how we are able to establish these results, as well as several other interesting results on the perverse $t$-exactness of Hecke operators on $\mathrm{Bun}_{G}$ (Theorem 1.11) and the splitting of Mantovan’s filtration on the cohomology of the Shimura variety (Theorem 1.17) along the way.
1.0.2. Proof Sketch of the Main Theorem
One of the basic ingredients used in verifying Conjecture 1.2 in all known instances is the perspective on Mantovan’s product formula provided by the Hodge-Tate period morphism. To explain this, we let $\mu\in\mathbb{X}_{*}(T_{\overline{\mathbb{Q}}_{p}})^{+}$ denote the minuscule geometric dominant cocharacter of $G$ determined by the inverse of the Hodge cocharacter of $X$ and an isomorphism $j:\mathbb{C}\simeq\overline{\mathbb{Q}}_{p}$ which we fix from now on. We consider the Kottwitz set $B(G)$, and with it the subset $B(G,\mu)\subset B(G)$ of $\mu$-admissible elements. Let $\mathfrak{p}|p$ be the prime dividing $p$ in the reflex field $E$ induced by the embedding $\overline{\mathbb{Q}}\rightarrow\overline{\mathbb{Q}}_{p}$ given by the isomorphism $j$. We let $E_{\mathfrak{p}}$ be the completion at $\mathfrak{p}$, $C:=\hat{\overline{E}}_{\mathfrak{p}}$ be the completion of the algebraic closure, and $\breve{E}_{\mathfrak{p}}$ be the compositum of $E_{\mathfrak{p}}$ with $\breve{\mathbb{Q}}_{p}$. We recall that, attached to each element $b\in B(G,\mu)$, there exists a diamond
$\mathrm{Sht}(G,b,\mu)_{\infty}\rightarrow\mathop{\rm Spd}(\breve{E}_{\mathfrak{p}})$
parametrizing modifications
$\mathcal{E}_{b}\dashrightarrow\mathcal{E}_{0}$
of meromorphy $\mu$ between the $G$-bundle $\mathcal{E}_{b}$ corresponding to $b$ and the trivial $G$-bundle. This space has an action by $\underline{G(\mathbb{Q}_{p})}=\mathrm{Aut}(\mathcal{E}_{0})$ and $\underline{J_{b}(\mathbb{Q}_{p})}\subset\mathrm{Aut}(\mathcal{E}_{b})$, where $J_{b}$ is the $\sigma$-centralizer of $b$. This allows us to consider the quotients
$\mathrm{Sht}(G,b,\mu)_{\infty}/\underline{K_{p}}\rightarrow\mathop{\rm Spd}(\breve{E}_{\mathfrak{p}})$
for varying compact open subgroups $K_{p}\subset G(\mathbb{Q}_{p})$. In certain cases, these quotients are representable by rigid analytic varieties called local Shimura varieties, but they are always representable as diamonds. We can consider the compactly supported cohomology
$R\Gamma_{c}(\mathrm{Sht}(G,b,\mu)_{\infty,C}/\underline{K_{p}^{\mathrm{hs}}},\overline{\mathbb{F}}_{\ell})$
at hyperspecial level with torsion coefficients. This has an action of $W_{E_{\mathfrak{p}}}\times J_{b}(\mathbb{Q}_{p})\times H_{K_{p}^{\mathrm{hs}}}$ [^1] It is clear that the space has an action of the inertia group $I_{E_{\mathfrak{p}}}$ of $E_{\mathfrak{p}}$ from the definition; however, the space $\mathrm{Sht}(G,b,\mu)_{\infty}\rightarrow\mathop{\rm Spd}(\breve{E}_{\mathfrak{p}})$ also has a canonical (non-effective) Frobenius descent datum giving a full action of $W_{E_{\mathfrak{p}}}$ on the cohomology (See [16, Section IX.3]). Now, the Mantovan product formula tells us that if we look at $R\Gamma(\mathrm{Sh}(\mathbf{G},X)_{K,\overline{E}},\overline{\mathbb{F}}_{\ell})$ then this should always admit a filtration in the derived category whose graded pieces are
$R\Gamma_{c}(\mathrm{Sht}(G,b,\mu)_{\infty,C}/\underline{K_{p}^{\mathrm{hs}}},\overline{\mathbb{F}}_{\ell}(d_{b}))[2d_{b}]\otimes^{\mathbb{L}}_{\mathcal{H}(J_{b})}R\Gamma(\mathrm{Ig}^{b},\overline{\mathbb{F}}_{\ell})$
for varying $b\in B(G,\mu)$, where the objects are as follows.
  1. (1)
    $\mathrm{Ig}^{b}$ is the perfect Igusa variety attached to an element $b\in B(G,\mu)$ in the $\mu$-admissible locus inside $B(G)$ and $d_{b}:=\mathrm{dim}(\mathrm{Ig}^{b})=\langle 2\rho_{G},\nu_{b}\rangle$, where $\rho_{G}$ is the half sum of all positive roots and $\nu_{b}$ is the slope cocharacter of $b$.
  2. (2)
    $\mathcal{H}(J_{b}):=C^{\infty}_{c}(J_{b}(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})$ is the usual smooth Hecke algebra.
  3. (3)
    $\overline{\mathbb{F}}_{\ell}(d_{b})$ is the sheaf on $\mathrm{Sht}(G,b,\mu)_{\infty,C}/\underline{K_{p}^{\mathrm{hs}}}$ with trivial Weil group action and $J_{b}(\mathbb{Q}_{p})$ action as defined in [37, Lemma 7.4].
Such a filtration should always exist, but is not currently proven in general. In the case that the Shimura datum $(\mathbf{G},X)$ is PEL of type AC, a modern proof of this result can be found in [37, Theorem 7.1].
This filtration on the complex $R\Gamma(\mathrm{Sh}(\mathbf{G},X)_{K,\overline{E}},\overline{\mathbb{F}}_{\ell})$ allows us to roughly split the verification of Conjecture 1.2 into two parts.
  1. (1)
    Controlling the cohomology of the shtuka spaces $R\Gamma_{c}(\mathrm{Sht}(G,b,\mu)_{\infty,C}/\underline{K_{p}^{\mathrm{hs}}},\overline{\mathbb{F}}_{\ell}(d_{b}))_{\mathfrak{m}}$.
  2. (2)
    Controlling the cohomology of the Igusa varieties $R\Gamma(\mathrm{Ig}^{b},\overline{\mathbb{F}}_{\ell})$.
We first discuss point (1). One of the key observations underlying Koshikawa’s method was that the cohomology of the space $\mathrm{Sht}(G,b,\mu)_{\infty}$ computes the action of a Hecke operator $T_{\mu}$ corresponding to $\mu$ on $\mathrm{Bun}_{G}$ the moduli stack of $G$-bundles of the Fargues-Fontaine curve. The Hecke operators commute with the action of the excursion algebra on $\mathrm{Bun}_{G}$, and the action of the excursion algebra on a smooth irreducible representation $\rho$, viewed as a sheaf on $\mathrm{Bun}_{G}$, determines the Fargues-Scholze parameter of $\rho$. It follows that $R\Gamma_{c}(\mathrm{Sht}(G,b,\mu)_{\infty,C}/\underline{K_{p}^{\mathrm{hs}}},\overline{\mathbb{F}}_{\ell}(d_{b}))_{\mathfrak{m}}$ as a complex of smooth $J_{b}(\mathbb{Q}_{p})$-modules will have irreducible constituents $\rho$ with Fargues-Scholze parameter $\phi_{\rho}^{\mathrm{FS}}$ equal to $\phi_{\mathfrak{m}}$ as conjugacy classes of parameters. When $\mathbf{G}_{\mathbb{Q}_{p}}=G$ is a product of $\mathrm{GL}_{n}$s, as in Theorem 1.3 (by the assumption that $p$ splits in $F$), it follows from the work of Hansen-Kaletha-Weinstein [31, Theorem 1.0.3] that the Fargues-Scholze correspondence for $J_{b}(\mathbb{Q}_{p})$ with rational coefficients agrees with the semi-simplification of the Harris-Taylor correspondence, where we recall that $J_{b}$ is a product of inner forms of $\mathrm{GL}_{n}$s in this case. In particular, using that $\mathfrak{m}$ is generic, it follows that $\phi_{\rho}^{\mathrm{FS}}=\phi_{\mathfrak{m}}$ must lift to a $\overline{\mathbb{Z}}_{\ell}$ parameter which is also generic in the analogous sense, and the condition of generic implies that the lift cannot come from the semi-simplification of a parameter with non-trivial monodromy. Using this, one can deduce that such a $\rho$ only exists if the group $J_{b}$ is quasi-split. In this particular case ($G$ is a product of $\mathrm{GL}_{n}$s), this can only happen if $b\in B(G,\mu)$ corresponds to the ordinary element (cf. Remark 4.28).
This argument of Koshikawa was formalized and generalized further in work of the first author [24]. In particular, it was noted that, for a general quasi-split $G$ and $\mathfrak{m}$ generic, the cohomology $R\Gamma_{c}(\mathrm{Sht}(G,b,\mu)_{\infty,C}/\underline{K_{p}^{\mathrm{hs}}},\overline{\mathbb{F}}_{\ell}(d_{b}))_{\mathfrak{m}}$ will only be non-trivial if $b\in B(G,\mu)_{\mathrm{un}}:=B(G)_{\mathrm{un}}\cap B(G,\mu)$, where $B(G)_{\mathrm{un}}$ is the set of elements lying in the image of the map $B(T)\rightarrow B(G)$, assuming that the Fargues-Scholze local Langlands correspondence has certain expected properties (Assumption 4.5). These unramified elements will be precisely the elements for which $J_{b}$ is quasi-split. The set $B(G,\mu)_{\mathrm{un}}$ corresponds to Weyl group orbits of weights in the representation $V_{\mu}$ of $\hat{G}$ restricted to $\hat{G}^{\Gamma}$. In particular, if $G$ is split then, since $\mu$ is minuscule, $B(G,\mu)_{\mathrm{un}}$ consists of only one element, corresponding to the unique Weyl group orbit of the highest weight. This is the situation occurring in the previous paragraph. Moreover, the contribution of the cohomology of this shtuka space is easily understood, and the problem completely reduces to controlling the cohomology of $\mathrm{Ig}^{b}$ when $b\in B(G,\mu)_{\mathrm{un}}$ is the $\mu$-ordinary element. However, if $G$ is not split then the restriction of $V_{\mu}$ to $\hat{G}^{\Gamma}$ may have multiple Weyl group orbits of weights. In particular, one needs to control the cohomology groups
$R\Gamma_{c}(\mathrm{Sht}(G,b,\mu)_{\infty,C}/\underline{K^{\mathrm{hs}}_{p}},\overline{\mathbb{F}}_{\ell}(d_{b}))_{\mathfrak{m}}$
for all possible $b\in B(G,\mu)_{\mathrm{un}}$. This makes the situation much more complicated; in fact, for non-split $G$, the basic element could be unramified, and in this case the Igusa variety $\mathrm{Ig}^{b}$ is just a profinite set, hence the problem of torsion vanishing for the contribution of the basic locus is completely reduced to controlling the generic part of the torsion cohomology of the local shtuka space attached to the basic element.
Such control of the cohomology of shtuka spaces with torsion coefficients for these more general situations was attained in [24]. In order to understand this, it is helpful to move away from the language of isotypic parts of shtuka spaces and consider the action of Hecke operators on $\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$, the derived category of étale $\overline{\mathbb{F}}_{\ell}$-sheaves on $\mathrm{Bun}_{G}$. Since we are interested in cohomology localized at a generic maximal ideal $\mathfrak{m}$, Hansen constructs in appendix A a full-subcategory $\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi_{\mathfrak{m}}}\subset\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$ together with an idempotent localization map $(-)_{\phi_{\mathfrak{m}}}:\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})\rightarrow\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi_{\mathfrak{m}}}$ such that, on smooth irreducible representations, the localization map is either an isomorphism or 0 depending on if the representation has Fargues-Scholze parameter conjugate to $\phi_{\mathfrak{m}}$ or not (Lemma 4.3 (1)). We let $\mathrm{D}^{\mathrm{ULA}}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$ denote the full subcategory of ULA objects, where we recall by [16, Theorem V.7.1], that this is equivalent to insisting that the restrictions to all the HN-strata indexed by $b\in B(G)$ are valued in the full subcategories $\mathrm{D}^{\mathrm{adm}}(J_{b}(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})$ of admissible complexes (i.e the invariants under all open pro-$p$ $K\subset J_{b}(\mathbb{Q}_{p})$ is a perfect complex). Using the results of [24], we show (Corollary 4.29) that, under some regularity and genericity hypothesis on $\mathfrak{m}$, that one has a direct sum decomposition:
(2) $\mathrm{D}^{\mathrm{ULA}}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi_{\mathfrak{m}}}\simeq\bigoplus_{b\in B(G)_{\mathrm{un}}}\mathrm{D}^{\mathrm{adm}}(J_{b}(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})_{\phi_{\mathfrak{m}}}.$
More precisely, we show that the !! and * push-forwards with respect to the inclusion of HN-strata agree on this sub-category, and so the excision semi-orthogonal decomposition splits on $\mathrm{D}^{\mathrm{ULA}}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi_{\mathfrak{m}}}$. This decomposition is a refinement of the fact mentioned above that only the shtuka spaces corresponding to the unramified elements $b\in B(G,\mu)_{\mathrm{un}}$ can contribute to the generic localization of the cohomology of the Shimura variety. The desired control of the shtuka spaces is now in turn encoded in understanding how Hecke operators interact with a perverse $t$-structure on $\mathrm{Bun}_{G}$ after restricting to the localized category $\mathrm{D}^{\mathrm{ULA}}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi_{\mathfrak{m}}}$.
We recall $\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$ has an action by Hecke operators. In particular, for each geometric dominant cocharacter $\mu$, we have a correspondence
$\mathrm{Hck}_{G,\leq\mu}$$\mathrm{Bun}_{G}$$\mathrm{Bun}_{G}\times\mathop{\rm Spd}(C),$$\scriptstyle{h_{\mu}^{\leftarrow}}$$\scriptstyle{h_{\mu}^{\rightarrow}}$
where $\mathrm{Hck}_{G,\leq\mu}$ is the stack parametrizing modifications $\mathcal{E}_{1}\dashrightarrow\mathcal{E}_{2}$ of a pair of $G$-bundles with meromorphy bounded by $\mu$ at the closed Cartier divisor defined by the fixed untilt given by $C$, and $h_{\mu}^{\rightarrow}$ (resp. $h_{\mu}^{\leftarrow}$) remembers $\mathcal{E}_{1}$ (resp. $\mathcal{E}_{2}$). We define
$T_{\mu}:\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})\rightarrow\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})^{BW_{E_{\mu}}}$
$A\mapsto h_{\mu*}^{\rightarrow}(h_{\mu}^{\leftarrow*}(A)\otimes^{\mathbb{L}}\mathcal{S}_{\mu})$
[…]
$T_{\mu}:\mathrm{D}^{\mathrm{ULA}}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi_{\mathfrak{m}}}\rightarrow\mathrm{D}^{\mathrm{ULA}}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})^{BW_{E_{\mu}}}_{\phi_{\mathfrak{m}}}$
on the localized category (See Lemma 4.3 (2)).
We are almost ready to state the result on Hecke operators we will need. To do this, we recall that $\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$ has a natural perverse $t$-structure, which can be defined as follows. The $v$-stack $\mathrm{Bun}_{G}$ is cohomologically smooth of $\ell$-dimension 0. Moreover, it has a locally closed HN-stratification $j_{b}:\mathrm{Bun}_{G}^{b}\hookrightarrow\mathrm{Bun}_{G}$, where each of the strata $\mathrm{Bun}_{G}^{b}$ are isomorphic to $\mathop{\rm Spd}(C)/\mathcal{J}_{b}]$, which is cohomologically smooth of $\ell$-dimension $d_{b}=-\mathrm{dim}(\mathrm{Ig}^{b})$. Therefore, we can define a perverse $t$-structure $\phantom{}{}^{\mathrm{p}}\mathrm{D}^{\geq 0}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$ (resp. $\phantom{}{}^{\mathrm{p}}\mathrm{D}^{\leq 0}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$) on $\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$ given by insisting that the !! (resp. *) restrictions to $\mathrm{Bun}_{G}^{b}$ are concentrated in degrees $\geq\langle 2\rho_{G},\nu_{b}\rangle$ (resp. $\leq\langle 2\rho_{G},\nu_{b}\rangle$). The key result that follows from the work of [24] and various compatibility results for the Fargues-Scholze correspondence is as follows.
Theorem 1.11.
(Corollary 4.27) Let $\mu$ be a minuscule geometric dominant cocharacter and $G$ a product of groups satisfying the conditions of Table (1) with $p$ and $\ell$ satisfying the corresponding conditions. Let $\mathfrak{m}$ be a generic maximal ideal with associated semi-simple $L$-parameter $\phi_{\mathfrak{m}}$. The restriction of the Hecke operator
$j_{1}^{*}T_{\mu}:\mathrm{D}^{\mathrm{ULA}}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi_{\mathfrak{m}}}\rightarrow\mathrm{D}^{\mathrm{adm}}(G(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})^{BW_{E_{\mu}}}_{\phi_{\mathfrak{m}}}$
to the neutral strata $j_{1}:[\ast/\underline{G(\mathbb{Q}_{p})}]\simeq\mathrm{Bun}_{G}^{1}\hookrightarrow\mathrm{Bun}_{G}$ is perverse $t$-exact (in fact much more is true (See Corollary 4.27)), where we note that the perverse $t$-structure on $\mathrm{D}(\mathrm{Bun}_{G}^{1},\overline{\mathbb{F}}_{\ell})\simeq\mathrm{D}(G(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})$ coincides with the usual $t$-structure.
Remark 1.12.
Assuming the Fargues-Scholze correspondence for $G$ behaves as expected with rational coefficients, the analysis in [24] allows one to verify this for any $\mu$ after imposing some additional conditions on the toral parameter $\phi_{T}^{\mathfrak{m}}$ attached to the maximal ideal $\mathfrak{m}$ such as being regular and satisfying Definition 4.16 for suitably chosen $\mu$, where the latter conditions are often implied by the genericity in the relevant use cases. We emphasize however that the above Theorem should always be true just under the condition that $\mathfrak{m}$ is generic, as formally stated in Conjecture 1.2.
These local torsion vanishing results would allow us to prove Conjecture 1.2 in several new cases if one could get control over the Igusa varieties $\mathrm{Ig}^{b}$. In Koshikawa’s argument, this is done by using a semi-perversity result proven by Caraiani-Scholze [8, Theorem 4.6.1], which was further generalized in work of Santos [54]. Roughly speaking, we want to show that $R\Gamma(\mathrm{Ig}^{b},\overline{\mathbb{F}}_{\ell})$ is concentrated in degrees $\geq d_{b}$, so that the complex of $J_{b}(\mathbb{Q}_{p})$-representations $R\Gamma(\mathrm{Ig}^{b},\overline{\mathbb{F}}_{\ell})$ defines the stalk of a semi-perverse sheaf on $\mathrm{Bun}_{G}$ at $b\in B(G)$, to which we can apply the previous result. In the case that the Shimura vareities $\mathrm{Sh}(\mathbf{G},X)_{K}$ are compact, there is a simpler way of seeing this. In particular, $\mathrm{Ig}^{b}$ is known to be a perfect affine scheme in this case, and so the desired semi-perversity just follows by applying Artin vanishing and then using Poincaré duality on the global Shimura variety. It turns out that this style of argument can be made to work even in the non-compact case. In [7, 8, 37, 54], the non-compactly supported cohomology $R\Gamma(\mathrm{Sh}(\mathbf{G},X)_{K},\overline{\mathbb{F}}_{\ell})_{\mathfrak{m}}$ is studied together with its filtration involving $R\Gamma(\mathrm{Ig}^{b},\overline{\mathbb{F}}_{\ell})$ coming from Mantovan’s formula, and shown to be concentrated in degrees $\geq d$. However, one could also study the compactly supported cohomology $R\Gamma_{c}(\mathrm{Sh}(\mathbf{G},X)_{K},\overline{\mathbb{F}}_{\ell})_{\mathfrak{m}}$ and show that it is concentrated in degrees $\leq d$, à la Poincaré duality. To do this, we recall [8, Section 3.3] that, in the non-compact case, the perfect scheme $\mathrm{Ig}^{b}$ is not affine, but it admits a partial minimal compactification $g_{b}:\mathrm{Ig}^{b}\hookrightarrow\mathrm{Ig}^{b,*}$ which is affine, as proven in this more general setting of PEL type AC by Santos [54]. We define
$V_{b}:=R\Gamma_{c-\partial}(\mathrm{Ig}^{b},\overline{\mathbb{F}}_{\ell}):=R\Gamma(\mathrm{Ig}^{b,*},g_{b!}(\overline{\mathbb{F}}_{\ell}))$
the partially compactly supported cohomology, which is supported in degrees $\leq d_{b}$ by Artin-vanishing (Proposition 3.6). Now, for $K\subset\mathbf{G}(\mathbb{A}_{f})$ a sufficiently small open compact, we define $\mathcal{S}(\mathbf{G},X)_{K}:=(\mathrm{Sh}(\mathbf{G},X)_{K}\otimes_{E}E_{\mathfrak{p}})^{\mathrm{ad}}$ to be the adic space over $\mathop{\rm Spa}(E_{\mathfrak{p}})$ attached to the Shimura variety. We can define the infinite level perfectoid Shimura varieties $\mathcal{S}(\mathbf{G},X)_{K^{p}}$ by taking the inverse limit of $\mathcal{S}(\mathbf{G},X)_{K^{p}K_{p}}$ as $K_{p}\rightarrow\{1\}$ in the category of diamonds. The base-change $\mathcal{S}(\mathbf{G},X)_{K^{p},C}$ is representable by a perfectoid space if $(\mathbf{G},X)$ is of pre-abelian type, and in general it is diamond. By the results of [57, 28], we have a Hodge-Tate period map
$\pi_{\mathrm{HT}}:[\mathcal{S}(\mathbf{G},X)_{K^{p},C}/\underline{G(\mathbb{Q}_{p})}]\rightarrow[\mathcal{F}\ell_{G,\mu^{-1}}^{\Diamond}/\underline{G(\mathbb{Q}_{p})}]$
recording the Hodge-Tate filtration on the abelian varieties with additional structure that $\mathcal{S}(\mathbf{G},X)_{K^{p},C}$ parametrizes. Here $\mathcal{F}\ell_{G,\mu^{-1}}^{:}=(G_{C}/P_{\mu^{-1}})^{\mathrm{ad}}$ is the adic flag variety attached to the parabolic in $G_{C}$ given by a dominant inverse of $\mu$ and the dynamical method (i.e we look at the set $g\in G_{C}$ such that $\mathop{\rm lim}\nolimits_{t\rightarrow 0}\mathrm{ad}(\mu(t))g$ exists, see also Remark 3.13 for a discussion of the sign conventions we use) and $\mathcal{F}\ell_{G,\mu^{-1}}^{\Diamond}$ denotes the associated diamond, and we have passed to the associated $v$-stack quotients by the $G(\mathbb{Q}_{p})$-action. We recall that the flag variety $\mathcal{F}\ell_{G,\mu^{-1}}^{\Diamond}/\underline{G(\mathbb{Q}_{p})}]$ admits a locally closed stratification $i_{b}:[\mathcal{F}\ell_{G,\mu^{-1}}^{b,\Diamond}/\underline{G(\mathbb{Q}_{p})}]\hookrightarrow[\mathcal{F}\ell_{G,\mu^{-1}}^{\Diamond}/\underline{G(\mathbb{Q}_{p})}]$ indexed by $b\in B(G,\mu)$, given by pulling the HN-stratification along the natural map $h^{\leftarrow}:[\mathcal{F}\ell_{G,\mu^{-1}}^{\Diamond}/\underline{G(\mathbb{Q}_{p})}]\rightarrow\mathrm{Bun}_{G}$. We will now impose the following very mild assumption in what follows.
Assumption 1.13.
Write $\partial\mathrm{Ig}^{b,*}\subset\mathrm{Ig}^{b,*}$ for the closed complement of $\mathrm{Ig}^{b}$ in $\mathrm{Ig}^{b,*}$. We assume that $(\mathbf{G},X)$ is a PEL datum of type AC such that, for all $b\in B(G,\mu)$, the perfect scheme $\partial\mathrm{Ig}^{b,*}$ is empty or has codimension in $\mathrm{Ig}^{b,*}$ greater than or equal to $2$.
Remark 1.14.
If $\mathbf{G}^{\mathrm{ad}}$ is simple then it is easy to show that this assumption will be satisfied if $\mathrm{dim}(\mathcal{S}(\mathbf{G},X)_{K^{p},C})\geq 2$, by using that the boundary of the partially minimally compactified Igusa varieties is expressible as the Igusa varieties of Shimura varieties attached to Levis of $\mathbb{Q}$-rational parabolics of $\mathbf{G}$, as we will explain in §2.2.2. Moreover, if $\mathcal{S}(\mathbf{G},X)_{K^{p},C}$ is compact then it is automatic that $\partial\mathrm{Ig}^{b,*}$ is empty. Therefore, if $\mathbf{G}^{\mathrm{ad}}$ is simple, this is excluding the cases where $\mathrm{dim}(\mathcal{S}(\mathbf{G},X)_{K^{p},C})=1$ and $\mathcal{S}(\mathbf{G},X)_{K^{p},C}$ is non-compact. There are two possibilities; either $(\mathbf{G},X)$ is the Shimura datum attached to the modular curve, or it is the Shimura datum attached to the unitary Shimura curve (See [68, Proposition 1.9]). In the latter case, we have that the connected components are given by modular curves. In these cases, the results of [37] are sufficient to prove Conjecture 1.2.
Now, assuming this, one can show that the stalk of $R\pi_{\mathrm{HT}!}(\overline{\mathbb{F}}_{\ell})$ at a geometric point $x:\mathop{\rm Spa}(C,C^{+})\rightarrow\mathcal{F}\ell_{G,\mu^{-1}}^{\Diamond}$ which lies in the adic Newton strata $\mathcal{F}\ell_{G,\mu^{-1}}^{b,\Diamond}$ is given by $V_{b}$. Moreover, if we write $h_{b}^{\leftarrow}:[\mathcal{F}\ell_{G,\mu^{-1}}^{b,\Diamond}/\underline{G(\mathbb{Q}_{p})}]\rightarrow[\mathop{\rm Spd}(C)/\mathcal{J}_{b}]\simeq\mathrm{Bun}_{G}^{b}$ for the pullback of $h^{\leftarrow}$ to $\mathrm{Bun}_{G}^{b}$ then one can deduce that the complex $i_{b!}i_{b}^{*}R\pi_{\mathrm{HT}!}(\overline{\mathbb{F}}_{\ell})$ is isomorphic to $h^{\leftarrow*}j_{b!}(V_{b})$. Therefore, by excision, we deduce that the complex of $G(\mathbb{Q}_{p})\times W_{E_{\mathfrak{p}}}$-representations
$h_{*}^{\rightarrow}R\pi_{\mathrm{HT}!}(\overline{\mathbb{F}}_{\ell})\simeq R\Gamma_{c}(\mathcal{S}(\mathbf{G},X)_{K^{p},C},\overline{\mathbb{F}}_{\ell})$ $\simeq\mathop{\rm colim}\nolimits_{K_{p}\rightarrow\{1\}}R\Gamma_{c}(\mathcal{S}(\mathbf{G},X)_{K^{p}K_{p},C},\overline{\mathbb{F}}_{\ell})$
$\simeq\mathop{\rm colim}\nolimits_{K_{p}\rightarrow\{1\}}R\Gamma_{c}(\mathrm{Sh}(\mathbf{G},X)_{K^{p}K_{p},C},\overline{\mathbb{F}}_{\ell})$
has a filtration with graded pieces isomorphic to $Rh^{\rightarrow}_{*}h^{\leftarrow*}j_{b!}(V_{b})$ for varying $b\in B(G,\mu)$, where $h^{\rightarrow}:[\mathcal{F}\ell_{G,\mu^{-1}}^{\Diamond}/\underline{G(\mathbb{Q}_{p})}]\rightarrow[\mathop{\rm Spd}(C)/\underline{G(\mathbb{Q}_{p})}]$ is the structure map quotiented by $G(\mathbb{Q}_{p})$, which we note is proper. Here the second isomorphism follows since taking compactly supported cohomology respects taking limits of spaces, and the third isomorphism is a standard comparison result due to Huber [32, Theorem 3.5.13].
Now, via the Bialynicki-Birula isomorphism, the flag variety $\mathcal{F}\ell_{G,\mu^{-1}}^{\Diamond}/\underline{G(\mathbb{Q}_{p})}]$ identifies with an open substack of $\mathrm{Hck}_{G,\leq\mu}$ for the fixed minuscule $\mu$. In particular, under this relationship the maps $h^{\rightarrow}_{\mu}$ and $h^{\leftarrow}_{\mu}$ identify with $h^{\rightarrow}$ and $h^{\leftarrow}$, and therefore we can relate the graded pieces of the excision filtration to Hecke operators. We write
$R\Gamma_{c}(G,b,\mu):=\mathop{\rm colim}\nolimits_{K_{p}\rightarrow\{1\}}R\Gamma_{c}(\mathrm{Sht}(G,b,\mu)/\underline{K_{p}},\overline{\mathbb{F}}_{\ell}(d_{b}))$
for the complex of $G(\mathbb{Q}_{p})\times J_{b}(\mathbb{Q}_{p})\times W_{E_{\mathfrak{p}}}$-modules defined by the compactly supported cohomology of this tower. Here $\overline{\mathbb{F}}_{\ell}(d_{b})$ is the sheaf with $J_{b}(\mathbb{Q}_{p})$-action defined as in [37, Lemma 7.4].
We deduce the following variant of the Mantovan product formula for the compactly supported cohomology.
Theorem 1.15.
Suppose that $(\mathbf{G},X)$ is a Shimura datum of PEL type AC satisfying Assumption 1.13 then the complex $R\Gamma_{c}(\mathcal{S}(\mathbf{G},X)_{K^{p},C},\overline{\mathbb{F}}_{\ell})$ has a filtration as a complex of $G(\mathbb{Q}_{p})$-representations with graded pieces isomorphic to $j_{1}^{*}T_{\mu}j_{b!}(V_{b})[-d]$. More specifically, the graded pieces are isomorphic to
$(R\Gamma_{c}(G,b,\mu)\otimes^{\mathbb{L}}_{\mathcal{H}(J_{b})}V_{b})[2d_{b}].$
as $G(\mathbb{Q}_{p})$-modules.
Remark 1.16.
When the Shimura variety is compact, so that $\mathrm{Ig}^{b,*}=\mathrm{Ig}^{b}$ and correspondingly that $R\Gamma_{c-\partial}(\mathrm{Ig}^{b},\overline{\mathbb{F}}_{\ell})\simeq R\Gamma(\mathrm{Ig}^{b},\overline{\mathbb{F}}_{\ell})$, and this recovers precisely [37, Theorem 7.1].
We now apply our localization functor $(-)_{\phi_{\mathfrak{m}}}:\mathrm{D}(\mathrm{Bun}_{G})\rightarrow\mathrm{D}(\mathrm{Bun}_{G})_{\phi_{\mathfrak{m}}}$ for a generic maximal ideal $\mathfrak{m}$ to get a complex $R\Gamma_{c}(\mathcal{S}(\mathbf{G},X)_{K^{p},C},\overline{\mathbb{F}}_{\ell})_{\phi_{\mathfrak{m}}}\in\mathrm{D}^{\mathrm{adm}}(G(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})_{\phi_{\mathfrak{m}}}$, which we view as a sheaf on $\mathrm{Bun}_{G}$ by !! extending along the neutral strata. Now we know, by the direct sum decomposition of $\mathrm{D}^{\mathrm{ULA}}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi_{\mathfrak{m}}}$ described in (2), that the natural map $j_{b!}(V_{b})\rightarrow Rj_{b*}(V_{b})$ is an isomorphism after applying $(-)_{\phi_{\mathfrak{m}}}$. Moreover, one only has interesting contributions coming from the unramified elements $B(G,\mu)_{\mathrm{un}}$. In particular, we can deduce the following Corollary of Theorem 1.15.
Theorem 1.17.
(Theorem 5.1) Suppose $(\mathbf{G},X)$ is a PEL datum of type AC satisfying Assumption 1.13 such that $\mathbf{G}_{\mathbb{Q}_{p}}$ is a product of groups as in Table (1) with $p$ and $\ell$ satisfying the corresponding conditions and let $\phi_{\mathfrak{m}}$ be a generic parameter of the form described in Theorem 1.11. Then the complex $R\Gamma_{c}(\mathcal{S}(\mathbf{G},X)_{K^{p},C},\overline{\mathbb{F}}_{\ell})_{\phi_{\mathfrak{m}}}$ breaks up as a direct sum
$\bigoplus_{b\in B(G,\mu)_{\mathrm{un}}}(R\Gamma_{c}(\mathrm{Sht}(G,b,\mu)_{\infty,C},\overline{\mathbb{F}}_{\ell}(d_{b}))_{\phi_{\mathfrak{m}}}\otimes^{\mathbb{L}}_{\mathcal{H}(J_{b})}R\Gamma_{c-\partial}(\mathrm{Ig}^{b},\overline{\mathbb{F}}_{\ell}))[2d_{b}]$
of $G(\mathbb{Q}_{p})$-modules.
Remark 1.18.
As we will explain more in §6.1, in the case that the unique basic element $b\in B(G,\mu)_{\mathrm{un}}$ is unramified, the contribution of the corresponding summand to middle degree cohomology serves as a generic fiber analogue of the description of the middle degree cohomology on the special fiber of the integral model at hyperspecial level, as provided in [67, Theorem 1.1.4].
It is now easy to see that we obtain from this Theorem 1.8, by combining Theorem 1.11 with the fact that $R\Gamma_{c-\partial}(\mathrm{Ig}^{b},\overline{\mathbb{F}}_{\ell})\in\mathrm{D}^{\leq d_{b}}(J_{b}(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})$, by Artin vanishing applied to the inverse limit defining the partial minimal compactification $\mathrm{Ig}^{b,*}$ (See Proposition 3.6 for details) using a Hochschild-Serre style argument.
Acknowledgements
We would like to thank Ana Caraiani, Jean-François Dat, David Hansen, Naoki Imai, Teruhisa Koshikawa, and Chris Skinner for helpful discussions pertaining to this work. Special thanks go to Mafalda Santos for sharing with us the results of her thesis, Peter Scholze for encouraging us to avoid working with the good reduction locus by directly describing the fibers of the Hodge-Tate period morphism, Matteo Tamiozzo for comments and corrections on an earlier draft, as well as suggestions for the arguments in §5.2, and Mingjia Zhang for very helpful discussions and filling in several gaps in the arguments used in §3, as well as several comments and corrections. We would also like the referees for carefully reading the paper and supplying several useful comments and corrections. This project was carried out while the second author was at the Max Planck Institute for Mathematics in Bonn and she thanks them for their hospitality and financial support.
Notation
  • Fix distinct primes $\ell\neq p$.
  • We write $\mathbb{Q}_{p}$ for the $p$-adic numbers, and $\breve{\mathbb{Q}}_{p}$ for the completion of the maximal unramified extension with Frobenius $\sigma$.
  • We let $\overline{\mathbb{F}}_{\ell}$ denote the algebraic closure of the finite field $\mathbb{F}_{\ell}$. We fix a choice of square root of $p$ in $\overline{\mathbb{F}}_{\ell}$ and normalize all half Tate twists and square roots of the norm character with respect to this choice. In particular, this how we normalize the geometric Satake correspondence of [16, Chapter VI] and in turn the Fargues-Scholze correspondence.
  • For $L/\mathbb{Q}_{p}$ a finite extension, we write $\breve{L}:=L\breve{\mathbb{Q}}_{p}$ for the compositum of $L$ with the maximal unramified extension and $W_{L}$ for the Weil group of $L$.
  • We let $\mathbb{A}$ (resp. $\mathbb{A}_{f}$) denote the adeles (resp. finite adeles) over $\mathbb{Q}$.
  • A pair $(\mathbf{G},X)$ will denote a Shimura datum. We will use $E$ to denote the reflex field. For $K\subset\mathbf{G}(\mathbb{A}_{f})$ a sufficiently small open compact, we write $\mathrm{Sh}(\mathbf{G},X)_{K}\rightarrow\mathop{\rm Spec}{E}$ for the attached Shimura variety of level $K$.
  • We fix an isomorphism $j:\overline{\mathbb{Q}}_{p}\xrightarrow{\simeq}\mathbb{C}$. Consider the induced embedding $\overline{\mathbb{Q}}\rightarrow\overline{\mathbb{Q}}_{p}$ this gives a finite place $\mathfrak{p}$ of $E$. We write $E_{\mathfrak{p}}$ for the completion at $\mathfrak{p}$.
  • We let $C:=\hat{\overline{E}}_{\mathfrak{p}}$ be the completed algebraic closure of $E_{\mathfrak{p}}$.
  • We use the symbol $G$ to always denote a connected reductive group over $\mathbb{Q}_{p}$, usually taken to be $\mathbf{G}_{\mathbb{Q}_{p}}$. We will always assume that $G$ is quasi-split with a fixed choice $T\subset B\subset G$ of maximal torus and Borel, respectively.
[…]
  • We let $\mathbb{X}_{*}(T_{\overline{\mathbb{Q}}_{p}})^{+}$ denote the set of geometric dominant cocharacters of $G$ and let $\mathbb{X}_{*}(T_{\overline{\mathbb{Q}}_{p}})^{+}/\Gamma$ denote the set of Galois orbits, where $\Gamma:=\mathrm{Gal}(\overline{\mathbb{Q}}_{p}/\mathbb{Q}_{p})$.
  • Let $B(G):=G(\breve{\mathbb{Q}}_{p})/(g\sim hg\sigma(h)^{-1})$ denote the Kottwitz set of $G$.
  • For $b\in B(G)$, we write $J_{b}$ for the $\sigma$-centralizer of $b$. We note that this depends on a choice of representative in the $\sigma$-conjugacy class attached to $b$, but the isomorphism class does not, as a result we will suppress this technicality.
  • For $\mu\in\mathbb{X}_{*}(T_{\overline{\mathbb{Q}}_{p}})^{+}$, we let $B(G,\mu)$ be the $\mu$-admissible locus (as defined in [52, Definition 2.3]).
  • Let $\mathrm{Perf}$ denote the category of affinoid perfectoid spaces in characteristic $p$ over $\ast:=\mathop{\rm Spd}(\overline{\mathbb{F}}_{p})$ endowed with the $v$-topology. For a perfectoid space $S$, let $\mathrm{Perf}_{S}$ denote the category of affinoid perfectoid spaces over the tilt $S^{\flat}$.
  • For $S\in\mathrm{Perf}$, let $X_{S}$ denote the relative schematic Fargues-Fontaine curve over $S$, as defined in the discussion proceeding [7, Definition 3.3.2].
  • For $\mathop{\rm Spa}{(F,\mathcal{O}_{F})}\in\mathrm{Perf}$ a geometric point, we will often drop the subscript on $X_{F}$ and just write $X$ for the associated Fargues-Fontaine curve.
  • For $b\in B(G)$, we write $\mathcal{E}_{b}$ for the associated $G$-bundle on $X$, as defined in [17, Definition 1.1]). As with the $\sigma$-centralizer $J_{b}$, this depends on a choice of representative in $G(\breve{\mathbb{Q}}_{p})$ of the $\sigma$-conjugacy class attached to $b$, but the isomorphism class does not, as a result we will suppress this technicality.
  • For $S\in\mathrm{Perf}$, we let $\mathcal{E}_{0}$ denote the trivial $G$-bundle on $X_{S}$.
  • To a diamond or $v$-stack $X$ over $\ast$, we write $\mathrm{D}(X,\overline{\mathbb{F}}_{\ell})$ for the category of étale $\overline{\mathbb{F}}_{\ell}$-sheaves, as defined in [56, Definition 14.13]. We let $\mathrm{D}^{\mathrm{ULA}}(X,\overline{\mathbb{F}}_{\ell})$ denote the full subcategory of ULA sheaves over $\ast$.
  • For an Artin $v$-stack $X$ and $\Lambda\in\{\overline{\mathbb{F}}_{\ell},\overline{\mathbb{Z}}_{\ell},\overline{\mathbb{Q}}_{\ell}\}$, we write $\mathrm{D}_{\blacksquare}(X,\Lambda)$ for the condensed $\infty$-category of solid $\overline{\mathbb{F}}_{\ell}$-sheaves on $X$, and write $\mathrm{D}_{\mathrm{lis}}(X,\Lambda)\subset\mathrm{D}_{\blacksquare}(X,\Lambda)$ for the full sub-category of $\Lambda$-lisse-étale sheaves, as defined in [16, Chapter VII].
  • If $X$ is an Artin $v$-stack ([16, Definition IV.V.1]) admitting a separated cohomologically smooth surjection $U\rightarrow X$ from a locally spatial diamond $U$ such that the étale site has a basis with bounded $\ell$-cohomological dimension (which will always be the case for our applications) then we will regard $\mathrm{D}(X,\overline{\mathbb{F}}_{\ell})$ as a condensed $\infty$-category via the identification $\mathrm{D}_{\mathrm{lis}}(X,\overline{\mathbb{F}}_{\ell})\simeq\mathrm{D}(X,\overline{\mathbb{F}}_{\ell})$ when viewed as objects in $\mathrm{D}_{\blacksquare}(X,\overline{\mathbb{F}}_{\ell})$ [16, Proposition VII.6.6].
  • We let $\hat{G}$ denote the Langlands dual group of $G$ with fixed splitting $(\hat{T},\hat{B},\{X_{\alpha}\})$.
  • If $F$ denotes the splitting field of $G$ then the action of $W_{\mathbb{Q}_{p}}$on $\hat{G}$ factors through $Q:=W_{\mathbb{Q}_{p}}/W_{F}$. We let $\phantom{}{}^{L}G:=\hat{G}\rtimes Q$ denote the $L$-group.
  • For $I$ a finite index set, we let $\mathrm{Rep}_{\overline{\mathbb{F}}_{\ell}}(\phantom{}^{L}G^{I})$ (resp. $\mathrm{Rep}_{\overline{\mathbb{F}}_{\ell}}(\hat{G}^{I})$) denote the category of finite-dimensional algebraic representations of $\phantom{}{}^{L}G^{I}$ (resp. $\hat{G}^{I}$) over $\overline{\mathbb{F}}_{\ell}$.
[…]
  • To any condensed $\infty$-category $\mathcal{C}$, we write $\mathcal{C}^{BW^{I}_{\mathbb{Q}_{p}}}$ for the category of objects with continuous $W_{\mathbb{Q}_{p}}^{I}$-action, as defined in [16, Section IX.1].
  • For any separated $v$-stack, $X\rightarrow\mathop{\rm Spa}(K,\mathcal{O}_{K})$ where $\mathop{\rm Spa}(K,\mathcal{O}_{K})$ is a non-archimedean field, we write $\overline{X}$ for the canonical compactification of $X$ with respect to the structure map ([56, Proposition 18.6], [32, Theorem 5.15]).
  • For a reductive group $H/\mathbb{Q}_{p}$, we write $\mathrm{D}(H(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})$ for the (left-completed) unbounded derived category of smooth $\overline{\mathbb{F}}_{\ell}$-representations.
2. Preliminaries on Shimura Varieties
In this section, we will recall some facts about Shimura varieties and Igusa varieties which we will use to study the geometry of the Hodge-Tate period morphism in the next section. More specifically, in §2.1, we start with reviewing the definition of a Shimura datum of PEL type AC that we will be interested in (§2.1.1), and then we introduce the minimal and toroidal compactifications of the Shimura varieties attached to these datum (§2.1.3). In the next section §2.2, we carry out an analogous discussion for Igusa varieties, first reviewing their definition and moduli interpretation (§2.2.1), and then studying their toroidal and minimal compactifications (§2.2.4, §2.2.3). This will be important in our discussion in §3.1 for describing the fibers of the Hodge-Tate period morphism on the open Shimura variety.
2.1. Shimura Varieties
We will mainly work with the following two types of Shimura varieties. Our main reference here is [41, §1].
2.1.1. PEL type AC
Let $(\mathcal{O}_{B},\ast,L,\langle\cdot,\cdot\rangle,h)$ be an (integral) PEL datum, as defined in [41, §1.2.1]. Note here that by tensoring $O_{B}$ and $L$ by $\mathbb{Q}$, we obtain the (rational) PEL data as introduced by Kottwitz [39, §1-4]. More precisely, we require that $B$ is a finite-dimensional semisimple $\mathbb{Q}$-algebra, $\ast$ is a $\mathbb{Q}$-linear involution of $B$, with fixed field $F$, $\mathcal{O}_{B}$ is a $\ast$-stable $\mathbb{Z}$-order of $B$, $L$ is a lattice with $\mathcal{O}_{B}$-action, and $\langle\cdot,\cdot\rangle:L\times L\rightarrow\mathbb{Z}(1)$ is a non-degenerate alternating form such that $\langle bv,v^{\prime}\rangle=\langle v,b^{*}v^{\prime}\rangle$, for all $b\in\mathcal{O}_{B}$ and $v,v^{\prime}\in L$. Moreover $h$ is an $\mathbb{R}$-algebra homomorphism
$h:\mathbb{C}\rightarrow\mathrm{End}_{O_{B}\otimes_{\mathbb{Z}}\mathbb{R}}(L\otimes_{\mathbb{Z}}\mathbb{R}),$
such that
  1. (1)
    $\langle h(z)v,w\rangle=\langle v,h(\bar{z})w\rangle$ for all $v,w\in L\otimes_{\mathbb{Z}}\mathbb{R}$ and all $z\in\mathbb{C}$
  2. (2)
    the symmetric bilinear form $(v,h(i)w)$ on $L\otimes_{\mathbb{Z}}\mathbb{R}$ is positive definite.
To our integral PEL datum, we can associate the following group scheme $\mathbf{G}$ over $\mathbb{Z}$ whose $R$-points, for each $\mathbb{Z}$-algebra $R$, are given by
$\mathbf{G}(R):=\{(g,r)\in\mathrm{GL}_{\mathcal{O}_{B}\otimes_{\mathbb{Z}}R}(L\otimes_{\mathbb{Z}}R)\times R^{\times}|\langle gv,gw\rangle=r\langle v,w\rangle\text{ for all }v,w\in L\otimes_{\mathbb{Z}}R\}.$
We will assume from now on that $B$ only has simple factors of type AC. Here, we follow [41, Definition 1.2.1.15] classifying simple $\mathbb{Q}$-algebras with a positive involution. Observe that from [41, Proposition 1.2.3.11], this is equivalent to $\mathbf{G}^{\mathrm{ad}}$ having simple factors only of type A and C.
We now further assume the data is unramified at $p$; namely, that each term in the decomposition $B_{\mathbb{Q}_{p}}=\prod_{\mathfrak{p}|p}B\otimes_{F}F_{\mathfrak{p}}$ is a matrix algebra over an unramified extension of $\mathbb{Q}_{p}$. We will moreover assume that $\mathcal{O}_{B}\otimes\mathbb{Z}_{p}$ is a maximal order in $B_{\mathbb{Q}_{p}}$, and $L$ is self-dual after localization at $p$. This can be arranged following [41, Remark 1.3.4.8]. Note that these conditions equivalently ensure that the group $\mathbf{G}_{\mathbb{Z}_{p}}$ is a smooth reductive group scheme.
We will now briefly discuss what conditions we may need to impose on the prime $p$ so that the form of the local group $G:=\mathbf{G}_{\mathbb{Q}_{p}}$ satisfies the conditions in Table (1).
Type A: In this case, the center $Z(B)=F_{c}$ is a quadratic imaginary extension of $F$. Let $n$ be the $\mathcal{O}_{B}$ rank of $L$. Observe that we will need to assume that the prime $p$ satisfies for all primes $\mathfrak{p}$ of $F$ above $p$,
  1. (1)
    $\mathfrak{p}$ is split in $F_{c}$; or
  2. (2)
    $F_{\mathfrak{p}}=\mathbb{Q}_{p}$, and $n$ is odd.
These conditions imply that $G$ will be a similitude subgroup of $\prod_{\mathfrak{p}}G_{\mathfrak{p}}$ where $G_{\mathfrak{p}}$ is either $\mathrm{Res}_{F_{\mathfrak{p}}/\mathbb{Q}_{p}}(\mathbb{G}_{m}\times\mathrm{GL}_{n})$ or $\mathrm{GU}_{n}$ for either $n=2$ or an odd unitary group.
Type C: Since the PEL data is unramified at $p$, we see that $B\otimes_{F}F_{\mathfrak{p}}$ is indefinite for all primes $\mathfrak{p}$ of $F$ above $p$, and thus $G$ will be a similitude subgroup of
$\prod_{\mathfrak{p}}\mathrm{Res}_{F_{\mathfrak{p}}/\mathbb{Q}_{p}}(\mathrm{GSp}_{2n}),$
where here we define $n$ such that $2n$ is the $O_{B}$ rank of $L$, so $n$ is half the $O_{B}$ rank of $L$. In this case, we will need that $n$ is either $1$ or $2$ to satisfy the conditions in Table (1).
Finally, note that in the case of mixed type A and C, we can decompose $B=B_{1}\times B_{2}$, there $B_{1}$ is of type A, $B_{2}$ is of type C, and hence the $O_{B}$-lattice also decomposes as $L=L_{1}\oplus L_{2}$, where $L_{1}$ is an $O_{B_{1}}$-lattice, and $L_{2}$ is an $O_{B_{2}}$-lattice. We impose the relevant conditions on $n_{1}$, the $O_{B_{1}}$-rank of $L_{1}$, and $n_{2}$, half the $O_{B_{2}}$-rank of $L_{2}$. In this case, the group is a product of groups of the types listed above, and we let $n=n_{1}+n_{2}$.
To each integral PEL data we can associate a moduli space of abelian varieties with extra structures. Let $K^{p}\subset G(\mathbb{A}_{f}^{p})$ be an open compact subgroup, and let $K=K_{p}^{\mathrm{hs}}K^{p}$, where we recall $K_{p}^{\mathrm{hs}}\subset G(\mathbb{Q}_{p})$ denotes a hyperspecial level. To any PEL data, we let $\mathcal{M}_{K}$ over $\mathcal{O}_{E_{\mathfrak{p}}}$ be the scheme which represents the functor that associates to each scheme $S$ over $\mathcal{O}_{E_{\mathfrak{p}}}$ the set of isomorphism classes of tuples $(A,\lambda,\iota,\eta^{p})$ consisting of
  1. (1)
    An abelian scheme $A/S$ of dimension $n[F:\mathbb{Q}]$ up to prime to $p$-isogeny,
  2. (2)
    A prime-to-$p$ polarization $\lambda:A\rightarrow A^{\vee}$,
  3. (3)
    An embedding $\iota:\mathcal{O}_{B}\otimes\mathbb{Z}_{(p)}\hookrightarrow\mathrm{End}(A)\otimes_{\mathbb{Z}}\mathbb{Z}_{(p)}$ of $\mathbb{Z}_{(p)}$-algebras such that
    $\lambda\circ\iota(b^{*})=\iota(b)^{\vee}\circ\lambda,$
  4. (4)
    A section $\eta^{p}\in\Gamma(S,\mathrm{Isom}_{B}(\underline{L\otimes_{\mathbb{Z}}\mathbb{A}_{f}^{p}},\underline{H^{1}(A,\mathbb{A}_{f}^{p}})/\underline{K^{p}}))$, where $\mathrm{Isom}_{B}(\underline{L\otimes_{\mathbb{Z}}\mathbb{A}_{f}^{p}},\underline{H^{1}(A,\mathbb{A}_{f}^{p}}))$ is the $\mathbf{G}(\mathbb{A}_{f}^{p})$ pro-étale torsor of $O_{B}\otimes\mathbb{A}_{f}^{p}$-equivariant isomorphisms that maps $\langle,\rangle$ to a $\underline{\mathbb{A}_{f}^{p\times}}$-multiple of the pairing on $H^{1}(A,\mathbb{A}_{f}^{p})$ defined by the Weil pairing.
satisfying the Kottwitz determinant condition that $\mathrm{det}(b|\mathrm{Lie}(A))=\mathrm{det}(b|V^{-1,0})$ as polynomial functions on $\mathcal{O}_{B}$, where $V=L\otimes\mathbb{Q}$ and $V_{\mathbb{C}}=V^{-1,0}\oplus V^{0,-1}$ is the Hodge decomposition. Here, the polynomial functions on $O_{B}$ for the action of $O_{B}$ on a finite locally free $O_{S}$-module $M$ for a scheme $S$ are defined in [39, §5].
Following [39, §8], we know that the generic fiber of $\mathcal{M}_{K}$ is a finite union of Shimura varieties for groups $\mathbf{G}^{\prime}$ with $\mathbf{G}(\mathbb{A})\simeq\mathbf{G}^{\prime}(\mathbb{A})$. In particular, the canonical integral model of the Shimura variety $\mathscr{S}(\mathbf{G},X)_{K}$ attached to $\mathbf{G}$ and the Hermitian space $X$ consisting of $\mathbf{G}(\mathbb{R})$-conjugates of $h$ is a union of connected components of $\mathcal{M}_{K}$. As such, we can still associate to $S$-points of $\mathscr{S}(\mathbf{G},X)_{K}$ an abelian scheme over $S$ with extra structures.
2.1.2. $p^{m}$-level
We will also consider normal integral models of Shimura varieties with deeper level at $p$, defined as follows. Assume from now on that the level structure $K$ is a principal congruence subgroup for some $N\geq 3$; namely, we have that
$K=K(N)=\{g\in\mathbf{G}(\hat{\mathbb{Z}})|g\equiv 1\pmod{N}\}$.
We consider Shimura varieties of level $K^{\prime}=K(p^{n}N)$ where $p\nmid N$. In this case, we consider $M_{K^{\prime}}$ the functor which assigns to every scheme $S$ over $E$ the set of isomorphism classes of tuples $(A,\lambda,\iota,\eta)$, where
  1. (1)
    AA is an abelian scheme over $S$, $\lambda$ a prime-to-$p$ polarization
  2. (2)
    $\iota:B\hookrightarrow\mathrm{End}(A)\otimes\mathbb{Q}$
  3. (3)
    $\eta$ is a section of the pro-étale torsor $\Gamma(S,\mathrm{Isom}_{B}(\underline{L\otimes_{\mathbb{Z}}\mathbb{A}_{f}},\underline{H^{1}(A,\mathbb{A}_{f}})/\underline{K^{\prime}}))$.
As in the case of the integral model at hyperspecial level, this moduli problem is representable for $N>3$ by a scheme over $E$. We can thus consider in this situation the composition
$M_{K^{\prime}}\rightarrow M_{K}\rightarrow\mathcal{M}_{K},$
and following [42] we let $\mathcal{M}_{K^{\prime}}$ denote the normalization of $\mathcal{M}_{K}$ in $M_{K^{\prime}}$, which by definition is a flat normal scheme over $\mathcal{O}_{E_{\mathfrak{p}}}$. Similar to before, $\mathcal{M}_{K^{\prime}}$ is a normal integral model for a finite union of Shimura varieties $\mathrm{Sh}_{K^{\prime}}(\mathbf{G},X)$ for groups $\mathbf{G}^{\prime}$ with $\mathbf{G}(\mathbb{A})\simeq\mathbf{G}^{\prime}(\mathbb{A})$, and we thus obtain a normal integral model $\mathscr{S}(\mathbf{G},X)_{K^{\prime}}$ for $\mathrm{Sh}_{K^{\prime}}(\mathbf{G},X)$.
2.1.3. Compactifications
We will now recall some constructions from the theory of toroidal compactifications of PEL type Shimura varieties from [41, 42].
We first recall the definition of a split, symplectic and admissible filtration from [41, §5.2.1]. Let $R$ be a commutative ring.
Definition 2.1.
A split, symplectic and admissible filtration on $L\otimes_{\mathbb{Z}}R$ is a two-step filtration on $L\otimes_{\mathbb{Z}}R$ by $(\mathcal{O}_{B}\otimes_{\mathbb{Z}}R)$-submodules, i.e. there is a flag
$0=Z_{-3}\subset Z_{-2}\subset Z_{-1}\subset L\otimes_{\mathbb{Z}}R=Z_{0},$
such that if we put $\mathrm{Gr}^{Z}_{-i}=Z_{-i}/Z_{-i-1}$ for $0\leq i\leq 2$, and $\mathrm{Gr}^{Z}=\oplus_{0\leq i\leq 2}\mathrm{Gr}^{Z}_{-i}$, we have
  1. (1)
    $\mathrm{Gr}^{Z}_{-i}$ is isomorphic to $M\otimes_{\mathbb{Z}}R$ for some $\mathcal{O}_{B}$-lattice $M$
  2. (2)
    There is some isomorphism of $\mathcal{O}_{B}$-lattices
    $L\otimes_{\mathbb{Z}}R\simeq\mathrm{Gr}^{Z}$
  3. (3)
    $Z_{-2}$ and $Z_{-1}$ are annihilators of each other under the pairing $\langle,\rangle$ induced from $L$.
Let $R=\hat{\mathbb{Z}}$ and suppose that we have a split symplectic admissible filtration $Z=Z_{\bullet}$ as above. The following is [41, Definition 5.4.1.3]:
Definition 2.2.
A torus argument $\Phi$ for $Z$ is a tuple $\Phi=(X,Y,\phi,\varphi_{-2},\varphi_{0})$, where
  1. (1)
    $X$ and $Y$ are $\mathcal{O}_{B}$-lattices of the same $B$-multi-rank, and $\phi:Y\hookrightarrow X$ is an $\mathcal{O}_{B}$-linear embedding
  2. (2)
    We have isomorphisms $\varphi_{-2}:\mathrm{Gr}^{Z}_{-2}\simeq\mathrm{Hom}_{R}(X\otimes_{\mathbb{Z}}R,R(1))$ and $\varphi_{0}:\mathrm{Gr}^{Z}_{0}\simeq Y\otimes_{\mathbb{Z}}R$ such that the pairing $\langle,\rangle_{20}:\mathrm{Gr}^{Z}_{-2}\times\mathrm{Gr}^{Z}_{0}\rightarrow R(1)$ is the pullback under these isomorphisms of the pairing
    $\langle\cdot,\cdot\rangle^{\phi}:\mathrm{Hom}_{R}(X\otimes R,R(1))\times(Y\otimes R)\xrightarrow{\mathrm{id}\times\phi}\mathrm{Hom}_{R}(X\otimes R,R(1))\times(X\otimes R)\rightarrow R(1),$
    where the last arrow is the tautological pairing.
In the above definition, we use the notion of $B$-multi-rank as defined in [41, Definition 1.2.1.20].
Definition 2.3.
A cusp label is a pair $(Z,\Phi)$, where $Z$ is a split symplectic admissible filtration on $L\otimes_{\mathbb{Z}}\hat{\mathbb{Z}}$, and $\Phi$ is a torus argument for $Z$.
There is an action of $\mathbf{G}(\mathbb{A}_{f})$ on the set of possible cusp labels, defined as follows. We can consider rational cusp labels $(Z\otimes\mathbb{Q},\Phi\otimes\mathbb{Q})$, where $Z\otimes\mathbb{Q}$ is the filtration on $L\otimes\mathbb{A}_{f}$ obtained by tensoring with $\mathbb{Q}$ over $\mathbb{Z}$, and $\Phi\otimes\mathbb{Q}$ is similarly the tuple obtained by tensoring $\varphi_{-2},\varphi_{0}$ with $\mathbb{Q}$. Observe that this association with rational cusp labels is a bijection, since we can recover the integral filtration by intersecting with $L\otimes_{\mathbb{Z}}\hat{\mathbb{Z}}$ and the integral isomorphisms $\varphi_{-2}$ and $\varphi_{0}$ by restriction, since $\mathrm{Gr}_{-i}^{Z}$ is a lattice in $\mathrm{Gr}_{-i}^{Z\otimes\mathbb{Q}}$. Every $g\in\mathbf{G}(\mathbb{A}_{f})$ defines an automorphism of $L\otimes\mathbb{A}_{f}$, so given a filtration $Z\otimes\mathbb{Q}$, we let $g(Z\otimes\mathbb{Q})$ be the filtration obtained as the image under $g$, and $g(\Phi\otimes\mathbb{Q})$ the torus argument obtained by precomposing $\varphi_{-2},\varphi_{0}$ with the isomorphisms $\mathrm{Gr}_{-i}^{g(Z\otimes\mathbb{Q})}\simeq\mathrm{Gr}_{-i}^{Z\otimes\mathbb{Q}}$ induced by $g$.
We define a cusp label of level $K(N)$ to be a $K(N)$-orbit of cusp labels.
Remark 2.4.
Note that this is the generalization of the cusp labels $(Z,X)$ considered in [8, §2.5.2], as for the PEL type A Shimura data they considered, the assumption of principal polarization means we can set $X=Y$, and the torus argument $\Phi$ is determined by the $\mathcal{O}_{F}$-isomorphism in part (2) of loc. cit. This definition is not the same as [41, Definition 5.4.1.9] for level $N$ (Lan uses $n$ as notation for the level). However, if we take a $K(N)$-orbit of cusps labels as defined here, we can look at the reduction mod $N$ of $\varphi_{-2}$ and $\varphi_{0}$ for any element in the orbit, and similarly taking the filtration mod $N$ gives a filtration on $L/NL$, and this recovers Lan’s definition.
To each cusp label $(Z,\Phi)$, we can associate a split torus $E_{\Phi}$ over $\mathbb{Z}$, as constructed by Lan in [41, §6.2.2]. Let $S_{\Phi}=\mathbb{X}^{*}(E_{\Phi})$, and note that from the construction it is a quotient of $Y\otimes_{\mathbb{Z}}X$. Let $S_{\Phi}^{\vee}:=\mathrm{Hom}_{\mathbb{Z}}(S_{\Phi},\mathbb{Z})$ be the $\mathbb{Z}$-dual of $S_{\Phi}$, and let $(S_{\Phi})^{\vee}_{\mathbb{R}}:=S^{\vee}_{\Phi}\otimes_{\mathbb{Z}}\mathbb{R}$. The $\mathbb{R}$-vector space $(S_{\Phi})^{\vee}_{\mathbb{R}}$ is isomorphic to the space of Hermitian pairings $|\cdot,\cdot|:(Y\otimes\mathbb{R})\times(Y\otimes\mathbb{R})\rightarrow\mathcal{O}_{B}\otimes\mathbb{R}$ with isomorphism constructed by sending a Hermitian pairing $|\cdot,\cdot|$ to the function $y\otimes\phi(y^{\prime})\mapsto\mathrm{Tr}_{B/\mathbb{Q}}(|y,y^{\prime}|)$ in $\mathrm{Hom}_{\mathbb{Z}}(S_{\Phi},\mathbb{R})$ for any $y,y^{\prime}\in Y$ (c.f. [41, §6.2.5]).
Thus, we have an $\mathbb{R}$-vector space $(S_{\Phi})_{\mathbb{R}}^{\vee}$ of Hermitian pairings, and we define $P_{\Phi}$ to be the subset of $(S_{\Phi})_{\mathbb{R}}^{\vee}$ corresponding to positive semi-definite Hermitian pairings with admissible radicals (see [41, Definition 6.2.5.4] and subsequent discussion for the precise definition of admissible radical). $P_{\Phi}$ will be a rational polyhedral cone in $(S_{\Phi})_{\mathbb{R}}^{\vee}$. Moreover, to every cusp $(Z,\Phi)$ we can also associate a stabilizer group $\Gamma_{Z}$ defined as follows. $\Gamma_{Z}\subset\mathrm{GL}_{O_{B}}(X)\times\mathrm{GL}_{O_{B}}(Y)$ is the group of pairs of isomorphisms $\gamma_{X}$ and $\gamma_{Y}$ of $X$ and $Y$ respectively satisfying $\phi=\gamma_{X}\circ\phi\circ\gamma_{Y}$, and mod $N$ we have $\varphi_{-2,N}=\gamma_{X,N}^{\vee}\circ\varphi_{-2,N}$ and $\varphi_{0,N}=\gamma_{Y,N}\circ\varphi_{0,N}$. Finally, let $\Sigma_{Z}$ be a $\Gamma_{Z}$-admissible rational polyhedral cone decomposition of $P_{Z}$, as in [41, Definition 6.1.1.14].
From now on, we will assume that we have fixed a compatible choice of admissible smooth rational polyhedral cone decomposition data (rpcd) $\Sigma$ for $K$; namely, we have
  1. (1)
    A complete set of representatives $(Z,\Phi)$ of cusp labels at level $K$,
  2. (2)
    A $\Gamma_{Z}$ -admissible smooth rational polyhedral cone decomposition $\Sigma_{\Phi}$ for each cusp $(Z,\Phi)$ so that the cone decompositions are pairwise compatible.
The precise definition and proof of existence of such smooth admissible rpcd is [41, §6.3.3.2,§6.6.3.3]. Associated to this admissible smooth rpcd, we have a toroidal compactification of $\mathscr{S}(G,X)_{K}$ for $K=K(N)$ as constructed by Lan [41, Theorem 6.4.1.1] (when $p\nmid N$), and [42, §7-10] (when $p\mid N$).
The toroidal compactification may be described as follows. To a cusp label $(Z,\Phi)$ of level $K$, there is a smaller-dimensional Shimura variety, denoted $\mathscr{S}_{Z}$, which we will sketch briefly below, and whose precise definition may be found in [41, §6.2] $(p\nmid N)$ and [42, §8] ($p\mid N$). Then, we have the following:
Theorem 2.5.
To each compatible choice $\Sigma=\{\Sigma_{\Phi}\}$ of admissible smooth rational polyhedral cone decomposition data, there is an associated flat proper normal algebraic scheme $\mathscr{S}(G,X)_{K}^{\mathrm{tor}}$ over $\mathcal{O}_{E,\mathfrak{p}}$ (which is even smooth when $p\nmid N$) containing $\mathscr{S}(G,X)_{K}$ as an open dense subscheme, together with a semiabelian family $\mathcal{A}$ over $\mathscr{S}(G,X)_{K}^{\mathrm{tor}}$. Moreover, we have the following
  1. (1)
    We have a decomposition into reduced locally closed subvarieties
    $\mathscr{S}(\mathbf{G},X)_{K}^{\mathrm{tor}}=\bigsqcup_{[Z]}\mathscr{S}(\mathbf{G},X)^{\mathrm{tor}}_{K,Z},$
    where $Z=(Z,\Phi)$ is a cusp label of level $K$, and where each $\mathscr{S}(\mathbf{G},X)^{\mathrm{tor}}_{K,Z}$ is flat over $O_{E_{\mathfrak{p}}}$.
  2. (2)
    There exists a scheme
    $C_{Z}\rightarrow\mathscr{S}_{Z}$
    (which when $p\nmid N$ is an abelian scheme torsor). Moreover, there exists a torsor under the torus $E_{\Phi}$, $\Xi_{Z}\rightarrow C_{Z}$, and a relative toroidal embedding $\Xi_{Z}\hookrightarrow\Xi_{Z,\Sigma_{Z}}$ induced by the cone $\Sigma_{Z}$ (see [41, Definition 6.1.2.3] for the definition of relative toroidal embedding), as depicted in the following diagram:
    $\Xi_{Z}$$C_{Z}$$\mathscr{S}_{Z}$$\Xi_{Z,\Sigma_{Z}}$}
    and $\Xi_{Z}$ has an action of $\Gamma_{Z}$ which extends to $\Xi_{Z,\Sigma_{Z}}$. We denote $\Xi_{Z,\Sigma_{Z}}\backslash\Xi_{Z}$ by $\partial\Xi_{Z,\Sigma_{Z}}$.
    If we let
    $\left(\mathscr{S}(\mathbf{G},X)^{\mathrm{tor}}_{K}\right)^{\wedge}_{[Z]},$
    denote the completion of $\mathscr{S}(\mathbf{G},X)_{K}^{\mathrm{tor}}$ along the locally closed component indexed by [Z][Z] [^2] By convention, by the formal completion of a scheme $X$ along a locally closed subscheme $Z$ we mean the formal completion of the open subscheme $X-(\overline{Z}-Z)$ along its closed subscheme $Z$., then we have an isomorphism
    $\left(\mathscr{S}(\mathbf{G},X)^{\mathrm{tor}}_{K}\right)^{\wedge}_{[Z]}\simeq\mathfrak{X}_{Z,\Sigma_{Z}}/\Gamma_{Z},$
    where $\mathfrak{X}_{Z,\Sigma_{Z}}$ is the completion of $\Xi_{Z,\Sigma_{Z}}$ along $\partial\Xi_{Z,\Sigma_{Z}}$.
We note that $C_{Z},\mathscr{S}_{Z},\Xi_{Z}$ parameterize certain degeneration structures on semiabelian schemes, which we will briefly sketch.
We first recall some facts about degenerations of abelian schemes, from [41, §3.3] and [8, §2.5.1]. Let $R$ be a normal local ring with fraction field $H$. Consider a polarized abelian variety $(A,\lambda)$ over $H$ with $\mathcal{O}_{B}$-structure, and $\mathcal{A}$ a semiabelian scheme over $R$ which is a degeneration of $A$ (that is, $\mathcal{A}_{H}\simeq A$). Then this uniquely determines a short exact sequence
$0\rightarrow T\rightarrow\mathcal{G}\rightarrow\mathcal{B}\rightarrow 0$
where $T$ is a torus, $\mathcal{B}$ is an abelian scheme over $R$, and $\mathcal{G}$ is the Raynaud extension (see [41, Definition 3.3.3.9] for definition). Let $X=\mathbb{X}^{*}(T)$, this is a free module over $R$. The lattice $X$ has an action of $\mathcal{O}_{B}$, and so does $\mathcal{B}$. Then, $\mathcal{G}$ determines, and is uniquely determined by, an $\mathcal{O}_{B}$-linear map $c:X\rightarrow\mathcal{B}^{\vee}$. Similarly, we can consider a degeneration $\mathcal{A}^{\vee}$ over $R$ of the dual $A^{\vee}/H$, which gives us a short exact sequence
$0\rightarrow T^{\vee}\rightarrow\mathcal{G}^{\vee}\rightarrow\mathcal{B}^{\vee}\rightarrow 0,$
and if we similarly let $Y=\mathbb{X}^{*}(T^{\vee})=\mathbb{X}_{*}(T)$, the extension $\mathcal{G}$ determines, and is uniquely determined by an $\mathcal{O}_{B}$-linear map $c^{\vee}:Y\rightarrow\mathcal{B}$.
Moreover, note that, as shown in [41, §3.4], the polarization on $A$ extends uniquely to a homomorphism $\lambda:\mathcal{G}\rightarrow\mathcal{G}^{\vee}$. The homomorphism $\lambda$ determines and is uniquely determined by the data of the polarization on $\mathcal{B}$, and an injective $\mathcal{O}_{B}$-linear map $\phi:Y\rightarrow X$.
Then, given a cusp label $Z$ of level $K(N)$, we can define a smaller-dimensional Shimura variety $\mathscr{S}_{Z}$ which will be a union of connected components of a PEL moduli space which parameterizes the data of abelian schemes $\mathcal{B}$ as described above, over which we consider another moduli space $C_{Z}$, additionally parametrizing the maps $c,c^{\vee}$, as well as extensions of maps $f_{0}:\frac{1}{N}X\rightarrow\mathcal{B}^{\vee}$ and $f_{0}^{\vee}:\frac{1}{N}Y\rightarrow\mathcal{B}$. The space $\Xi_{Z}$ parameterizes liftings of the map $f^{\vee}:\frac{1}{N}Y\rightarrow\mathcal{B}$ to a map $f^{\vee}:\frac{1}{N}Y\rightarrow\mathcal{G}$.
Finally, we also want to consider the integral model over $\mathcal{O}_{E_{\mathfrak{p}}}$ of the minimal compactification which was constructed in [41, §7.2.3], and we denote this by $\mathscr{S}(\mathbf{G},X)_{K}^{*}$. We will denote the generic fibers of the toroidal and minimal compactifications by $\mathrm{Sh}(\mathbf{G},X)_{K}^{\mathrm{tor}}$ and $\mathrm{Sh}(\mathbf{G},X)^{*}_{K}$, respectively.
2.2. Igusa Varieties
We now turn our attention to studying Igusa varieties. Throughout this subsection the level $K=K(N)$ satisfies $p\nmid N$. We start with the basic definition.
2.2.1. The Basic Definition
Definition 2.6.
A $p$-divisible group $\mathbb{X}$ with $\mathbf{G}_{\mathbb{Z}_{p}}$-structure is a tuple $(\mathbb{X},\lambda,\iota)$ where $\mathbb{X}$ is a $p$-divisible group of dimension $n[F:\mathbb{Q}]$, $\lambda$ is a quasi-polarization (i.e. a quasi-isogeny $\lambda:\mathbb{X}\dashrightarrow\mathbb{X}^{\vee}$), and $\iota:\mathcal{O}_{B_{\mathbb{Q}_{p}}}\hookrightarrow\mathrm{End(\mathbb{X})}$, where $\mathcal{O}_{B_{\mathbb{Q}_{p}}}$ is the $p$-adic completion of $\mathcal{O}_{B}$ in $B_{\mathbb{Q}_{p}}$, such that $\lambda\circ\iota(b^{*})=\iota(b)^{\vee}\circ\lambda$.
For notational simplicity, we will often drop $\lambda,\iota$ from the notation.
Let $\mu\in\mathbb{X}_{*}(T_{\overline{\mathbb{Q}}_{p}})^{+}$ be the dominant cocharacter induced by the inverse of the Hodge cocharacter associated with the conjugacy class $X$. We fix now some $b\in B(G,\mu)$, and consider a geometric point $x\in\mathscr{S}(\mathbf{G},X)_{K}(\overline{\mathbb{F}}_{p})$ lying in the Newton strata for $b$. This defines an abelian variety $\mathcal{A}_{x}/\mathop{\rm Spec}(\overline{\mathbb{F}}_{p})$. Consider now the associated $p$-divisible group $\mathbb{X}:=\mathcal{A}_{x}[p^{\infty}]$, the $p$-divisible group $\mathbb{X}$ has $\mathbf{G}_{\mathbb{Z}_{p}}$-structure acquired from the additional structures on $\mathcal{A}_{x}$.
Up to replacing $x$ by another element in its isogeny class, we can assume $\mathbb{X}$ is completely slope divisible, following a result of Oort and Zink [49, Theorem 2.1]. Thus, we can write $\mathbb{X}=\oplus_{i=1}^{r}\mathbb{X}_{i}$, where the $\mathbb{X}_{i}$ are isoclinic $p$-divisible groups of strictly decreasing slopes.
We consider the following subset
$\mathscr{C}_{\mathbb{X}}:=\{x\in\mathscr{S}(\mathbf{G},X)_{K,\overline{\mathbb{F}}_{p}}:\exists\text{ isomorphism }\rho:\mathcal{A}_{x}[p^{\infty}]\times k(\overline{x})\simeq\mathbb{X}\times k(\overline{x})\text{ preserving $\mathbf{G}_{\mathbb{Z}_{p}}$-structure}\},$
known as the central leaf, where we denote by $\mathscr{S}(\mathbf{G},X)_{K,\overline{\mathbb{F}}_{p}}$ the (geometric) special fiber of $\mathscr{S}(\mathbf{G},X)_{K}$. It is shown in [46, Proposition 1] that this is a locally closed subset of $\mathscr{S}(\mathbf{G},X)_{K,\overline{\mathbb{F}}_{p}}$, and that we can give this subset the induced reduced scheme structure which makes the associated scheme $\mathscr{C}_{\mathbb{X}}$ smooth.
Let $\mathscr{G}$ be the $p$-divisible group of the restriction to $\mathscr{C}_{\mathbb{X}}$ of the universal abelian variety over $\mathscr{S}(\mathbf{G},X)_{K}$. We further define $\mathrm{Ig}^{b}$ to be the functor parametrizing, for any perfect $\mathscr{C}_{\mathbb{X}}$-scheme $\mathscr{T}$, isomorphisms $\mathscr{G}\times_{\mathscr{C}_{\mathbb{X}}}\mathscr{T}\simeq\mathbb{X}\times_{\overline{\mathbb{F}}_{p}}\mathscr{T}$ which preserve $\mathbf{G}_{\mathbb{Z}_{p}}$-structure. Equivalently, we can define $\mathrm{Ig}^{b}$ as the functor sending an $\overline{\mathbb{F}}_{p}$-algebra $R$ to the set
(3) $\mathrm{Ig}^{b}(R)=\{(\rho,x):x\in\mathscr{S}(\mathbf{G},X)_{K}(R),\rho:\mathcal{A}_{x}[p^{\infty}]\xrightarrow{\sim}\mathbb{X}_{R}\text{ preserving $\mathbf{G}_{\mathbb{Z}_{p}}$-structure}\}.$
By [7, Proposition 4.3.3], we know that this functor is representable, and by [7, Corollary 4.3.5] the scheme is perfect, and hence it lifts uniquely to a flat $p$-adic formal scheme, which we denote by $\mathfrak{Ig}^{b}$ over $\mathrm{Spf}(W(\overline{\mathbb{F}}_{p}))$. Moreover, from [8, Corollary 2.3.2], we know that $\mathrm{Ig}^{b}\rightarrow\mathscr{C}_{\mathbb{X}}$ is a torsor for $\mathrm{Aut}_{G}(\mathbb{X})_{\overline{\mathbb{F}}_{p}}$, the formal group scheme of automorphisms of $\mathbb{X}$ compatible with extra structures viewed as an fpqc sheaf over $\overline{\mathbb{F}}_{p}$.
We write $\mathfrak{Ig}_{C}^{b}$ for the perfectoid space attached to the adic generic fiber of $\mathfrak{Ig}^{b}$ over $C$. These spaces are supposed to model the fibers of the Hodge-Tate period morphism, a connection we will elaborate upon in §3.1.
2.2.2. Compactifications
In order to understand (partial) minimal and toroidal compactifications of $\mathrm{Ig}^{b}$, we must first consider compactifications of the central leaf $\mathscr{C}_{\mathbb{X}}$. As constructed [54, §3.2.1], the central leaf $\mathscr{C}_{\mathbb{X}}$ admits partial toroidal and minimal compactifications, which we will denote by $\mathscr{C}_{\mathbb{X}}^{\mathrm{tor}}$ and $\mathscr{C}_{\mathbb{X}}^{*}$ respectively. Moreover, it is also shown in [54, §3.2.1] that $\mathscr{C}_{\mathbb{X}}^{\mathrm{tor}}$ is a locally closed subset of $\mathscr{S}(\mathbf{G},X)_{K,\overline{\mathbb{F}}_{p}}^{\mathrm{tor}}$, and $\mathscr{C}_{\mathbb{X}}^{*}$ is a locally closed subset of $\mathscr{S}(\mathbf{G},X)_{K,\overline{\mathbb{F}}_{p}}^{*}$. Let $Z$ be a cusp label at level $K(N)$. This determines a locally closed boundary stratum $\mathscr{C}_{\mathbb{X},Z}\subset\mathscr{C}_{\mathbb{X}}^{\mathrm{tor}}$.
The Igusa variety $\mathrm{Ig}^{b}$ over $\mathscr{C}_{\mathbb{X}}$ extends to a perfect scheme $\mathrm{Ig}^{b,\mathrm{tor}}$ over $\mathscr{C}_{\mathbb{X}}^{\mathrm{tor}}$. More precisely, we can define $\mathrm{Ig}^{b,\mathrm{tor}}$ as follows. Let $\mathcal{A}$ denote the universal semi-abelian scheme over $\mathscr{C}_{\mathbb{X}}^{\mathrm{tor}}$. This is the restriction to $\mathscr{C}_{\mathbb{X}}^{\mathrm{tor}}$ of the universal semi-abelian scheme over $\mathscr{S}(\mathbf{G},X)^{\mathrm{tor}}_{K,\overline{\mathbb{F}}_{p}}$. Then, we know from [54, Proposition 3.2.6] (which is exactly the same argument as in [8, Proposition 3.2.1]) that both the connected part $\mathcal{A}[p^{\infty}]^{\circ}$ of $\mathcal{A}[p^{\infty}]$ and the multiplicative part $\mathcal{A}[p^{\infty}]^{\mu}$ are $p$-divisible groups. We thus let $\mathcal{A}[p^{\infty}]^{(0,1)}=\mathcal{A}[p^{\infty}]^{\circ}/\mathcal{A}[p^{\infty}]^{\mu}$ be the biconnected part. We can similarly define $\mathbb{X}^{\circ},\mathbb{X}^{(0,1)}$ as the connected and biconnected parts of $\mathbb{X}$.
Following [8, Definition 3.2.9] we can define $\mathrm{Ig}^{b,\mathrm{tor}}$ to be the functor which, for a perfect $\mathscr{C}_{\mathbb{X}}^{\mathrm{tor}}$-scheme $\mathscr{T}$, parametrizes
  1. (1)
    $\mathcal{O}_{B_{\mathbb{Q}_{p}}}$ -linear isomorphisms $\rho:\mathcal{A}[p^{\infty}]^{\circ}\times_{\mathscr{C}_{\mathbb{X}}^{\mathrm{tor}}}\mathscr{T}\xrightarrow{\sim}\mathbb{X}^{\circ}\times_{\bar{\mathbb{F}}_{p}}{\mathscr{T}}$
  2. (2)
    A scalar in $\mathbb{Z}^{\times}_{p}(\mathscr{T})$ such that the induced isomorphism $\rho^{(0,1)}:\mathcal{A}[p^{\infty}]^{(0,1)}\times_{\mathscr{C}_{\mathbb{X}}^{\mathrm{tor}}}\mathscr{T}\xrightarrow{\sim}\mathbb{X}_{\mathscr{T}}^{(0,1)}$ obtained by quotienting by the multiplicative parts commutes with the polarizations up to the given element of $\mathbb{Z}_{p}^{\times}(\mathscr{T})$.
Here, $\mathbb{Z}_{p}^{\times}(\mathscr{T})$ is the set of $\mathscr{T}$-points of the group scheme $\mathbb{Z}_{p}^{\times}\times\mathscr{C}_{\mathbb{X}}^{\mathrm{tor}}$. As explained in [8, §3.2.3-§3.2.7], this is representable by a perfect scheme.
We define the partial minimal compactification $\mathrm{Ig}^{b,*}$ as the normalization of $\mathscr{C}_{\mathbb{X}}^{*}$ in $\mathrm{Ig}^{b}$. Since we have $\mathrm{Ig}^{b}\subset\mathrm{Ig}^{b,*},\mathrm{Ig}^{b,\mathrm{tor}}$, we will denote the boundaries by $\partial\mathrm{Ig}^{b,*}$ and $\partial\mathrm{Ig}^{b,\mathrm{tor}}$, respectively. These schemes are all perfect, and we can lift them to $p$-adic formal schemes $\mathfrak{Ig}^{b,*}$, $\mathfrak{Ig}^{b,\mathrm{tor}}$, $\partial\mathfrak{Ig}^{b,*}$, and $\partial\mathfrak{Ig}^{b,\mathrm{tor}}$ over $\mathrm{Spf}(W(\overline{\mathbb{F}}_{p}))$. We similarly denote by $\mathfrak{Ig}_{C}^{b,*}$, $\mathfrak{Ig}_{C}^{b,\mathrm{tor}}$, $\partial\mathfrak{Ig}_{C}^{b,*}$, and $\partial\mathfrak{Ig}_{C}^{b,\mathrm{tor}}$ the associated perfectoid spaces over $C$.
2.2.3. Igusa Cusp Labels
In order to understand the boundary components $\partial\mathrm{Ig}^{b,*}$ and $\partial\mathrm{Ig}^{b,\mathrm{tor}}$, we will recall the notion of Igusa cusp labels, as in [54, Definition 3.2.19] (which we have slightly modified to match the definition of cusp labels previously introduced, and which is similar to [8, Definition 3.3.10]). We let $\mathbb{X}_{b}:=\mathbb{X}$ be the completely slope divisible $p$-divisible group attached to $b$ defined above. We reintroduce $b$ in the notation to emphasise that all constructions here depend on $b$. Finally, observe that, since from the moduli problem the polarization on $\mathcal{A}_{x}$ is prime-to-$p$, the $p$-divisible group $\mathbb{X}_{b}=\mathcal{A}_{x}[p^{\infty}]$ is principally polarized.
Definition 2.7.
We define an Igusa cusp label as a tuple $(Z_{b},Z^{p},X,Y,\phi,\varphi_{0},\varphi_{-2},\tilde{\varphi}_{0},\delta_{b})$ where
  1. (1)
    $Z_{b}$ is an $\mathcal{O}_{B_{\mathbb{Q}_{p}}}$ -stable filtration of $\mathbb{X}_{b}$ by $p$-divisible subgroups of the form
    $0=Z_{b,-3}\subset Z_{b,-2}\subset Z_{b,-1}\subset\mathbb{X}_{b},$
    with quotients which are also $p$-divisible groups, where $\mathrm{Gr}_{-2}^{Z_{b}}=Z_{b,-2}$ is multiplicative, and $\mathrm{Gr}_{0}^{Z_{b}}=\mathbb{X}_{b}/Z_{b,-1}$ is étale, and $Z_{b,-1}$, $Z_{b,-2}$ are Cartier dual to each other under the principal polarization on $\mathbb{X}_{b}$.
  2. (2)
    $\delta_{b}$ is an $\mathcal{O}_{B_{\mathbb{Q}_{p}}}$-linear isomorphism
    $\delta_{b}:\mathrm{Gr}^{Z_{b}}\simeq\mathbb{X}_{b}$
  3. (3)
    $Z^{p}$ is an $\mathcal{O}_{B}$ -stable split, symplectic and admissible filtration
    $0=Z^{p}_{-3}\subset Z^{p}_{-2}\subset Z^{p}_{-1}\subset L\otimes_{\mathbb{Z}}\hat{\mathbb{Z}}^{p}$
  4. (4)
    $X,Y$ are $\mathcal{O}_{B}$-lattices of the same $B$-multirank, together with an $\mathcal{O}_{B}$-linear embedding $\phi:Y\hookrightarrow X$, and we have isomorphisms
    $\varphi_{0}:Y\otimes_{\mathbb{Z}}\hat{\mathbb{Z}}^{p}\simeq\mathrm{Gr}_{0}^{Z^{p}}$
    $\varphi_{-2}:\mathop{\rm Hom}\nolimits(X\otimes_{\mathbb{Z}}\hat{\mathbb{Z}}^{p},\hat{\mathbb{Z}}^{p}(1))\simeq\mathrm{Gr}_{-2}^{Z^{p}}$
    $\tilde{\varphi}_{0}:Y\otimes(\mathbb{Q}_{p}/\mathbb{Z}_{p})\simeq\mathrm{Gr}_{0}^{Z_{b}}$
    such that the pairing $\langle,\rangle_{20}:\mathrm{Gr}^{Z^{p}}_{-2}\times\mathrm{Gr}^{Z^{p}}_{0}\rightarrow\hat{\mathbb{Z}}^{p}(1)$ induced from the one on $L$ is the pullback via $\varphi_{-2},\varphi_{0}$ of the one defined on $X,Y$.
There is an action of $J_{b}(\mathbb{Q}_{p})\times\mathbf{G}(\mathbb{A}^{p}_{f})$ on Igusa cusp labels. For the action of $\mathbf{G}(\mathbb{A}^{p}_{f})$ this works exactly as for the Shimura variety (see discussion after Definition 2.3) since we observe that in this case there is also a bijection between tuples $(Z^{p},X,Y,\phi,\varphi_{0},\varphi_{-2})$ and rational versions $(Z^{p}\otimes\mathbb{Q},X,Y,\phi,\varphi_{0}\otimes\mathbb{Q},\varphi_{-2}\otimes\mathbb{Q})$. For the action of $J_{b}(\mathbb{Q}_{p})$, we first observe that by Dieudonné theory we may replace all $p$-divisible groups with their Dieudonné modules, ie. we ask for a filtration
$Z_{b}:0\subset\mathbb{D}(Z_{b,-2})\subset\mathbb{D}(Z_{b,-1})\subset\mathbb{D}(\mathbb{X}_{b}),$
by Dieudonné submodules with $\mathcal{O}_{B_{\mathbb{Q}_{p}}}$-action such that the quotients are also free $W(\overline{\mathbb{F}}_{p})$-modules with $\mathcal{O}_{B_{\mathbb{Q}_{p}}}$-action and isomorphisms
$\delta_{b}:\mathbb{D}(\mathrm{Gr}^{Z_{b}})\simeq\mathbb{D}(\mathbb{X}_{b})\qquad\text{and}\qquad\tilde{\varphi_{0}}:Y\otimes_{\mathbb{Z}}\mathbb{D}(\mathbb{Q}_{p}/\mathbb{Z}_{p})\simeq\mathbb{D}(\mathrm{Gr}^{Z_{b}}_{0}).$
We can consider rational versions obtained by inverting $p$, i.e.
$Z_{b}\left[\frac{1}{p}\right]:0\subset\mathbb{D}(Z_{b,-2})\left[\frac{1}{p}\right]\subset\mathbb{D}(Z_{b,-1})\left[\frac{1}{p}\right]\subset\mathbb{D}(\mathbb{X}_{b})\left[\frac{1}{p}\right],$
and isomorphisms $\delta_{b}\left[\frac{1}{p}\right],\tilde{\varphi_{0}}\left[\frac{1}{p}\right]$, and we see that to obtain a bijection we need to additionally require that $\delta_{b}\left[\frac{1}{p}\right],\tilde{\varphi}_{0}\left[\frac{1}{p}\right]$ restricts to an isomorphism of Dieudonné modules after restricting to the lattice $\mathbb{D}(\mathbb{X}_{b})$. We now see that $J_{b}(\mathbb{Q}_{p})$ clearly acts on rational data with this extra condition.
If $H\subset J_{b}(\mathbb{Q}_{p})\times\mathbf{G}(\mathbb{A}^{p}_{f})$ is a compact open subgroup then an Igusa cusp label at level $H$ is a $H$-orbit of Igusa cusp labels. For a general closed subgroup $H^{\prime}\subset J_{b}(\mathbb{Q}_{p})\times\mathbf{G}(\mathbb{A}^{p}_{f})$, an Igusa cusp label at level $H^{\prime}$ is a compatible family of Igusa cusp labels at level $H$ for all $H^{\prime}\supset H$.
Additionally, we have the following relationship between Igusa cusp labels and cusp labels for the Shimura variety. Given a cusp label $(Z,\Phi)$, we say that an Igusa cusp label $(Z_{b},Z^{p},X,Y,\phi,\varphi_{0},\varphi_{-2},\tilde{\varphi}_{0},\delta_{b})$ lives over $(Z,\Phi)$ if the lattices $X,Y$ are the same, and the base change of $Z$ to $\hat{\mathbb{Z}}^{p}$ recovers the filtration $Z^{p}$ and the isomorphisms $\varphi_{0},\varphi_{-2}$.
Remark 2.8.
The definition in [54, Definition 3.2.15] of an Igusa cusp label is different from ours in that the definition in loc. cit. involves (using our terminology) first fixing a cusp label of level $K=K(N)$, and then considering Igusa cusp labels of level $\{1\}\times K^{p}$ living over this cusp label. The definition of cusp labels in loc. cit. is adapted from Lan’s definition, c.f. the discussion in 2.4. In particular, the Igusa cusp labels in [54] contains extra data of lattices and isomorphisms in $L\otimes\mathbb{Z}_{p}$. This slight variation does not affect any of the results quoted here, since we will always be in a situation where we fix a cusp label $Z$ and are looking at Igusa cusp labels $\tilde{Z}$ living over $Z$.
2.2.4. Boundary components
We now explain how we can understand the boundary $\partial\mathrm{Ig}^{b,\mathrm{tor}}$ according to Igusa cusp labels following [8] and [54]. Note that since $K=K(N)$ admits a decomposition as $K_{p}K^{p}$, where $K^{p}\subseteq\mathbf{G}(\mathbb{A}_{f}^{p})$, we henceforth use $K^{p}(N)$ to denote this prime-to-$p$ level.
Recall that for a cusp label $Z=(Z,\Phi)$, we have a locally closed boundary stratum $\mathscr{C}_{\mathbb{X}_{b},Z}\subset\mathscr{C}_{\mathbb{X_{b}}}^{\mathrm{tor}}$, and we denote by $\mathrm{Ig}^{b,\mathrm{tor}}_{Z}$ the preimage of $\mathscr{C}_{\mathbb{X}_{b},Z}^{\mathrm{perf}}$ under the map $f:\mathrm{Ig}^{b,\mathrm{tor}}\rightarrow\mathscr{C}_{\mathbb{X_{b}}}^{\mathrm{tor},\mathrm{perf}}$. We want to understand the decomposition of $\mathrm{Ig}^{b,\mathrm{tor}}_{Z}$ according to Igusa cusp labels.
We first recall the following definition in [54] of the perfect boundary components corresponding to the Igusa cusp labels. Set $\mathscr{C}_{C_{Z}}:=\mathscr{C}_{\mathbb{X}_{b},Z}\times_{\mathscr{S}_{Z}}C_{Z}$, and recall that over $\mathscr{C}_{C_{Z}}$ there is a universal Raynaud extension
$0\rightarrow T\rightarrow\mathcal{G}\rightarrow\mathcal{B}\rightarrow 0,$
where $\mathcal{B}$ is an abelian scheme. Similar to above, $\mathcal{B}[p^{\infty}]$ we let the multiplicative part be denoted by $\mathcal{B}[p^{\infty}]^{\mu}$ and the connected part be denoted by $\mathcal{B}[p^{\infty}]^{(0,1)}$.
Definition 2.9.
[54, Definition 3.2.34] Let $\tilde{Z}$ be an Igusa cusp label of level $\{1\}\times K^{p}(N)$ living over a cusp label $Z$ of level $K(N)$. Let $\mathrm{Ig}_{\tilde{Z},C_{Z}}^{b}$ be the functor on perfect schemes $\mathscr{T}$ over $\mathscr{C}_{C_{Z}}$ parametrizing:
  1. (1)
    An isomorphism $\rho:\mathcal{B}[p^{\infty}]^{\mu}\times_{\mathscr{C}_{C_{Z}}}\mathscr{T}\xrightarrow{\sim}\bigl(\mathrm{Gr}^{Z_{b}}_{-1}\bigr)^{\mu}\times_{\overline{\mathbb{F}}_{p}}\mathscr{T}$ compatible with the $\mathcal{O}_{B_{\mathbb{Q}_{p}}}$-actions;
  2. (2)
    A splitting $\delta:\;\mathcal{B}[p^{\infty}]^{\mu}\times_{\mathscr{C}_{C_{Z}}}\mathscr{T}\longrightarrow\mathcal{G}[p^{\infty}]^{\mu}\times_{\mathscr{C}_{C_{Z}}}\mathscr{T}$ compatible with the $\mathcal{O}_{B_{\mathbb{Q}_{p}}}$-actions;
  3. (3)
    An $\mathcal{O}_{B_{\mathbb{Q}_{p}}}$ -linear isomorphism $\rho^{(0,1)}:\mathcal{B}[p^{\infty}]^{(0,1)}\times_{\mathscr{C}_{\mathbb{X}}^{\mathrm{tor}}}\mathscr{T}\xrightarrow{\sim}\mathbb{X}^{(0,1)}\times_{\bar{\mathbb{F}}_{p}}{\mathscr{T}}$ which commutes with the polarizations up to the given element of $\mathbb{Z}_{p}^{\times}(\mathscr{T})$.
This functor is representable by a perfect scheme, as implicitly seen from [54]. More precisely, we can first consider a level $p^{m}$-variant $\mathrm{Ig}_{m,\tilde{Z}_{m},C_{Z}}$ as in [54, Definition 3.2.20] for each $m\in\mathbb{N}$ (this is denoted $\mathrm{Ig}^{\mathbb{X}_{b}}_{\tilde{Z}^{m}_{n},C_{\Phi_{n}},\delta_{n}}$ in loc. cit.). Using [8, Theorem 2.2.4] we see that each of these functors is representable by a scheme over $\mathscr{C}_{Z}$, where we further note that in light of (1) the data in (2) of the splitting is equivalent to asking for an isomorphism
$T[p^{\infty}]\oplus\mathcal{B}^{\mu}[p^{\infty}]\simeq\mathcal{G}^{\mu}[p^{\infty}]$
compatible with $\mathcal{O}_{B_{\mathbb{Q}_{p}}}$ actions. Now, taking the perfection of the inverse limit over $m$ recovers the functor defined above, and hence $\mathrm{Ig}_{\tilde{Z},C_{Z}}^{b}$ [^3] This is denoted by $\mathfrak{Ig}^{\mathbb{X}_{b}}_{Z_{n},C_{\Phi_{n}},\delta_{n}}$ in [54], but in this paper we reserve the fraktur font for formal schemes. is also representable by a perfect scheme.
We have a decomposition of $\mathrm{Ig}_{Z}^{b,\mathrm{tor}}$ into locally closed boundary strata indexed by Igusa cusp labels $\tilde{Z}$ at level $K^{p}(N)$ living over a fixed cusp label $Z$. These boundary components are related to $\mathrm{Ig}_{\tilde{Z},C_{Z}}^{b}$ in the following way: Denote by $\widehat{\mathrm{Ig}}^{b,\mathrm{tor}}_{Z}$ the completion of $\mathrm{Ig}^{b,\mathrm{tor}}$ along the locally closed perfect subscheme $\mathrm{Ig}^{b,\mathrm{tor}}_{Z}$. We have the following:
Theorem 2.10.
[54, Theorem 3.2.36]
  1. (1)
    $\widehat{\mathrm{Ig}}^{b,\mathrm{tor}}_{Z}$ admits a decomposition into open and closed formal subschemes
    $\widehat{\mathrm{Ig}}^{b,\mathrm{tor}}_{Z}=\bigsqcup_{\tilde{Z}}\widehat{\mathrm{Ig}}^{b,\mathrm{tor}}_{\tilde{Z}},$
    where $\tilde{Z}$ runs over all Igusa cusp labels of level $\{1\}\times K^{p}(N)$ living over $Z$.
  2. (2)
    Each $\widehat{\mathrm{Ig}}^{b,\mathrm{tor}}_{\tilde{Z}}$ admits the following description: We have a commutative diagram with both squares Cartesian:
    (4) $\mathrm{Ig}^{b}_{\tilde{Z},\Xi}$$\Xi_{Z}^{\mathrm{perf}}$$\mathrm{Ig}^{b}_{\tilde{Z},\Xi,\Sigma_{Z}}$$\Xi_{Z,\Sigma_{Z}}^{\mathrm{perf}}$$\mathrm{Ig}^{b}_{\tilde{Z},C_{Z}}$$C_{Z}^{\mathrm{perf}},$}
    where the composition of the vertical maps are torsors for the torus $E_{\Phi}^{\mathrm{perf}}$, and the inclusion maps in the top square define a relative toroidal embedding.
[…]
    $\widehat{\mathrm{Ig}}^{b,\mathrm{tor}}_{\tilde{Z}}\simeq\widehat{\mathrm{Ig}}^{b}_{\tilde{Z},\Xi,\Sigma_{Z}}/\Gamma_{\tilde{Z}}$,
    Here, we define $\Gamma_{\tilde{Z}}$ to be the subgroup of $\Gamma_{Z}$ consisting isomorphisms $(\gamma_{X},\gamma_{Y})$ such that $\tilde{\varphi}_{0}=\gamma_{Y}\circ\tilde{\varphi}_{0}$.
For later use, we give also an alternate moduli description of $\mathrm{Ig}^{b}_{\tilde{Z},\Xi}$.
Proposition 2.11.
We have the following:
  1. (1)
    1. (a)
      Over $\mathrm{Ig}^{b}_{\tilde{Z},C_{Z}}$, we have universal Raynaud extensions
      $0\rightarrow T\rightarrow\mathcal{G}\rightarrow\mathcal{B}\rightarrow 0$
      and its dual
      $0\rightarrow T^{\vee}\rightarrow\mathcal{G}^{\vee}\rightarrow\mathcal{B}^{\vee}\rightarrow 0,$
      equivalently maps $c:X\rightarrow\mathcal{B}$ and $c^{\vee}:Y\rightarrow\mathcal{B}$.
    2. (b)
      Over $\mathrm{Ig}^{b}_{\tilde{Z},C_{Z}}$, there exists a universal $\mathcal{O}_{B_{\mathbb{Q}_{p}}}$-equivariant embedding $\rho:\mathcal{G}[p^{\infty}]\hookrightarrow\mathbb{X}_{b}\times_{\overline{\mathbb{F}}_{p}}\mathrm{Ig}^{b}_{\tilde{Z},C_{Z}}$ which is compatible with polarizations in the following sense: The induced filtration
      $Z_{\mathbb{X}_{b}}:T[p^{\infty}]\subset\mathcal{G}[p^{\infty}]\subset\mathbb{X}_{b}\times_{\overline{\mathbb{F}}_{p}}\mathrm{Ig}^{b}_{\tilde{Z},C_{Z}}$
      is the base change of $Z_{b}$, and the induced isomorphism
      $\mathrm{Gr}^{Z_{b}}_{-1}\simeq\mathcal{G}[p^{\infty}]/T[p^{\infty}]=\mathcal{B}[p^{\infty}]$
      is compatible with the polarizations
    3. (c)
      There is a lift of $c^{\vee}:Y\rightarrow\mathcal{B}$ to $\tilde{c}^{\vee}:Y[\frac{1}{p}]\rightarrow\mathcal{B}$ corresponding to a symplectic splitting of the filtration $Z_{\mathbb{X}_{b}}$.
  2. (2)
    For $\tilde{c}^{\vee}$ as above, $\mathrm{Ig}^{b}_{\tilde{Z},\Xi}$ is the space of symmetric lifts $\iota:Y[\frac{1}{p}]\rightarrow\mathcal{G}$ lifting $\tilde{c}^{\vee}$.
Proof.
This follows from [54, Lemma 3.2.32,Proposition 3.2.33]. More precisely, loc. cit. gives a moduli description for $\mathrm{Ig}^{b}_{Z}$, by considering a moduli functor of degeneration perfect Igusa-level structures [54, Definition 3.2.30], and a torsor for $E_{\Phi}^{\mathrm{perf}}$ over this moduli space. If we restrict now to the boundary component indexed by some $\tilde{Z}$ living over the fixed $Z$, as in the decomposition of Theorem 2.10, (c.f. the discussion after [54, Definition 3.2.34]), we see that this amounts to identifying the filtration with $Z_{b}$. ∎
Let $\widehat{\mathscr{S}}(\mathbf{G},X)^{\mathrm{tor}}_{K^{p}}$ denote the inverse limit over $n$ of the $p$-adic completion of toroidal integral models with level $K(p^{n}N)$,
$\widehat{\mathscr{S}}(\mathbf{G},X)^{\mathrm{tor}}_{K^{p}}=\mathop{\rm lim}\nolimits\limits_{\begin{subarray}{c}\longleftarrow\\ n\end{subarray}}\widehat{\mathscr{S}}(\mathbf{G},X)^{\mathrm{tor}}_{K(p^{n}N)}.$
Fix a lift $\mathbb{X}_{O_{C}}$ of $\mathbb{X}=\mathbb{X}_{b}$ to a $p$-divisible group with $\mathbf{G}_{\mathbb{Z}_{p}}$-structure over $O_{C}$, and an isomorphism $\alpha:T_{p}(\mathbb{X}_{O_{C}})\simeq L\otimes_{\mathbb{Z}}\mathbb{Z}_{p}$. Following [54, Theorem 4.3.8], we have a map of formal schemes
$g:\mathfrak{Ig}^{b,\mathrm{tor}}\rightarrow\widehat{\mathscr{S}}(\mathbf{G},X)^{\mathrm{tor}}_{K^{p}}.$
This map $g$ is related to the natural map $f:\mathrm{Ig}^{b,\mathrm{tor}}\rightarrow\mathscr{C}^{\mathrm{tor},\mathrm{perf}}\rightarrow\mathscr{S}(\mathbf{G},X)^{\mathrm{tor}}_{K}$ in that on the mod $p$ fiber $g$ factors $f$, i.e. the composition of the mod $p$ fiber of $g$ with the natural projection $\mathscr{S}(\mathbf{G},X)_{K^{p}}\rightarrow\mathscr{S}(\mathbf{G},X)_{K}$ is $f$.
For later use, we will need to understand this map $g$ on toroidal boundary components. Moreover, we also want to describe the lifting of $\widehat{\mathrm{Ig}}^{b,\mathrm{tor}}_{\tilde{Z}}$ to a formal scheme over $O_{C}$. Let $\mathfrak{Ig}_{Z,O_{C}}^{b,\mathrm{tor}}$ be the base change to $O_{C}$ of the canonical flat formal scheme lifting $\mathrm{Ig}^{b,\mathrm{tor}}_{Z}$. Note that an Igusa cusp label $\tilde{Z}$ of level $K^{p}(N)$ is the same as the data of a compatible system of Igusa cusp labels $\tilde{Z}_{m}$ of level $\Gamma_{b}(p^{m})K^{p}(N)$ for all $m\geq 0$, thus we define
$C_{\tilde{Z}}:=\mathop{\rm lim}\nolimits\limits_{\begin{subarray}{c}\longleftarrow\\ m\end{subarray}}C_{\tilde{Z}_{m}}\qquad\Xi_{\tilde{Z}}:=\mathop{\rm lim}\nolimits\limits_{\begin{subarray}{c}\longleftarrow\\ m\end{subarray}}\Xi_{\tilde{Z}_{m}}\qquad\Xi_{\tilde{Z},\Sigma_{Z}}:=\mathop{\rm lim}\nolimits\limits_{\begin{subarray}{c}\longleftarrow\\ m\end{subarray}}\Xi_{\tilde{Z}_{m},\Sigma_{Z}},$
while $\widehat{C}_{\tilde{Z}}$, $\widehat{\Xi}_{\tilde{Z}}$, $\widehat{\Xi}_{\tilde{Z},\Sigma_{Z}}$ denote the $p$-adic completions.
We have the following:
Proposition 2.12.
There is a flat formal lift of $\widehat{\mathrm{Ig}}^{b,\mathrm{tor}}_{\tilde{Z}}$ over $\mathrm{Spf}O_{C}$, denoted $\widehat{\mathfrak{Ig}}^{b,\mathrm{tor}}_{\tilde{Z},O_{C}}$, which may be described as follows. We have a commutative diagram of formal schemes
(5) $\mathfrak{Ig}^{b}_{\tilde{Z},\Xi,O_{C}}$$\widehat{\Xi}_{\tilde{Z},O_{C}}$$\mathfrak{Ig}^{b}_{\tilde{Z},\Xi,\Sigma_{Z},O_{C}}$$\widehat{\Xi}_{\tilde{Z},\Sigma_{Z,O_{C}}}$$\mathfrak{Ig}^{b}_{\tilde{Z},C_{Z},O_{C}}$$\widehat{C}_{\tilde{Z},O_{C}},$}
and $\widehat{\mathfrak{Ig}}^{b,\mathrm{tor}}_{\tilde{Z},O_{C}}$ is isomorphic to $\mathfrak{X}_{\tilde{Z}}/\Gamma_{\tilde{Z}}$, where $\mathfrak{X}_{\tilde{Z}}$ is the completion of $\mathfrak{Ig}^{b}_{\tilde{Z},\Xi,\Sigma_{Z},O_{C}}$ along the boundary.
Moreover, this gives a stratification of $\mathfrak{Ig}^{b,\mathrm{tor}}_{Z,O_{C}}$ into formal toroidal boundary neighborhoods, such that $\widehat{\mathfrak{Ig}}^{b,\mathrm{tor}}_{\tilde{Z},O_{C}}$ cover $\widehat{\mathfrak{Ig}}^{b,\mathrm{tor}}_{Z,O_{C}}$ for $\tilde{Z}$ running over the Igusa cusp labels $\tilde{Z}$ at level $K^{p}(N)$ over $Z$.
Finally, the map $g$ respects the stratification given by cusps, i.e. for every $\tilde{Z}$ we have maps
$g_{\tilde{Z}}:\widehat{\mathfrak{Ig}}^{b,\mathrm{tor}}_{\tilde{Z},O_{C}}\rightarrow\widehat{\mathscr{S}}(\mathbf{G},X)^{\mathrm{tor}}_{K^{p},\tilde{Z},O_{C}},$
induced by the maps in (5) on toroidal boundary charts.
In the above proposition, we used $\tilde{Z}$ to denote both an Igusa cusp label at level $K^{p}(N)$ and a usual cusp label at level $K^{p}(N)$, since we observe the following. To define a cusp label at level $K^{p}$, it suffices to define the filtration on $L\otimes_{\mathbb{Z}}\mathbb{Z}_{p}$ and the isomorphisms $\varphi_{0},\varphi_{-2}$, but given the isomorphism $\alpha:T_{p}(\mathbb{X}_{O_{C}})\simeq L\otimes_{\mathbb{Z}}\mathbb{Z}_{p}$ which we fixed before, we obtain this from $Z_{b}$ and $\tilde{\varphi_{0}}$.
Proof.
This result can be extracted from the construction of $g$ in the proof of [54, Theorem 4.3.8], which we explicate here for the reader’s convenience.
The left column is the formal $p$-adic lift of the left column of (4) in Theorem 2.10, and in particular the composition is given by a $E^{\mathrm{perf}}_{\Phi,O_{C}}$-torsor, where $E^{\mathrm{perf}}_{\Phi,O_{C}}$ is the lift of the perfection of the torus $E^{\mathrm{perf}}_{\Phi}$.
We will now give a more concrete description of the functors defining these lifts. Recall from Proposition 2.11 that the $E^{\mathrm{perf}}_{\Phi}$-torsor $\mathrm{Ig}^{b}_{\tilde{Z},\Xi}\rightarrow\mathrm{Ig}^{b}_{\tilde{Z},C_{Z}}$ can be described as the space parametrizing symmetric lifts $\iota:Y[\frac{1}{p}]\rightarrow\mathcal{G}$ of $\tilde{c}^{\vee}:Y[\frac{1}{p}]\rightarrow\mathcal{B}$. Now, we want to lift these to $O_{C}$. Fix a map $k\rightarrow O_{C}/p$. Recall that we have fixed a lifting $\mathbb{X}_{O_{C}}$ of $\mathbb{X}$, and we also fix as well as an $O_{B}$-linear splitting
$\delta_{\mathbb{X}_{O_{C}}}:\mathbb{X}_{O_{C}}\simeq\mathbb{X}^{\mu}_{O_{C}}\oplus\mathbb{X}^{(0,1)}_{O_{C}}\oplus\mathbb{X}^{\acute{e}t}_{O_{C}}.$
Firstly, by [8, Proposition 4.3.1] there exists some rational $\epsilon\in[0,1]$ such that we have an isomorphism $\mathbb{X}\times_{k}O_{C}/p^{\epsilon}\simeq\mathbb{X}_{O_{C}}\times_{O_{C}}O_{C}/p^{\epsilon}$ of $p$-divisible groups with $G$-structures, lifting the identity, such that
$\delta_{\mathbb{X}_{O_{C}}}\times_{O_{C}}O_{C}/p\simeq\delta_{b}\times_{k}{O_{C}/p^{\epsilon}}.$
Thus, we first base change to $O_{C}/p^{\epsilon}$. Observe that since the quotient $\mathbb{X}/\mathcal{G}[p^{\infty}]$ is étale, it deforms uniquely to a $p$-divisible group over $O_{C}$. Thus, we have a unique lift of $\mathcal{G}[p^{\infty}]$ to $\mathcal{G}[p^{\infty}]_{O_{C}}$ such that the embedding $\rho$ lifts to an embedding $\rho_{O_{C}}:\mathcal{G}[p^{\infty}]_{O_{C}}\hookrightarrow\mathbb{X}_{O_{C}}$. By Serre-Tate theory, this lift $\mathcal{G}[p^{\infty}]$ induces a unique lift of the Raynaud extension
$0\rightarrow T_{O_{C}}\rightarrow\mathcal{G}_{O_{C}}\rightarrow\mathcal{B}_{O_{C}}\rightarrow 0,$
and by similarly considering duals, also a unique lift of the dual Raynaud extension
$0\rightarrow T_{O_{C}}^{\vee}\rightarrow\mathcal{G}^{\vee}_{O_{C}}\rightarrow\mathcal{B}^{\vee}_{O_{C}}\rightarrow 0.$
In particular, we have a map $c_{O_{C}}^{\vee}:Y\rightarrow\mathcal{B}_{O_{C}}$ lifting the map $c^{\vee}$ from Proposition 2.11.
We now lift a splitting of
$T[p^{\infty}]\subset\mathcal{G}[p^{\infty}]\subset\mathbb{X}$
over $O_{C}$, this will define a lifting of $c_{O_{C}}^{\vee}$ to a map $\tilde{c}_{O_{C}}^{\vee}:Y[\frac{1}{p}]\rightarrow\mathcal{B}_{O_{C}}$. Note that given the splitting $\delta_{\mathbb{X}_{O_{C}}}$, to get a splitting of
$T[p^{\infty}]_{O_{C}}\subset\mathscr{G}[p^{\infty}]_{O_{C}}\subset\mathbb{X}_{O_{C}}$
we only need to defined a splitting on the étale parts. This splitting can be defined as a lift of the restriction to étale parts of the splitting mod $p$ as constructed in Proposition 2.11 (1)(a), and thus there is a unique such lift to $O_{C}$.
Thus, we see that the $E^{\mathrm{perf}}_{\Phi,O_{C}}$-torsor $\mathfrak{Ig}^{b}_{\tilde{Z},\Xi,O_{C}}\rightarrow\mathfrak{Ig}^{b}_{\tilde{Z},C_{Z},O_{C}}$ can be described as the torsor of liftings of $\tilde{c}_{O_{C}}^{\vee}$ to $\iota:Y[\frac{1}{p}]\rightarrow\mathcal{G}_{O_{C}}$. Now, we can consider a relative (perfect) toroidal embedding $\mathfrak{Ig}^{b}_{\tilde{Z},\Xi,O_{C}}\hookrightarrow\mathfrak{Ig}^{b}_{\tilde{Z},\Xi,\Sigma_{Z},O_{C}}$ defined using the polyhedral cone $\Xi_{Z}\subset S_{\Phi}[\frac{1}{p}]\otimes\mathbb{R}$, where $S_{\Phi}[\frac{1}{p}]$ is the character lattice of $E_{\Phi}^{\mathrm{perf}}$, and thus also by construction the character lattice of $E^{\mathrm{perf}}_{\Phi,O_{C}}$. Observe that from the uniqueness of relative (perfect) toroidal embeddings, since $\mathfrak{Ig}^{b}_{\tilde{Z},\Xi,\Sigma_{Z},O_{C}}$ is flat over $O_{C}$ it is the unique lift of $\mathrm{Ig}^{b}_{\tilde{Z},\Xi,\Sigma_{Z}}$, since the (perfect) toroidal embedding $\mathrm{Ig}^{b}_{\tilde{Z},\Xi}\hookrightarrow\mathrm{Ig}^{b}_{\tilde{Z},\Xi,\Sigma_{Z}}$ is exactly defined also using the polyhedral cone $\Xi_{Z}$.
Now fix any $m\geq 0$. Let $\tilde{Z}_{m}$ be the associated Igusa cusp label of level $\Gamma_{b}(p^{m})K^{p}(N)$. Note that from $\tilde{Z}_{m}$ we obtain a usual cusp label of level $K(p^{m}N)$ using the isomorphism $X_{\mathcal{O}_{C}}[p^{m}](\mathbb{C})\cong L/p^{m}$. Moreover, the data of an Igusa cusp label of level $\Gamma_{b}(p^{m})K^{p}(N)$ $\tilde{Z}_{m}$ gives us a filtration $Z_{b,m}$ on $\mathbb{X}[p^{m}]$, with an $\mathcal{O}_{B_{\mathbb{Q}_{p}}}$ -splitting $\delta_{b,m}$ of $Z_{b,m}$, and an isomorphism $\psi_{m}:\mathrm{Gr}^{Z_{b,m}}_{0}\simeq Y/p^{m}$.
Thus, we first want to understand a map to $\widehat{\mathscr{S}}^{\mathrm{tor}}_{K(p^{m}N),\,\tilde{Z}_{m},\,O_{C}}.$ For this, firstly note that the prime-to-$p$ structures on both sides are identical, so it suffices to understand the structures at $p$. Consider the symplectic splitting $L/p^{m}\cong\bigoplus_{i=-2}^{0}\mathrm{Gr}^{Z_{p^{m}}}_{i}$ of the filtration we obtain on $L/p^{m}$ using the trivialization $\alpha$ and the splitting $\delta_{b,m}$ of the filtration $Z_{b,m}$, and we assume it is compatible with the splitting $\delta$. Thus, we see that we have a splitting $\mathrm{Gr}^{Z_{b,m}}_{-1}\to Z_{b,m,-1}$, which then induces a splitting of
$0\longrightarrow T[p^{m}]\longrightarrow\mathcal{G}_{{O}_{C}}[p^{m}]\longrightarrow\mathcal{B}_{O_{C}}[p^{m}]\longrightarrow 0,$
and also by duality, a splitting of the dual sequence
$0\rightarrow T^{\vee}[pm]\rightarrow\mathcal{G}^{\vee}_{O_{C}}[p^{m}]\rightarrow\mathcal{B}^{\vee}_{O_{C}}[p^{m}].$
We thus obtain lifts of the maps defining the Raynaud extension and its dual, $c$ and $c^{\vee}$ to maps $c_{m}:\frac{1}{p^{m}}X\rightarrow\mathcal{B}^{\vee}_{O_{C}}$ and $c_{m}^{\vee}:\frac{1}{p^{m}}Y\rightarrow\mathcal{B}_{O_{C}}$. Thus, we have a map to the abelian scheme $C_{\tilde{Z}_{m},O_{C}}$ on the toroidal boundary chart defining $\mathscr{S}^{\mathrm{tor}}_{K(p^{m}N),\tilde{Z}_{m}}$. Now, the torus torsor $\Xi_{\tilde{Z}_{m},O_{C}}$ is defined as the moduli space parametrizing lifts $\iota_{m}:\frac{1}{p^{m}}Y\rightarrow\mathcal{G}_{O_{C}}$.
Since $\mathfrak{Ig}^{b}_{\tilde{Z},\Xi,O_{C}}\rightarrow\mathfrak{Ig}^{b}_{\tilde{Z},C_{Z},O_{C}}$ can be described as the torsor of liftings of $\tilde{c}_{O_{C}}^{\vee}$ to $\iota:Y[\frac{1}{p}]\rightarrow\mathcal{G}_{O_{C}}$, we obtain a map from $\mathfrak{Ig}^{b}_{\tilde{Z},\Xi,O_{C}}$ to $\Xi_{\tilde{Z}_{m},O_{C}}$. Observe that this extends to a map of relative toroidal embeddings $\mathfrak{Ig}^{b}_{\tilde{Z},\Xi,\Sigma_{Z},O_{C}}$ to $\Xi_{\tilde{Z}_{m},\Sigma_{Z},O_{C}}$, since a map of torus torsors extends exactly when the induced map on polyhedral cones maps every face in one into a face of the other. This last condition clearly holds true for the map $\mathfrak{Ig}^{b}_{\tilde{Z},\Xi,O_{C}}\rightarrow\Xi_{\tilde{Z}_{m},O_{C}}$, since we in fact get the identity map on the polyhedral cone $\Sigma_{Z}$.
By taking limits over $m$, we get the desired maps. ∎
Finally, we want to understand the inclusion of adic generic fibers $\partial\mathfrak{Ig}^{b,\mathrm{tor}}_{C}\subset\mathfrak{Ig}^{b,\mathrm{tor}}_{C}$. Let $\widehat{\partial\mathscr{S}}(\mathbf{G},X)^{\mathrm{tor}}_{K,O_{C}}$ denote the (base-change over to $O_{C}$ of the) closed formal subscheme of $\widehat{\mathscr{S}}(\mathbf{G},X)^{\mathrm{tor}}_{K,O_{C}}$ given by taking the formal completion of the boundary component $\partial\mathscr{S}(\mathbf{G},X)_{K}$ (which is flat over $O_{E_{\mathfrak{p}}}$) along its special fiber, then we may consider the formal subscheme
$\tilde{\partial}\mathfrak{Ig}^{b,\mathrm{tor}}_{O_{C}}:=\mathfrak{Ig}^{b,\mathrm{tor}}_{O_{C}}\times_{\widehat{\mathscr{S}}(\mathbf{G},X)_{K^{p},O_{C}}}\widehat{\partial\mathscr{S}}(\mathbf{G},X)^{\mathrm{tor}}_{K^{p},O_{C}}.$
Since the inclusion $\widehat{\partial\mathscr{S}}(\mathbf{G},X)^{\mathrm{tor}}_{K,O_{C}}\hookrightarrow\widehat{\mathscr{S}}(\mathbf{G},X)_{K,O_{C}}$ is an adic morphism of $p$-adic formal schemes, taking limits over $K_{p}$ and base-changing we see that the induced map $\tilde{\partial}\mathfrak{Ig}^{b,\mathrm{tor}}_{O_{C}}\hookrightarrow\mathfrak{Ig}^{b,\mathrm{tor}}_{O_{C}}$ is also an adic morphism, and in particular $\tilde{\partial}\mathfrak{Ig}^{b,\mathrm{tor}}_{O_{C}}$ is also a $p$-adic formal scheme.
Finally, observe that by the previous Proposition we know that image of an Igusa boundary component labelled by $\tilde{Z}$ maps to a boundary component of the Shimura variety labelled by $\tilde{Z}$, hence the perfection of the special fiber of $\tilde{\partial}\mathfrak{Ig}^{b,\mathrm{tor}}_{O_{C}}$ is
$\left(\mathrm{Ig}^{b,\mathrm{tor}}\times_{\mathscr{S}(\mathbf{G},X)_{K^{p},\overline{\mathbb{F}}_{p}}}\partial\mathscr{S}(\mathbf{G},X)^{\mathrm{tor}}_{K^{p},\overline{\mathbb{F}}_{p}}\right)^{\mathrm{perf}}\simeq\partial\mathrm{Ig}^{b,\mathrm{tor}}.$
Proposition 2.13.
We have a closed inclusion of $p$-adic formal schemes
$\partial\mathfrak{Ig}^{b,\mathrm{tor}}_{O_{C}}\subset\tilde{\partial}\mathfrak{Ig}^{b,\mathrm{tor}}_{O_{C}},$
such that both have the same reduced special fiber. Moreover, this inclusion induces an equality of $\mathrm{Spf}O_{C^{\prime}}$-points, for $C^{\prime}$ any complete algebraically closed non-archimedean field containing $C$.
Proof.
We will first show that every $\overline{\mathbb{F}}_{p}$-point of the special fiber of $\tilde{\partial}\mathfrak{Ig}^{b,\mathrm{tor}}_{O_{C}}$ lifts. To see this, we may consider some toroidal boundary component $\mathrm{Ig}^{b,\mathrm{tor}}_{\tilde{Z}}$, and we want to show that in fact this entire component admits a flat lift over $\mathrm{Spf}O_{C}$, to a locally closed $p$-adic formal subscheme of $\tilde{\partial}\mathfrak{Ig}^{b,\mathrm{tor}}$.
To see this, we claim that the map
$g_{\tilde{Z}}:\mathfrak{Ig}^{b,\mathrm{tor}}_{\tilde{Z},O_{C}}\rightarrow\widehat{\mathscr{S}}(\mathbf{G},X)^{\mathrm{tor}}_{K^{p},\tilde{Z},O_{C}},$
respects boundary components. For this, we will look at the moduli interpretations of both sides, and we want to show that the map in the middle row of (5) respects the boundary. On the left hand side, from the previous Proposition and the proof of [54, Theorem 4.3.10], for any $p$-complete $O_{C}$-algebra $R$, we see that $\mathrm{Spf}R$-points of this are prescribed by the following data:
  1. (1)
    An abelian scheme $\mathcal{B}$ over $R$ with $\mathcal{O}_{B}$-action and a prime-to-$p$ polarization
  2. (2)
    An $\mathcal{O}_{B}$ -linear extension
    $0\rightarrow T\rightarrow\mathcal{G}\rightarrow\mathcal{B}\rightarrow 0$
    where $\mathbb{X}_{*}(T)=X$; equivalently, an $\mathcal{O}_{B}$ -linear map $c:X\rightarrow\mathcal{B}^{\vee}$.
  3. (3)
    An $\mathcal{O}_{B_{\mathbb{Q}_{p}}}$-linear isomorphism
    $\rho:\mathcal{G}[p^{\infty}]\simeq Z_{b,-1}$
    that extends the isomorphism $T[p^{\infty}]\simeq Z_{b,-2}$ obtained by applying Cartier duality to $\tilde{\varphi}_{0}$. By the splitting $\delta_{b}$, this induces a splitting of
    $0\rightarrow T[p^{\infty}]\rightarrow\mathcal{G}[p^{\infty}]\rightarrow\mathcal{B}[p^{\infty}]\rightarrow 0,$
    and in particular $c$ extends to a map $c:X[1/p]\rightarrow\mathcal{B}^{\vee}$.
  4. (4)
    An $\mathcal{O}_{B}$ -linear extension
    $0\rightarrow T^{\vee}\rightarrow\mathcal{G}^{\vee}\rightarrow\mathcal{B}^{\vee}\rightarrow 0$
    where $\mathbb{X}^{*}(T)=Y$. Equivalently, an $\mathcal{O}_{B}$ -linear map $c^{\vee}:Y\rightarrow\mathcal{B}$. By the splitting $\delta_{b}$, as well as by duality (using that $\mathcal{B}[p^{\infty}]$ will be principally polarized), we have a splitting of
    $0\rightarrow T^{\vee}[p^{\infty}]\rightarrow\mathcal{G}^{\vee}[p^{\infty}]\rightarrow\mathcal{B}^{\vee}[p^{\infty}]\rightarrow 0$
    and in particular we extend $c^{\vee}$ to a map $c^{\vee}:Y[1/p]\rightarrow\mathcal{B}$.
  5. (5)
    An $R$ -point of $\mathcal{P}^{\prime}_{\Sigma_{Z}}$. Here, we note that away from the boundary we have a torsor $\mathcal{P}^{\prime}$ over $R$ for the torus $E_{\Phi}$ with character group $S_{\Phi}$, parametrizing lifts of $c^{\vee}$ to $\iota:Y[1/p]\rightarrow\mathcal{G}$. $\mathcal{P}^{\prime}_{\Sigma_{Z}}\supset\mathcal{P}^{\prime}$ is the toroidal embedding defined by the admissible rational polyhedral cone decomposition $\Sigma_{Z}$ for the cusp $(Z,\Phi)$.
Moreover, we see that the boundary component is given by $\mathrm{Spf}R$-points which in (5) lie entirely in the boundary $\mathcal{P}^{\prime}_{\Sigma_{Z}}\backslash\mathcal{P}^{\prime}$. By construction, the image of these boundary points under $g$ lies in $\Xi_{Z,\Sigma_{\tilde{Z}},O_{C}}\backslash\Xi_{Z,O_{C}}$, since being in the boundary of $\Xi_{Z,\Sigma_{\tilde{Z}},O_{C}}$ is also exactly determined by the data in (5) of the $R$-point of $\mathcal{P}^{\prime}_{\Sigma_{Z}}$.
Thus, observe that for the formal toroidal boundary components $\partial\mathfrak{Ig}^{b}_{\tilde{Z},\Xi,\Sigma_{Z},O_{C}}$ and $\partial\Xi_{Z,\Sigma_{Z},O_{C}}:=\Xi_{Z,\Sigma_{Z},O_{C}}\backslash\Xi_{Z,O_{C}}$ we have a commutative diagram
$\partial\mathfrak{Ig}^{b}_{\tilde{Z},\Xi,\Sigma_{Z},O_{C}}$$\widehat{\partial\Xi}_{\tilde{Z},\Sigma_{Z},O_{C}}$$\mathfrak{Ig}^{b}_{\tilde{Z},C_{\tilde{Z}},O_{C}}$$\widehat{C}_{\tilde{Z},O_{C}},$}
where on the right column we take completions along the special fiber. We see that the quotient
$\partial\mathfrak{Ig}^{b}_{\tilde{Z},\Xi,\Sigma_{Z},O_{C}}/\Gamma_{\tilde{Z}}$
is a flat $p$-adic formal subscheme of $\tilde{\partial}\mathfrak{Ig}^{b,\mathrm{tor}}_{O_{C}}$. Moreover, since the image of $\partial\mathfrak{Ig}^{b}_{\tilde{Z},\Xi,\Sigma_{Z},O_{C}}$ under $g$ by construction lies in $\widehat{\partial\Xi}_{Z,\Sigma_{Z},O_{C}}$, and under the identification of formal toroidal neighborhoods of $\widehat{\mathscr{S}}(\mathbf{G},X)^{\mathrm{tor}}_{K^{p},\tilde{Z},O_{C}}$ with $\widehat{\Xi}_{Z,\Sigma_{Z},O_{C}}/\Gamma_{\tilde{Z}}$ this lies in the boundary $\widehat{\partial\mathscr{S}}(\mathbf{G},X)^{\mathrm{tor}}_{K^{p},\tilde{Z},O_{C}}$.
To see the second part of the proposition, we may consider the closed formal subscheme $\mathfrak{X}\subset\tilde{\partial}\mathfrak{Ig}^{b,\mathrm{tor}}_{O_{C}}$ obtained by quotienting by the $p^{\infty}$-torsion ideal of $\mathcal{O}_{\tilde{\partial}\mathfrak{Ig}^{b,\mathrm{tor}}_{O_{C}}}$, this is a flat closed $p$-adic formal subscheme of $\tilde{\partial}\mathfrak{Ig}^{b,\mathrm{tor}}_{O_{C}}$ whose reduced special fiber is necessarily $\partial\mathrm{Ig}^{b,\mathrm{tor}}$, since it contains every geometric point of $\partial\mathrm{Ig}^{b,\mathrm{tor}}$. Observe that every $\mathrm{Spf}O_{C^{\prime}}$-point of $\tilde{\partial}\mathfrak{Ig}^{b,\mathrm{tor}}_{O_{C}}$ necessarily factors through $\mathfrak{X}$, since $O_{C^{\prime}}$ is $p$-torsionfree. Finally, observe that since $\mathfrak{X}$ is flat, and the special fiber of $\mathfrak{X}$ contains as a reduced closed subscheme $\partial\mathrm{Ig}^{b,\mathrm{tor}}$, $\mathfrak{X}$ contains the lift $\partial\mathfrak{Ig}^{b,\mathrm{tor}}_{O_{C}}$ as a closed formal subscheme. Now, any $\mathrm{Spf}O_{C^{\prime}}$-point of $\mathfrak{X}$ must also necessarily factor through $\partial\mathfrak{Ig}^{b,\mathrm{tor}}_{O_{C}}$, since its image mod $p$ is reduced. Thus, the set of $\mathrm{Spf}O_{C^{\prime}}$-points of $\tilde{\partial}\mathfrak{Ig}^{b,\mathrm{tor}}_{O_{C}}$ and $\partial\mathfrak{Ig}^{b,\mathrm{tor}}_{O_{C}}$ are the same. ∎
3. Mantovan’s Formula and the Hodge-Tate Period Morphism
We now review some basic facts about the geometry of the Hodge-Tate period morphism, establishing some facts about the fiber of the Hodge-Tate period morphism in §3.1. In §3.2, we then apply these basic geometric facts to establish a version of Mantovan’s product formula for non-compact Shimura varieties (Theorem 5.1). This will in particular allow us to bound the degrees of the torsion cohomology of the Shimura variety in terms of the semiperversity of certain sheaves coming from Igusa varieties and the perverse t-exactness of Hecke operators on $\mathrm{Bun}_{G}$.
We recall that $(\mathbf{G},X)$ denotes a Shimura datum with reflex field $E/\mathbb{Q}$. We let $\mathrm{Sh}(\mathbf{G},X)_{K}/\mathop{\rm Spec}(E)$ denote the associated Shimura variety for a sufficiently small open compact subgroup $K\subset\mathbf{G}(\mathbb{A}_{f})$. We will assume throughout that the Shimura datum $(\mathbf{G},X)$ is of PEL type AC.
We recall that $\ell\neq p$ are distinct primes, $E_{\mathfrak{p}}$ is the completion at the place $\mathfrak{p}|p$ determined by a fixed choice of isomorphism $j:\overline{\mathbb{Q}}_{p}\simeq\mathbb{C}$, and $C$ is the completed algebraic closure of $E_{\mathfrak{p}}$. The group $G:=\mathbf{G}_{\mathbb{Q}_{p}}$ will denote the local group, and we let $K=K^{p}K_{p}$ be the decomposition into the level $K^{p}\subset\mathbf{G}(\mathbb{A}_{f}^{p})$ away from $p$ (resp. the level $K_{p}\subset G(\mathbb{Q}_{p})$ at $p$).
3.1. The Hodge-Tate Period Morphism
For $K\subset\mathbf{G}(\mathbb{A}_{f})$ a sufficiently small open compact, we define $\mathcal{S}(\mathbf{G},X)_{K}:=(\mathrm{Sh}(\mathbf{G},X)_{K}\otimes_{E}E_{\mathfrak{p}})^{\mathrm{ad}}$ to be the adic space over $\mathop{\rm Spa}(E_{\mathfrak{p}})$ attached to the Shimura variety. For brevity, in this section we will often denote this simply by $\mathcal{S}_{K}$. When $K=K_{p}^{\mathrm{hs}}K^{p}$ with $K_{p}^{\mathrm{hs}}$ a hyperspecial subgroup, the space $\mathcal{S}_{K}$ has a canonical integral model $\mathscr{S}_{K}$ over $\mathcal{O}_{E,\mathfrak{p}}$. Let $\mathcal{S}_{K}^{\circ}\subset\mathcal{S}_{K}$ be the good reduction locus, i.e. the open subspace of $\mathcal{S}_{K}$ obtained from the adic generic fiber of the $p$-adic completion $\mathscr{S}^{\wedge}_{K}$ of the scheme $\mathscr{S}_{K}$. We define $\mathcal{S}_{K^{\prime}}^{\circ}\subset\mathcal{S}_{K^{\prime}}$ for $K^{\prime}\subset K$ by taking the preimage under the natural map from $\mathcal{S}_{K^{\prime}}$ to $\mathcal{S}_{K}$. We also consider the adic spaces $\mathcal{S}^{*}_{K}:=(\mathrm{Sh}(\mathbf{G},X)^{*}_{K}\otimes_{E}E_{\mathfrak{p}})^{\mathrm{ad}}$ and $\mathcal{S}^{\mathrm{tor}}_{K}:=(\mathrm{Sh}(\mathbf{G},X)^{\mathrm{tor}}_{K}\otimes_{E}E_{\mathfrak{p}})^{\mathrm{ad}}$ attached to the minimal and toroidal compactifications of the Shimura variety $\mathrm{Sh}(\mathbf{G},X)_{K}$ over $E$, respectively, as described in §2.1.3, where, for the former, we fix a compatible choice $\Sigma=\{\Sigma_{\Phi}\}$ of admissible smooth rational polyhedral cone decomposition data, as described in loc.cit. We use the superscript $(-)^{\Diamond}$ for the diamond attached to these analytic adic spaces, as defined in [55, Lecture X].
Associated to the $\mathbf{G}(\mathbb{R})$-conjugacy class $X$, we have a conjugacy class of a minuscule cocharacter $\mu$ of $G_{C}$ with reflex field $E_{\mathfrak{p}}$ attached to the inverse of the Hodge cocharacter of $X$. Let $\mathcal{F}\ell_{G,\mu^{-1}}$ be the adic flag variety over $\mathop{\rm Spa}(C)$ associated to the conjugacy class of $\mu^{-1}$, the dominant inverse of $\mu$ (i.e we quotient out by the parabolic defined by the set of $g\in G_{C}$ such that $\mathop{\rm lim}\nolimits_{t\rightarrow 0}\mathrm{ad}(\mu(t))g$ exists). Since $\mu$ is minuscule, via the Bialynicki-Birula isomorphism ([7, Theorem 3.4.5]), the diamond attached to the flag variety $\mathcal{F}\ell_{G,\mu^{-1}}^{\Diamond}$ represents the following functor on $\mathrm{Perf}_{C}$. Given any $S\in\mathrm{Perf}_{C}$, $\mathcal{F}\ell_{G,\mu^{-1}}^{\Diamond}(S)$ is the set of modifications of vector bundles $\mathcal{E}\dashrightarrow\mathcal{E}_{0}$ of meromorphy $\mu$ on $X_{S}$, the relative Fargues-Fontaine curve over $S$, such that the modification occurs over the degree one Cartier divisor in $X_{S}$ corresponding to the untilt of $S$ defined by the map $S\rightarrow\mathop{\rm Spd}(C)$.
Let
$\mathcal{S}_{K^{p},C}^{\circ}:=\mathop{\rm lim}\nolimits\limits_{\begin{subarray}{c}\longleftarrow\\ K_{p}\end{subarray}}\mathcal{S}_{K^{p}K_{p},C}^{\circ,\Diamond}\subset\mathcal{S}_{K^{p},C}:=\mathop{\rm lim}\nolimits\limits_{\begin{subarray}{c}\longleftarrow\\ K_{p}\end{subarray}}\mathcal{S}_{K^{p}K_{p},C}^{\Diamond}\subset\mathcal{S}_{K^{p},C}^{\mathrm{tor}}:=\mathop{\rm lim}\nolimits\limits_{\begin{subarray}{c}\longleftarrow\\ K_{p}\end{subarray}}\mathcal{S}^{\mathrm{tor},\Diamond}_{K^{p}K_{p},C}\rightarrow\mathcal{S}_{K^{p},C}^{*}:=\mathop{\rm lim}\nolimits\limits_{\begin{subarray}{c}\longleftarrow\\ K_{p}\end{subarray}}\mathcal{S}^{*,\Diamond}_{K^{p}K_{p},C}$
be the associated perfectoid Shimura varieties [^4] If $(\mathbf{G},X)$ is more generally of Hodge type then this is representable by a perfectoid space by [57, Theorem IV.1.1], and in general they have the structure of diamonds (See [28] and [50, Section 2]), where this limit is computed in the category of sheaves on $\mathrm{Perf}_{C}$ with respect to the $v$-topology on $\mathrm{Perf}_{C}$, as defined in [56, Definition 8.1] and we recall that any diamond defines such a $v$-sheaf in light of [56, Proposition 11.9]. Throughout, we identify adic spaces attached to perfectoids and their associated $v$-sheaves under the Yoneda embedding, recalling that the $v$-topology is sub-canonical, by [56, Theorem 1.2].
We also consider $\overline{\mathcal{S}}^{\circ}_{K^{p},C}$, the canonical compactification of the good reduction locus over $\mathop{\rm Spd}(C)$ in the sense of [56, Proposition 18.6]. This will be a subspace of $\mathcal{S}_{K^{p},C}$, since $\mathcal{S}_{K^{p},C}$ is partially proper. Caraiani-Scholze [7, §2.1] consider the Hodge-Tate period morphism on $\mathcal{S}_{K^{p}}$
$\pi_{\mathrm{HT}}:\mathcal{S}_{K^{p},C}\rightarrow\mathcal{F}\ell_{G,\mu^{-1}}^{\Diamond},$
which records the relative position of the Hodge-Tate filtration associated with the abelian varieties that $\mathcal{S}_{K^{p},C}$ parametrizes. This extends [8, §4.1] to a Hodge-Tate period morphism on the minimal compactification
$\pi_{\mathrm{HT}}^{*}:\mathcal{S}_{K^{p},C}^{*}\rightarrow\mathcal{F}\ell_{G,\mu^{-1}}^{\Diamond}$
and toroidal compactification
$\pi_{\mathrm{HT}}^{\mathrm{tor}}:\mathcal{S}_{K^{p},C}^{\mathrm{tor}}\rightarrow\mathcal{F}\ell_{G,\mu^{-1}}^{\Diamond}.$
We write $\pi_{\mathrm{HT}}^{\circ}$ for the restriction to the good reduction locus, and $\overline{\pi}_{\mathrm{HT}}^{\circ}$ for the canonical compactification of $\pi_{\mathrm{HT}}^{\circ}$, where we note that this again maps to $\mathcal{F}\ell_{G,\mu^{-1}}^{\Diamond}$ as this is proper over $\mathop{\rm Spa}(C)$.
These maps have the following properties:
  1. (1)
    $\overline{\pi}_{\mathrm{HT}}^{\circ}$, $\pi_{\mathrm{HT}}^{*}$, and $\pi_{\mathrm{HT}}^{\mathrm{tor}}$ are partially proper and qcqs; hence, proper.
  2. (2)
    $\pi_{\mathrm{HT}}$ is partially proper, but not always qcqs (Unless the Shimura variety is compact).
  3. (3)
    $\pi_{\mathrm{HT}}^{\circ}$ is qcqs, but not partially proper (Unless the Shimura variety is compact).
With these properties in mind, let us study the fibers of these maps. For our purposes, we will focus on the compactly supported cohomology of $\mathcal{S}_{K^{p},C}$ and in turn the sheaf $R\pi_{\mathrm{HT}!}(\overline{\mathbb{F}}_{\ell})$ on $\mathcal{F}\ell_{G,\mu^{-1}}^{\Diamond}$.
Our goal is to describe the stalks of $R\pi_{\mathrm{HT}!}(\overline{\mathbb{F}}_{\ell})$ at a geometric rank $1$ point $x:\mathop{\rm Spa}(C,O_{C})\rightarrow\mathcal{F}\ell^{\Diamond}_{G,\mu^{-1}}$. We assume the geometric point $x$ factors through the adic Newton strata $\mathcal{F}\ell_{G,\mu^{-1}}^{b,\Diamond}$ for $b\in B(G,\mu)$, and choose a completely slope divisible $p$-divisible group $\mathbb{X}_{b}$ over $\overline{\mathbb{F}}_{p}$ corresponding to $b\in B(G,\mu)$. Let $\mathrm{Ig}^{b}$ be the associated perfect Igusa variety as defined in §2.2.1, with toroidal compactification $\mathrm{Ig}^{b,\mathrm{tor}}$ and minimal compactification $\mathrm{Ig}^{b,*}$. Recall that we have associated perfectoid Igusa varieties, $\mathfrak{Ig}^{b}_{C}$$\mathfrak{Ig}^{b,\mathrm{tor}}_{C}$, and $\mathfrak{Ig}^{b,*}_{C},$ which should model the fibers of $\overline{\pi}_{\mathrm{HT}}^{\circ}$, $\pi_{\mathrm{HT}}^{\mathrm{tor}}$, and $\pi^{*}_{\mathrm{HT}}$, respectively. We let $\partial\mathfrak{Ig}_{C}^{b,*}$ and $\partial\mathfrak{Ig}_{C}^{b,\mathrm{tor}}$ be the Zariski closed subspaces attached to the boundaries $\partial\mathrm{Ig}^{b,*}$ and $\partial\mathrm{Ig}^{b,\mathrm{tor}}$, respectively.
Let $g_{b}:\mathrm{Ig}^{b}\hookrightarrow\mathrm{Ig}^{b,*}$ be the natural open immersion of $\overline{\mathbb{F}}_{p}$-schemes. We define the partially compactly supported cohomology
$R\Gamma_{c-\partial}(\mathrm{Ig}^{b},\overline{\mathbb{F}}_{\ell}):=R\Gamma(\mathrm{Ig}^{b,*},g_{b!}(\overline{\mathbb{F}}_{\ell})).$
Our goal is to show that this computes the fibers of $R\pi_{\mathrm{HT}!}(\overline{\mathbb{F}}_{\ell})$ at geometric points of $\mathcal{F}\ell_{G,\mu^{-1}}^{b,\Diamond}$.
To get a clearer picture of how these spaces interact with each other, we have the following theorem.
Theorem 3.1.
[8, 54] For a geometric rank $1$ point, $x:\mathop{\rm Spa}(C,O_{C})\rightarrow\mathcal{F}\ell_{G,\mu^{-1}}^{\Diamond}$, there exists a diagram of spaces of the form
$(\pi_{\mathrm{HT}}^{\circ})^{-1}(x)$$\pi_{\mathrm{HT}}^{-1}(x)$$(\pi_{\mathrm{HT}}^{\mathrm{tor}})^{-1}(x)$$(\pi_{\mathrm{HT}}^{*})^{-1}(x)$$\mathfrak{Ig}^{b}_{C}$$\mathfrak{Ig}^{b,\mathrm{tor}}_{C},$$\mathfrak{Ig}^{b,*}_{C}$$\scriptstyle{i}$$\scriptstyle{\phantom{}{}^{\mathrm{tor}}i}$$\scriptstyle{\phantom{}{}^{*}i}$
where the maps $i$, $\phantom{}{}^{*}i$, and $\phantom{}{}^{\mathrm{tor}}i$ are open immersions whose image contains all rank $1$ points. Moreover, the fibers $(\pi_{\mathrm{HT}}^{*})^{-1}(x)$ and $(\pi_{\mathrm{HT}}^{\mathrm{tor}})^{-1}(x)$ are partially proper, so in particular $\phantom{}{}^{*}i$ and $\phantom{}{}^{\mathrm{tor}}i$ are canonical compactifications in the sense of [56, Proposition 18.6].
Proof.
This theorem in the case where $(\mathbf{G},X)$ is of PEL type $A$ attached to a globally quasi-split unitary group of even dimension is [8, Theorems 2.7.2,Theorem 4.5.1], and the general case of PEL type AC is proven in [54, Theorems 4.3.10,4.3.12]. ∎
We will also combine this with the following result.
Theorem 3.2.
[8, 54] The partially minimally compactified Igusa variety $\mathrm{Ig}^{b,*}$ is affine; in particular, the attached adic space $\mathfrak{Ig}^{b,*}_{C}$ is affinoid perfectoid. Moreover, there exists a proper map
$\mathfrak{Ig}^{b,\mathrm{tor}}_{C}\rightarrow\mathfrak{Ig}^{b,*}_{C}$
which induces an isomorphism on global sections.
Proof.
For the affineness, the case of PEL type $A$ attached to a global quasi-split unitary group of even dimension is covered by [8, Theorem 1.7] and [8, Lemma 4.5.2]. The general case of PEL type AC is covered in [54, Lemma 3.3.7]. To see the properness, we note that the map
$\mathfrak{Ig}_{C}^{b,\mathrm{tor}}\rightarrow\mathfrak{Ig}_{C}^{b,*}$
is the one appearing in the Stein factorization described in [8, Proposition 3.3.4] and [54, Proposition 3.3.5]. This also shows that one has an isomorphism on global sections. ∎
We will also need the following Corollary.
Corollary 3.3.
The boundary $\partial\mathfrak{Ig}_{C}^{b,\mathrm{tor}}$ is quasi-compact and $\partial\mathfrak{Ig}_{C}^{b,*}$ is affinoid perfectoid (In particular, quasi-compact).
Proof.
The fact that $\mathfrak{Ig}_{C}^{b,*}$ is affinoid perfectoid follows from the previous Theorem. Moreover, $\partial\mathfrak{Ig}_{C}^{b,*}$ is a Zariski closed subspace, since it came from considering the adic generic fiber of a formal model of the perfect closed subscheme $\partial\mathrm{Ig}^{b,*}\subset\mathrm{Ig}^{b,*}$, and a Zariski closed subspace of a perfectoid is still perfectoid by [3, Remark 7.5]. The claim for the toroidal compactification follows since the map $\mathfrak{Ig}_{C}^{b,\mathrm{tor}}\rightarrow\mathfrak{Ig}_{C}^{b,*}$ is proper, and maps the boundary $\partial\mathfrak{Ig}_{C}^{b,\mathrm{tor}}$ to $\partial\mathfrak{Ig}_{C}^{b,*}$. ∎
It is natural to wonder how one could describe the fiber of $\pi_{\mathrm{HT}}$ in terms of the spaces described above. In particular, we now deduce the following Corollary.
Corollary 3.4.
For $x:\mathop{\rm Spa}(C,O_{C})\rightarrow\mathcal{F}\ell_{G,\mu^{-1}}^{b,\Diamond}$ a geometric point, we have isomorphisms
$\pi_{\mathrm{HT}}^{-1}(x)\simeq\overline{\mathfrak{Ig}}_{C}^{b,*}\setminus\overline{\partial\mathfrak{Ig}}_{C}^{b,*}\simeq\overline{\mathfrak{Ig}}_{C}^{b,\mathrm{tor}}\setminus\overline{\partial\mathfrak{Ig}}_{C}^{b,\mathrm{tor}}$
induced by the natural open immersions $\pi_{\mathrm{HT}}^{-1}(x)\hookrightarrow(\pi_{\mathrm{HT}}^{*})^{-1}(x)\simeq\overline{\mathfrak{Ig}}_{C}^{b,*}$ (resp. $\pi_{\mathrm{HT}}^{-1}(x)\hookrightarrow(\pi_{\mathrm{HT}}^{\mathrm{tor}})^{-1}(x)\simeq\overline{\mathfrak{Ig}}_{C}^{b,\mathrm{tor}}$), as given by Theorem 3.1. Here $\overline{\partial\mathfrak{Ig}}_{C}^{b,*}$ (resp. $\overline{\partial\mathfrak{Ig}}_{C}^{b,\mathrm{tor}}$) is the Zarisiki closed subset of $\overline{\mathfrak{Ig}}_{C}^{b,*}$ (resp. $\overline{\mathfrak{Ig}}_{C}^{b,\mathrm{tor}}$) defined by the canonical compactification of the boundary $\partial\mathfrak{Ig}^{b}_{C}\subset\mathfrak{Ig}^{b}_{C}$ (resp. $\partial\mathfrak{Ig}_{C}^{b,\mathrm{tor}}\subset\mathfrak{Ig}^{b,\mathrm{tor}}_{C}$).
Proof.
We first establish the claim for the toroidal compactification. We consider the closed immersion
$\overline{\mathfrak{Ig}}_{C}^{b,\mathrm{tor}}\times_{\mathcal{S}^{\mathrm{tor}}_{K^{p},C}}\partial\mathcal{S}^{\mathrm{tor}}_{K^{p},C}\hookrightarrow\overline{\partial\mathfrak{Ig}}_{C}^{b,\mathrm{tor}}$
obtained by base-changing the closed immersion $\partial\mathcal{S}^{\mathrm{tor}}_{K^{p},C}\hookrightarrow\mathcal{S}^{\mathrm{tor}}_{K^{p},C}$ to the fiber $(\pi_{\mathrm{HT}}^{\mathrm{tor}})^{-1}(x)$ and applying Theorem 3.1. To show the desired claim, it suffices to show this is an isomorphism. We note that both the RHS and LHS are perfectoid by [3, Remark 7.5], since they are Zariski closed in the perfectoid space $\overline{\mathfrak{Ig}}_{C}^{b,\mathrm{tor}}$. Moreover, both the LHS and RHS are partially proper; therefore, to show this map is an isomorphism, it suffices to show it induces an isomorphism on rank one geometric points using [56, Lemma 5.4] (where we note that condition (iv) of [56, Lemma 5.4] will be implied by the claim on rank one points and the partial properness). In particular, given $C^{\prime}/C$ a complete algebraically closed non-archimedean field, we claim that there exists a Cartesian diagram of the form
[…]
Observe that since both $\widehat{\mathscr{S}}(\mathbf{G},X)_{K^{p}}^{\mathrm{tor}}$ and the boundary $\widehat{\partial\mathscr{S}}(\mathbf{G},X)_{K^{p}}^{\mathrm{tor}}$ are proper, we have that the $\mathop{\rm Spa}(C^{\prime},O_{C^{\prime}})$ points of the adic generic fiber $\widehat{\mathscr{S}}(\mathbf{G},X)_{K^{p}}^{\mathrm{tor}}$ and $\mathcal{S}^{\mathrm{tor}}_{K^{p}}$ are the same, and similarly for the adic generic fiber of $\widehat{\partial\mathscr{S}}(\mathbf{G},X)^{\mathrm{tor}}_{K^{p}}$ and $\partial\mathcal{S}^{\mathrm{tor}}_{K^{p},C}$. Moreover, the $\mathop{\rm Spa}(C^{\prime},O_{C^{\prime}})$-points of $\widehat{\mathscr{S}}(\mathbf{G},X)_{K^{p}}^{\mathrm{tor}}$ are also the same as the $\mathrm{Spf}O_{C^{\prime}}$-points of the formal scheme $\widehat{\mathscr{S}}(\mathbf{G},X)_{K^{p},O_{C}}^{\mathrm{tor}}$ . A similar result also holds for the boundary. Thus, we can apply Proposition 2.13 to conclude, since $\tilde{\partial}\mathfrak{Ig}^{b,\mathrm{tor}}_{O_{C}}$ is defined as the fiber product of formal schemes $\mathfrak{Ig}^{b,\mathrm{tor}}_{O_{C}}\times_{\widehat{\mathscr{S}}(\mathbf{G},X)_{K^{p},O_{C}}}\widehat{\partial\mathscr{S}}(\mathbf{G},X)^{\mathrm{tor}}_{K^{p},O_{C}}$.
It remains to see the analogous claim for the minimal compactification. This follows easily using Theorem 3.1 and the fact that the proper surjective map $\mathfrak{Ig}_{C}^{b,\mathrm{tor}}\rightarrow\mathfrak{Ig}_{C}^{b,*}$ sends $\partial\mathfrak{Ig}_{C}^{b,\mathrm{tor}}$ to $\partial\mathfrak{Ig}_{C}^{b,*}$ by construction. ∎
We have the following corollary.
Corollary 3.5.
For a geometric point $x:\mathop{\rm Spa}(C,O_{C})\rightarrow\mathcal{F}\ell_{G,\mu^{-1}}^{b,\Diamond}$, we have an identification:
$R\Gamma_{c-\partial}(\mathrm{Ig}^{b},\overline{\mathbb{F}}_{\ell})\simeq R\pi_{\mathrm{HT}!}(\overline{\mathbb{F}}_{\ell})_{x}.$
Proof.
We have an identification $\pi_{\mathrm{HT}}^{-1}(x)\simeq\overline{\mathfrak{Ig}}^{b,*}_{C}\setminus\overline{\partial\mathfrak{Ig}}^{b,*}_{C}$ by the previous corollary, so, by proper base-change, we are tasked with computing the compactly supported cohomology of this space. We note, by Theorem 3.2, the adic spaces $\mathfrak{Ig}^{b,*}_{C}$ are affinoid perfectoid. It follows that the canonical compactification $\overline{\mathfrak{Ig}}^{b,*}_{C}\simeq(\pi_{\mathrm{HT}}^{*})^{-1}(x)$ is also affinoid perfectoid, by [56, Proposition 18.7 (iv)]. In particular, it is quasi-compact and partially proper, so in particular proper. It therefore follows by excision [^5] One easily checks that the excision sequence is exact on geometric points, and this is sufficient by [56, Proposition 14.3]. that we have a distinguished triangle
(6) $R\Gamma_{c}(\pi_{\mathrm{HT}}^{-1}(x),\overline{\mathbb{F}}_{\ell})\rightarrow R\Gamma(\overline{\mathfrak{Ig}}^{b,*}_{C},\overline{\mathbb{F}}_{\ell})\rightarrow R\Gamma(\overline{\partial\mathfrak{Ig}}^{b,*}_{C},\overline{\mathbb{F}}_{\ell})\xrightarrow{+1}.$
Applying Theorem 3.1 again, we know that $k:\mathfrak{Ig}^{b,*}\hookrightarrow\overline{\mathfrak{Ig}}^{b,*}_{C}$ is a qcqs open immersion of perfectoid spaces inducing an isomorphism on rank $1$ points, and it follows that the same is true for the induced map on the Zariski closed subspaces $\phantom{}{}^{\partial}k:\partial\mathfrak{Ig}^{b,*}\hookrightarrow\overline{\partial\mathfrak{Ig}}^{b,*}_{C}$. Therefore, we can apply [7, Lemma 4.4.2], this tells us that the natural maps
$\overline{\mathbb{F}}_{\ell}\rightarrow k_{*}(\overline{\mathbb{F}}_{\ell})$
$\overline{\mathbb{F}}_{\ell}\rightarrow\phantom{}^{\partial}k_{*}(\overline{\mathbb{F}}_{\ell})$
are isomorphisms, giving identifications $R\Gamma(\overline{\mathfrak{Ig}}_{C}^{b,*},\overline{\mathbb{F}}_{\ell})\simeq R\Gamma(\mathfrak{Ig}^{b,*}_{C},\overline{\mathbb{F}}_{\ell})$ and $R\Gamma(\overline{\partial\mathfrak{Ig}}_{C}^{b,*},\overline{\mathbb{F}}_{\ell})\simeq R\Gamma(\partial\mathfrak{Ig}^{b,*}_{C},\overline{\mathbb{F}}_{\ell})$. Now, by [7, Lemma 4.4.3], we have further identifications of $R\Gamma(\partial\mathfrak{Ig}^{b,*}_{C},\overline{\mathbb{F}}_{\ell})$ and $R\Gamma(\mathfrak{Ig}^{b,*}_{C},\overline{\mathbb{F}}_{\ell})$ with the cohomology of the perfect schemes $\partial\mathrm{Ig}^{b,*}$ and $\mathrm{Ig}^{b,*}$, respectively. Substituting this into the triangle (6), we get a distinguished triangle
$R\Gamma_{c}(\pi_{\mathrm{HT}}^{-1}(x),\overline{\mathbb{F}}_{\ell})\rightarrow R\Gamma(\mathrm{Ig}^{b,*},\overline{\mathbb{F}}_{\ell})\rightarrow R\Gamma(\partial\mathrm{Ig}^{b,*},\overline{\mathbb{F}}_{\ell})\xrightarrow{+1}.$
By applying quasi-compact base-change [56, Proposition 17.6] and then using that the inclusion $\partial\mathfrak{Ig}_{C}^{b,*}\subset\mathfrak{Ig}_{C}^{b,*}$ is induced from taking the rigid generic fiber over $C$ of Witt vectors applied to $\partial\mathrm{Ig}^{b,*}\subset\mathrm{Ig}^{b,*}$, we identify the last map with the natural restriction map on the cohomology. However, this identifies the first term with precisely the partially compactly supported cohomology, as desired. ∎
We will combine this with the following Proposition, which already hints at our expectation that $R\pi_{\mathrm{HT}!}(\overline{\mathbb{F}}_{\ell})$ is connective in some suitable perverse $t$-structure.
Proposition 3.6.
Let $d_{b}:=\langle 2\rho_{G},\nu_{b}\rangle$. We have $d_{b}=\mathrm{dim}(\mathrm{Ig}^{b,*})=\mathrm{dim}(\mathrm{Ig}^{b})$, and the cohomology of the complex
$R\Gamma_{c-\partial}(\mathrm{Ig}^{b},\overline{\mathbb{F}}_{\ell})\simeq R\pi_{\mathrm{HT}!}(\overline{\mathbb{F}}_{\ell})_{x}$
is concentrated in degrees $\leq d_{b}$.
Proof.
From [23, Corollary 7.8], we know that the dimension of $\mathscr{C}_{\mathbb{X}_{b}}$ is $d_{b}$, and since
$\mathrm{Ig}^{b}\rightarrow\mathscr{C}_{\mathbb{X}_{b}}^{\mathrm{perf}},$
is a pro-étale cover with Galois group $\mathrm{Aut}_{G}(\mathbb{X})(\overline{\mathbb{F}}_{p})$ over the perfection of the central leaf attached to $\mathbb{X}_{b}$, we may conclude the first part about dimensions. We also saw in Theorem 3.2 that $\mathrm{Ig}^{b,*}$ is an affine scheme. So we would like to apply Artin vanishing [1, Exposé XIV, Corrolary 3.2]; however, $\mathrm{Ig}^{b,*}$ is also a perfect scheme so not in general of finite type. Recall from [7, Proposition 4.3.8] that $\mathrm{Ig}^{b}$ is obtained as the perfection of the limit of the finite étale covers
$\mathrm{Ig}^{b}_{m}\rightarrow\mathscr{C}_{\mathbb{X}_{b}},$
where $\mathrm{Ig}^{b}_{m}$ is as defined in [8, Remark 2.3.6]. These spaces are of finite type over $\overline{\mathbb{F}}_{p}$. Define $\mathrm{Ig}^{b,*}_{m}$ to be the normalization of $\mathscr{C}_{\mathbb{X}}^{*}$ in $\mathrm{Ig}^{b}_{m}$ of the finite étale cover $\mathrm{Ig}^{b}_{m}\rightarrow\mathscr{C}_{\mathbb{X}_{b}}$. By [54, Theorem 3.33], $\mathscr{C}_{\mathbb{X}}^{*}$ is affine of dimension $d_{b}$; therefore, it follows that $\mathrm{Ig}^{b,*}_{m}$ is a normal and affine scheme of dimension $d_{b}$ which will be of finite type, since $\mathrm{Ig}^{b}_{m}$ is. Let $g_{b,m}:\mathrm{Ig}^{b}_{m}\rightarrow\mathrm{Ig}^{b,*}_{m}$ be the natural map, then Artin vanishing implies that the cohomology of the complex $g_{b,m!}(\overline{\mathbb{F}}_{\ell})$ is concentrated in degrees $\leq d_{b}$. By definition of $\mathrm{Ig}^{b,*}$, it is the perfection of $\mathop{\rm lim}\nolimits_{m\geq 1}\mathrm{Ig}_{m}^{b,*}$. Therefore, since étale cohomology is invariant under perfection, we can conclude by applying [61, Tag 09QY] to the system of sheaves $g_{b,m!}(\overline{\mathbb{F}}_{\ell})$. ∎
Now we would like to link this analysis with the semi-perversity of certain sheaves on $\mathrm{Bun}_{G}$ via a version of the Mantovan product formula.
3.2. Mantovan’s Product Formula
We consider the Hodge-Tate period morphism
$\pi_{\mathrm{HT}}:[\mathcal{S}_{K^{p}}/\underline{G(\mathbb{Q}_{p})}]\rightarrow[\mathcal{F}\ell_{G,\mu^{-1}}^{\Diamond}/\underline{G(\mathbb{Q}_{p})}]$
from the previous section quotiented out by $G(\mathbb{Q}_{p})$, where we abusively use $\pi_{\mathrm{HT}}$ for this quotiented out map. We let $h^{\rightarrow}:[\mathcal{F}\ell_{G,\mu^{-1}}^{\Diamond}/\underline{G(\mathbb{Q}_{p})}]\rightarrow[\mathop{\rm Spd}(C)/\underline{G(\mathbb{Q}_{p})}]\simeq\mathrm{Bun}_{G,C}^{1}$ be the structure map quotiented out by $G(\mathbb{Q}_{p})$. Note this is a proper map, since $\mathcal{F}\ell_{G,\mu^{-1}}^{\Diamond}$ is proper over $\mathop{\rm Spd}(C)$.
Then we have an identification
(7) $R\Gamma_{c}(\mathcal{S}_{K^{p},C},\overline{\mathbb{F}}_{\ell})\simeq Rh^{\rightarrow}_{*}R\pi_{\mathrm{HT}!}(\overline{\mathbb{F}}_{\ell})$
of $G(\mathbb{Q}_{p})$-representations.
Remark 3.7.
We note that it is always true that the compactly supported cohomology at infinite level is the colimit of the compactly supported cohomology at finite levels, but, for usual cohomology one needs to assume the spaces are qcqs for this to be true (e.g the tower defined by the good reduction locus or the minimal/toroidal compactifications).
Similarly, we have a map
$h^{\leftarrow}:[\mathcal{F}\ell_{G,\mu^{-1}}^{\Diamond}/\underline{G(\mathbb{Q}_{p})}]\rightarrow\mathrm{Bun}_{G}$
remembering the isomorphism class of the bundle $\mathcal{E}$ in the moduli interpretation of $\mathcal{F}\ell_{G,\mu^{-1}}^{\Diamond}$ as a diamond described in §3.1. This defines a cohomologically smooth map by [16, Theorem IV.1.19], and the image identifies with the open subset $B(G,\mu)\subset B(G)$ under the identification $|\mathrm{Bun}_{G}|\simeq B(G)$ of topological spaces [65], where $|\mathrm{Bun}_{G}|$ denotes the underlying topological space of the $v$-stack $\mathrm{Bun}_{G}$ and $B(G)$ has the topology given by its natural partial ordering (i.e it is generated by the opens $U_{b}:=\{x\in B(G)|x\leq b\}$, where $x\leq b$ means that their Kottwitz invariants $\kappa(x)=\kappa(b)$ agree and the differences of slopes $\nu_{b}-\nu_{x}$ is given by a $\mathbb{Q}_{\geq 0}$-linear combination of positive coroots (See [16, Section I.5] for details)). For each $b$, we have a locally closed Harder-Narasimhan stratum $j_{b}:\mathrm{Bun}_{G,C}^{b}\hookrightarrow\mathrm{Bun}_{G,C}$, and we can define the locally closed subset $\mathcal{F}\ell_{G,\mu^{-1}}^{b,\Diamond}/\underline{G(\mathbb{Q}_{p})}]$, by pulling back this HN-strata along $h^{\leftarrow}$. This defines a locally closed stratification of $\mathcal{F}\ell_{G,\mu^{-1}}^{\Diamond}/\underline{G(\mathbb{Q}_{p})}]$. Let $i_{b}:[\mathcal{F}\ell_{G,\mu^{-1}}^{b,\Diamond}/\underline{G(\mathbb{Q}_{p})}]\hookrightarrow[\mathcal{F}\ell_{G,\mu^{-1}}^{\Diamond}/\underline{G(\mathbb{Q}_{p})}]$ denote the associated locally closed immersion. We write $\pi_{\mathrm{HT}}^{b}:[\mathcal{S}^{b}_{K^{p},C}/\underline{G(\mathbb{Q}_{p})}]\rightarrow[\mathcal{F}\ell_{G,\mu^{-1}}^{b,\Diamond}/\underline{G(\mathbb{Q}_{p})}]$ (resp. $\pi_{\mathrm{HT}}^{b,*}:[\mathcal{S}^{b,*}_{K_{p},C}/\underline{G(\mathbb{Q}_{p})}]\rightarrow[\mathcal{F}\ell_{G,\mu^{-1}}^{b,\Diamond}/\underline{G(\mathbb{Q}_{p})}]$) for the pullbacks of $\pi_{\mathrm{HT}}$ (resp. $\pi^{*}_{\mathrm{HT}}$) along $i_{b}$. On the good reduction locus, we also have an additional stratification coming from pulling back the Newton stratification on the special fiber along the specialization map. We recall the rather subtle point that this does not agree with the pullback of the locally closed strata $\mathcal{F}\ell_{G,\mu^{-1}}^{b,\Diamond}$ (namely, the closure relationships are opposite with respect to the partial ordering on $B(G)$). We write $\mathcal{S}^{b,\circ,\mathrm{rd}}_{K^{p},C}$ for these Newton strata coming from the special fiber. There exists a natural map
$\mathcal{S}^{b,\circ,\mathrm{rd}}_{K^{p},C}\times_{\mathcal{F}\ell_{G,\mu^{-1}}^{\Diamond}}\mathcal{F}\ell_{G,\mu^{-1}}^{b,\Diamond}\hookrightarrow\mathcal{S}^{b,\circ}_{K^{p},C}$
which is a qcqs open immersion containing all rank $1$ points ([7, Page 68], [37, Page 8]). We write $\pi_{\mathrm{HT}}^{b,\circ}:[(\mathcal{S}^{b,\circ,\mathrm{rd}}_{K^{p},C}\times_{\mathcal{F}\ell^{\Diamond}_{G,\mu^{-1}}}\mathcal{F}\ell_{G,\mu^{-1}}^{b,\Diamond})/\underline{G(\mathbb{Q}_{p})}]\rightarrow[\mathcal{F}\ell_{G,\mu^{-1}}^{b,\Diamond}/\underline{G(\mathbb{Q}_{p})}]$ for the Hodge-Tate period map on this locus, and similarly we write $\overline{\pi}_{\mathrm{HT}}^{b,\circ}:[\overline{\mathcal{S}}^{b,\circ}_{K^{p},C}/\underline{G(\mathbb{Q}_{p})}]\rightarrow[\mathcal{F}\ell_{G,\mu^{-1}}^{b,\Diamond}/\underline{G(\mathbb{Q}_{p})}]$ for the induced map on the canonical compactification, where we note that this agrees with the canonical compactification of $(\mathcal{S}^{b,\circ,\mathrm{rd}}_{K^{p},C}\times_{\mathcal{F}\ell^{\Diamond}_{G,\mu^{-1}}}\mathcal{F}\ell_{G,\mu^{-1}}^{b,\Diamond})/\underline{G(\mathbb{Q}_{p})}]$ by the previous remark on rank $1$ points and the fact that $\mathcal{F}\ell_{G,\mu^{-1}}^{b,\Diamond}$ is partially proper [^6] The partial properness of these strata follows directly from the moduli interpretation, since the category of vector bundles on the Fargues-Fontaine curve is insensitive to the choice of $R^{+}$ for a Tate Huber pair $(R,R^{+})$..
Define the group diamond $\mathcal{J}_{b}:=\mathrm{Aut}(\mathcal{E}_{b})$, as in [16, Proposition III.5.1] parametrizing automorphism of the $G$-bundle $\mathcal{E}_{b}$ on the Fargues-Fontaine curve attached to $b$. We have an isomorphism $j_{b}:\mathrm{Bun}_{G}^{b}\simeq[\mathop{\rm Spd}(C)/\mathcal{J}_{b}]\hookrightarrow\mathrm{Bun}_{G}$ with the locally closed HN-strata in $\mathrm{Bun}_{G}$ defined by $b$. There is a $\mathcal{J}_{b}$-torsor over the adic Newton strata $\mathcal{F}\ell_{G,\mu^{-1}}^{b,\Diamond}$ given by rigidifying the bundle $\mathcal{E}$ to be isomorphic to $\mathcal{E}_{b}$. This gives a map
(8) $h^{\leftarrow}_{b}:[\mathcal{F}\ell_{G,\mu^{-1}}^{b,\Diamond}/\underline{G(\mathbb{Q}_{p})}]\rightarrow[\mathop{\rm Spd}(C)/\mathcal{J}_{b}]\simeq\mathrm{Bun}_{G,C}^{b}$
such that $h^{\leftarrow}\circ i_{b}=j_{b}\circ h^{\leftarrow}_{b}$.
The diamond attached to the perfectoid Igusa variety $\mathfrak{Ig}^{b}_{C}$ comes equipped with an action of $\mathcal{J}_{b}$. In particular, we consider the formal group scheme $\mathrm{Aut}_{G}(\tilde{\mathbb{X}}_{b})$ over $\mathrm{Spf}(W(\overline{\mathbb{F}}_{p}))$ described in [7, Definition 4.2.9] parametrizing automorphisms of the universal cover of $\mathbb{X}_{b}$. We write $\mathrm{Aut}_{G}(\tilde{\mathbb{X}}_{b})_{\overline{\mathbb{F}}_{p}}$ for the special fiber of this formal group scheme, which appeared in 2.2.1. This defines a formal group over $\overline{\mathbb{F}}_{p}$, which we regard as an fpqc sheaf on schemes over $\overline{\mathbb{F}}_{p}$. The formal group scheme $\mathrm{Aut}_{G}(\tilde{\mathbb{X}}_{b})$ over $\mathrm{Spf}(W(\overline{\mathbb{F}}_{p}))$ is then given by taking Witt vectors of this formal group.
As is done in [7, Corollary 4.3.5], one can, using the moduli description of $\mathrm{Ig}^{b}$ provided in (3), show that the fpqc sheaf $\mathrm{Aut}_{G}(\tilde{\mathbb{X}}_{b})_{\overline{\mathbb{F}}_{p}}$ acts on $\mathrm{Ig}^{b}$, by noting that quasi-isogenies of $\mathbb{X}_{b}$ are the same as automorphisms of the universal cover.
Functorality of Witt vectors then gives rise to an action of $\mathrm{Aut}_{G}(\tilde{\mathbb{X}}_{b})$ on the Witt vector lift of $\mathrm{Ig}^{b}$, and passing to rigid generic fibers over $C$ gives an action of the rigid generic fiber of $\mathrm{Aut}_{G}(\tilde{\mathbb{X}}_{b})$ on $\mathfrak{Ig}^{b}_{C}$, which we denote by $\mathrm{Aut}_{G}(\tilde{\mathbb{X}}_{b})_{C}$. By [7, Lemma 4.2.10], $\mathrm{Aut}_{G}(\tilde{\mathbb{X}}_{b})_{C}$ is a perfectoid space which we identify with its associated $v$-sheaf on $\mathrm{Perf}_{C}$ using the Yoneda embedding. The following seems to be well known, but we sketch a proof as we could not find it explicitly written down in the literature.
Proposition 3.8.
There is an identification
(9) $\mathrm{Aut}_{G}(\tilde{\mathbb{X}}_{b})_{C}\simeq\mathcal{J}_{b}$
of group-valued functors on $\mathrm{Perf}_{C}$, where $\mathcal{J}_{b}$ is the group diamond parametrizing automorphisms of the $G$-bundle $\mathcal{E}_{b}$ on the Fargues-Fontaine curve attached to $b$.
Proof.
(Sketch) We fix $(\mathbb{X}_{b})_{O_{C}}$ to be a lift of $\mathbb{X}_{b}$ up to quasi-isogeny (with its $G$-structure). For an object $S=\mathop{\rm Spa}(R,R^{+})\in\mathrm{Perf}_{C}$, the $S$-points of LHS identifies with the set of formal quasi-isogenies
$(\mathbb{X}_{b})_{O_{C}}\times_{\mathrm{Spf}(O_{C})}\mathrm{Spf}(R^{+})\dashrightarrow(\mathbb{X}_{b})_{O_{C}}\times_{\mathrm{Spf}(O_{C})}\mathrm{Spf}(R^{+})$
in the sense of [10, Section 2.5], respecting the $G$-structure. In other words, this is a quasi-isogeny
$(\mathbb{X}_{b})_{O_{C}}\times_{O_{C}}R^{+\flat}/\varpi\dashrightarrow(\mathbb{X}_{b})_{O_{C}}\times_{O_{C}}R^{+\flat}/\varpi$
respecting the additional structures, where $\varpi\in R^{+\flat}$ is a suitable choice of pseudo-uniformizer in the tilt $R^{+\flat}$ of $R^{+}$, and the quotient map is given by the map $R^{+}\rightarrow R^{+\flat}/\varpi$ induced by Fontaine’s map $\theta_{(R,R^{+})}:W(R^{+\flat})\rightarrow R^{+}$, and the quasi-isogeny doesn’t depend on the choice of pseudo-uniformizer by Serre-Tate theory.
As described in [10, Section 2.5.3], to a $p$-divisible $G_{0}$ group over $R^{+}/\varpi$, one can associated a vector bundle $\mathcal{E}(G_{0})$ on $X_{S}$ using crystalline Dieudonné theory. As explained in [10, Section 2.5.6], a quasi-isogeny $G_{0}\dashrightarrow G_{0}$ induces an isomorphism of vector bundles
$\mathcal{E}(G_{0})\xrightarrow{\simeq}\mathcal{E}(G_{0})$
on $X_{S}$. By applying the Tannakian-formalism, it follows that for a $p$-divisible group $G_{0}$ with $G$-structure and a quasi-isogeny $G_{0}\dashrightarrow G_{0}$ respecting the additional structures, that one obtains a $G$-bundle $\mathcal{E}_{G}(G_{0})$ and an isomorphism of $G$-bundles $\mathcal{E}_{G}(G_{0})\xrightarrow{\simeq}\mathcal{E}_{G}(G_{0})$. Applying this to $G_{0}=(\mathbb{X}_{b})_{O_{C}}\times_{O_{C}}R^{+\flat}/\varpi$, one can check by our assumption that $(\mathbb{X}_{b})_{O_{C}}$ was a lift of $\mathbb{X}_{b}$ and [16, Theorem II.0.4], that one has pro-étale locally an identification $\mathcal{E}_{G}(G_{0})=\mathcal{E}_{b,S}$, where $\mathcal{E}_{b,S}$ is the pullback of $\mathcal{E}_{b}$ the vector bundle attached to the isocrystal $b\in B(G)$ on $X_{C}$ along the natural map $X_{S}\rightarrow X_{C}$. This defines the desired natural transformation from the functor of points of the LHS to the functor of points of the RHS of (9). In order to check it is an equivalence, we use that the LHS and RHS of (9) are partially proper over $\mathop{\rm Spa}(C,O_{C})$ to reduce to checking it on $(R,R^{\circ})$-points. For the RHS, this follows from the definition, as the category of vector bundles on $X_{(R,R^{+})}$ and $X_{(R,R^{\circ})}$ are equivalent, and for the LHS it follows from [7, Proposition 4.2.22]. The desired equivalence is then reduced to the fully faithfullness statement in [50, Proposition 2.2.7], via [10, Lemma 2.5.5].
Using the previous Proposition, one sees that the action of $\mathrm{Aut}_{G}(\tilde{\mathbb{X}}_{b})_{C}$ on $\mathfrak{Ig}^{b}_{C}$ described above gives rise to an action of the group diamond $\mathcal{J}_{b}$ parametrizing automorphism of $\mathcal{E}_{b}$.
This allows us to form the $v$-stack quotient $\mathfrak{Ig}_{C}^{b}/\mathcal{J}_{b}]$, which is equipped with a natural map:
$\pi_{\mathfrak{Ig}}^{b}:[\mathfrak{Ig}_{C}^{b}/\mathcal{J}_{b}]\rightarrow[\mathop{\rm Spd}(C)/\mathcal{J}_{b}].$
We would like to say $\pi_{\mathfrak{Ig}}^{b}$ pulls back to the map $\pi^{b}_{\mathrm{HT}}$ along the map (8). However, as seen in Corollary 3.5, we need to account for the additional points in the fiber of the Hodge-Tate period morphism that are not seen by the perfectoid Igusa varieties $\mathfrak{Ig}^{b}_{C}$. To capture this, we need to show that $\pi_{\mathfrak{Ig}}^{b}$ also extends to the partial minimal compactification in a suitable sense. We recall that we have a decomposition
(10) $\mathcal{J}_{b}\simeq\underline{J_{b}(\mathbb{Q}_{p})}\ltimes\mathcal{J}_{b}^{>0},$
where $\underline{J_{b}(\mathbb{Q}_{p})}$ is the constant group scheme coming from the automorphisms of $\mathcal{E}_{b}$ which preserve the graded pieces of the Harder-Narasimhan filtration and $\mathcal{J}_{b}^{>0}$ is an $\ell$-adically contractible unipotent group diamond of dimension $\langle 2\rho_{G},\nu_{b}\rangle$. Analogously, we have a decomposition of fpqc sheaves
(11) $\mathrm{Aut}_{G}(\tilde{\mathbb{X}}_{b})_{\overline{\mathbb{F}}_{p}}=\underline{J_{b}(\mathbb{Q}_{p})}\ltimes\widetilde{\mathcal{U}}_{b},$
where $J_{b}(\mathbb{Q}_{p})$ can be identified with the group of self-quasi-isogenies of $\mathbb{X}_{b}$ over $\overline{\mathbb{F}}_{p}$ that respect the $G$-structure, $\underline{J_{b}(\mathbb{Q}_{p})}$ is the associated constant group scheme, and $\tilde{\mathcal{U}_{b}}$ is a unipotent formal group scheme, see [7, Proposition 4.2.11]. We now have the following.
Proposition 3.9.
Assuming 1.13, the action of $\mathcal{J}_{b}$ on the perfectoid Igusa variety $\mathfrak{Ig}^{b}_{C}$ extends uniquely to an action on $\mathfrak{Ig}^{b,*}_{C}$. In particular, by functorality of the formation of the canonical compactification ([56, Proposition 18.6]), we have a map
$\pi_{\mathfrak{Ig}}^{b,*}:[\overline{\mathfrak{Ig}}_{C}^{b,*}/\mathcal{J}_{b}]\rightarrow[\mathop{\rm Spd}(C)/\mathcal{J}_{b}]$
extending $\pi_{\mathfrak{Ig}}^{b}$. This restriction of the action to the subgroup $\underline{J_{b}(\mathbb{Q}_{p})}\subset\mathcal{J}_{b}$ also preserves the boundary $\partial\mathfrak{Ig}^{b,*}$, so, in particular, we also get a map
$\tilde{\pi}_{\mathfrak{Ig}}^{b,\partial}:[(\overline{\mathfrak{Ig}}^{b,*}_{C}\setminus\overline{\partial\mathfrak{Ig}}^{b,*}_{C})/\underline{J_{b}(\mathbb{Q}_{p})}]\rightarrow[\mathop{\rm Spd}(C)/\underline{J_{b}(\mathbb{Q}_{p})}]$
by pulling back $\pi_{\mathfrak{Ig}}^{b,*}$ along the map $\mathop{\rm Spd}(C)/\underline{J_{b}(\mathbb{Q}_{p})}]\rightarrow[\mathop{\rm Spd}(C)/\mathcal{J}_{b}]$ induced by the decomposition (10) and restricting.
Proof.
We consider the open immersion
$g_{b}:\mathrm{Ig}^{b}\rightarrow\mathrm{Ig}^{b,*}$
of perfect schemes, which we claim induces an isomorphism on global sections. To show this, we write $g_{b}$ as the perfection of the limit of the corresponding maps at finite level
$g_{b,m}:\mathrm{Ig}^{b}_{m}\hookrightarrow\mathrm{Ig}^{b,*}_{m},$
as explained in the proof of Proposition 3.6. Under Assumption 1.13, we can apply the algebraic form of Hartogs’ principle (See for example [57, Proposition III.2.9]) to the open inclusion $g_{b,m}$ to conclude an isomorphism of global sections via restriction. This gives the corresponding claim for the map $g_{b}$ of perfect schemes, by taking inverse limits and perfectifying. In particular, we have an isomorphism
(12) $\mathcal{O}(\mathrm{Ig}^{b})\simeq\mathcal{O}(\mathrm{Ig}^{b,*})$
on global sections. Since the perfect scheme $\mathrm{Ig}^{b,*}$ is affine, the isomorphism of global sections implies that the action of the formal group scheme $\mathrm{Aut}_{G}(\tilde{\mathbb{X}}_{b})$ regarded as a fpqc sheaf on the perfect scheme $\mathrm{Ig}^{b}$ extends to an action on the perfect affine scheme $\mathrm{Ig}^{b,*}$. As before, by taking Witt vectors and passing to the rigid generic fiber over $C$, this induces an action of $\mathrm{Aut}_{G}(\tilde{\mathbb{X}}_{b})_{C}$ on the perfectoid space $\mathfrak{Ig}^{b,*}_{C},$ extending the action on $\mathfrak{Ig}^{b}_{C}$. Passing through the identification $\mathrm{Aut}_{G}(\tilde{\mathbb{X}}_{b})_{C}\simeq\mathcal{J}_{b}$ of Proposition 3.8, this gives us the desired action of $\mathcal{J}_{b}$ on $\mathfrak{Ig}^{b,*}_{C}$.
Now, to understand the preservation of the boundary, we understand the construction of the action of $\mathcal{J}_{b}$ in a more abstract way. In particular, we consider the small diamond functor $(-)^{\diamond}$ of [22, Section 3.1]. Recall $\mathrm{Perf}$ is the category of perfectoid spaces in characteristic $p$ over $\overline{\mathbb{F}}_{p}$, the $v$-sheaf $\mathrm{Aut}_{G}(\tilde{\mathbb{X}}_{b})^{\diamond}$ on $\mathrm{Perf}$ attached to the formal group $\mathrm{Aut}_{G}(\tilde{\mathbb{X}}_{b})^{\diamond}$ is defined by the functor sending $(R,R^{+})\in\mathrm{Perf}$ to the group
[…]
for some pseudo-uniformizer $\varpi$, where the last equivalence follows from Serre-Tate lifting of $p$-divisible groups along $p$-nilpotent rings, as in [8, Theorem 2.4.1]. The $v$-sheaves $\mathrm{Ig}^{b,\diamond}$ (resp. $\mathrm{Ig}^{b,*,\diamond}$) are defined by sending $(R,R^{+})$ to $\mathrm{Ig}^{b}(\mathop{\rm Spec}R^{+})$ and $\mathrm{Ig}^{b,*}(\mathop{\rm Spec}R^{+})$. We note that, since $\mathrm{Ig}^{b}$ and $\mathrm{Ig}^{b,*}$ are perfect and $R^{+}$ is perfectoid, we have identifications $\mathrm{Ig}^{b}(\mathop{\rm Spec}R^{+})\simeq\mathrm{Ig}^{b}(\mathop{\rm Spec}R^{+}/\varpi)$ and $\mathrm{Ig}^{b,*}(\mathop{\rm Spec}R^{+})\simeq\mathrm{Ig}^{b,*}(\mathop{\rm Spec}R^{+}/\varpi)$. In particular, we may understand the action of $\mathrm{Aut}_{G}(\tilde{\mathbb{X}}_{b})^{\diamond}$ on the $v$-sheaves $\mathrm{Ig}^{b,\diamond}$ and $\mathrm{Ig}^{b,*,\diamond}$ in terms of the action of $\mathrm{Aut}_{G}(\tilde{\mathbb{X}}_{b})(\mathop{\rm Spec}R^{+}/\varpi)$ on $\mathrm{Ig}^{b}(\mathop{\rm Spec}R^{+}/\varpi)$ and $\mathrm{Ig}^{b,*}(\mathop{\rm Spec}R^{+}/\varpi)$.
The action of $\mathcal{J}_{b}$ on $\mathfrak{Ig}^{b}_{C}$ and $\mathfrak{Ig}^{b,*}_{C}$ described above may now be given by base-changing $\mathrm{Aut}_{G}(\tilde{\mathbb{X}}_{b})^{\diamond}$, $\mathrm{Ig}^{b}$, $\mathrm{Ig}^{b,*}_{C}$ and base-changing along $\mathop{\rm Spd}(C)\rightarrow\mathop{\rm Spd}(\overline{\mathbb{F}}_{p})$ in the category of $v$-sheaves and identifying the resulting functors on $\mathrm{Perf}_{C}$ with $\mathcal{J}_{b}\simeq\mathrm{Aut}_{G}(\tilde{\mathbb{X}}_{b})_{C}$, $\mathfrak{Ig}^{b}_{C}$, and $\mathfrak{Ig}^{b,*}_{C}$. In particular, the relationship between the construction described above in terms of taking Witt vectors and the passing to the rigid generic fiber over $C$ follows from from the relationship between untilts of a characteristic $p$ perfectoid $(R,R^{+})$ and Cartier divisors in the Witt ring $W(R^{+})$, as described in [55, Lemma 6.28].
We now use this perspective to show that the action of sub group-sheaf $\underline{J_{b}(\mathbb{Q}_{p})}\subset\mathrm{Aut}_{G}(\tilde{\mathbb{X}}_{b})^{\diamond}$ given by the decomposition (11) preserves the closed (by the valuative criterion of properness) subsheaf $\partial\mathrm{Ig}^{b,*,\diamond}\subset\mathrm{Ig}^{b,*,\diamond}$. By base-changing along $\mathop{\rm Spd}(C)\rightarrow\mathop{\rm Spd}(\overline{\mathbb{F}}_{p})$, this will give the desired claim.
[…]
Lemma 3.10.
Suppose we have a group-valued fpqc sheaf $G/\overline{\mathbb{F}}_{p}$ acting on a scheme $X/\overline{\mathbb{F}}_{p}$ via $m:G\times X\rightarrow X$ and suppose that $Z\subset X$ is a closed subscheme which is reduced. We write $U:=X\setminus Z$ for the open complement and assume the action of $G$ preserves $U\subset X$. Then, for all $T/\overline{\mathbb{F}}_{p}$ reduced, the action of the $T$-points $G(T)$ on $X(T)$ preserves $Z(T)$.
Proof.
For $T$ as in the statement, consider a $T$-point $x:T\rightarrow Z\rightarrow X$ and an element $g\in G(T)$. We need to show that $m(g,x):T\rightarrow X$ factors through $Z\rightarrow X$. Since $T$ and $Z$ are reduced, it suffices by [61, Lemma 0356], to show that the induced map $|m(g,x)|:|T|\rightarrow|X|$ on the underlying topological space factors through $|Z|\rightarrow|X|$. Suppose for the sake of contradiction that it does not, then there exists a geometric point $\mathop{\rm Spec}(k)\rightarrow T$ factoring through $U\rightarrow X$. Now if we consider $m(g^{-1},m(g,x))=x:T\rightarrow X$, then we arrive at a contradiction since the precomposition of this map with $\mathop{\rm Spec}(k)\rightarrow T$ must factor through $U\subset X$, by the assumption that $G$ preserves $U$. This however contradicts the decomposition $|X|=|U|\sqcup|Z|$ on topological spaces. ∎
Remark 3.11.
It is true that the full action of $\mathcal{J}_{b}$ preserves the Zariski closed boundary $\partial\mathfrak{Ig}^{b,*}_{C}\subset\mathfrak{Ig}^{b,*}_{C}$. This is shown in [6, Section 4-5] (In fact, a stronger claim is shown that $\mathcal{J}_{b}$ preserves all the closed boundary strata of $\partial\mathfrak{Ig}^{b,*}_{C}$). However, this requires a much more detailed argument involving lifting to the toroidal compactification. As a result, we will not address this here.
Lastly, we will consider the map $\overline{\pi}^{b}_{\mathfrak{Ig}}:[\overline{\mathfrak{Ig}}^{b}_{C}/\mathcal{J}_{b}]\rightarrow[\mathop{\rm Spd}(C)/\mathcal{J}_{b}]$, given by taking the canonical compactification of $\pi^{b}_{\mathfrak{Ig}}$, where we note, by [7, Proposition 4.2.22], the $v$-stack $\mathop{\rm Spd}(C)/\mathcal{J}_{b}]$ is partially proper over $\mathop{\rm Spd}(C)$. We now have the following Proposition.
Proposition 3.12.
The maps constructed above fit into the following Cartesian squares [^7] We emphasize that these are really diagrams of $v$-stacks and that all fiber products are formed in this category.
(13) $\left[(\mathcal{S}^{b,\circ,\mathrm{rd}}_{K^{p},C}\times_{\mathcal{F}\ell^{\Diamond}_{G,\mu^{-1}}}\mathcal{F}\ell_{G,\mu^{-1}}^{b,\Diamond})/\underline{G(\mathbb{Q}_{p})}\right]$$\left[\mathcal{F}\ell^{b,\Diamond}_{G,\mu^{-1}}/\underline{G(\mathbb{Q}_{p})}\right]$$\left[\mathfrak{Ig}_{C}^{b}/\mathcal{J}_{b}\right]$$\left[\mathop{\rm Spd}(C)/\mathcal{J}_{b}\right]$$\scriptstyle{\pi_{\mathrm{HT}}^{b,\circ}}$$\scriptstyle{\tilde{h}_{b}^{\leftarrow}}$$\scriptstyle{h^{\leftarrow}_{b}}$$\scriptstyle{\pi_{\mathfrak{Ig}}^{b}}$
and
(14) $\left[\overline{\mathcal{S}}^{b,\circ}_{K^{p},C}/\underline{G(\mathbb{Q}_{p})}\right]$$\left[\mathcal{F}\ell^{b,\Diamond}_{G,\mu^{-1}}/\underline{G(\mathbb{Q}_{p})}\right]$$\left[\overline{\mathfrak{Ig}}_{C}^{b}/\mathcal{J}_{b}\right]$$\left[\mathop{\rm Spd}(C)/\mathcal{J}_{b}\right].$$\scriptstyle{\overline{\pi}_{\mathrm{HT}}^{b,\circ}}$$\scriptstyle{\tilde{h}_{b}^{\leftarrow}}$$\scriptstyle{h^{\leftarrow}_{b}}$$\scriptstyle{\overline{\pi}_{\mathfrak{Ig}}^{b}}$
Proof.
Consider the moduli space of local shtukas $\mathrm{Sht}(G,b,\mu)_{\infty,C}$, as defined in [55, Definition 23.1.1]. This represents the functor sending $S\in\mathrm{Perf}_{C}$ to the set of all pairs $(S^{\#},\alpha)$ where $S^{\#}$ is the untilt of $S$ coming from the map $S\rightarrow\mathop{\rm Spd}(C)$, and $\alpha$ is a modification from $\mathcal{E}_{b}\dashrightarrow\mathcal{E}_{0}$ with meromorphy along $S^{\#}$ and bounded by $\mu$. We have a local Hodge-Tate period morphism
$\mathrm{Sht}(G,b,\mu)_{\infty,C}\rightarrow\mathcal{F}\ell_{G,\mu^{-1}}^{b,\Diamond},$
which fits into the following Cartesian diagram coming from the definition of $\mathrm{Sht}(G,b,\mu)_{\infty,C}$.
(15) $\mathrm{Sht}(G,b,\mu)_{\infty,C}$$\mathop{\rm Spd}(C)$$\mathcal{F}\ell_{G,\mu^{-1}}^{b,\Diamond}$$\left[\mathop{\rm Spd}(C)/\mathcal{J}_{b}\right]$}
Let $\mathrm{Sht}(G,b,\mu)_{\infty,C}\times^{\mathcal{J}_{b}}\mathfrak{Ig}_{C}^{b}$ denote the quotient of $\mathrm{Sht}(G,b,\mu)_{\infty,C}\times_{C}\mathfrak{Ig}_{C}^{b}$ by $\{(jx,j^{-1}y):j\in\mathcal{J}_{b},x\in\mathrm{Sht}(G,b,\mu)_{\infty,C},y\in\mathfrak{Ig}_{C}^{b}\}$. To see that the diagram (13) above is Cartesian, observe that (15) implies we have an isomorphism
$\mathcal{F}\ell^{b,\Diamond}_{G,\mu^{-1}}\times_{[\mathop{\rm Spd}(C)/\mathcal{J}_{b}]}[\mathfrak{Ig}_{C}^{b}/\mathcal{J}_{b}]\simeq\mathrm{Sht}(G,b,\mu)_{\infty,C}\times^{\mathcal{J}_{b}}\mathfrak{Ig}_{C}^{b}.$
Moreover, we see, by [7, Corollary 4.3.19,Lemma 4.3.20], that we have an isomorphism
(16) $\mathrm{Sht}(G,b,\mu)_{\infty,C}\times_{C}\mathfrak{Ig}_{C}^{b}\simeq\mathcal{S}^{b,\circ,\mathrm{rd}}_{K^{p}}\times_{\mathcal{F}\ell^{\Diamond}_{G,\mu^{-1}}}\mathrm{Sht}(G,b,\mu)_{\infty,C}.$
Moreover, this map is $\mathcal{J}_{b}$-equivariant. Here $\mathcal{J}_{b}$ acts on the LHS via $(x,y)\mapsto(jx,j^{-1}y)$, and $\mathcal{J}_{b}$ acts on the RHS by the trivial action on $\mathcal{S}^{b,\circ,\mathrm{rd}}_{K^{p}}$ and the natural action on $\mathrm{Sht}(G,b,\mu)_{\infty,C}$ by automorphisms of bundles.
Now applying (15) implies that $\mathcal{S}^{b,\circ,\mathrm{rd}}_{K^{p}}\times_{\mathcal{F}\ell^{\Diamond}_{G,\mu^{-1}}}\mathcal{F}\ell_{G,\mu^{-1}}^{b,\Diamond}$ is isomorphic to the quotient of $\mathcal{S}^{b,\circ,\mathrm{rd}}_{K^{p}}\times_{\mathcal{F}\ell^{\Diamond}_{G,\mu^{-1}}}\mathrm{Sht}(G,b,\mu)_{\infty,C}$ by the action of $\mathcal{J}_{b}$ (here $\mathcal{J}_{b}$ acts via the action on the second factor). Thus, we have an isomorphism:
$\mathrm{Sht}(G,b,\mu)_{\infty,C}\times^{\mathcal{J}_{b}}\mathfrak{Ig}_{C}^{b}\simeq\mathcal{S}^{b,\circ,\mathrm{rd}}_{K^{p}}\times_{\mathcal{F}\ell^{\Diamond}_{G,\mu^{-1}}}\mathcal{F}\ell_{G,\mu^{-1}}^{b,\Diamond}.$
This gives us the Cartesian diagram (14). Now, the natural map,
$\mathcal{S}_{K^{p},C}^{b,\circ,\mathrm{rd}}\times_{\mathcal{F}\ell^{\Diamond}_{G,\mu^{-1}}}\mathcal{F}\ell_{G,\mu^{-1}}^{b,\Diamond}\hookrightarrow\mathcal{S}^{b,\circ}_{K^{p},C}$
is a qcqs open immersion, which is an isomorphism on rank $1$ points. In turn, it induces an isomorphism of canonical compactifications over the partially proper strata $\mathcal{F}\ell_{G,\mu^{-1}}^{b,\Diamond}$. Therefore, by passing to canonical compactifications over $\mathcal{F}\ell_{G,\mu^{-1}}^{b,\Diamond}$, we deduce that the diagram (14) is also Cartesian, using that the formation of canonical compactifications commutes with limits [56, Proposition 18.7 (v)]. ∎
Remark 3.13.
We note that our conventions for moduli spaces of shtukas differs from that in [55, Definition 23.1.1]; in particular, there the space $\mathrm{Sht}(G,b,\mu)_{\infty}$ denotes the moduli space $\mathcal{E}_{0}\dashrightarrow\mathcal{E}_{b}$ parametrizes modifications of meromorphy $\mu$, which would identify more naturally with the space we denote by $\mathrm{Sht}(G,b,\mu^{-1})_{\infty}$. This is consistent with [27, 24, 47], where this convention is nice in that the Hecke operator $j_{1}^{*}T_{\mu}$ is related to $\mathrm{Sht}(G,b,\mu)_{\infty}$ for $b\in B(G,\mu)$ instead of $\mathrm{Sht}(G,b,\mu^{-1})$, as in the usual convention. Here we also note that we have uniformized the Schubert stratification of the $\mathbb{B}_{\mathrm{dR}}^{+}$-Grassmannian so that $G(\mathbb{B}_{\mathrm{dR},C}):=\bigsqcup_{\mu\in\mathbb{X}_{*}(T_{\overline{\mathbb{Q}}_{p}})^{+}}G(\mathbb{B}^{+}_{\mathrm{dR},C})\mu(\xi)^{-1}G(\mathbb{B}_{\mathrm{dR},C}^{+})$, where the coset $G(\mathbb{B}^{+}_{\mathrm{dR},C})\mu(\xi)^{-1}G(\mathbb{B}_{\mathrm{dR},C}^{+})$ defines $\mathrm{Gr}_{G,\mu}$, as in the discussion proceeding [7, Definition 3.4.1]. We warn the reader that this differs by an inverse from the uniformization used in [55] and [10], which is why the index set for the adic Newton stratification of $\mathcal{F}\ell_{G,\mu^{-1}}^{\Diamond}$ is for us $B(G,\mu)$ instead of $B(G,\mu^{-1})$.
We now invoke a result of Zhang [68].
Theorem 3.14.
[68, Theorem 1.3, Section 9.4] Assuming 1.13, for all $b\in B(G,\mu)$ the Cartesian diagram (14) extends to a Cartesian diagram
(17) $\left[\mathcal{S}^{b,*}_{K^{p},C}/\underline{G(\mathbb{Q}_{p})}\right]$$\left[\mathcal{F}\ell^{b,\Diamond}_{G,\mu^{-1}}/\underline{G(\mathbb{Q}_{p})}\right]$$\left[\overline{\mathfrak{Ig}}_{C}^{b,*}/\mathcal{J}_{b}\right]$$\left[\mathop{\rm Spd}(C)/\mathcal{J}_{b}\right]$$\scriptstyle{\pi_{\mathrm{HT}}^{b,*}}$$\scriptstyle{\phantom{}{}^{*}\tilde{h}_{b}^{\leftarrow}}$$\scriptstyle{h^{\leftarrow}_{b}}$$\scriptstyle{\pi_{\mathfrak{Ig}}^{b,*}}$
of $v$-stacks.
Remark 3.15.
In fact Zhang shows a much stronger claim, that there exists a series of larger Cartesian diagrams living over $\mathrm{Bun}_{G.C}$ such that the diagrams (14) and (17), are the base-change along the inclusions $j_{b}:\mathrm{Bun}_{G,C}^{b}\hookrightarrow\mathrm{Bun}_{G,C}$ of HN-strata for $b\in B(G,\mu)$ varying.
Remark 3.16.
The rough idea behind proving this is to apply a relative $\mathop{\rm Spa}$ construction in the category of $v$-sheaves to the horizontal maps of the diagram (17) by invoking Hartogs’ principle, as in Proposition 3.9.
We now state the key Corollary that we will need. We write $\tilde{s}_{b}:[\widetilde{\mathcal{F}\ell}_{G,\mu^{-1}}^{b}/\underline{G(\mathbb{Q}_{p})}]\rightarrow[\mathcal{F}\ell_{G,\mu^{-1}}^{b,\Diamond}/\underline{G(\mathbb{Q}_{p})}]$ for the $\mathcal{J}_{b}^{>0}$-torsor defined by pulling back $h^{\leftarrow}_{b}$ along the natural map $s_{b}:[\mathop{\rm Spd}(C)/\underline{J_{b}(\mathbb{Q}_{p})}]\rightarrow[\mathop{\rm Spd}(C)/\mathcal{J}_{b}]$ induced by (10). We write $\tilde{\pi}_{\mathrm{HT}}^{b}:[\widetilde{\mathcal{S}}^{b}_{K^{p},C}/\underline{G(\mathbb{Q}_{p})}]\rightarrow[\widetilde{\mathcal{F}\ell}_{G,\mu^{-1}}^{b,\Diamond}/\underline{G(\mathbb{Q}_{p})}]$ and $\tilde{\pi}_{\mathrm{HT}}^{b,*}:[\widetilde{\mathcal{S}}^{b,*}_{K^{p},C}/\underline{G(\mathbb{Q}_{p})}]\rightarrow[\widetilde{\mathcal{F}\ell}^{\Diamond}_{G,\mu^{-1}}/\underline{G(\mathbb{Q}_{p})}]$ for the pullbacks of $\pi_{\mathrm{HT}}^{b}$ and $\pi_{\mathrm{HT}}^{b,*}$ along the map $\widetilde{\mathcal{F}\ell}_{G,\mu^{-1}}^{b,\Diamond}/\underline{G(\mathbb{Q}_{p})}]\rightarrow[\mathcal{F}\ell_{G,\mu^{-1}}^{b,\Diamond}/\underline{G(\mathbb{Q}_{p})}]$.
We now have the following.
Corollary 3.17.
Assuming 1.13, for all $b\in B(G,\mu)$ we have a Cartesian diagram
(18) $\left[\widetilde{\mathcal{S}}^{b,*}_{K^{p},C}/\underline{G(\mathbb{Q}_{p})}\right]$$\left[\widetilde{\mathcal{F}\ell}^{b,\Diamond}_{G,\mu^{-1}}/\underline{G(\mathbb{Q}_{p})}\right]$$\left[(\overline{\mathfrak{Ig}}_{C}^{b,*}\setminus\overline{\partial\mathfrak{Ig}}_{C}^{b,*})/\underline{J_{b}(\mathbb{Q}_{p})}\right]$$\left[\mathop{\rm Spd}(C)/\underline{J_{b}(\mathbb{Q}_{p})}\right].$$\scriptstyle{\tilde{\pi}_{\mathrm{HT}}^{b,*}}$$\scriptstyle{\phantom{}{}^{\partial}\tilde{h}_{b}^{\leftarrow}}$$\scriptstyle{\tilde{\pi}_{\mathfrak{Ig}}^{b,\partial}}$
with maps as defined above
Proof.
This follows from the Cartesian diagram (17) and Corollary 3.4. More precisely, consider the open substack
$\widetilde{\mathcal{S}}_{K^{p},C}^{b,*}/\underline{G(\mathbb{Q}_{p})}]\times_{[\overline{\mathfrak{Ig}}^{b,*}_{C}/\underline{J_{b}(\mathbb{Q}_{p})}]}\left[(\overline{\mathfrak{Ig}}_{C}^{b,*}\setminus\overline{\partial\mathfrak{Ig}}_{C}^{b,*})/\underline{J_{b}(\mathbb{Q}_{p})}\right]\hookrightarrow[\widetilde{\mathcal{S}}^{b,*}_{K^{p},C}/\underline{G(\mathbb{Q}_{p})}]$
obtained by considering the base-change of $\left[\widetilde{\mathcal{F}\ell}^{b,\Diamond}_{G,\mu^{-1}}/\underline{G(\mathbb{Q}_{p})}\right]\rightarrow[\mathop{\rm Spd}(C)/\underline{J_{b}(\mathbb{Q}_{p})}]$ along the open immersion $\overline{\mathfrak{Ig}}_{C}^{b,*}\setminus\overline{\partial\mathfrak{Ig}}^{b,*}_{C}/\underline{J_{b}(\mathbb{Q}_{p})}]\hookrightarrow[\overline{\mathfrak{Ig}}^{b,*}_{C}/\underline{J_{b}(\mathbb{Q}_{p})}]$. The claim follows from showing that this open substack identifies with the open substack
$\widetilde{\mathcal{S}}^{b}_{K_{p},C}/\underline{G(\mathbb{Q}_{p})}]\hookrightarrow[\widetilde{\mathcal{S}}^{b,*}_{K^{p},C}/\underline{G(\mathbb{Q}_{p})}].$
Using [56, Proposition 11.15], it suffices to check that the open subspaces of the underlying topological space of $\widetilde{\mathcal{S}}^{b,*}_{K^{p},C}/\underline{G(\mathbb{Q}_{p})}]$ that these two substacks define agree with one another. Since both subspaces are partially proper (as they are $\mathcal{J}_{b}^{>0}$-torsors over something partially proper), it suffices to show that they have the same rank $1$ points, and therefore it suffices to show that they identify with the same subspace after taking the fiber of $\tilde{\pi}_{\mathrm{HT}}^{b,*}$ over a rank $1$ point of $\widetilde{\mathcal{F}\ell}_{G,\mu^{-1}}^{b,\Diamond}$, and this reduces us to Corollary 3.4. ∎
By the Cartesian diagram (18), if we look at the sheaf
$\tilde{s}_{b}^{*}i_{b}^{*}R\pi_{\mathrm{HT}!}(\overline{\mathbb{F}}_{\ell})$
we can see that this is canonically identified with $R\tilde{\pi}_{\mathfrak{Ig}!}^{b,\partial}(\overline{\mathbb{F}}_{\ell})\in\mathrm{D}(J_{b}(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})$ pulled back along the natural map $\widetilde{\mathcal{F}\ell}_{G,\mu^{-1}}^{b,\Diamond}/\underline{G(\mathbb{Q}_{p})}]\rightarrow[\mathop{\rm Spd}(C)/\underline{J_{b}(\mathbb{Q}_{p})}]$. Moreover, we can identify $R\tilde{\pi}_{\mathfrak{Ig}!}^{b,\partial}(\overline{\mathbb{F}}_{\ell})$ simply with the complex $V_{b}:=R\Gamma_{c-\partial}(\mathrm{Ig}^{b},\overline{\mathbb{F}}_{\ell})$ of $J_{b}(\mathbb{Q}_{p})$-modules, as in Corollary 3.5. We recall that we have an identification $\mathrm{D}(\mathrm{Bun}_{G}^{b},\overline{\mathbb{F}}_{\ell})\simeq\mathrm{D}(J_{b}(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})$, which is induced via pullback along $p_{b}:\mathrm{Bun}_{G}^{b}\simeq[\mathop{\rm Spd}(C)/\mathcal{J}_{b}]\rightarrow[\ast/\underline{J_{b}(\mathbb{Q}_{p})}]$, as in [16, Proposition V.2.2]. The map $s_{b}$ is a section to $p_{b}$ induced by the splitting of the Harder-Narasimhan filtration and $s_{b}^{*}$ is inverse to $p_{b}^{*}$. If we write $\tilde{p}_{b}:[\mathcal{F}\ell^{b,\Diamond}_{G,\mu^{-1}}/\underline{G(\mathbb{Q}_{p})}]\rightarrow[\widetilde{\mathcal{F}\ell}^{b,\Diamond}_{G,\mu^{-1}}/\underline{G(\mathbb{Q}_{p})}]$ for the map obtained by the base-change of $p_{b}$ then $\tilde{s}_{b}$ will be a section of $\tilde{p}_{b}$. In particular, we conclude that
$\tilde{p}_{b}^{*}\tilde{s}_{b}^{*}i_{b}^{*}R\pi_{\mathrm{HT}!}(\overline{\mathbb{F}}_{\ell})\simeq i_{b}^{*}R\pi_{\mathrm{HT}!}(\overline{\mathbb{F}}_{\ell})$
is the pullback of $V_{b}$ along the map $\widetilde{\mathcal{F}\ell}^{b,\Diamond}_{G,\mu^{-1}}/\underline{G(\mathbb{Q}_{p})}]\rightarrow[\mathop{\rm Spd}(C)/\underline{J_{b}(\mathbb{Q}_{p})}]$. Abusively identifying $V_{b}$ and $p_{b}^{*}(V_{b})$ as in the identification $\mathrm{D}(\mathrm{Bun}_{G}^{b},\overline{\mathbb{F}}_{\ell})\simeq\mathrm{D}(J_{b}(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})$, this allows us to deduce an isomorphism
$h^{\leftarrow*}_{b}(V_{b})\simeq i_{b}^{*}R\pi_{\mathrm{HT}!}(\overline{\mathbb{F}}_{\ell}).$
We can further refine this using the following Lemma.
Lemma 3.18.
We have isomorphisms
$i_{b!}h_{b}^{\leftarrow*}(V_{b})\simeq h^{\leftarrow*}j_{b!}(V_{b})$
and
$i_{b*}h_{b}^{\leftarrow*}(V_{b})\simeq h^{\leftarrow*}Rj_{b*}(V_{b})$
of sheaves on $\mathcal{F}\ell_{G,\mu^{-1}}^{\Diamond}$.
Proof.
The first isomorphism follows from proper base-change ([56, Theorem 1.9 (ii)]). For the second isomorphism, we note that
$h^{\leftarrow}:[\mathcal{F}\ell_{G,\mu^{-1}}^{\Diamond}/\underline{G(\mathbb{Q}_{p})}]\rightarrow\mathrm{Bun}_{G,C}$
is cohomologically smooth separated and representable in locally spatial diamonds; therefore, the result follows by smooth base-change [56, Proposition 23.16 (2)]. ∎
Remark 3.19.
In particular, the graded pieces of the filtration
$R\Gamma([\mathcal{F}\ell_{G,\mu^{-1}}^{\Diamond}/\underline{G(\mathbb{Q}_{p})}],i_{b!}i_{b}^{*}(R\pi_{\mathrm{HT}!}(\overline{\mathbb{F}}_{\ell})))$
on the cohomology of the Shimura variety coming from excision applied to $R\pi_{\mathrm{HT}!}(\overline{\mathbb{F}}_{\ell})$ and the Newton stratification of the flag variety are identified with $Rh^{\rightarrow}_{*}h^{\leftarrow*}j_{b!}(V_{b})\in\mathrm{D}(G(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})$, and similarly for $R\Gamma([\mathcal{F}\ell_{G,\mu^{-1}}^{\Diamond}/\underline{G(\mathbb{Q}_{p})}],i_{b*}i_{b}^{*}(R\pi_{\mathrm{HT}!}(\overline{\mathbb{F}}_{\ell})))$ and $Rh^{\rightarrow}_{*}h^{\leftarrow*}Rj_{b*}(V_{b})$.
All in all, we get the following.
Proposition 3.20.
Assuming 1.13, we have a filtration on $R\Gamma_{c}(\mathcal{S}(\mathbf{G},X)_{K^{p},C},\overline{\mathbb{F}}_{\ell})$ by complexes of smooth representations of $G(\mathbb{Q}_{p})$, with graded pieces isomorphic to $Rh^{\rightarrow}_{*}h^{\leftarrow*}j_{b!}(V_{b})$, where $V_{b}:=R\Gamma_{c-\partial}(\mathrm{Ig}^{b},\overline{\mathbb{F}}_{\ell})\in\mathrm{D}(J_{b}(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})$ is regarded as a sheaf on $\mathrm{Bun}_{G}^{b}$ via the natural identificaiton $\mathrm{D}(\mathrm{Bun}_{G}^{b},\overline{\mathbb{F}}_{\ell})\simeq\mathrm{D}(J_{b}(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})$.
The functor $h^{\rightarrow}_{*}h^{\leftarrow*}(-)$ appearing on the graded pieces is manifestly related to the action on $\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$ by Hecke operators. In particular, for each geometric dominant cocharacter $\mu\in\mathbb{X}_{*}(T_{\overline{\mathbb{Q}}_{p}})^{+}$, we have a correspondence
$\mathrm{Hck}_{G,\leq\mu}$$\mathrm{Bun}_{G}$$\mathrm{Bun}_{G}\times\mathop{\rm Spd}(C)$$\scriptstyle{h_{\mu}^{\leftarrow}}$$\scriptstyle{h_{\mu}^{\rightarrow}}$
where $\mathrm{Hck}_{G,\leq\mu}$ is the stack parametrizing modifications $\mathcal{E}_{1}\rightarrow\mathcal{E}_{2}$ of a pair of $G$-bundles with meromorphy bounded by $\mu$ over the fixed untilt defined by $C$. We define the Hecke operator [16, Section IX.2]
$T_{\mu}:\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})\rightarrow\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})^{BW_{E_{\mu}}}$
$A\mapsto h_{\mu*}^{\rightarrow}(h_{\mu}^{\leftarrow*}(A)\otimes^{\mathbb{L}}\mathcal{S}_{\mu}),$
where $E_{\mu}$ is the reflex field of $\mu$ and $\mathcal{S}_{\mu}$ is a sheaf on $\mathrm{Hck}_{G,\leq\mu}$ attached to the highest weight tilting module $\mathcal{T}_{\mu}$ by geometric Satake.
Remark 3.21.
To see why this operator gives a functor of the desired shape, we note that we have an isomorphism $\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})\simeq\mathrm{D}(\mathrm{Bun}_{G}\times\mathop{\rm Spd}(C),\overline{\mathbb{F}}_{\ell})$ by [16, Corollary V.2.3.], and, under this isomorphism, the operator $T_{\mu}$ a priori only defines a functor $\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})\rightarrow\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$; however, it follows from [16, Corollary IX.2.3] and the surrounding discussion that this factors through the forgetful functor $\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})^{BW_{E_{\mu}}}\rightarrow\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$ forgetting the continuous $W_{E_{\mu}}$-action.
If we now let $\mu$ be the minuscule cocharacter appearing above then the Bialynicki-Birula map gives an isomorphism of diamonds between the open locus of $\mathrm{Hck}_{G,\leq\mu}$ where $\mathcal{E}_{1}$ is isomorphic to the trivial bundle and $\mathcal{F}\ell_{G,\mu^{-1}}^{\Diamond}/\underline{G(\mathbb{Q}_{p})}]$, which identifies $h_{\mu}^{\rightarrow}$ (resp. $h_{\mu}^{\leftarrow})$) with $h^{\rightarrow}$ (resp. $h^{\leftarrow}$). Moreover, this is a cohomologically smooth space of dimension $d:=\langle 2\rho_{G},\mu\rangle$, and we have an isomorphism $\mathcal{S}_{\mu}\simeq\overline{\mathbb{F}}_{\ell}[d](\frac{d}{2})$ [^8] This is true for Satake sheaf attached to the highest weight module $V_{\mu}$ by definition and this agrees with the highest weight tilting module $\mathcal{T}_{\mu}$, since $\mu$ is minuscule.. It follows, by proper base-change, that we have an isomorphism
$h^{\rightarrow}_{*}h^{\leftarrow*}j_{b!}(V_{b})\simeq j_{1}^{*}T_{\mu}j_{b!}(V_{b})[-d](-\frac{d}{2})$
of $G(\mathbb{Q}_{p})$-modules, where $1\in B(G)$ denotes the trivial element. This gives us Theorem 1.15.
Corollary 3.22.
Assuming 1.13, the complex $R\Gamma_{c}(\mathcal{S}(\mathbf{G},X)_{K^{p},C},\overline{\mathbb{F}}_{\ell})$ has a $G(\mathbb{Q}_{p})$-equivariant filtration with graded pieces given by $j_{1}^{*}T_{\mu}j_{b!}(V_{b})[-d](-\frac{d}{2})$ for varying $b\in B(G,\mu)$, where $V_{b}\simeq R\Gamma_{c-\partial}(\mathrm{Ig}^{b},\overline{\mathbb{F}}_{\ell})$ and we implicitly identify $T_{\mu}$ with its postcomposition with the forgetful functor $\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{p})^{BW_{E_{\mathfrak{p}}}}\rightarrow\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{p})$.
Moreover, we obtain that each graded piece is isomorphic to
$(R\Gamma_{c}(G,b,\mu)\otimes^{\mathbb{L}}_{\mathcal{H}(J_{b})}V_{b})[2d_{b}]$
as $G(\mathbb{Q}_{p})$-modules. Here
$R\Gamma_{c}(G,b,\mu):=\mathop{\rm colim}\nolimits_{K_{p}\rightarrow\{1\}}R\Gamma_{c}(\mathrm{Sht}(G,b,\mu)_{\infty,C}/\underline{K_{p}},\overline{\mathbb{F}}_{\ell}(d_{b}))$
is a complex of $G(\mathbb{Q}_{p})\times J_{b}(\mathbb{Q}_{p})\times W_{E_{\mathfrak{p}}}$-modules, where $\mathrm{Sht}(G,b,\mu)_{\infty,C}$ is as defined above, and $\overline{\mathbb{F}}_{\ell}(d_{b})$ is the sheaf with $J_{b}(\mathbb{Q}_{p})$-action given as in [37, Lemma 7.4].
Proof.
It remains to explain the description of $j_{1}^{*}T_{\mu}j_{b!}(V_{b})[-d](-\frac{d}{2})$. By applying the second part of [24, Proposition 10.3] and noting that $\mathcal{S}_{\mu}\simeq\overline{\mathbb{F}}_{\ell}[d](\frac{d}{2})$ since $\mu$ is minuscule (where we recall that, since $\mu$ is minuscule, the representation $\mathcal{T}_{\mu}$ agrees with the usual highest weight representation), we obtain that the graded pieces are isomorphic to
$\mathop{\rm colim}\nolimits_{K_{p}\rightarrow\{1\}}(R\Gamma_{c}(\mathrm{Sht}(G,b,\mu)_{\infty,C}/\underline{K_{p}},\overline{\mathbb{F}}_{\ell})\otimes^{\mathbb{L}}_{\mathcal{H}(J_{b})}V_{b}\otimes\kappa^{-1})[2d_{b}]$
as desired, where $\kappa$ is the character of $J_{b}(\mathbb{Q}_{p})$ defined by the action of $J_{b}(\mathbb{Q}_{p})$ on the compactly supported cohomology of the $\ell$-adically contractible group diamond $\mathcal{J}_{b}^{>0}$, where $\mathcal{J}_{b}\simeq\mathcal{J}_{b}^{>0}\ltimes J_{b}(\mathbb{Q}_{p})$ is the semi-direct product structure given by allowing $\mathrm{Aut}(\mathcal{E}_{b})$ to act on its canonical reduction. However, by combining this with [37, Lemma 7.6] and its proof, we can rewrite this as
$(\mathop{\rm colim}\nolimits_{K_{p}\rightarrow\{1\}}R\Gamma_{c}(\mathrm{Sht}(G,b,\mu)_{\infty,C}/\underline{K_{p}},\overline{\mathbb{F}}_{\ell}(d_{b}))\otimes^{\mathbb{L}}_{\mathcal{H}(J_{b})}V_{b})[2d_{b}],$
as desired (See [25, Corollary 1.9 (1)]). ∎
Remark 3.23.
The analogous statement involving the full Hecke operator $T_{\mu}:\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})\rightarrow\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})^{BW_{E_{\mu}}}$ and the natural Weil group action on $R\Gamma_{c}(\mathcal{S}(\mathbf{G},X)_{K^{p},C})$ also holds. The process of matching the actions of the inertia group $I_{E_{\mathfrak{p}}}\subset W_{E_{\mathfrak{p}}}$ is fairly straight forward; in particular, one just needs to check that all the above Cartesian diagrams descend to $\breve{E}_{\mathfrak{p}}$. The process of matching the Frobenius descent datum is a bit tricky however so we do not address this here; for a relevant discussion of this see [10, Sections 8.2,8.4.9].
4. The Local Results
In this section, we establish the key local results (Corollaries 4.27,4.29) necessary for the proof of Theorem 1.8. We will let $G/\mathbb{Q}_{p}$ be a quasi-split connected reductive group with a choice of Borel $B$ and maximal torus $T$ as before. For the application of these local results to Theorem 1.8, we will in practice take $G$ to be the base-change of the global group $\mathbf{G}$ to $\mathbb{Q}_{p}$.
More specifically, for a semi-simple $L$-parameter $\phi$, in §4.1, we review the spectral action of the category of quasi-coherent sheaves on the stack of L-parameters on $\mathrm{D}(\mathrm{Bun}_{G})$, as well as its relationship to the action of Hecke operators on $\mathrm{D}(\mathrm{Bun}_{G})$. This will allow us to introduce a certain localization functor $\mathrm{D}(\mathrm{Bun}_{G})\rightarrow\mathrm{D}(\mathrm{Bun}_{G})_{\phi}$ that Hansen will study more extensively in Appendix A. We establish some basic properties of the localization functor that will be used in the later sections, and then turn our attention to studying the derived category $\mathrm{D}(\mathrm{Bun}_{G})_{\phi}$ for $\phi$ induced from a generic toral L-parameter. In particular, in §4.2, we discuss how, assuming the Fargues-Scholze local Langlands correspondence is compatible with a suitably nice form of local Langlands, one can show that every object in $\mathrm{D}(\mathrm{Bun}_{G})_{\phi}$ can be described in terms of constituents of certain explicit parabolic inductions. We then collect in §4.2.2 various instances for which this compatibility holds, using the results of [47, 24, 31].
In §4.3, we turn to the perverse $t$-exactness of the Hecke operators on the subcategory $\mathrm{D}(\mathrm{Bun}_{G})_{\phi}$, explaining how the theory of geometric Eisenstein series allows one to verify the perverse t-exactness for the subcategory spanned by the inductions described in §4.2, under some possible constraints including the genericity and regularity of a toral parameter $\phi_{T}$ inducing $\phi$. This will allow us to prove our main results under some possible additional constraints. In §4.4, we study the relationship between the genericity and these additional constraints, for the groups for which compatibility of the Fargues-Scholze local Langlands is known, showing that the additional constraints always follow from genericity in the relevant cases. This will allow us to conclude the precise form of our main local results.
4.1. The Spectral Action
[…]
$\mathrm{Perf}(\mathfrak{X}_{\hat{G}})\rightarrow\mathrm{End}(\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})^{\omega})$
$C\mapsto\{A\mapsto C\star A\}$
which, extending by filtered colimits, gives
$\mathrm{Ind}(\mathrm{Perf}(\mathfrak{X}_{\hat{G}}))\rightarrow\mathrm{End}(\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})),$
where the notation $\mathrm{Ind}(-)$ is the operation described in [45, Section 5.3]. We recall the following basic properties of this action [^9] These properties follow from the construction, since it is given by [16, Corollary X.I.3] applied to the functorial in $I$ Hecke action of $\mathrm{Rep}_{\overline{\mathbb{F}}_{\ell}}(\phantom{}^{L}G^{I})$ on $\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})^{\omega}$, as at the end of [16, Section X.3]. In particular, property (1) immediately follows from this, and property (2) also follows since the action appearing in [16, Corollary X.I.3] is monoidal..
  1. (1)
    For $V=\boxtimes_{i\in I}V_{i}\in\mathrm{Rep}_{\overline{\mathbb{F}}_{\ell}}(\phantom{}^{L}G^{I})$, there is an attached vector bundle $C_{V}\in\mathrm{Perf}(\mathfrak{X}_{\hat{G}})^{BW_{\mathbb{Q}_{p}}^{I}}$ whose pullback to a $\overline{\mathbb{F}}_{\ell}$-point of $\mathfrak{X}_{\hat{G}}$ corresponding to a (not necessarily semi-simple) $L$-parameter $\tilde{\phi}:W_{\mathbb{Q}_{p}}\rightarrow\phantom{}^{L}G(\overline{\mathbb{F}}_{\ell})$ is the vector space $V$ with $W_{\mathbb{Q}_{p}}^{I}$-action given by $\boxtimes_{i\in I}r_{V_{i}}\circ\tilde{\phi}$, where $r_{V_{i}}:\phantom{}^{L}G\rightarrow\mathrm{GL}(V_{i})$ is morphism defined by $V_{i}$. More specifically, these bundles are given by pulling back $V|_{\hat{G}^{I}}$ along the natural map $\mathfrak{X}_{\hat{G}}\rightarrow[\mathop{\rm Spec}(\overline{\mathbb{F}}_{\ell})/\hat{G}$], where we note that these have a natural $W_{\mathbb{Q}_{p}}^{I}$-equivariant structure coming from the full action of $\phantom{}{}^{L}G$ on $V$. The endomorphism
    $C_{V}\star(-):\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})\rightarrow\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})^{BW_{\mathbb{Q}_{p}}^{I}}$
    is the Hecke operator $T_{V}$ attached to $V$.
  2. (2)
[…]
    $(C_{1}\otimes^{\mathbb{L}}C_{2})\star(-)\simeq C_{1}\star(C_{2}\star(-)).$
Remark 4.1.
The Hecke operators appearing here are slightly different then the ones discussed in §3.2. In particular, for $\mu$ a geometric dominant cocharacter of $G$, we recall that there is a canonical away of attaching a representation of $\phantom{}{}^{L}G$. Namely, by the procedure of [38, Lemma 2.1.2], the highest weight representation $V_{\mu}$ of $\hat{G}$ descends to a representation of $\hat{G}\rtimes W_{E_{\mu}}$ where $E_{\mu}$ is the reflex field of the cocharacter $\mu$. This can then be induced along the inclusion $W_{E_{\mu}}\subset W_{\mathbb{Q}_{p}}$ to get a full representation of $\phantom{}{}^{L}G$ which we denote by $V$. Correspondingly, there is a Hecke operator
$T_{V}:\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})\rightarrow\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})^{BW_{\mathbb{Q}_{p}}}$
attached to the representation of $\phantom{}{}^{L}G$. This Hecke operator factors as
$T_{\mu}:\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})\rightarrow\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})^{BW_{E_{\mu}}}$
composed with the map $\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})^{BW_{E_{\mu}}}\rightarrow\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})^{BW_{\mathbb{Q}_{p}}}$ induced by induction along the inclusion $W_{E_{\mu}}\subset W_{\mathbb{Q}_{p}}$ (cf. the discussion on [16, Page 322]). The Hecke operator $T_{\mu}:\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})\rightarrow\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})^{BW_{E_{\mu}}}$ was the one appearing in §3.2. We will often cite results for Hecke operators $T_{V}$ for $V\in\mathrm{Rep}_{\overline{\mathbb{F}}_{\ell}}(\phantom{}^{L}G)$ to deduce results about the $T_{\mu}$, implicitly using this dictionary in the process.
[…]
$A\simeq\mathcal{O}_{\mathfrak{X}_{\hat{G}}}\star A\rightarrow\mathcal{O}_{\mathfrak{X}_{\hat{G}}}\star A\simeq A$
[…]
$\phi_{A}^{\mathrm{FS}}:W_{\mathbb{Q}_{p}}\rightarrow\phantom{}^{L}G(\overline{\mathbb{F}}_{\ell})$
[…]
With this in hand, we can make our key definition.
Definition 4.2.
We define $\iota_{\phi}:\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi}\hookrightarrow\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$ to be the full-subcategory of objects $A$ for which the endomorphisms $A\rightarrow A$ induced by $f\in\mathcal{O}_{\mathfrak{X}_{\hat{G}}}(\mathfrak{X}_{\mathfrak{X}_{\hat{G}}})\setminus\mathfrak{m}_{\phi}$ are isomorphisms.
It is easy to check that the subcategory $\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi}\subset\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$ is preserved under colimits and limits, and therefore, by the $\infty$-categorical adjoint functor Theorem [45, Corollary 5.5.2.9], there exists a left adjoint to the inclusion $\iota_{\phi}$ denoted by $\mathcal{L}_{\phi}$ (See Appendix A for details). We define $(-)_{\phi}:=\iota_{\phi}\mathcal{L}_{\phi}(-)$. This, by the fully faithfulness of $\iota_{\phi}$, will define an idempotent functor.
We now have the following key Lemma.
Lemma 4.3.
The following is true.
  1. (1)
    Any Schur irreducible object $A\in\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi}$ has Fargues-Scholze parameter equal to $\phi$ as conjugacy classes of parameters.
  2. (2)
    Given $V\in\mathrm{Rep}_{\overline{\mathbb{F}}_{\ell}}(\phantom{}^{L}G^{I})$, the Hecke operator $T_{V}:\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})\rightarrow\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})^{BW_{\mathbb{Q}_{p}}^{I}}$ takes the subcategory $\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi}$ to $\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi}^{BW_{\mathbb{Q}_{p}}^{I}}$, and there is a natural isomorphism $T_{V}((-)_{\phi})\simeq(T_{V}(-))_{\phi}$.
  3. (3)
    Given $A\in\mathrm{D}(G(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})$, we have an isomorphism
    $R\Gamma(K_{p}^{\mathrm{hs}},A)_{\mathfrak{m}}\simeq R\Gamma(K_{p}^{\mathrm{hs}},A_{\phi_{\mathfrak{m}}}),$
    where the LHS is the usual localization under the smooth unramified Hecke algebra $H_{K_{p}^{\mathrm{hs}}}$ at some maximal ideal $\mathfrak{m}\subset H_{K_{p}^{\mathrm{hs}}}$. Here we abuse notation and identify $A\in\mathrm{D}(G(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})$ with its image under the fully faithful embedding $\mathrm{D}(G(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})\xrightarrow{j_{1!}}\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$.
  4. (4)
    If $A\in\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$ is ULA then one has a direct sum decomposition
    $A\simeq\bigoplus_{\phi}A_{\phi}$
    ranging over all semi-simple $L$-parameters.
Proof.
Claims (2) and (4) follow from Proposition A.2 and Proposition A.5, respectively, where for claim (2) we use the relationship between Hecke operators and the spectral action described above.
[…]
For (3), we use the arguments in Koshikawa [37, Page 6]. Consider the map
$\mathcal{O}_{\mathfrak{X}_{\hat{G}}}(\mathfrak{X}_{\hat{G}})\rightarrow\mathrm{End}_{G(\mathbb{Q}_{p})}(\mathrm{cInd}_{K_{p}^{\mathrm{hs}}}^{G(\mathbb{Q}_{p})}(\overline{\mathbb{F}}_{\ell}))\simeq H_{K_{p}^{\mathrm{hs}}}^{\mathrm{op}}$
[…]
$R\mathrm{Hom}(\mathrm{cInd}_{K_{p}^{\mathrm{hs}}}^{G(\mathbb{Q}_{p})}(\overline{\mathbb{F}}_{\ell}),A_{\phi_{\mathfrak{m}}})\simeq R\mathrm{Hom}(\mathrm{cInd}_{K_{p}^{\mathrm{hs}}}^{G(\mathbb{Q}_{p})}(\overline{\mathbb{F}}_{\ell}),A)_{\mathfrak{m}_{\phi_{\mathfrak{m}}}}.$
Using Frobenius reciprocity, this gives an identification:
$R\Gamma(K_{p}^{\mathrm{hs}},A_{\phi_{\mathfrak{m}}})\simeq R\Gamma(K_{p}^{\mathrm{hs}},A)_{\mathfrak{m}_{\phi_{\mathfrak{m}}}},$
but the RHS identifies with $R\Gamma(K_{p}^{\mathrm{hs}},A)_{\mathfrak{m}}$, as explained above. ∎
We note that we get the following Corollary of this.
Corollary 4.4.
Let $A$ be a complex of smooth $G(\mathbb{Q}_{p})$-representations which is admissible (i.e $A^{K}$ is a perfect complex for all open pro-$p$ $K\subset G(\mathbb{Q}_{p})$). We then have a decomposition
$A\simeq\bigoplus_{\phi}A_{\phi}$
running over semisimple $L$-parameters, where any irreducible constituent $\pi$ of $A_{\phi}$ has Fargues-Scholze parameter equal to $\phi$, as conjugacy classes of parameters.
Proof.
This follows by applying to Lemma 4.3 (1) and (4) to the full subcategory $\mathrm{D}(G(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})\subset\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$, where we recall that a sheaf $A\in\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$ lies in the ULA subcategory if and only if its pullbacks $j_{b}^{*}(A)\in\mathrm{D}(\mathrm{Bun}_{G}^{b},\overline{\mathbb{F}}_{\ell})\simeq\mathrm{D}(J_{b}(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})$ for all $b\in B(G)$ are admissible as in the statement of the Corollary ([16, Theorem V.7.1]) . ∎
We now turn our attention to studying the subcategory $\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$ in certain cases coming from generic toral parameters, by using compatibility of the Fargues-Scholze local Langlands correspondence with more classical and well understood instances of local Langlands.
4.2. Local-Global Compatibility of the Fargues-Scholze Local Langlands
Now our goal is to describe the subcategory $\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi}$ more explicitly, using the results of [24] in the case that $\phi$ is induced from a generic toral parameter $\phi_{T}$, as in Definition 1.1. To do this, we will need to have some information about the Fargues-Scholze local Langlands correspondence. We first introduce the main result in this direction. We keep our running assumption that $\ell\nmid|\pi_{0}(Z(G))|$, so that the localized category $\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi}$ described in the previous section is well-defined.
4.2.1. Implications of Compatibility for Generic Parameters
We let $B(G)_{\mathrm{un}}:=\text{Im}{(B(T)\rightarrow B(G))}$. We recall that these are precisely the elements $b\in B(G)$ such that the $\sigma$-centralizer $J_{b}$ is quasi-split ([24, Lemma 2.12]). In particular, the fixed choice of Borel $B\subset G$ transfers to a Borel subgroup $B_{b}$ for all $b\in B(G)_{\mathrm{un}}$, and $J_{b}\simeq M_{b}$ under the inner twisting, where $M_{b}\subset G$ is the Levi subgroup of $G$ determined by the centralizer of the slope homomorphism of $b$ in $G$. We let $\delta_{P_{b}}$ denote the modulus character of the standard parabolic $P_{b}$ of $G$ with Levi factor $M_{b}$ transferred to $J_{b}$ under the inner twisting (See [25, Definition 3.14] for a more precise discussion, where we use the same conventions for the modulus character as loc.cit). We set $W_{b}:=W_{G}/W_{M_{b}}$ to be the quotient of the relative Weyl group of $G$ by the relative Weyl group of $M_{b}$. We identify $W_{b}$ with a choice of representatives in $w\in W_{G}$ of minimal length. We set $\rho_{b,w}:=i_{B_{b}}^{J_{b}}(\chi^{w})\otimes\delta_{P_{b}}^{-1/2}$ to be the normalized parabolic induction of $\chi^{w}$, where $\chi$ is the character of $T(\mathbb{Q}_{p})$ attached to a toral parameter $\phi_{T}$ under local class field theory and $\delta_{P_{b}}$ is the modulus character of $M_{b}\simeq J_{b}$ (Here we have slightly abused notation by omitting $\chi$ from the notation, where it will usually be clear from context. If further clarifications is necessary then we will write $\rho_{b,w}^{\chi}$.).
We will need to assume the following properties of the Fargues-Scholze local Langlands correspondence, as in [24, Assumption 6.5].
Assumption 4.5.
For a connected reductive group $H/\mathbb{Q}_{p}$, we have
  • the set $\Pi(H)$ of smooth irreducible $\overline{\mathbb{Q}}_{\ell}$-representations of $H(\mathbb{Q}_{p})$,
  • the set $\Phi(H)$ of conjugacy classes of continuous maps
    $W_{\mathbb{Q}_{p}}\times\mathrm{SL}(2,\overline{\mathbb{Q}}_{\ell})\rightarrow\phantom{}^{L}H(\overline{\mathbb{Q}}_{\ell})$
    where $\overline{\mathbb{Q}}_{\ell}$ has the discrete topology, $\mathrm{SL}(2,\overline{\mathbb{Q}}_{\ell})$ acts via an algebraic representation, $W_{\mathbb{Q}_{p}}$ acts through a semi-simple homorphism (in the sense that if it factors through $\phantom{}{}^{L}P$ for a proper parabolic $P\subset G$ then it also factors through the associated Levi $\phantom{}{}^{L}M$), and the map respects the action of $W_{\mathbb{Q}_{p}}$ on $\phantom{}{}^{L}H(\overline{\mathbb{Q}}_{\ell})$, the $L$-group of $H$,
  • the set $\Phi^{\mathrm{ss}}(H)$ of continuous (where again $\overline{\mathbb{Q}}_{\ell}$ has the discrete topology) semi-simple homomorphisms,
    $W_{\mathbb{Q}_{p}}\rightarrow\phantom{}^{L}H(\overline{\mathbb{Q}}_{\ell}),$
  • and the semi-simplification map $(-)^{\mathrm{ss}}:\Phi(H)\rightarrow\Phi^{\mathrm{ss}}(H)$ defined by precomposition with
    $W_{\mathbb{Q}_{p}}\rightarrow W_{\mathbb{Q}_{p}}\times\mathrm{SL}(2,\overline{\mathbb{Q}}_{\ell})$
    $g\mapsto(g,\begin{pmatrix}|g|^{1/2}&0\\ 0&|g|^{-1/2}\end{pmatrix})$.
Then, we assume that there exists a collection of maps
$\mathrm{LLC}_{b}:\Pi(J_{b})\rightarrow\Phi(J_{b})$
$\rho\mapsto\phi_{\rho},$
for all $b\in B(G)$, satisfying the following properties:
  1. (1)
    The diagram
    $\Pi(J_{b})$$\Phi(J_{b})$$\Phi^{\mathrm{ss}}(J_{b})$$\scriptstyle{\mathrm{LLC}_{b}}$$\scriptstyle{\mathrm{LLC}^{\mathrm{FS}}_{b}}$$\scriptstyle{(-)^{\mathrm{ss}}}$
    commutes, where $\mathrm{LLC}_{b}^{\mathrm{FS}}$ is the Fargues-Scholze local Langlands correspondence for $J_{b}$.
  2. (2)
    For $\rho$ a non-zero smooth irreducible representation of $J_{b}(\mathbb{Q}_{p})$ for some $b\in B(G)$, consider $\phi_{\rho}$ as an element of $\Phi(G)$ given by composing with the twisted embedding $\phantom{}{}^{L}J_{b}(\overline{\mathbb{Q}}_{\ell})\simeq\phantom{}^{L}M_{b}(\overline{\mathbb{Q}}_{\ell})\rightarrow\phantom{}^{L}G(\overline{\mathbb{Q}}_{\ell})$ (as defined in [16, Section IX.7.1]). Then $\phi_{\rho}$ factors through the natural (up to $\hat{G}-$conjugacy) embedding $\phantom{}{}^{L}T\rightarrow\phantom{}^{L}G$ if and only if $b\in B(G)_{\mathrm{un}}$.
  3. (3)
    If $\rho$ is a non-zero representation of $J_{b}(\mathbb{Q}_{p})$ such that $W_{\mathbb{Q}_{p}}\times\mathrm{SL}(2,\overline{\mathbb{Q}}_{\ell})\xrightarrow{\phi_{\rho}}\phantom{}^{L}J_{b}(\overline{\mathbb{Q}}_{\ell})\rightarrow\phantom{}^{L}G(\overline{\mathbb{Q}}_{\ell})$ factors through the embedding $\phantom{}{}^{L}T\rightarrow\phantom{}^{L}G$ with induced parameter $\phi_{T}$, then, by (2), the element $b$ is unramified, and we require that $\rho$ is isomorphic to an irreducible subquotient of $i_{B_{b}}^{J_{b}}(\chi^{w})\otimes\delta_{P_{b}}^{-1/2}$ for $w\in W_{b}$, where $\chi:T(\mathbb{Q}_{p})\rightarrow\overline{\mathbb{Q}}_{\ell}^{*}$ is the character attached to $\phi_{T}$ by class field theory and the notation is as above. Here the map $\phantom{}{}^{L}T\rightarrow\phantom{}^{L}G$ is the non-twisted embedding and the map $\phantom{}{}^{L}J_{b}\rightarrow\phantom{}^{L}G$ is the twisted embedding.
The importance of this assumption is that it allows us to deduce the following Proposition.
Proposition 4.6.
Assuming 4.5, we have that the following is true for a parameter $\phi$ induced from a generic parameter $\phi_{T}$. Given any Schur-irreducible object $A\in\mathrm{D}(\mathrm{Bun}_{G}^{b},\overline{\mathbb{F}}_{\ell})_{\phi}\subset\mathrm{D}(\mathrm{Bun}_{G}^{b},\overline{\mathbb{F}}_{\ell})\simeq\mathrm{D}(J_{b}(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})$ then $A$ is non-zero if and only if $b\in B(G)_{\mathrm{un}}$, and in this case it must be an irreducible sub-quotient of $\rho_{b,w}$, for some $w\in W_{b}$.
Proof.
This follows from combining the proof of [24, Corollary 6.7] with Lemma 4.3 (1). ∎
Since we want some flexibility in the groups for which we have the above results, we discuss how Assumption 4.5 behaves under central isogenies, and then spell out the main relevant cases for which this assumption holds.
4.2.2. Assumption 4.5 under Central Isogenies
We consider an injective map $\psi:G^{\prime}\hookrightarrow G$ of connected reductive groups which induces an isomorphism of derived groups, and the induced map $\psi_{\mathrm{Bun}}:B(G^{\prime})\rightarrow B(G)$ on the associated Kottwitz sets. We now have the following Lemma.
Lemma 4.7.
If $\psi:G^{\prime}\rightarrow G$ is an injective map which induces an isomorphism on derived groups then it follows that $\psi_{\mathrm{Bun}}:B(G^{\prime})\rightarrow B(G)$ induces an injection $\psi_{b^{\prime}}:J_{b^{\prime}}\rightarrow J_{b}$ for all $b=\psi_{\mathrm{Bun}}(b^{\prime})$ and $b^{\prime}\in B(G^{\prime})$, which is an isomorphism on the derived groups.
Proof.
Since $\psi$ is an inclusion it easily follows that it induces an inclusion $J_{b^{\prime}}\rightarrow J_{b}$ of $\sigma$-centralizers. To see that it induces an isomorphism on derived groups, recall that $J_{b}$ is an inner form of a Levi subgroup $M_{b}$ of $G$ given by the centralizer of the slope homomorphism of $b$. The preimage of $M_{b}$ under $\psi$ defines a Levi subgroup $M_{b^{\prime}}$ of $G$ which will be the centralizer of the slope homomorphism of $b^{\prime}$, since $\psi$ induces an isomorphism on derived groups and in turn also on adjoint groups (since $(G^{\mathrm{der}})^{\mathrm{ad}}\simeq G^{\mathrm{ad}}$)). Moreover, the inner twisting from $J_{b}$ to $M_{b}$ and $J_{b^{\prime}}$ to $M_{b^{\prime}}$ are compatible with the inclusion in the sense that the inclusion $J_{b^{\prime}}\rightarrow J_{b}$ is given by applying the inner twist of $M_{b}$ to the inclusion $M_{b^{\prime}}\rightarrow M_{b}$. Since the formation of derived groups respects inner twists, this reduces us to showing that the map $M_{b^{\prime}}\rightarrow M_{b}$ on Levi subgroups induces an isomorphism on the derived groups, and this is clear. ∎
We now consider a map $\mathrm{LLC}_{\mathrm{Bun}_{G}}:\bigsqcup_{b\in B(G)}\Pi(J_{b})\rightarrow\Phi(G)$ determined by components $\mathrm{LLC}_{b}:\Pi(J_{b})\rightarrow\Phi(J_{b})$ and satisfying Assumption 4.5. We now wish to define $\mathrm{LLC}_{\mathrm{Bun}_{G^{\prime}}}:\bigsqcup_{b^{\prime}\in B(G^{\prime})}\Pi(J_{b})\rightarrow\Phi(G^{\prime})$ in terms of $\mathrm{LLC}_{\mathrm{Bun}_{G}}$, and show that it also satisfies Assumption 4.5. To do this, we note that, for varying $b^{\prime}\in B(G^{\prime})$, we define $\mathrm{LLC}_{b^{\prime}}:\Pi(J_{b^{\prime}})\rightarrow\Phi(J_{b^{\prime}})$ to be the correspondence that makes the following diagram
$\Pi(J_{b})$$\Phi(J_{b})$$\Pi(J_{b^{\prime}})$$\Phi(J_{b^{\prime}})$$\scriptstyle{\mathrm{LLC}_{b}}$$\scriptstyle{\mathrm{LLC}_{b^{\prime}}}$
commute, where $b:=\psi_{\mathrm{Bun}}(b^{\prime})$. Here the right vertical arrow is given by composing a parameter $\phi:W_{\mathbb{Q}_{p}}\times\mathrm{SL}_{2}(\overline{\mathbb{Q}}_{\ell})\rightarrow\phantom{}^{L}J_{b}(\overline{\mathbb{Q}}_{\ell})$ with the induced map $\phantom{}{}^{L}J_{b}\rightarrow\phantom{}^{L}J_{b^{\prime}}$ on the dual groups, and the left vertical arrow is not a map at all, it is a correspondence defined by the subset of $\Pi(J_{b})\times\Pi(J_{b^{\prime}})$ consisting of pairs $(\pi_{b},\pi_{b^{\prime}})$ such that $\pi_{b^{\prime}}$ is a constituent of the restriction of $\pi_{b}$ to $J_{b^{\prime}}(\mathbb{Q}_{p})$. We will now show that this gives rise to a well-defined map under our assumptions on $\psi$. Given a representation $\pi_{b^{\prime}}\in\Pi(J_{b^{\prime}})$, it follows by [63, Proposition 2.2] and the previous Lemma that we can find a lift $\pi_{b}\in\Pi(J_{b})$ such that $\pi_{b^{\prime}}$ is an irreducible constituent of $\pi_{b}|_{J_{b^{\prime}}(\mathbb{Q}_{p})}$. It also follows from [62, Lemma 2.1] that the set $\Pi_{\pi_{b}}(J_{b^{\prime}})$ of representations of $J_{b^{\prime}}$ occurring in the restriction of $\pi_{b}$ is finite. Now, using the previous Lemma, we have the following.
Lemma 4.8.
[21, Corollary 2.5] For the map $J_{b^{\prime}}\rightarrow J_{b}$ of $\sigma$-centralizers induced by a map $\psi$ as above and $\pi_{b}^{1},\pi_{b}^{2}\in\Pi(J_{b})$, the following are equivalent.
  1. (1)
    There exists a character $\chi\in(J_{b}(\mathbb{Q}_{p})/J_{b^{\prime}}(\mathbb{Q}_{p}))^{\vee}$ such that $\pi_{b}^{1}\simeq\pi_{b}^{2}\otimes\chi$, where $(-)^{\vee}$ denotes the Pontryagin dual.
  2. (2)
    $\Pi_{\pi_{b}^{1}}(J_{b^{\prime}})\cap\Pi_{\pi_{b}^{2}}(J_{b^{\prime}})\neq\emptyset$
  3. (3)
    $\Pi_{\pi_{b}^{1}}(J_{b^{\prime}})=\Pi_{\pi_{b}^{2}}(J_{b^{\prime}})$.
Now we can use this to define $\mathrm{LLC}_{b^{\prime}}:\Pi(J_{b^{\prime}})\rightarrow\Phi(J_{b^{\prime}})$ in terms of $\mathrm{LLC}_{b}:\Pi(J_{b})\rightarrow\Phi(J_{b})$ for $\mathrm{LLC}_{\mathrm{Bun}_{G}}$ satisfying Assumption 4.5. Namely, for $\pi_{b^{\prime}}\in\Pi(J_{b^{\prime}})$, we let $\pi_{b}\in\Pi(J_{b})$ be a representation such that $\pi_{b^{\prime}}$ occurs as an irreducible constituent of $\pi_{b}|_{J_{b^{\prime}}(\mathbb{Q}_{p})}$. We set $\phi_{\pi_{b^{\prime}}}$ to be the parameter $\phi_{\pi_{b}}$ attached to $\pi_{b}$ under $\mathrm{LLC}_{b}$ composed with the map $\phantom{}{}^{L}J_{b}\rightarrow\phantom{}^{L}J_{b^{\prime}}$ on dual groups induced by $\psi$. By the previous Lemma, any two choices of lifts $\pi^{1}_{b}$ and $\pi^{2}_{b}$ of $\pi_{b^{\prime}}$ will differ by a character twist of $\chi\in(J_{b}(\mathbb{Q}_{p})/J_{b^{\prime}}(\mathbb{Q}_{p}))^{\vee}$. We note that the Fargues-Scholze local Langlands correspondence is compatible with character twists [16, Theorem I.9.6 (ii)]. Since $\mathrm{LLC}_{b}$ is compatible with the Fargues-Scholze local Langlands after semi-simplification by assumption, it follows that the same is true for $\mathrm{LLC}_{b^{\prime}}$. Therefore, $\phi_{\pi^{1}_{b}}$ and $\phi_{\pi^{2}_{b}}$ differ by a character twist that becomes trivial after composing with $\phantom{}{}^{L}J_{b}\rightarrow\phantom{}^{L}J_{b^{\prime}}$, and so $\phi_{\pi_{b}}$ does not depend on the choice of lift. We let $\mathrm{LLC}_{\mathrm{Bun}_{G^{\prime}}}:\bigsqcup_{b^{\prime}\in B(G^{\prime})}\Pi(J_{b^{\prime}})\rightarrow\Phi(G)$ be the local Langlands correspondence defined by the $\mathrm{LLC}_{b^{\prime}}$ for $b^{\prime}$ varying. We now prove that our assumption is compatible with central isogenies.
Proposition 4.9.
Suppose we have an injective map $\psi:G^{\prime}\hookrightarrow G$ of quasi-split connected reductive groups inducing an isomorphism on derived groups. Assume we have a local Langlands correspondence $\mathrm{LLC}_{\mathrm{Bun}_{G}}:\bigsqcup_{b\in B(G)}\Pi(J_{b})\rightarrow\Phi(G)$ such that Assumption 4.5 holds. If we let $\mathrm{LLC}_{\mathrm{Bun}_{G^{\prime}}}:\bigsqcup_{b^{\prime}\in B(G^{\prime})}\Pi(J_{b^{\prime}})\rightarrow\Phi(G^{\prime})$ be the local Langlands correspondence induced by $\mathrm{LLC}_{\mathrm{Bun}_{G}}$ and $\psi$ as above then $\mathrm{LLC}_{\mathrm{Bun}_{G^{\prime}}}$ satisfies Assumption 4.5 as well.
Proof.
We note, since the Fargues-Scholze local Langlands correspondence is compatible with maps $G^{\prime}\rightarrow G$ that induce an isomorphism of adjoint groups [16, Theorem I.9.6 (v)] (which follows from the map inducing an isomorphism on derived groups), it follows by the above construction that if Assumption 4.5 (1) holds true for $\mathrm{LLC}_{\mathrm{Bun}_{G}}$ then it also holds true for $\mathrm{LLC}_{\mathrm{Bun}_{G^{\prime}}}$. Suppose we have $b^{\prime}\in B(G^{\prime})$ mapping to $b\in B(G)_{\mathrm{un}}$. We let $B_{b}\subset J_{b}$ be the corresponding Borel then, since the map $J_{b^{\prime}}\rightarrow J_{b}$ induces an isomorphism on adjoint groups by Lemma 4.7, it follows that $B\cap J_{b^{\prime}}=:B_{b^{\prime}}\subset J_{b^{\prime}}$ is a Borel of $J_{b^{\prime}}$. In particular, $b^{\prime}$ must be an unramified element of $B(G^{\prime})_{\mathrm{un}}$. Now, the map $J_{b^{\prime}}\rightarrow J_{b}$ induces an isomorphism
$J_{b^{\prime}}/B^{\prime}\simeq J_{b}/B.$
If we let $T$ be the maximal split torus of $J_{b}$ then the preimage $T^{\prime}$ under $\psi_{b^{\prime}}$ the map induced by $\psi$ is a maximal torus of $J_{b^{\prime}}$, and the previous isomorphism of flag varieties implies that, given a character $\chi:T(\mathbb{Q}_{p})\rightarrow\overline{\mathbb{Q}}_{\ell}^{*}$, we have an isomorphism:
$i_{B_{b}}^{J_{b}}(\chi)|_{J_{b^{\prime}}(\mathbb{Q}_{p})}\simeq i_{B_{b^{\prime}}}^{J_{b^{\prime}}}(\chi|_{T^{\prime}(\mathbb{Q}_{p})}).$
Indeed, recall that the left hand side is described by the set of compactly supported smooth functions on $J_{b}(\mathbb{Q}_{p})$ which transform under $B_{b}(\mathbb{Q}_{p})$ via the character $\chi$ of $T(\mathbb{Q}_{p})$ inflated to $B_{b}(\mathbb{Q}_{p})$. In particular, by the previous isomorphism of flag varieties, restricting to $J_{b^{\prime}}(\mathbb{Q}_{p})$ gives the collection of compactly supported smooth (since $J_{b^{\prime}}(\mathbb{Q}_{p})\rightarrow J_{b}(\mathbb{Q}_{p})$ will be a closed subgroup, by Lemma 4.7) functions on $J_{b^{\prime}}(\mathbb{Q}_{p})$ which transform under $B_{b^{\prime}}(\mathbb{Q}_{p})$ by the restriction of the inflated character $\chi$ to $B_{b^{\prime}}(\mathbb{Q}_{p})$, but this is just the inflation of $\chi|_{T^{\prime}(\mathbb{Q}_{p})}$ to $B_{b^{\prime}}(\mathbb{Q}_{p})$, giving the desired claim.
Given $\pi_{b^{\prime}}$ and a lift $\pi_{b}$ to $J_{b}$ then, by definition of $\phi_{\pi_{b}}$, we have that it is equal to
$W_{\mathbb{Q}_{p}}\times\mathrm{SL}_{2}(\overline{\mathbb{Q}}_{\ell})\xrightarrow{\phi_{\pi_{b}}}\phantom{}^{L}J_{b}(\overline{\mathbb{Q}}_{\ell})\rightarrow\phantom{}^{L}J_{b^{\prime}}(\overline{\mathbb{Q}}_{\ell})$
as a conjugacy class of parameters for $J_{b^{\prime}}$. Therefore, $\phi_{\pi_{b^{\prime}}}$ factors through $\phantom{}{}^{L}T^{\prime}$ if and only if $\phi_{\pi_{b}}$ factors through the preimage of $\phantom{}{}^{L}T^{\prime}$ under the map $\phantom{}{}^{L}J_{b}(\overline{\mathbb{Q}}_{\ell})\rightarrow\phantom{}^{L}J_{b^{\prime}}(\overline{\mathbb{Q}}_{\ell})$ of $L$-groups, but this is precisely $\phantom{}{}^{L}T$, and so Assumption 4.5 (2) holds for $\mathrm{LLC}_{\mathrm{Bun}_{G^{\prime}}}$. Moreover, by Assumption 4.5 (3) for $\mathrm{LLC}_{\mathrm{Bun}_{G}}$, we have that, in the above situation, $\pi_{b}$ is an irreducible sub-quotient of $i_{B_{b}}^{J_{b}}(\chi^{w})\otimes\delta_{P_{b}}^{-1/2}$, but this implies that $\pi_{b^{\prime}}$ is an irreducible constituent of the restriction $i_{B_{b^{\prime}}}^{J_{b^{\prime}}}(\chi^{w}|_{T^{\prime}(\mathbb{Q}_{p})})\otimes\delta_{P_{b^{\prime}}}^{-1/2}$. From this, it follows that Assumption 4.5 (3) also holds for holds for $\mathrm{LLC}_{\mathrm{Bun}_{G^{\prime}}}$. ∎
Now that we have shown this compatibility assumption is somewhat flexible, we can state the groups we know to satisfy Assumption 4.5. This result is largely contained in [27, 16, 31, 47], but we also want to consider an additional group $\mathrm{GU}_{2}$, where we have the following construction of $\mathrm{LLC}_{\mathrm{Bun}_{\mathrm{GU}_{2}}}$. For $L/\mathbb{Q}_{p}$ a finite extension, recall that $\mathrm{GU}_{2}/L$ can be written as
$\mathrm{GU}_{2}:=(\mathrm{GL}_{2}\times\mathrm{Res}_{L^{\prime}/L}\mathbb{G}_{m})/\mathbb{G}_{m},$
where $\mathbb{G}_{m}$ is embedded in $H:=\mathrm{GL}_{2}\times\mathrm{Res}_{L^{\prime}/L}(\mathbb{G}_{m})$ via $a\mapsto(\mathrm{diag}(a,a),a^{-1})$, and $L^{\prime}/L$ is an unramified quadratic extension. Let $\psi:B(H)\rightarrow B(\mathrm{GU}(2))$ and let $\tilde{\psi}:B(H)\rightarrow B(\mathrm{GL}_{2})$ be the map of Kottwitz sets. Given $b\in B(H)$, let $b^{\prime}=\psi(b)$, $\tilde{b}=\tilde{\psi}(b)$.
Lemma 4.10.
[…]
Proof.
We will show that we have an isomorphism
$J_{b^{\prime}}\simeq(J_{\tilde{b}}\times\mathrm{Res}_{L^{\prime}/L}\mathbb{G}_{m})/\mathbb{G}_{m}$
of groups over $L$. Assuming this, we see that since by Hilbert’s Theorem 90 we have $H^{1}(L,\mathbb{G}_{m})=0$, the projection map from $(J_{\tilde{b}}\times\mathrm{Res}_{L^{\prime}/L}\mathbb{G}_{m})$ to $J_{b^{\prime}}$ induces a short exact sequence on $L$-points
$0\rightarrow L^{\times}\rightarrow J_{\tilde{b}}(L)\times(L^{\prime})^{\times}\rightarrow J_{b^{\prime}}(L)\rightarrow 0,$
so we see that to give a smooth irreducible $\overline{\mathbb{Q}}_{\ell}$-representation of $J_{b^{\prime}}(L)$ is exactly equivalent to giving a smooth irreducible $\overline{\mathbb{Q}}_{\ell}$-representation of $J_{\tilde{b}}(L)$ and $(L^{\prime})^{\times}$ which agree on the restriction to the central copy of $L^{\times}$.
To see this isomorphism, recall that $J_{b}$ (resp. $J_{b^{\prime}},J_{\tilde{b}}$) is an inner form of $M_{b}$ (resp. $M_{b^{\prime}},M_{\tilde{b}}$), the Levi subgroup of $H$ (resp. $\mathrm{GU}_{2},\mathrm{GL}_{2}$) given by the centralizer of the slope homomorphism of $b$ (resp. $b^{\prime},\tilde{b}$). Note that we have $J_{b}=J_{\tilde{b}}\times\mathrm{Res}_{L^{\prime}/L}\mathbb{G}_{m}\subset H$, and the quotient map $J_{b}\rightarrow J_{b^{\prime}}$ induces an isomorphism on adjoint and derived subgroups of $J_{b^{\prime}}$. Moreover, we see that we have an isomorphism
$M_{b^{\prime}}\simeq(M_{\tilde{b}}\times\mathrm{Res}_{L^{\prime}/L}\mathbb{G}_{m})/\mathbb{G}_{m}$
and thus we have a surjective map $M_{b}\rightarrow M_{b^{\prime}}$, since $M_{b}=M_{\tilde{b}}\times\mathrm{Res}_{L^{\prime}/L}\mathbb{G}_{m}$. Moreover, we see that under these maps, we have isomorphisms $M_{b}^{\mathrm{ad}}\simeq M_{b^{\prime}}^{\mathrm{ad}}\simeq M_{\tilde{b}}^{\mathrm{ad}}$, and the inner twist $H^{1}(L,M_{b}^{\mathrm{ad}})$ corresponding to $J_{b}$ is, under this identification, the inner twist inducing $J_{b^{\prime}}$ and $J_{\tilde{b}}$. The identification of $J_{b^{\prime}}$ then follows. ∎
Moreover, we observe that we have an exact sequence of dual groups
$1\rightarrow\widehat{J}_{b^{\prime}}\rightarrow\widehat{J}_{b}\rightarrow\mathbb{G}_{m}\rightarrow 1,$
where the map $p:\widehat{J}_{b}\rightarrow\mathbb{G}_{m}$ is defined as follows. We can write $\widehat{J}_{b}=\widehat{J}_{\tilde{b}}\times\mathbb{G}_{m}^{2}$, and we have maps $\hat{i}_{1}:\widehat{J}_{\tilde{b}}\rightarrow\mathbb{G}_{m}$, $\hat{i}_{2}:\widehat{\mathrm{Res}_{L^{\prime}/L}\mathbb{G}_{m}}=\mathbb{G}_{m}^{2}\rightarrow\mathbb{G}_{m}$ induced from the inclusion maps $i_{1}:\mathbb{G}_{m}\hookrightarrow J_{\tilde{b}}$ and $i_{2}:\mathbb{G}_{m}\hookrightarrow\mathrm{Res}_{L^{\prime}/L}\mathbb{G}_{m}$, and $p(g,h)=\hat{i}_{1}(g)\hat{i}_{2}(h)^{-1}$.
Now, we want to define $\mathrm{LLC}_{\mathrm{Bun}_{\mathrm{GU}_{2}}}$ in terms of $\mathrm{LLC}_{\mathrm{Bun}_{H}}$. More precisely, for any $b^{\prime}\in B(\mathrm{GU}_{2})$ we define $\mathrm{LLC}_{b^{\prime}}:\Pi(J_{b^{\prime}})\rightarrow\Phi(J_{b^{\prime}})$ in terms of $\mathrm{LLC}_{b}:\Pi(J_{b})\rightarrow\Phi(J_{b})$ for $\mathrm{LLC}_{\mathrm{Bun}_{H}}$. For $\pi_{b^{\prime}}=(\tilde{\pi},\chi)\in\Pi(J_{b^{\prime}})$, we consider the image $\phi=\mathrm{LLC}_{\mathrm{Bun}_{H}}((\tilde{\pi},\chi))$, and we want to show that $\phi:W_{L}\times\mathrm{SL}_{2}(\overline{\mathbb{Q}}_{\ell})\rightarrow\phantom{}^{L}J_{b}(\overline{\mathbb{Q}}_{\ell})$ factors through $\phantom{}{}^{L}J_{b^{\prime}}(\overline{\mathbb{Q}}_{\ell})$. To see this, from the exact sequence above, it suffices to show that the composition of $\phi$ with the map $\phantom{}{}^{L}J_{b}(\overline{\mathbb{Q}}_{\ell})\rightarrow\phantom{}^{L}\mathbb{G}_{m}(\overline{\mathbb{Q}}_{\ell})$ is trivial. However, we observe that the condition that $\chi|_{L^{\times}}=\omega_{\tilde{\pi}}|_{L^{\times}}$ exactly implies that this image is trivial, since the composition is the $L$-parameter associated with the character $\omega_{\tilde{\pi}}\chi^{-1}|_{L^{\times}}$. Thus, we have an $L$-parameter $\phi^{\prime}:W_{L}\times\mathrm{SL}_{2}(\overline{\mathbb{Q}}_{\ell})\rightarrow\phantom{}^{L}J_{b^{\prime}}(\overline{\mathbb{Q}}_{\ell})$. We thus define the map $\mathrm{LLC}_{b^{\prime}}$ to take $\pi_{b^{\prime}}$ to $\phi^{\prime}$.
For the various groups discussed above, we will also want to check some assumptions on the prime $\ell$ relative to the group $G$ that we will need to assume for the proof our main results, since they are required for the arguments in [24]. In particular, we define the following.
Definition 4.11.
We say that $\ell$ is very decent with respect to $G$ if the following holds.
  1. (1)
    We have that $\ell\nmid|\pi_{0}(Z(G))|$, where $Z(G)$ denotes the center of $G$.
  2. (2)
    We have that $\ell\nmid|Q|$, where we recall that $Q$ is the smallest quotient through which the action of $W_{\mathbb{Q}_{p}}$ on $\hat{G}$ factors.
Remark 4.12.
The motivation for the first condition was already explained in §4.1; in particular, it is required for the construction of the spectral action and in turn the Fargues-Scholze local Langlands correspondence and the localized derived category $\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi}$. The second assumption comes from results on the theory of geometric Eisenstein series introduced in [24]. In particular, it will guarantee that the theory of tilting modules for the group $\hat{G}/\overline{\mathbb{F}}_{\ell}$ extends to a nice theory of tilting modules for the group $\phantom{}{}^{L}G:=\hat{G}\ltimes Q$, as discussed in [24, Section 9], which is used to establish the main results of [24]. In particular, the semisimplicity of the category of tilting modules for the full $L$-group $\phantom{}{}^{L}G$ requires this.
We now have the following Theorem about the groups we know to satisfy Assumption 4.5, preempting our discussion in the next section.
Theorem 4.13.
[27, 16, 31, 47] Assumption 4.5 is true and $\ell$ is very decent in the following cases.
  1. (1)
    The group $\mathrm{Res}_{L/\mathbb{Q}_{p}}(\mathrm{GSp}_{4})$ for $L/\mathbb{Q}_{p}$ an unramified extension with $p>2$ , $L:\mathbb{Q}_{p}]\geq 2$, and $\ell\nmid[L:\mathbb{Q}_{p}]$ or $L=\mathbb{Q}_{p}$ for all $p$ and $\ell$.
  2. (2)
    The groups $\mathrm{GU}_{n}$ or $\mathrm{U}_{n}$ for $n$ odd and defined with respect to an unramified quadratic extension $E/\mathbb{Q}_{p}$, and $\ell\neq 2$.
  3. (3)
    The group $\mathrm{Res}_{L/\mathbb{Q}_{p}}(\mathrm{GU}_{2})$ defined with respect to an unramified quadratic extension $L^{\prime}/L$, and $\ell$ such that $\ell\nmid 2[L:\mathbb{Q}_{p}]$.
  4. (4)
    The group $\mathrm{Res}_{L/\mathbb{Q}_{p}}(\mathrm{GL}_{n})$ for all $p$ and $\ell$ such that $\ell\nmid[L:\mathbb{Q}_{p}]$.
Proof.
We first start with the conditions on $\ell$. For all the groups described above the center is connected, and therefore condition (1) of Definition 4.11 is automatically satisfied. For condition (2) of Definition 4.11, it is straightforward. In particular, in the unramified restriction of scalar cases discussed in (1) and (4), the quotient $Q$ will be isomorphic to the Galois group of $L/\mathbb{Q}_{p}$, and in the unitary group case in (2) it will be isomorphic to the Galois group of the quadratic extension $E/\mathbb{Q}_{p}$. In case (3), it will be given by the Galois group of the extension $L^{\prime}/\mathbb{Q}_{p}$.
Now, we turn to Assumption 4.5 (1). For $\mathrm{GL}_{n}$, this follows from [16, Theorem I.9.6] and [31, Theorem 1.0.3], where $\mathrm{LLC}_{b}$ is given by the Harris-Taylor correspondence precomposed with Badulescu’s Jacquet Langlands map [2]. For $\mathrm{Res}_{L/\mathbb{Q}_{p}}(\mathrm{GSp}_{4})$ and $L/\mathbb{Q}_{p}$ as described above, this follows from [27, Theorem 1.1], where $\mathrm{LLC}_{b}$ is given by Harris-Taylor for the non-basic $b$ and Gan-Takeda [19] and Gan-Tantono [20] for the basic elements [^10] In the current version of [27], the assumption that $p>2$ is only used to invoke basic uniformization of abelian type Shimura varieties, but this condition is actually unnecessary if $L=\mathbb{Q}_{p}$, following the construction of Pappas-Rapoport [50] of the uniformization in Hodge type case when $p=2$ and the group is unramified.. For $\mathrm{U}_{n}$ or $\mathrm{GU}_{n}$, this is [47, Theorem 1.1], where $\mathrm{LLC}_{b}$ for $b\in B(G)$ was constructed by Mok [48] and Kaletha-Minguez-Shin-White [34].
Now we explain why Assumption 4.5 (2) is satisfied. We recall that if $J_{b}$ is a non quasi-split group then the fibers of the $\mathrm{LLC}_{b}$ over an $L$-parameter $\phi:W_{\mathbb{Q}_{p}}\times\mathrm{SL}_{2}(\overline{\mathbb{Q}}_{\ell})\rightarrow\phantom{}^{L}J_{b}(\overline{\mathbb{Q}}_{\ell})$ should be empty if $\phi$ factors through $\phantom{}{}^{L}M(\overline{\mathbb{Q}}_{\ell})\rightarrow\phantom{}^{L}G(\overline{\mathbb{Q}}_{\ell})$ via the the natural (equivalently twisted) embedding for a Levi subgroup $M\subset G$ which does not transfer to a Levi subgroup of $J_{b}$ (recalling that $G$ is quasi-split in this case, so that $J_{b}$ is an inner form of the quasi-split Levi of $G$). In particular, such parameters are called irrelevant, and we expect the fiber to be empty if and only if $\phi$ is irrelevant [35, Conjecture A.2]. Assumption 4.5 (2) reduces to checking that the fibers over the irrelevant $L$-parameters are empty for all $J_{b}$ and the $\mathrm{LLC}_{b}$ in question. This is because $b\in B(G)_{\mathrm{un}}$ if and only if $J_{b}$ is quasi-split or a torus (See [24, Lemma 2.12]) implying that if $b\notin B(G)_{\mathrm{un}}$ then all $L$-parameters $\phi\in\Phi(J_{b})$ which factor through $\phantom{}{}^{L}T$ are irrelevant. Since every representation in $\Pi(J_{b})$ occurs as the constituent of a parabolic induction of a supercuspidal representation of a Levi subgroup (where we recall that supercuspidals are in particular discrete series) to show this claim, it suffices to show that the semi-simplification of $\mathrm{LLC}_{b}$ is compatible with respect to parabolic induction and the semi-simplification of a local Langlands correspondence on each Levi $L_{b}\subset J_{b}$ (as in [16, Theorem I.9.6 (viii)]), and that the non-semisimplified local Langlands correspondence on the Levi sends discrete series representations to discrete $L$-parameters of $L_{b}$ (in the sense that they do not factor through some smaller Levi of $\phantom{}{}^{L}L_{b}$), as expected [35, Conjecture A.4]).
For the Harris-Taylor correspondence, it is know that the fibers over irrelevant parameters are empty by the standard properties of Jacquet-Langlands. In particular, it follows from the discussion before [40, Theorem 4.2.1] that the Harris-Taylor parameter for a discrete series representation of $\mathrm{GL}_{n}$ always defines a discrete L-parameter (in the sense that its L-parameter does not factor through a proper Levi subgroup) and is compatible with parabolic induction. The claim for $\mathrm{GL}_{m}(D_{\frac{1}{n}})$ now follows from the fact that the Jacquet Langlands transfer of any smooth irreducible representation of $D_{\frac{1}{n}}^{*}$ is discrete, and the compatibility of parabolic induction of Baduluescu’s Jacquet Langlands transfer discussed in [2, Section 3.1]. This implies the claim on irrelevant parameters for $J_{b}$ in the case of $G=\mathrm{GL}_{n}$, since $J_{b}$ in this case will always be a product of groups of the form $\mathrm{GL}_{m}(D_{\frac{1}{n}})$.
For $G=\mathrm{GU}_{n}$ or $\mathrm{U}_{n}$, we note that odd unitary groups are always quasi-split, so that the inner forms of such a group are always quasi-split; in particular, there are no irrelevant $L$-parameters for these groups and there is nothing to check. Therefore in this case, we can reduce to the claim on irrelevant $L$-parameters to $\mathrm{GL}_{n}$ and its inner forms, since these occur as direct factors of $J_{b}$ for such $G$, which was already explained above.
For $\mathrm{GSp}_{4}$, one needs to show the claim on irrelevant $L$-parameters for $\mathrm{LLC}_{\mathrm{GU}_{2}(D)}$, where $\mathrm{GU}_{2}(D)$ is the unique non-split inner form of $\mathrm{GSp}_{4}$. Here the claim on irrelevant $L$-parameters follows directly from the construction of Gan-Tantono (See the discussion before the main Theorem in [20]).
For $\mathrm{GU}_{2}$, we observe that since $\mathrm{GU}_{2}$$H:=\mathrm{GL}_{2}\times\mathrm{Res}_{L^{\prime}/L}(\mathbb{G}_{m})$ and, $\mathrm{GL}_{2}$ all have the same adjoint group, $b^{\prime}\in B(\mathrm{GU}_{2})_{\mathrm{un}}$ is unramified exactly when $b$ is unramified with notation as in Lemma 4.10. Now, let $\tilde{T}$ be a maximal split torus of $\mathrm{GL}_{2}$, and observe that $T^{\prime}=(\tilde{T}\times\mathrm{Res}_{L^{\prime}/L}\mathbb{G}_{m})/\mathbb{G}_{m}$ is a maximal torus of $\mathrm{GU}_{2}$ over $L$. Given $\pi_{b^{\prime}}=(\tilde{\pi},\chi)$, and $\phi_{\pi_{b^{\prime}}}:W_{L}\times\mathrm{SL}_{2}(\overline{\mathbb{Q}}_{\ell})\rightarrow\phantom{}^{L}J_{b^{\prime}}(\overline{\mathbb{Q}}_{\ell})$, we see from the construction that this factors through $\phantom{}{}^{L}T^{\prime}(\overline{\mathbb{Q}}_{\ell})$ exactly when the associated $L$-parameter for $H$, $\phi_{(\tilde{\pi},\chi)}:W_{L}\times\mathrm{SL}_{2}(\overline{\mathbb{Q}}_{\ell})\rightarrow\phantom{}^{L}J_{b}(\overline{\mathbb{Q}}_{\ell})$, factors through $\phantom{}{}^{L}T(\overline{\mathbb{Q}}_{\ell})$, where $T=\tilde{T}\times\mathrm{Res}_{L^{\prime}/L}\mathbb{G}_{m}$. Since Assumption 4.5 (2) is clearly compatible with taking products, the group $H$ via $a\mapsto(\mathrm{diag}(a,a),a^{-1})$ satisfies this assumption. Thus, we see that so does $\mathrm{GU}_{2}$, since if we let $T_{H}$ denote the maximal torus of $T$ such that its image under the projection to $\mathrm{GU}_{2}$ is $T$, and we let $\rho_{H}$ be a smooth irreducible representation of $H(L)$ lifting $\rho$, we see that by definition of $\mathrm{LLC}_{\mathrm{GU}_{2}}$ we have $\phi_{\rho}$ factors through $\phantom{}{}^{L}T$ if and only if $\phi_{\rho_{H}}$ factors through $\phantom{}{}^{L}T_{H}$.
Now we explain why Assumption 4.5 (3) is satisfied. First, note that any parameter $\phi:W_{\mathbb{Q}_{p}}\times\mathrm{SL}_{2}(\overline{\mathbb{Q}}_{\ell})\rightarrow\phantom{}^{L}G(\overline{\mathbb{Q}}_{\ell})$ induced from a toral parameter $\phi_{T}$ has necessarily trivial monodromy, since $\phantom{}{}^{L}T(\overline{\mathbb{Q}}_{\ell})$ consists only of semi-simple elements. Moreover, since $J_{b}$ is an inner form of $M_{b}$, it follows that the set of all distinct conjugacy classes of parameters $\phi^{\prime}:W_{\mathbb{Q}_{p}}\times\mathrm{SL}_{2}(\overline{\mathbb{Q}}_{\ell})\rightarrow\phantom{}^{L}J_{b}(\overline{\mathbb{Q}}_{\ell})$, which can give rise to $\phi$ under the twisted embedding $\phantom{}{}^{L}J_{b}(\overline{\mathbb{Q}}_{\ell})\rightarrow\phantom{}^{L}G(\overline{\mathbb{Q}}_{\ell})$ are parameterized by a set of minimal length representatives of $W_{b}=W_{G}/W_{M_{b}}$ via conjugating $\phi^{\prime}$. We expect (See [35, Conjecture A.5]) that the fiber of $\mathrm{LLC}_{b}$ over such a $\phi^{\prime}$ inducing $\phi$ to be the irreducible constituents of the normalized induction of the $L$-packet of $\phi_{T}^{w}$, which is just $\chi^{w}$ by local class field theory for $w\in W_{b}$. For cases (1),(2),and (4), the required compatibility with parabolic induction above was already discussed above for the Harris-Taylor correspondence and the relevant claim for the Kaletha-Minguez-Shin-White/Mok correspondence follows from [47, Section 2.2.3]. For the Gan-Takeda/Gan-Tantono correspondence, it follows directly from the construction and the analogous claim for the Harris-Taylor correspondence. We note that the twists by $\delta_{P_{b}}^{-1/2}$ appear to account for the half Tate twists appearing in the definition of the twisted embedding $\phantom{}{}^{L}J_{b}(\overline{\mathbb{Q}}_{\ell})\rightarrow\phantom{}^{L}G(\overline{\mathbb{Q}}_{\ell})$ (cf. [25, Another proof of Proposition 4.1]). For $\mathrm{GU}_{2}$, we see that when $b^{\prime}$ is unramified, we have isomorphisms of flag varieties
$J_{b^{\prime}}/B_{b^{\prime}}\simeq J_{\tilde{b}}/B_{\tilde{b}}\simeq J_{b}/B_{b}.$
In the above situation where $\phi_{\pi_{b^{\prime}}}$ factors through $\phantom{}{}^{L}T^{\prime}$, we see that since $H,\mathrm{GL}_{2}$ satisfy Assumption 4.5 (3), the corresponding representation of $H(L)$ is of the form $(\tilde{\pi},\chi)$, where $\tilde{\pi}$ is an irreducible constituent of $i_{B_{\tilde{b}}}^{J_{\tilde{b}}}(\chi_{1}^{w})\otimes\delta_{P_{\tilde{b}}}^{-1/2}$, for the associated character $\chi_{1}$ of $\tilde{T}$. In particular, we see that $\pi_{b^{\prime}}$ is a constituent of $i_{B_{b^{\prime}}}^{J_{b^{\prime}}}(\chi_{1}^{w}\otimes\chi)\otimes\delta_{P_{b^{\prime}}}^{-1/2}$, as desired. ∎
We now turn our attention to deriving the key local results from Assumption 4.5.
4.3. Perverse $t$-exactness
We recall that $\mathrm{Bun}_{G}^{b}\simeq[\mathop{\rm Spd}(C)/\mathcal{J}_{b}]$, where $\mathcal{J}_{b}:=\mathrm{Aut}(\mathcal{E}_{b})$ is the group diamond parameterizing automorphisms of the bundle $\mathcal{E}_{b}$ attached to $b\in B(G)$ on $X$, as in (9) though here we view it as a diamond over $\mathop{\rm Spd}(\overline{\mathbb{F}}_{p})$. The diamond $\mathcal{J}_{b}$ has pure cohomological $\ell$-dimension over the base (in the sense of [16, Definition IV.1.17]) equal to $\langle 2\rho_{G},\nu_{b}\rangle$, where $\nu_{b}$ is the slope homomorphism of $b$. Moreover, we have that $\mathrm{Bun}_{G}$ is cohomologically smooth of pure $\ell$-dimension equal to 0 over the base. This motivates the following definition.
Definition 4.14.
We define a perverse $t$-structure $(\phantom{}^{\mathrm{p}}\mathrm{D}^{\leq 0}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell}),\phantom{}^{\mathrm{p}}\mathrm{D}^{\geq 0}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell}))$ on $\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$ such that $A\in\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$ lies in $\phantom{}{}^{\mathrm{p}}\mathrm{D}^{\leq 0}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$ (resp. $\phantom{}{}^{\mathrm{p}}\mathrm{D}^{\geq 0}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$) if and only if $j_{b}^{*}(A)$ (resp. $j_{b}^{!}(A)$) sits in cohomological degrees $\leq\langle 2\rho_{G},\nu_{b}\rangle$ (resp. $\geq\langle 2\rho_{G},\nu_{b}\rangle$) (See [10, Proposition 8.1.5] for a detailed existence proof).
For the rest of this section, we assume that $\ell$ is very decent in the sense of Definition 4.11 so that we may invoke the results of [24] and §4.1.
We let $\phi$ be a semisimple L-parameter, and we write $(\phantom{}^{\mathrm{p}}\mathrm{D}^{\geq 0}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi},\phantom{}^{\mathrm{p}}\mathrm{D}^{\leq 0}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi})$ for the restriction of this $t$-structure to the full subcategory $\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi}$. One can check that this gives rise to a well-defined $t$-structure on $\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi}$ such that the localization map $(-)_{\phi}$ is $t$-exact. Indeed, this follows by applying Proposition A.3 to the two set of compact generators $\{j_{b!}(\mathrm{cInd}_{K}^{J_{b}(\mathbb{Q}_{p})}(\Lambda))\}$ and $\{j_{b\sharp}(\mathrm{cInd}_{K}^{J_{b}(\mathbb{Q}_{p})}(\Lambda))\}$ for $K\subset J_{b}(\mathbb{Q}_{p})$ a varying open pro-$p$ subgroup of $\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$ (See [16, Theorem V.4.1]) and using that localization is an exact functor. Here $j_{b\sharp}$ denotes the exceptional left adjoint to $j_{b}^{*}$ considered in [16, Proposition VII.7.2]. Moreover, one easily checks that this also gives rise to a $t$-structure after further intersecting with $\mathrm{D}^{\mathrm{ULA}}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})\subset\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$ the full subcategory of sheaves $A$ such that $j_{b}^{*}A$ satisfies the property that its invariants under any pro-$p$ group $K\subset J_{b}(\mathbb{Q}_{p})$ is a perfect complex, by an analogous argument where one uses that the category of perfect complexes of $\overline{\mathbb{F}}_{\ell}$-vector spaces is preserved under standard truncation (more generally this holds for coefficients in a regular Noetherian ring). We let $\mathrm{Perv}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi}$ and $\mathrm{Perv}^{\mathrm{ULA}}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi}$ denote the hearts of these $t$-structures.
One of the technical conditions introduced in [24] is the following.
Definition 4.15.
We say that a toral parameter $\phi_{T}:W_{\mathbb{Q}_{p}}\rightarrow\phantom{}^{L}T(\overline{\mathbb{F}}_{\ell})$ is generic regular if it is generic in the sense of Definition 1.1 and $\phi_{T}$ is also regular in the sense that, if $\chi$ denotes the character attached to $\phi_{T}$ under local class field theory, we have that $\chi$ is regular, i.e for all $w\in W_{G}$ non-trivial, we have that
(19) $\chi\not\simeq\chi^{w}.$
We let $\phi:W_{\mathbb{Q}_{p}}\rightarrow\phantom{}^{L}G(\overline{\mathbb{F}}_{\ell})$ be the L-parameter induced by a toral parameter $\phi_{T}$. To motivate the condition, we recall that, if $\phi_{T}$ is generic regular, we have by [24, Theorem 9.10] an object $\mathrm{nEis}(\mathcal{S}_{\phi_{T}})\in\mathrm{Perv}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$, which is a perverse filtered Hecke eigensheaf on $\mathrm{Bun}_{G}$, assuming 4.5 holds. Moreover, it is supported on the set of unramified elements and, for $b\in B(G)_{\mathrm{un}}$, its stalks are given by
$\mathrm{Red}_{b,\phi}^{\mathrm{tw}}:=\bigoplus_{w\in W_{b}}\rho_{b,w}^{\chi}[-\langle 2\rho_{G},\nu_{b}\rangle],$
where we recall that $\rho_{b,w}^{\chi}:=i_{B_{b}}^{J_{b}}(\chi^{w})\otimes\delta_{P_{b}}^{-1/2}$ and $\chi$ is the character attatched to $\phi_{T}$ via local class field theory. In particular, by compatability of the Fargues-Scholze correspondence with local Langlands for torii and parabolic induction [16, Theorem I.9.6 (i), (viii)], it defines an object in the localized category $\mathrm{Perv}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi}$. To show the desired perverse $t$-exactness property, we would like to use the Hecke eigensheaf property of $\mathrm{nEis}(\mathcal{S}_{\phi_{T}})$. Given a geometric dominant cocharacter $\mu$ of $G$, we consider the highest weight tilting module $\mathcal{T}_{\mu}$ attached to $\mu$, we let
$T_{\mu}:\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})\rightarrow\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})^{BW_{E_{\mu}}}$
be the Hecke operator attached to the representation $\mathcal{T}_{\mu}$, where $E_{\mu}$ denotes the reflex field of $\mu$. The sheaf $T_{\mu}(\mathrm{nEis}(\mathcal{S}_{\phi_{T}}))$ carries a filtration which, if it splits, guarantees an isomorphism $\mathrm{nEis}(\mathcal{S}_{\phi_{T}})\boxtimes r_{\mu}\circ\phi\simeq T_{\mu}(\mathrm{nEis}(\mathcal{S}_{\phi_{T}}))$, and we say that $\phi_{T}$ is $\mu$-regular ([24, Definition 9.11]) if such a splitting exists. Here $r_{\mu}:\hat{G}\rightarrow\mathrm{GL}(\mathcal{T}_{\mu})$ is the map defined by the tilting module $\mathcal{T}_{\mu}$. The condition of being $\mu$-regular is guaranteed by the following stronger condition, using [24, Theorem 9.10].
Definition 4.16.
We write $(-)^{\Gamma}:\mathbb{X}_{*}(T_{\overline{\mathbb{Q}}_{p}})\rightarrow\mathbb{X}_{*}(T_{\overline{\mathbb{Q}}_{p}})/\Gamma$ for the natural map from geometric cocharacters to their $\Gamma$-orbits. For a toral parameter $\phi_{T}:W_{\mathbb{Q}_{p}}\rightarrow\phantom{}^{L}T(\overline{\mathbb{F}}_{\ell})$ and a geometric dominant cocharacter $\mu$, we say $\phi_{T}$ is strongly $\mu$-regular if the Galois cohomology complexes
$R\Gamma(W_{\mathbb{Q}_{p}},(\nu-\nu^{\prime})^{\Gamma}\circ\phi_{T})$
are trivial for $\nu$$\nu^{\prime}$ defining distinct $\Gamma$-orbits of weights in the highest weight tilting module $\mathcal{T}_{\mu}$.
Remark 4.17.
In particular, strong $\mu$-regularity implies $\mu$-regularity, and if we know strong $\mu$-regularity then it implies $\mu^{\prime}$-regularity for any $\mathcal{T}_{\mu^{\prime}}$ which occurs as a direct summand of the tensor product $\mathcal{T}_{\mu}^{\otimes n}$, by [27, Proposition 10.12], and this uses the second part of Definition 4.11 (2) in a key way. Also, as we will see, strong $\mu$-regularity is often implied by genericity for some suitably chosen $\mu$.
More importantly, this condition can be used to deduce the following semi-simplicity result for the inductions $\rho_{b,w}$.
Proposition 4.18.
For any $\phi$ induced from a generic regular $\phi_{T}$ we assume, for all $b\in B(G)_{\mathrm{un}}$ and $w\in W_{b}$, that the representations $\rho_{b,w}$ are semi-simple, and that Assumption 4.5 is true. Then we have a direct sum decomposition
$\bigoplus_{b\in B(G)_{\mathrm{un}}}\mathrm{D}^{\mathrm{adm}}(\mathrm{Bun}_{G}^{b},\overline{\mathbb{F}}_{\ell})_{\phi}\simeq\mathrm{D}^{\mathrm{ULA}}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi},$
where $\mathrm{D}^{\mathrm{adm}}(\mathrm{Bun}_{G}^{b},\overline{\mathbb{F}}_{\ell})\subset\mathrm{D}(\mathrm{Bun}_{G}^{b},\overline{\mathbb{F}}_{\ell})\simeq\mathrm{D}(J_{b}(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})$ denotes the subcategory of admissible complexes regarded as a full subcategory of $\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$ by !!-pushforward.
Moreover, for any $A\in\mathrm{D}^{\mathrm{ULA}}(\mathrm{Bun}_{G}^{b},\overline{\mathbb{F}}_{\ell})_{\phi}\simeq\mathrm{D}^{\mathrm{adm}}(J_{b}(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})_{\phi}$, we have that the !! and * pushforwards agree with respect to the inclusion $j_{b}:\mathrm{Bun}_{G}^{b}\rightarrow\mathrm{Bun}_{G}$.
Proof.
The first part of the Proposition follows from the second part. To see this, we use the semi-orthogonal decomposition of $\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$ into $\mathrm{D}(\mathrm{Bun}_{G}^{b},\overline{\mathbb{F}}_{\ell})\simeq\mathrm{D}(J_{b}(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})$ via the excision triangles. Using that the !! and *-pushforwards agree for all objects $A\in\mathrm{D}(\mathrm{Bun}_{G}^{b},\overline{\mathbb{F}}_{\ell})_{\phi}$, we see that the excision spectral sequence degenerates and the first part of the claim follows. To see the second part, we now use Proposition 4.6 to see that an object $A\in\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi}$ can only be supported on the HN-strata $\mathrm{Bun}_{G}^{b}$ for $b\in B(G)_{\mathrm{un}}$, and that the restriction of $A$ to $\mathrm{Bun}_{G}^{b}$ has irreducible constituents valued in subquotients of the representations $\rho_{b,w}$ for $w\in W_{b}$ varying. For the representations $\rho_{b,w}$, we use the following claim of [24].
Proposition 4.19.
[24, Proposition 10.12] For all $b\in B(G)_{\mathrm{un}}$ and $w\in W_{b}$, the natural map
$j_{b!}(\rho_{b,w})\rightarrow Rj_{b*}(\rho_{b,w})$
is an isomorphism assuming $\phi_{T}$ is generic regular and Assumption 4.5 is true.
Therefore, we know the !! and * pushforwards agree on the $\rho_{b,w}$, and, since we are assuming the representations $\rho_{b,w}$ are semisimple, the claim follows for any constituent of $\rho_{b,w}$. This is enough to conclude the claim for any $A\in\mathrm{D}^{\mathrm{ULA}}(\mathrm{Bun}^{b}_{G},\overline{\mathbb{F}}_{\ell})_{\phi}\simeq\mathrm{D}^{\mathrm{adm}}(J_{b}(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})_{\phi}$ using the following claim.
Lemma 4.20.
Assuming 4.5, for $\phi$ a generic parameter and any $A\in\mathrm{D}^{\mathrm{adm}}(J_{b}(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})_{\phi}$ the cohomology of $A$ in the standard $t$-structure has finite length.
Proof.
By Proposition 4.6, we know that any irreducible subquotient of the cohomology of $A$ is an irreducible constituent of $\rho_{b,w}$ for some $w\in W_{b}$. It follows by [66, Section II.5.13] that there are only finitely many possibilities for the irreducible constituents of $\rho_{b,w}$. Therefore, by choosing $K\subset G(\mathbb{Q}_{p})$ a sufficiently small open compact such that all these representations have an invariant vector, we deduce, since $A^{K}$ is a perfect complex by assumption, that $A$ must have finite length cohomology. ∎
We note that the semi-simplicity of $\rho_{1,1}=i_{B}^{G}(\chi)$ is implied by the conditions discussed above.
Lemma 4.21.
Let $\phi_{T}:W_{\mathbb{Q}_{p}}\rightarrow\phantom{}^{L}T(\overline{\mathbb{F}}_{\ell})$ be a generic regular paramater, and set $\chi:T(\mathbb{Q}_{p})\rightarrow\overline{\mathbb{F}}_{\ell}^{*}$ to be the character attached to $\phi_{T}$ under class field theory. Suppose there exists a $\mu$ which is not fixed under any $w\in W_{G}$ and $\phi_{T}$ is $\mu$-regular. Then, $i_{B}^{G}(\chi)$ is irreducible.
Proof.
It follows, just as in the proof of [24, Corollary 10.22] and the assumed $\mu$-regularity, that we have an isomorphism $i_{B}^{G}(\chi)\simeq i_{B}^{G}(\chi^{w})=i_{B^{w}}^{G}(\chi)$ for all $w\in W_{G}$. Here $B^{w}$ is the conjugate of $B$ by $w$. We write $r_{B}^{G}$ for the normalized parabolic restriction functor. We recall that we are working with $\ell$-modular coefficients in possibly non-banal characteristic so $i_{B}^{G}(\chi)$ may have cuspidal constituents. In particular, we will need the following Lemma.
Lemma 4.22.
Let $w_{0}\in W_{G}$ be the element of longest length. For a character $\chi:T(\mathbb{Q}_{p})\rightarrow\overline{\mathbb{F}}_{\ell}^{*}$, if we have an isomorphism $i_{B}^{G}(\chi)\simeq i_{B}^{G}(\chi^{w_{0}})$ of $G(\mathbb{Q}_{p})$-modules then any non-zero quotient $\sigma^{\prime}$ of $i_{B}^{G}(\chi)$ satisfies $r_{B}^{G}(\sigma^{\prime})\neq 0$
Proof.
We apply second adjointness [12, Corollary 1.3] to the non-zero map
$i_{B^{w_{0}}}^{G}(\chi)\xrightarrow{\simeq}i_{B}^{G}(\chi)\rightarrow\sigma^{\prime}$
to conclude the existence of a non-zero map $\chi\rightarrow r_{B}^{G}(\sigma^{\prime})$, which implies the claim. ∎
Now suppose for the sake of contradiction that $i_{B}^{G}(\chi)$ is not irreducible. Then there exists an exact sequence
$0\rightarrow\sigma\rightarrow i_{B}^{G}(\chi)\rightarrow\sigma^{\prime}\rightarrow 0.$
Since parabolic restriction is exact ([66, Section II.2.1]), we get an exact sequence
$0\rightarrow r_{B}^{G}(\sigma)\rightarrow r_{B}^{G}i_{B}^{G}(\chi)\rightarrow r_{B}^{G}(\sigma^{\prime})\rightarrow 0.$
This allows us to conclude an equality of lengths of representations:
$\ell(r_{B}^{G}(\sigma))+\ell(r_{B}^{G}(\sigma^{\prime}))=\ell(r_{B}^{G}(i_{B}^{G}(\chi)))\leq|W_{G}|,$
where the inequality follows from the geometric Lemma [11, Section 2.8] [^11] Note that this bound however fails without taking normalized restriction because of the aforementioned cuspidal constituents of $i_{B}^{G}(\chi)$ in non-banal characteristic (cf. [11, Page 48]).. By the previous Lemma, we conclude that $\ell(r_{B}^{G}(\sigma))<|W_{G}|$. Now, since we know that $\sigma\subset i_{B}^{G}(\chi)\simeq i_{B}^{G}(\chi^{w})$ for all $w\in W_{G}$, Frobenius reciprocity implies that we have non-zero maps $r_{B}^{G}(\sigma)\rightarrow\chi^{w}$ for all $w\in W_{G}$. This gives a contradiction by the regularity of $\chi$, since we have exhibited $|W_{G}|$ quotients of $r_{B}^{G}(\sigma)$ given by $\chi^{w}$ which are all distinct representations, by the regularity of the character $\chi$. ∎
We now have the following key claim.
Theorem 4.23.
Let $\mu$ be a geometric dominant cocharacter. We write
$T_{\mu}:\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})\rightarrow\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})^{BW_{E_{\mu}}}$
for the Hecke operator attached to the highest weight tilting module $\mathcal{T}_{\mu}$ of highest weight $\mu$, where $E_{\mu}$ denotes the reflex field of $E$. Then the operator restricted to $\mathrm{D}^{\mathrm{ULA}}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi}$ is perverse $t$-exact if $\phi_{T}$ is generic regular, Assumption 4.5 is true, the $\rho_{b,w}$ are semi-simple for all $b\in B(G)_{\mathrm{un}}$ and $w\in W_{b}$, and $\phi_{T}$ is $\mu$-regular. Here $\phi:W_{\mathbb{Q}_{p}}\rightarrow\phantom{}^{L}G(\overline{\mathbb{F}}_{\ell})$ is the parameter induced by $\phi_{T}$.
Proof.
Using Lemma 4.20, the commutation of Hecke operators with colimits (which follows since it is a left adjoint (See the discussion preceding [16, Theorem I.7.2])), Proposition 4.6, Proposition 4.18, and semi-simplicity of the representations $\rho_{b,w}$, we can reduce to showing, for all $b\in B(G)_{\mathrm{un}}$, that if we consider the complex
$\mathrm{Red}^{\mathrm{tw}}_{b,\phi}:=\bigoplus_{w\in W_{b}}i_{B_{b}}^{J_{b}}(\chi^{w})\otimes\delta_{P_{b}}^{-1/2}[-\langle 2\rho_{G},\nu_{b}\rangle]\in\mathrm{Perv}^{\mathrm{ULA}}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi}$
then we have a containment
$T_{\mu}(j_{b!}(\mathrm{Red}^{\mathrm{tw}}_{b,\phi}))\in\mathrm{Perv}^{\mathrm{ULA}}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi}$
for the fixed $\mu$. However, $\mathrm{Red}^{\mathrm{tw}}_{b,\phi}$ are the stalks of the perverse filtered Hecke eigensheaf $\mathrm{nEis}(\mathcal{S}_{\phi_{T}})$ given by [24, Theorem 9.10] and, since $\phi_{T}$ is $\mu$-regular then, by definition ([24, Definition 9.11]), we have an isomorphism:
$T_{\mu}(\mathrm{nEis}(\mathcal{S}_{\phi_{T}}))\simeq\mathrm{nEis}(\mathcal{S}_{\phi_{T}})\boxtimes r_{\mu}\circ\phi\in\mathrm{Perv}^{\mathrm{ULA}}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi}^{BW_{E_{\mu}}}.$
This gives the desired claim. ∎
We are almost ready to deduce the result we need for torsion vanishing. To do this, we will first need to discuss when the additional assumptions of regularity and $\mu$-regularity are superfluous, possibly under certain assumptions on $\ell$.
4.4. Verification of additional assumptions
In this section, we will study how the conditions of regularity (Definition 4.15) and strong $\mu$-regularity (Definition 4.16) described in §4.3 are related to the condition of being generic in the particular case of the groups appearing in Theorem 4.13. In particular, we will see in many situations these extra conditions are in fact implied by the generic condition after imposing some additional constraints on $\ell$. In §4.4.1, we will study this for groups of type $A_{n}$, and then in §4.4.2 we will deal with the case of groups of type $C_{n}$. We then finally conclude with our main local results in §4.4.3 (Corollaries 4.27 and 4.29).
Before proceeding, we describe the following general Lemma which will allow us to base change to splitting fields.
Lemma 4.24.
Let $G$ be a quasi-split connected reductive group with splitting field $F$. Suppose that $\ell\nmid[F:\mathbb{Q}]$ (e.g if $\ell$ is very decent with respect to $G$). Then $\phi_{T}$ is generic if and only if $R\Gamma(W_{F},\tilde{\alpha}\circ\phi_{T}|_{W_{F}})$ is trivial for all absolute coroots $\tilde{\alpha}\in\mathbb{X}_{*}(T_{\overline{\mathbb{Q}}_{p}})$.
Proof.
We recall that, given a $\Gamma$-orbit of absolute coroots $\alpha\in\mathbb{X}_{*}(T_{\overline{\mathbb{Q}}_{p}})^{+}/\Gamma$, if $F_{\alpha}$ denotes the reflex field of $\alpha$ then the representation of $\phantom{}{}^{L}T$ defined by $\alpha$ is given by choosing a representative $\tilde{\alpha}\in\mathbb{X}_{*}(T_{\overline{\mathbb{Q}}_{p}})^{+}$ of $\alpha$, and inducing the representation of $\hat{T}\rtimes W_{F_{\alpha}}/W_{F}$ defined by it to $W_{\mathbb{Q}_{p}}/W_{F}$. By Schapiro’s Lemma, we deduce that genericity is equivalent to $R\Gamma(W_{F_{\alpha}},\tilde{\alpha}\circ\phi_{T}|_{W_{F_{\alpha}}})\simeq 0$. Now, for all $i\in\mathbb{N}_{\geq 0}$, we recall that we have the maps defined by restriction and corestriction
$H^{i}(W_{F_{\alpha}},\tilde{\alpha}\circ\phi_{T}|_{W_{F_{\alpha}}})\xrightarrow{\mathrm{res}}H^{i}(W_{F},\tilde{\alpha}\circ\phi_{T}|_{W_{F}})\xrightarrow{\mathrm{cores}}H^{i}(W_{F_{\alpha}},\tilde{\alpha}\circ\phi_{T}|_{W_{F_{\alpha}}}),$
and that the composite is given by multiplication by $F:F_{\alpha}]$ (See [58, Chapter VIII,Proposition 4]). Therefore, since $\ell\nmid[F:F_{\alpha}]$ by our assumption, it follows that the vanishing of $R\Gamma(W_{F_{\alpha}},\tilde{\alpha}\circ\phi_{T}|_{W_{F_{\alpha}}})$ is equivalent to the vanishing of $R\Gamma(W_{F},\tilde{\alpha}\circ\phi_{T}|_{W_{F}})$, since multiplication by $F:F_{\alpha}]$ is invertible on $\overline{\mathbb{F}}_{\ell}$-vector spaces. This shows the desired claim. ∎
We will also want to consider a few additional groups other than those listed in Theorem 4.13. There are a few more groups of interest to us. We will define them now.
Let $L/\mathbb{Q}_{p}$ be a finite extension. We have the similitude maps from $\mathrm{GL}_{n}$ (resp. $\mathrm{GSp}_{4}$)
$\nu:\mathrm{Res}_{L/\mathbb{Q}_{p}}\mathrm{GL}_{n}\rightarrow\mathrm{Res}_{L/\mathbb{Q}_{p}}\mathbb{G}_{m}$
(resp.
$\nu:\mathrm{Res}_{L/\mathbb{Q}_{p}}\mathrm{GSp}_{4}\rightarrow\mathrm{Res}_{L/\mathbb{Q}_{p}}\mathbb{G}_{m}).$
We thus define
$G(\mathrm{SL}_{n,L}):=\mathrm{Res}_{L/\mathbb{Q}_{p}}\mathrm{GL}_{n}\times_{\nu,\mathrm{Res}_{L/\mathbb{Q}_{p}}\mathbb{G}_{m}}\mathbb{G}_{m},$
$G(\mathrm{Sp}_{4,L}):=\mathrm{Res}_{L/\mathbb{Q}_{p}}\mathrm{GSp}_{4}\times_{\nu,\mathrm{Res}_{L/\mathbb{Q}_{p}}\mathbb{G}_{m}}\mathbb{G}_{m}.$
Now we will study the additional assumptions for the groups of type $A$.
4.4.1. Groups of type A
The first main result is of the following form.
Lemma 4.25.
Let $L/\mathbb{Q}_{p}$ be a finite extension and $G$ be one of the following groups:
  1. (1)
    $\mathrm{Res}_{L/\mathbb{Q}_{p}}\mathrm{U}_{n}$,
  2. (2)
    $\mathrm{Res}_{L/\mathbb{Q}_{p}}\mathrm{GU}_{n}$,
  3. (3)
    $\mathrm{Res}_{L/\mathbb{Q}_{p}}\mathrm{GL}_{n}$,
  4. (4)
    $G(\mathrm{SL}_{2,L})$.
Suppose that $\ell\nmid 2[L:\mathbb{Q}_{p}]$ in cases (1)-(2) and that $\ell\nmid[L:\mathbb{Q}_{p}]$ in cases (3)-(4).
If $\phi_{T}$ is a generic toral parameter for $G$ then $\phi_{T}$ is also regular. Moreover, for (1)(3)(1)-(3), $\phi_{T}$ will be $\mu$-regular for all $\mu$, while, for $(4)$, $\phi_{T}$ will be $\mu$-regular for $\mu$ which are of the form $\prod_{\tau:L\hookrightarrow\overline{\mathbb{Q}}_{p}}\mu^{\prime}$ for $\mu^{\prime}$ a cocharacter of $\mathrm{GL}_{2}$; in particular, it will be $\mu$-regular for all minuscule $\mu$.
Proof.
We first establish the implication that $\phi_{T}$ generic implies $\phi_{T}$ is regular in cases (1)-(3). We may assume for simplicity that $L=\mathbb{Q}_{p}$ with the proof in general essentially being the same in light of Lemma 4.24 and our assumption on $\ell$. If $G=\mathrm{GL}_{n}$ then we write $\chi(t_{1},\ldots,t_{n})$ as $\prod_{i=1}^{n}\chi_{i}(t_{i})$, where $t_{i}$ are the coordinates of $T(\mathbb{Q}_{p})\simeq(\mathbb{Q}_{p}^{*})^{n}$. It then easily follows from evaluating the relationship $\chi\simeq\chi^{w}$ for some non-trivial $w$ at any coordinate of $T(\mathbb{Q}_{p})$ not fixed by $w$ and setting the remaining coordinates equal to $1$ to derive the relation $\chi_{i}\chi_{j}^{-1}\simeq\mathbf{1}$ for some $i\neq j$, which contradicts genericity. We now consider the case of $G=\mathrm{U}_{n}$ defined with respect to a quadratic extension $E/\mathbb{Q}_{p}$. Suppose there exists a non-trivial $w\in W_{G}$ such that we have an isomorphism:
$\chi\simeq\chi^{w}$
of characters on $T(\mathbb{Q}_{p})$. We recall that $G_{E}\simeq\mathrm{GL}_{n,E}$ where $E/\mathbb{Q}_{p}$ denotes the quadratic extension defining the unitary group. We then precompose the isomorphism $\chi\simeq\chi^{w}$ with the norm map $T(E)\rightarrow T(\mathbb{Q}_{p})$ to obtain an analogous relationship of characters on the torus $T(E)$, which is the maximal torus of $\mathrm{GL}_{n,E}$. Then Lemma 4.24 and our assumpition on $\ell$ reduces us to the $\mathrm{GL}_{n}$ case discussed above.
The case of $\mathrm{GU}_{n}$ similarly reduces to the $\mathrm{U}_{n}$ case by setting the coordinate on $T(\mathbb{Q}_{p})$ corresponding to the similitude factor to be equal to $1$.
We now turn to case (4), let $d=[L:\mathbb{Q}_{p}]$. Observe that we have an isomorphism $G(\mathrm{SL}_{2,L})_{L}\simeq H_{L}$, where
$H=\left\{(g_{i})\in\prod_{i=1}^{d}\mathrm{GL}_{2}:\mathrm{det}(g_{i})=\mathrm{det}(g_{j})\penalty 10000\ \forall i,j\right\}.$
where $d$ is the size of the set of embeddings $L\hookrightarrow\overline{\mathbb{Q}}_{p}$. Applying Lemma 4.24 again and arguing as for unitary groups, it suffices to work with $H_{L}$. The maximal torus $T^{\prime}$ in $H$ can be identified with
$\mathbb{G}_{m}^{d}\times\mathbb{G}_{m},$
via the map $(t_{1},\dots,t_{d},t)\mapsto(\mathrm{diag}(t_{i},tt_{i}^{-1}))$. We have $W_{H}=\prod_{i=1}^{d}\mathbb{Z}/2\mathbb{Z}$, where the non-trivial element act on each of the factors by sending $t_{i}\mapsto t_{i}^{-1}t$. Write $\chi(t_{1},\ldots,t_{d})=\chi_{1}(t_{1})\otimes\cdots\otimes\chi_{d}(t_{d})\otimes\nu(t)$ for characters $\chi_{1},\dots,\chi_{d},\nu$. If we have a permutation $w\in W_{H}$ which is non-trivial on the $i$th factor then setting $t_{i}=x$ and $t=x$ and setting the remaining coordinates $t_{j}$ for $i\neq j$ equal to $1$ we obtain the relationship
$\chi_{i}(t)\simeq\mathbf{1},$
which contradicts genericity.
We now show $\mu$-regularity. As before, for cases (1), (2), and (3), observe that if $G$ is of the form $\mathrm{Res}_{L/\mathbb{Q}_{p}}G^{\prime}$ and $T^{\prime}$ denotes the maximal torus of $G^{\prime}$ then we have an isomorphism
$\mathbb{X}_{*}(T_{\overline{\mathbb{Q}}_{p}})\simeq\prod_{\phi\in\mathrm{Hom}_{\mathbb{Q}_{p}}(L,\overline{L})}\mathbb{X}_{*}(T^{\prime}_{\overline{L}}),$
where $\overline{L}$ is an algebraic closure of $L$. Using this, we can without loss of generality assume that $L=\mathbb{Q}_{p}$. In the case that $G=\mathrm{GL}_{n},\mathrm{U}_{n}$, or $\mathrm{GU}_{n}$, this follows as in the proof of [24, Corollary 9.16]. We recall briefly how this goes.
One can consider the geometric dominant cocharacter $\mu=(1,0,\ldots,0,0)$ of $\mathrm{GL}_{n}$. This defines the standard representation $V_{\mathrm{std}}$ of $\hat{G}\simeq\mathrm{GL}_{n}$. This cocharacter is in particular minuscule so the weights form a closed Weyl group orbit with representative $(1,0,\ldots,0,0)$. From here, it easily follows that the difference of the weights appearing in $V_{\mathrm{std}}$ define coroots of $G$. In particular, it follows that, if $\phi_{T}$ is generic then it is strongly $\mu$-regular for $\mu=(1,0,\ldots,0)$ in the sense of Definition 4.16, and this implies the filtration on $T_{\mu}(\mathrm{nEis}(\mathcal{S}_{\phi_{T}}))$ splits by [24, Theorem 9.10] for this $\mu$. Now, the tilting modules $\mathcal{T}_{\omega_{i}}=\Lambda^{i}(V_{\mathrm{std}})$ attached to the other fundamental coweights $\omega_{i}=(1^{i},0^{n-i})$ of $G$ can be realized as direct summands of $V_{\mathrm{std}}^{\otimes i}$ (See [24, Proposition 9.1,Theorem 9.2, and Proposition 9.3] for a review of how this works), and it follows that $\phi_{T}$ is $\mu$-regular for $\mu=\omega_{i}$ by [24, Corollary 9.12]. Since any dominant cocharacter can be written as a linear combination of fundamental coweights, the claim for any $\mu$ now follows from [24, Corollary 9.13]. The case of $\mathrm{GU}_{n}$ and $\mathrm{U}_{n}$ follows in a very similar way, using Lemma 4.24 and our assumption on $\ell$.
For case (4), observe that as before, we can base change to $L$, and since $H$ is a subgroup of $\prod{\mathrm{GL}_{2}}$, all cocharacters $\mu$ of $H$ define products of cocharacters for $\mathrm{GL}_{2}$. Now, consider a cocharacter of the form $\mu=\prod_{\tau}\mu_{\tau}$, where $\mu_{\tau^{\prime}}=\mu_{\tau}$ for all $\tau,\tau^{\prime}$ and $\mu_{\tau}$ is a cocharacter of $\mathrm{GL}_{2}$. Note that every dominant minuscule cocharacter of $H$ will be of this form. This is because a cocharacter $\mu=\prod_{\tau}\mu_{\tau}$ of $\prod\mathrm{GL}_{2}$ factors through $H$ exactly when the composition with the determinant is equal for all $\tau$. Moreover, we see that for $\mu$ to be minuscule, $\mu_{\tau}$ must be one of the fundamental coweights $\omega_{i}$, and if some $\mu_{\tau}\neq\mu_{\tau^{\prime}}$, then after composing with the determinant we will get different characters. Now, we observe that the same argument as above holds to show that the difference of Weyl conjugates define a coroot of $H$ for the cocharacter $\mu_{1}=\prod(1,0)$, while for all other cocharacters of the form $\mu=\prod\mu^{\prime}$, where $\mu^{\prime}$ is a fundamental weight of $\mathrm{GL}_{2}$, they appear as weights in some tensor power of the highest weight representation corresponding to $\mu_{1}$. The claim for any $\mu=\prod\mu^{\prime}$, where $\mu^{\prime}$ is a dominant cocharacter of $\mathrm{GL}_{2}$, follows by the same argument as above, using [24, Corollary 9.13]. ∎
We now turn our attention to the groups of type $C$.
4.4.2. Groups of type C
The main lemma is the following.
Lemma 4.26.
Let $L/\mathbb{Q}_{p}$ be a finite extension, and $G$ be one of the following groups:
  1. (1)
    $\mathrm{Res}_{L/\mathbb{Q}_{p}}\mathrm{GSp}_{4}$
  2. (2)
    $G(\mathrm{Sp}_{4,L})$.
If $\phi_{T}$ is a generic toral parameter for $G$ then $\phi_{T}$ is regular. Moreover, in case $(1)$, the parameter $\phi_{T}$ will be $\mu$-regular for all $\mu$, and, for $(2)$, $\phi_{T}$ will be $\mu$-regular for $\mu$ which are of the form $\prod_{\tau:L\hookrightarrow\overline{\mathbb{Q}}_{p}}\mu^{\prime}$ for $\mu^{\prime}$ a cocharacter of $\mathrm{GSp}_{4}$; in particular, it will be $\mu$-regular for all minuscule $\mu$.
Proof.
We will first establish regularity, and suppress giving the proof that $\phi_{T}$ is regular as it is strictly easier. Again, for (1), we assume that $L=\mathbb{Q}_{p}$ for this part with the proof in general being essentially the same in light of Lemma 4.24. We will show this by contradiction. Suppose on the contrary that there exists some $w\in W_{G}$ such that we have an isomorphism
(20) $\chi\simeq\chi^{w}.$
For case (1), consider the following parametrization of the maximal torus $T$
(21) $a:(\mathbb{Q}_{p}^{*})^{2}\times\mathbb{Q}_{p}^{*}\rightarrow T(\mathbb{Q}_{p})$
$(t_{1},t_{2},t)\mapsto\begin{pmatrix}t_{1}&0&0&0\\ 0&t_{2}&0&0\\ 0&0&tt_{2}^{-1}&0\\ 0&0&0&tt_{1}^{-1}\end{pmatrix}$,
as in [63, Page 135]. This allows us to write the character $\chi:T(\mathbb{Q}_{p})\rightarrow\overline{\mathbb{F}}_{\ell}^{*}$ as $\chi_{1}(t_{1})\chi_{2}(t_{2})\nu(t)$, for characters $\mathbb{Q}_{p}^{*}\rightarrow\overline{\mathbb{F}}_{\ell}^{*}$. We now check that (20) cannot hold for all seven non-trivial elements of the Weyl group.
Consider the Weyl group element corresponding to the translation:
$w_{1}:a(t_{1},t_{2},t)\mapsto a(t_{2},t_{1},t)$
If we consider equation (20) with respect to this element and evaluate on $(x,1,x)=(t_{1},t_{2},t)$ then we obtain the equation
$\chi_{1}(x)\simeq\chi_{2}(x)$
which gives an isomorphism $\chi_{1}\chi_{2}^{-1}(x)\simeq\mathbf{1}$ contradicting genericity.
Similarly, if we consider the simple Weyl group element
$w_{2}:a(t_{1},t_{2},t)\mapsto a(t_{1},t_{2}^{-1}t,t)$
then evaluating equation (20) becomes
$\chi_{1}(t_{1})\chi_{2}(t_{2})\nu(t)\simeq\chi_{1}(t_{1})\chi_{2}(t_{2}^{-1}t)\nu(t)$
cancelling terms we obtain that
$\chi_{2}(t)^{-1}\chi_{2}(t_{2})^{2}\simeq\mathbf{1}$
and setting $t=1$ this gives
$\chi_{2}^{-1}(t)\simeq\mathbf{1}$
which contradicts genericity (See [63, Page 167] for the enumeration of $1$-parameter subgroups attached to the coroots in the parametrization (21).
Consider now the Weyl group element
$w_{3}:a(t_{1},t_{2},t)\mapsto a(t_{2}^{-1}t,t_{1},t)$
if we evaluate equation (20) then we obtain
$\chi_{1}(t_{1})\chi_{2}(t_{2})\nu(t)\simeq\chi_{1}(t_{2})^{-1}\chi_{1}(t)\chi_{2}(t_{1})\nu(t)$
rearranging and canceling terms we obtain
$\chi_{1}\chi_{2}^{-1}(t_{1})\chi_{2}\chi_{1}(t_{2})\chi_{1}(t)^{-1}\simeq\mathbf{1}$
so if we evaluate at $(t_{1},t_{2},t)=(1,1,x)$ we obtain that
$\chi_{1}^{-1}(x)\simeq\mathbf{1}$
which contradicts genericity (See [63, Page 167] for the enumeration of $1$-parameter subgroups attached to the coroots in the parametrization (21).
Consider the reflection
$w_{4}:a(t_{1},t_{2},t)\mapsto a(t_{1}^{-1}t,t_{2},t)$
then equation (20) becomes
$\chi_{1}(t_{1})\chi_{2}(t_{2})\nu(t)\simeq\chi_{1}(t_{1}^{-1}t)\chi_{2}(t_{2})\nu(t)$
which gives
$\chi_{1}(t_{1}^{2})\chi_{1}(t)^{-1}\simeq\mathbf{1}$
so if we evaluate at $(t_{1},t_{2},t)=(1,1,x)$, this becomes
$\chi_{1}(x)^{-1}\simeq\mathbf{1}$
which contradicts genericity.
Now consider the Weyl group element
$w_{5}:a(t_{1},t_{2},t)\mapsto a(t_{2},t_{1}^{-1}t,t)$
then equation (20)
$\chi_{1}(t_{1})\chi_{2}(t_{2})\nu(t)\simeq\chi_{1}(t_{2})\chi_{2}(t_{1}^{-1}t)\nu(t)$
which simplifies to
$\chi_{2}\chi_{1}^{-1}(t_{2})\chi_{1}\chi_{2}(t_{1})\chi_{2}(t)^{-1}\simeq\mathbf{1}$
so if we evaluate at say $(t_{1},t_{2},t)=(x,1,x)$ then this gives
$\chi_{1}(x)\simeq\mathbf{1}$
which contradicts genericity.
Now consider the Weyl group element
$w_{6}:a(t_{1},t_{2},t)\mapsto a(t_{1}^{-1}t,t_{2}^{-1}t,t)$
then equation (20) becomes
$\chi_{1}(t_{1})\chi_{2}(t_{2})\nu(t)\simeq\chi_{1}(t_{1}^{-1}t)\chi_{2}(t_{2}^{-1}t)\nu(t)$
which simplifies to
$\chi_{1}^{2}(t_{1})\chi_{2}^{2}(t_{2})\chi_{1}\chi_{2}(t)^{-1}\simeq\mathbf{1}$
so if we evaluate at $(t_{1},t_{2},t)=(1,1,x)$ then this becomes
$\chi_{1}\chi_{2}(x)\simeq\mathbf{1}$
which contradicts genericity.
Now finally we consider
$w_{7}:a(t_{1},t_{2},t)\mapsto a(t_{2}^{-1}t,t_{1}^{-1}t,t)$
then equation (20) becomes
$\chi_{1}(t_{1})\chi_{2}(t_{2})\nu(t)\simeq\chi_{1}(t_{2}^{-1}t)\chi_{2}(t_{1}^{-1}t)\nu(t)$
which simplifies to
$\chi_{1}\chi_{2}(t_{1})\chi_{1}\chi_{2}(t_{2})\chi_{1}\chi_{2}(t^{-1})\simeq\mathbf{1}$
evaluated at $(t_{1},t_{2},t)=(1,1,x)$ simplifies to
$\chi_{1}\chi_{2}(x)\simeq\mathbf{1}$
which contradicts genericity.
This concludes our discussion of generic regularity for $\mathrm{Res}_{L/\mathbb{Q}_{p}}\mathrm{GSp}_{4}$.
We now turn to the case of $G(\mathrm{Sp}_{4,L})$. As in the proof of the previous Lemma, observe that if we let
(22) $H=\{(g_{i})\in\prod_{L\hookrightarrow\overline{\mathbb{Q}}_{p}}\mathrm{GSp}_{4}\text{ such that }\nu(g_{i})=\nu(g_{j}),\forall i,j\},$
then we have $H_{L}\simeq G(\mathrm{Sp}_{4,L})_{L}$. Thus, we may reduce to the case of $H$. Since $H\subset\prod\mathrm{GSp}_{4}$, we may also use the parametrization in [63, 135] to see that the maximal torus $T^{\prime}$ is given by a parametrization
$(\mathbb{Q}_{p}^{*})^{2d}\times\mathbb{Q}_{p}^{*}\rightarrow T^{\prime}(\mathbb{Q}_{p}),$
$((t_{\tau 1},t_{\tau 2})_{\tau:L\hookrightarrow\overline{\mathbb{Q}}_{p}},t)\mapsto\begin{pmatrix}t_{\tau 1}&0&0&0\\ 0&t_{\tau 2}&0&0\\ 0&0&tt_{\tau 2}^{-1}&0\\ 0&0&0&tt_{\tau 1}^{-1}\end{pmatrix}_{\tau:L\hookrightarrow\overline{\mathbb{Q}}_{p}}$
where we note that the common similitude factor is the last coordinate $t$.
Thus, if we wanted to argue by contradiction, using the notation of the proof above, when $w_{\tau}=w_{i}$ for $i=1,\dots,7$, we should substitute for $t_{\tau 1},t_{\tau 2},t$ the values we considered above, subject to the additional constraint that we must have $t$, the similitude factor, being equal for all $\tau$. However, in all of the above cases $t=x$, and we choose different values for $t_{1},t_{2}$ depending on the particular $w_{i}$. In particular, as in the previous proof, to derive a contradiction we just need to evaluate the relationship
$\chi\simeq\chi^{w}$
on $t=x$ and set $t_{\tau 1}=1$ or $x$ and $t_{\tau 2}$ or $x$ depending on $w_{\tau}$, as in the proof described above. Since the roots of $H$ are expressible as products of roots of $\mathrm{GSp}_{4}$, this will allow us to derive a contradiction to genericity by the same argument as above.
We now show $\mu$-regularity. As in the proof of the previous Lemma, for case (1) it suffices to check the claim when $G=\mathrm{GSp}_{4}$, the Langlands dual group is given by $\mathrm{GSpin}_{5}$, which is isomorphic to $\mathrm{GSp}_{4}$. The spin representation
$\mathrm{spin}:\mathrm{GSpin}_{5}\rightarrow\mathrm{GL}_{4}(V_{\mathrm{spin}})$
defines a minuscule highest weight representation, which, under the isomorphism $\mathrm{GSpin}_{5}\simeq\mathrm{GSp}_{4}$, identifies with the defining representation of $\mathrm{GSp}_{4}$. From here it is easy to see that the differences of the weights are roots of $\mathrm{GSp}_{4}$ (= coroots of $\mathrm{GSpin}_{5}$). For example, by using the parametrization of the maximal torus, as in (21), and the description of the roots in this parametrization provided on [63, Page 167]. Therefore, genericity guarantees strong $\mu$-regularity for this representation which implies $\mu$-regularity as before. The other fundamental tilting module of $\mathrm{GSpin}_{5}$ is given by the defining representation $\mathrm{GSpin}_{5}\rightarrow\mathrm{SO}_{5}\rightarrow\mathrm{GL}_{5}$ or its maximal irreducible submodule. Moreover, this occurs as a $5$-dimensional summand of $V_{\mathrm{spin}}\otimes V_{\mathrm{spin}}$, and it follows by [24, Proposition 9.12] that we know $\mu$-regularity for this representation as well. Therefore, since we know $\mu$-regularity for the fundamental coweights, we are now done by [24, Corollary 9.13] as before.
Now, for case (2), note that, as in the previous Lemma, all cocharacters of $H$ are of the form $\mu=\prod_{\tau}\mu_{\tau}$ for some cocharacters $\mu_{\tau}$ of $\mathrm{GSp}_{4}$. The argument given above for $\mathrm{GSp}_{4}$ shows that if we let $\mu^{\prime}$ be one of the fundamental coweights of $\mathrm{GSp}_{4}$, then if we take $\mu=\prod\mu^{\prime}$, (i.e. $\mu_{\tau}=\mu^{\prime}$ for all $\tau$) then we are $\mu$-regular for such $\mu$. Applying [24, Corollary 9.13] again shows that if $\mu=\prod_{\tau}\mu^{\prime}$, where $\mu^{\prime}$ is a dominant cocharacter of $\mathrm{GSp}_{4}$, then $\phi_{T}$ will be $\mu$-regular for such $\mu$. Note that, as in the case of $G(\mathrm{SL}_{2})$, every dominant minuscule cocharacter of $H$ is of this form. ∎
With all the extra conditions now properly understood in the relevant cases, we now turn our attention to formulating and proving the main local results.
4.4.3. The Local Results
We consider the following table, summarizing the groups and primes for which our results apply. We have left the entry blank if no constraint is imposed, and just mentioned the groups that appear as local constituents of global groups that admit a Shimura datum and for which $G$ is unramified.
$G$ Constraint on $G$ $\ell$ $p$
$\mathrm{Res}_{L/\mathbb{Q}_{p}}(\mathrm{GL}_{n})$ $L/\mathbb{Q}_{p}$ unramified $(\ell,[L:\mathbb{Q}_{p}])=1$
$\mathrm{Res}_{L/\mathbb{Q}_{p}}(\mathrm{GSp}_{4})$ $L=\mathbb{Q}_{p}$
$L/\mathbb{Q}_{p}$ unramified $(\ell,[L:\mathbb{Q}_{p}])=1$ $p\neq 2$
$\mathrm{Res}_{L/\mathbb{Q}_{p}}(\mathrm{GU}_{2})$ $L/\mathbb{Q}_{p}$ unramified, $(\ell,2[L:\mathbb{Q}_{p}])=1$
$G=\mathrm{U}_{n}(L/\mathbb{Q}_{p})$ $n$ odd $L$ unramified $\ell\neq 2$
$G=\mathrm{GU}_{n}(L/\mathbb{Q}_{p})$ $n$ odd $L$ unramified $\ell\neq 2$
$G(\mathrm{SL}_{2,L})$ $L/\mathbb{Q}_{p}$ unramified $(\ell,[L:\mathbb{Q}_{p}])=1$
$G(\mathrm{Sp}_{4,L})$ $L/\mathbb{Q}_{p}$ unramified, $L\neq\mathbb{Q}_{p}$ $(\ell,[L:\mathbb{Q}_{p}])=1$ $p\neq 2$
We now apply Theorem 4.23.
Corollary 4.27.
Assume $G$ is a product of the groups appearing in Table (23) with $p$ and $\ell$ satisfying the corresponding conditions. Let $\phi$ be a semisimple $L$-parameter which is induced from a generic toral parameter $\phi_{T}$. Let $\mu$ be a geometric dominant cocharacter of $G$ which satisfies the constraint that it is of the form $\prod_{\tau:L\hookrightarrow\overline{\mathbb{Q}}_{p}}\mu^{\prime}$, for $\mu^{\prime}$ a cocharacter of $\mathrm{GL}_{2}$ (resp. $\mathrm{GSp}_{4}$) on the direct factors of $G$ equal to $G(\mathrm{SL}_{2,L})$ or $G(\mathrm{Sp}_{4,L})$, then we have that the natural functor
$T_{\mu}:\mathrm{D}^{\mathrm{ULA}}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi}\rightarrow\mathrm{D}^{\mathrm{ULA}}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi}$
is perverse t-exact. In particular, it is perverse $t$-exact for all minuscule $\mu$, and if we look at the natural functor
$j_{1}^{*}T_{\mu}:\mathrm{D}^{\mathrm{ULA}}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi}\rightarrow\mathrm{D}^{\mathrm{adm}}(G(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})_{\phi}$
then this is exact for the perverse $t$-structure on the source and the perverse (= standard $t$-structure) on the target.
Proof.
First note that, using the decomposition $\mathrm{Bun}_{G_{1}\times G_{2}}:=\mathrm{Bun}_{G_{1}}\times\mathrm{Bun}_{G_{2}}$, we can assume that $G$ is isomorphic to one of the groups appearing in Table (23).
Observe that all the groups in Table (23) satisfy Assumption 4.5, where the first five rows follows from Theorem 4.13, and the last two from Proposition 4.9. We also note that the $\ell$ very decent assumption is also guaranteed by our assumptions on $\ell$ as in Theorem 4.13, so we may use the results of §4.3. We now apply Theorem 4.23. To do this, we also need to check that if $\phi_{T}$ is a generic toral parameter then it is also generic regular, and $\mu$-regular for all $\mu$ if $G\neq\mathrm{Res}_{L/\mathbb{Q}_{p}}H$ and $H=\mathrm{GL}_{n},\mathrm{GU}_{n}$, or $\mathrm{U}_{n}$. If $G=G(\mathrm{SL}_{2,L})$ (resp. $G(\mathrm{Sp}_{4,L})$), and we need to check if $\phi_{T}$ is generic then it is generic regular and $\mu$-regular for all $\mu$ of the form $\prod_{\tau:L\hookrightarrow\overline{\mathbb{Q}}_{p}}\mu^{\prime}$, for $\mu^{\prime}$ a cocharacter of $\mathrm{GL}_{2}$ (resp. $\mathrm{GSp}_{4}$) it is $\mu$-regular, where these give rise to all minuscule cocharacters of $G$ in this case. Finally, if $G=\mathrm{Res}_{L/\mathbb{Q}_{p}}\mathrm{GSp}_{4}$ we need to check if $\phi$ is generic then it is regular and $\mu$-regular for all $\mu$. These claims follow from Lemma 4.25 and Lemma 4.26.
Lastly, we need to check that the representations $\rho_{b,w}:=i_{B_{b}}^{J_{b}}(\chi^{w})\otimes\delta_{P_{b}}^{-1/2}$ are semi-simple for all $b\in B(G)_{\mathrm{un}}$ and $w\in W_{b}$. We claim that they are in fact irreducible. Recall that $J_{b}\simeq M_{b}\subset G$, where $M_{b}$ is a Levi of $G$. Moreover, we note that any such Levi $M_{b}$ is a product of groups also appearing in (23). Therefore, the desired irreducibility follows from the $\mu$-regularity and regularity of $\phi_{T}$ discussed above combined with Lemma 4.21. Note that in the case of $G(\mathrm{SL}_{2,L})$ and $G(\mathrm{Sp}_{4,L})$, we can always find a cocharacter $\mu$ of the form $\prod_{\tau}\mu^{\prime}$ which is not fixed by the Weyl group, since we can simply look at any cocharacter of $\mu^{\prime}$ of $\mathrm{GL}_{2}$ (resp. $\mathrm{GSp}_{4}$) which is not fixed by the Weyl group, and take $\mu$ to be the product of these $\mu^{\prime}$ in order to see that the assumptions of Lemma 4.21 are satisfied. ∎
Remark 4.28.
We note that if the local group $G$ is split then the set $B(G,\mu)_{\mathrm{un}}$ is a singleton for any minuscule cocharacter $\mu$, by invoking [24, Corollary 2.9]. We denote the unique element in this set as $b_{\mu}$. In this case, for a semisimple $L$-parameter $\phi$ induced from a generic $\phi_{T}$, by using Proposition 4.6 and that assumption 4.5 holds, we see that the adjunction map $j_{b_{\mu}!}j_{b_{\mu}}^{!}\rightarrow\mathrm{id}$ induces a natural isomorphism
$j_{1}^{*}T_{\mu}j_{b_{\mu}!}j_{b_{\mu}}^{*}\simeq j_{1}^{*}T_{\mu}$
of operators on the full subcategory $\mathrm{D}(\mathrm{Bun}_{G})_{\phi}\subset\mathrm{D}(\mathrm{Bun}_{G})$ (Here we have implicitly used the isomorphism $j_{1}^{*}T_{\mu}\simeq j_{1}^{*}T_{\mu}j_{\leq\mu!}j_{\leq\mu}^{*}$, by [53, Proposition A.9] , to rewrite $j_{b_{\mu}}^{!}$ as $j_{b_{\mu}}^{*}$ using Proposition 4.6, where $j_{\leq\mu}:\mathrm{Bun}_{G,\leq\mu}\hookrightarrow\mathrm{Bun}_{G}$ is the open immersion of the substack corresponding to $B(G,\mu)\subset B(G)$). We have that the $\sigma$-centralizer of $J_{b_{\mu}}$ identifies with a proper Levi of $M_{b_{\mu}}\subset G$ and the LHS of this isomorphism can be explicitly computed (up to a shift by the dimension of the Shimura variety) in terms of the parabolic induction from $J_{b_{\mu}}\simeq M_{b_{\mu}}$ to $G$, using the results of [18]. In this case, the second part of Corollary 4.27 is a simple consequence of the exactness of the parabolic induction functor (even without restricting to the ULA subcategory). In particular, for any of cases appearing in Table (23) where the group $G$ is split, Corollary 4.27 also holds immediately upon knowing the that an object in $\mathrm{D}(\mathrm{Bun}_{G})_{\phi}$ is supported on just the unramified elements for $\phi$ induced from a generic toral parameter $\phi_{T}$, which follows from Assumption 4.5 and Proposition 4.6. As described in [37, 54] in the case when $G$ is a product of $\mathrm{GL}_{n}$s, by an observation of Koshikawa one can actually weaken the condition on $\phi$ to just assuming that $H^{2}(W_{\mathbb{Q}_{p}},\alpha\circ\phi_{T})\simeq 0$ (cf. Remark 6.8) to deduce that $\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi}$ is supported on the unramified elements from assumption 4.5. The key point here is that non-trivial classes in the $H^{0}(W_{\mathbb{Q}_{p}},\alpha\circ\phi_{T})$ will only give rise to non Frobenius semi-simple deformations of $\phi_{T}$ as a $\phantom{}{}^{L}T(\overline{\mathbb{F}}_{\ell})$-valued parameter to a $\phantom{}{}^{L}B(\overline{\mathbb{F}}_{\ell})$-valued parameter, which cannot occur as mod $\ell$-reductions of the semi-simplification of parameters appearing in the classical Local Langlands correspondence (cf. Remark 6.5). By the reasoning described above, this will in turn imply Conjecture 1.2 under this weaker genericity condition in the case when $G$ is split (See [10, Proposition 10.2.4,10.2.5] for details). However, under this weaker hypothesis, the splitting of the semi-orthogonal decomposition (See Corollary 4.29 below) should not hold (cf. Remark 6.8) and this becomes important when $B(G,\mu)_{\mathrm{un}}$ has more than one element.
We also have the following.
Corollary 4.29.
Assume $G$ is a product of the groups appearing in Table (23) with $p$ and $\ell$ satisfying the corresponding conditions. Then, for $\phi$ a semisimple parameter induced from a generic $\phi_{T}$, we have that
$\mathrm{D}^{\mathrm{ULA}}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi}\simeq\bigoplus_{b\in B(G)_{\mathrm{un}}}\mathrm{D}^{\mathrm{adm}}(J_{b}(\mathbb{Q}_{p},\overline{\mathbb{F}}_{\ell})_{\phi}.$
Moreover, the !! and * pushforwards agree for any $A\in\mathrm{D}^{\mathrm{adm}}(J_{b}(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})_{\phi}\simeq\mathrm{D}^{\mathrm{ULA}}(\mathrm{Bun}_{G}^{b},\overline{\mathbb{F}}_{\ell})_{\phi}$
Proof.
This follows from Proposition 4.18, where the semisimplicity of the $\rho_{b,w}$ follows as in the proof of the Previous corollary. ∎
5. The Proof of the Main Theorems
We recall that the pair $(\mathbf{G},X)$ will denote a Shimura datum with reflex field $E/\mathbb{Q}$. We let $\ell\neq p$ be distinct prime numbers, and $E_{\mathfrak{p}}$ the completion at the place $\mathfrak{p}$ in $E$ dividing $p$ determined by the fixed isomorphism $j:\overline{\mathbb{Q}}_{p}\xrightarrow{\simeq}\mathbb{C}$ with completed algebraic closure denoted by $C$. We will let $G:=\mathbf{G}_{\mathbb{Q}_{p}}$ be the local group with maximal torus $T$, and we let $\mu$ be the conjugacy class of geometric dominant cocharacters of $G$ attached to the inverse of the Hodge cocharacter of $X$ and the isomorphism $j:\overline{\mathbb{Q}}_{p}\xrightarrow{\simeq}\mathbb{C}$. We assume that $G$ is unramified, and let $K_{p}^{\mathrm{hs}}\subset G(\mathbb{Q}_{p})$ denote a hyperspecial level. We let $H_{K_{p}^{\mathrm{hs}}}$ denote the spherical Hecke algebra with coefficients in $\overline{\mathbb{F}}_{\ell}$, and recall that $B(G)_{\mathrm{un}}:=\mathrm{Im}(B(T)\rightarrow B(G))$ denotes the set of unramified elements.
For $K\subset\mathbf{G}(\mathbb{A}_{f})$ a sufficiently small open compact, we write $\mathrm{Sh}(\mathbf{G},X)_{K}/\mathop{\rm Spec}(E)$ for the Shimura variety attatched to the Shimura datum, with its associated adic space $\mathcal{S}(\mathbf{G},X)_{K}$ over $\mathop{\rm Spa}(E_{\mathfrak{p}})$. and if $K=K^{p}K_{p}$, where $K_{p}$ (resp. $K^{p}$) denotes the level at $p$ (resp. away from $p$), we write $\mathcal{S}(\mathbf{G},X)_{K^{p}}$ for the associated Shimura variety at infinite level introduced in §3.1.
In this section, we will reap the fruit of our work in the previous section and finally prove our main results. More precisely, in §5.1, we will establish our main results on the splitting of Mantovan’s filtration (Theorem 5.1) for the torsion cohomology of a PEL type AC Shimura variety, as well as show the bounds for the torsion cohomology after generic localization (Theorem 5.2). In §5.2, we push our results a bit further and deduce additional applications to abelian type Shimura varieties (Corollaries 5.6, 5.7).
5.1. Proof of Theorems 1.8 and 1.17
We can now reap the fruit of our work in the previous sections and finally prove our main results. We start with Theorem 1.17.
Theorem 5.1.
Suppose $(\mathbf{G},X)$ is a PEL datum of type AC satisfying Assumption 1.13 such that $\mathbf{G}_{\mathbb{Q}_{p}}$ is a product of groups as in Table (1) with $p$ and $\ell$ satisfying the corresponding conditions and let $\phi_{T}$ be a generic toral parameter with associated semi-simple L-parameter $\phi$ of $G$. Then the complex $R\Gamma_{c}(\mathrm{Sh}(\mathbf{G},X)_{K,C},\overline{\mathbb{F}}_{\ell})_{\phi}$ breaks up as a direct sum
$\bigoplus_{b\in B(G,\mu)_{\mathrm{un}}}(R\Gamma_{c}(\mathrm{Sht}(G,b,\mu)_{\infty,C},\overline{\mathbb{F}}_{\ell}(d_{b}))_{\phi}\otimes^{\mathbb{L}}_{\mathcal{H}(J_{b})}R\Gamma_{c-\partial}(\mathrm{Ig}^{b},\overline{\mathbb{F}}_{\ell}))[2d_{b}]$
of $G(\mathbb{Q}_{p})$-modules.
Proof.
By Corollary 3.22, the complex $R\Gamma_{c}(\mathcal{S}(\mathbf{G},X)_{K^{p},C},\overline{\mathbb{F}}_{\ell})$ has a $G(\mathbb{Q}_{p})$-equivariant filtration with graded pieces isomorphic to $j_{1}^{*}T_{\mu}j_{b!}(V_{b})[-d](-\frac{d}{2})$. The cohomology of Igusa varieties $V_{b}$ and the global Shimura variety is admissible as a complex (For the $V_{b}$, we can write $V_{b}$ as the fiber of the natural excision map $R\Gamma(\mathrm{Ig}^{b,*},\overline{\mathbb{F}}_{\ell})\rightarrow R\Gamma(\partial\mathrm{Ig}^{b,*},\overline{\mathbb{F}}_{\ell})$ as in the proof of Corollary 3.5 and then analogously to [68, Proposition 8.21] show that $R\Gamma(\mathrm{Ig}^{b,*},\overline{\mathbb{F}}_{\ell})$ and $R\Gamma(\partial\mathrm{Ig}^{b,*},\overline{\mathbb{F}}_{\ell})$ are admissible, since the condition of admissibility is preserved under finite limits, the claim follows), and therefore it follows that $j_{b!}(V_{b})\in\mathrm{D}^{\mathrm{ULA}}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$, and so we can apply the results of the previous section to them. We consider the localization
$(j_{1}^{*}T_{\mu}j_{b!}(V_{b}))_{\phi}[-d](-\frac{d}{2}).$
This defines a filtration on $R\Gamma_{c}(\mathcal{S}(\mathbf{G},X)_{K^{p},C},\overline{\mathbb{F}}_{\ell})_{\phi}$. The filtration on $R\Gamma_{c}(\mathcal{S}(\mathbf{G},X)_{K^{p},C},\overline{\mathbb{F}}_{\ell})_{\phi_{\mathfrak{m}}}$ considered above comes from applying $(-)_{\phi}$ to $R\Gamma([\mathcal{F}\ell_{G,\mu^{-1}}^{\Diamond}/\underline{G(\mathbb{Q}_{p})}],i_{b!}i_{b}^{*}(R\pi_{\mathrm{HT!}}(\overline{\mathbb{F}}_{\ell})))$ viewed as a $G(\mathbb{Q}_{p})$-representation. Using Corollary 4.29, we see that these graded pieces are also isomorphic to
$(j_{1}^{*}T_{\mu}Rj_{b*}(V_{b}))_{\phi}[-d](-\frac{d}{2})$
via the natural transformation $j_{b!}\rightarrow Rj_{b*}$ and are trivial for $b\notin B(G,\mu)_{\mathrm{un}}$. However, using Lemma 3.18, this implies that the natural map
$R\Gamma([\mathcal{F}\ell_{G,\mu^{-1}}^{\Diamond}/\underline{G(\mathbb{Q}_{p})}],i_{b!}i_{b}^{*}(R\pi_{\mathrm{HT!}}(\overline{\mathbb{F}}_{\ell})))_{\phi_{\mathfrak{m}}}\rightarrow R\Gamma([\mathcal{F}\ell_{G,\mu^{-1}}^{\Diamond}/\underline{G(\mathbb{Q}_{p})}],i_{b*}i_{b}^{*}(R\pi_{\mathrm{HT!}}(\overline{\mathbb{F}}_{\ell})))_{\phi_{\mathfrak{m}}}$
is an isomorphism (See Remark 3.19). Therefore, we see that the edge maps in the excision spectral sequence actually degenerate after applying $(-)_{\phi}$, giving us a direct sum decomposition
$R\Gamma_{c}(\mathcal{S}(\mathbf{G},X)_{K^{p},C},\overline{\mathbb{F}}_{\ell})_{\phi}\simeq\bigoplus_{b\in B(G,\mu)_{\mathrm{un}}}j_{1}^{*}T_{\mu}j_{b!}(V_{b})[-d](-\frac{d}{2})_{\phi},$
as desired. ∎
We now turn our attention to Theorem 1.8.
Theorem 5.2.
Suppose $(\mathbf{G},X)$ is a PEL datum of type AC satisfying Assumption 1.13 such that $\mathbf{G}_{\mathbb{Q}_{p}}$ is a product of groups as in Table (1) with $p$ and $\ell$ satisfying the corresponding conditions, and let $\mathfrak{m}\subset H_{K_{p}^{\mathrm{hs}}}$ be a generic maximal ideal with associated semi-simple L-parameter $\phi_{\mathfrak{m}}$. Then, for a level $K=K^{p}K_{p}^{\mathrm{hs}}\subset\mathbf{G}(\mathbb{A}_{f})$, the cohomology of $R\Gamma(\mathrm{Sh}(\mathbf{G},X)_{K,\overline{E}},\overline{\mathbb{F}}_{\ell})_{\mathfrak{m}}$ (resp. $R\Gamma_{c}(\mathrm{Sh}(\mathbf{G},X)_{K,\overline{E}},\overline{\mathbb{F}}_{\ell})_{\mathfrak{m}}$) is concentrated in degrees $d\leq i\leq 2d$ (resp. $0\leq i\leq d$).
Proof.
We recall, by Proposition 3.6, that $V_{b}$ is a complex of smooth $J_{b}(\mathbb{Q}_{p})$-representations concentrated in degree $\leq d_{b}$. It follows that we have $\bigoplus_{b\in B(G,\mu)}j_{b!}(V_{b})_{\phi_{\mathfrak{m}}}\in\phantom{}^{\mathrm{p}}\mathrm{D}^{\leq 0,\mathrm{ULA}}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi_{\mathfrak{m}}}$, using Proposition A.5. Corollary 4.27 implies that
$j_{1}^{*}T_{\mu}j_{b!}(V_{b})[-d]_{\phi_{\mathfrak{m}}}\in\mathrm{D}^{\leq d}(G(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})_{\phi_{\mathfrak{m}}}$
after forgetting the Weil group action. Therefore, we conclude that
$R\Gamma_{c}(\mathcal{S}(\mathbf{G},X)_{K^{p},C},\overline{\mathbb{F}}_{\ell})_{\phi_{\mathfrak{m}}}$
is concentrated in degrees $0\leq i\leq d$. By applying Poincaré duality at the pro-$p$ finite levels and Corollary A.7, this allows us to conclude that the non-compactly supported cohomology
$R\Gamma(\mathcal{S}(\mathbf{G},X)_{K^{p},C},\overline{\mathbb{F}}_{\ell})_{\phi_{\mathfrak{m}}^{\vee}}$
localized at $\phi_{\mathfrak{m}}^{\vee}$ is concentrated in degrees $d\leq i\leq 2d$, where we define this to be the colimit over the non-compactly supported cohomology of finite levels (cf. Remark 3.7). Moreover, we note that genericity is preserved under taking duals and the same is true for regularity. It therefore follows that
$R\Gamma(K_{p}^{\mathrm{hs}},R\Gamma(\mathcal{S}(\mathbf{G},X)_{K^{p},C},\overline{\mathbb{F}}_{\ell})_{\phi_{\mathfrak{m}}^{\vee}})$
is also concentrated in degrees $\geq d$, but this is isomorphic to
$R\Gamma(\mathcal{S}(\mathbf{G},X)_{K^{p}K_{p}^{\mathrm{hs}},C},\overline{\mathbb{F}}_{\ell})_{\mathfrak{m}^{\vee}}$
by Lemma 4.3 (3). This establishes Theorem 1.8, by applying Poincaré duality to the Shimura variety at finite level $K_{p}^{\mathrm{hs}}$. ∎
We now turn our attention to the abelian type case.
5.2. Proof of Corollary 1.10
We would like to obtain the main Theorem for some Shimura varities of non-PEL type, especially Hilbert-Siegel modular varieties (attached to $\mathrm{Res}_{F/\mathbb{Q}}\mathrm{GL}_{2}$ or $\mathrm{Res}_{F/\mathbb{Q}}\mathrm{GSp}_{4}$). We will show this in a more general setup, as follows.
Let $(\mathbf{G},X)$, $(\mathbf{G}_{1},X_{1})$ be a pair of abelian type Shimura data such that $\mathbf{G},\mathbf{G}_{1}$ are centrally isogenous, and we have an isomorphism of derived subgroups
$\mathbf{G}^{\mathrm{der}}\xrightarrow{\sim}\mathbf{G}_{1}^{\mathrm{der}},$
(and hence of adjoint quotients). We further assume that both groups $G,G_{1}$ are unramified (observe that we can always arrange this, up to changing the prime $p$). Consider the associated Shimura varieties $\mathrm{Sh}(\mathbf{G},X)_{K}$ and $\mathrm{Sh}(\mathbf{G}_{1},X_{1})_{K_{1}}$, where we will choose the level $K,K_{1}$ such that the level at $p$ satisfies that we have an equality $K_{p}\cap G^{\mathrm{der}}(\mathbb{Q}_{p})=K_{1p}\cap G_{1}^{\mathrm{der}}(\mathbb{Q}_{p})$. In fact, we can and will take $K_{p}$ and $K_{1p}$ hyperspecial and hence $K^{\prime}_{p}=K_{p}\cap G^{\mathrm{der}}(\mathbb{Q}_{p})$ is also hyperspecial. This is possible because we observe that the group $G^{\mathrm{der}}$ is hence also unramified, and we have a natural map of Bruhat-Tits buildings
$\mathcal{B}(G^{\mathrm{der}},\breve{\mathbb{Q}}_{p})\rightarrow\mathcal{B}(G,\breve{\mathbb{Q}}_{p})$
which sends a hyperspecial point on $\mathcal{B}(G^{\mathrm{der}},\breve{\mathbb{Q}}_{p})$ to a hyperspecial point, so by fixing a choice of hyperspecial point of $\mathcal{B}(G^{\mathrm{der}},\breve{\mathbb{Q}}_{p})$ we can obtain hyperspecial subgroups $K_{p}^{\prime},K_{p},K_{1p}$ as desired.
Now, observe that by the Satake isomorphism, we have an isomorphism of $\overline{\mathbb{F}}_{\ell}$-algebras
$H_{K_{p}}\simeq\overline{\mathbb{F}}_{\ell}[X_{*}(T)]^{W_{G}},$
and, since $G,G^{\mathrm{der}}$ have isomorphic adjoint groups, the inclusion of cocharacters $X_{*}(T^{\prime})\subset X_{*}(T)$ induces an inclusion of Hecke algebras $H^{\prime}_{K^{\prime}_{p}}\subset H_{K_{p}}$, where $T^{\prime}$ denotes the torus $T\cap G^{\mathrm{der}}$, and $H^{\prime}_{K^{\prime}_{p}}$ denotes the spherical Hecke algebra for $G^{\mathrm{der}}$. Moreover, given a maximal ideal $\mathfrak{m}\subset H_{K_{p}}$, then $\mathfrak{m}^{\prime}=\mathfrak{m}\cap H^{\prime}_{K^{\prime}_{p}}$ is a maximal ideal of $H^{\prime}_{K^{\prime}_{p}}$.
We first want to understand how this passage to the derived subgroup works for localization at various $L$-parameters $\phi$. We have the following Lemma:
Lemma 5.3.
Let $G^{\prime}\rightarrow G$ be a map inducing an isomorphism on adjoint groups with $g:\phantom{}^{L}G\rightarrow\phantom{}^{L}G^{\prime}$, the induced map on dual groups. For $\phi:W_{\mathbb{Q}_{p}}\rightarrow\phantom{}^{L}G(\overline{\mathbb{F}}_{\ell})$ a L-parameter and $A$ an admissible complex of $G(\mathbb{Q}_{p})$-modules, then there is a natural isomorphism of $G^{\prime}(\mathbb{Q}_{p})$-modules
$(A|_{G^{\prime}(\mathbb{Q}_{p})})_{\phi^{\prime}}\simeq\bigoplus_{\begin{subarray}{c}\phi\\ \phi^{\prime}=g\circ\phi\end{subarray}}A_{\phi}|_{G^{\prime}(\mathbb{Q}_{p})}.$
where the LHS is the composite $\mathrm{D}(G^{\prime}(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})\xrightarrow{j_{1!}}\mathrm{D}(\mathrm{Bun}_{G^{\prime}},\overline{\mathbb{F}}_{\ell})\xrightarrow{(-)_{\phi^{\prime}}}\mathrm{D}(\mathrm{Bun}_{G^{\prime}},\overline{\mathbb{F}}_{\ell})_{\phi^{\prime}}$ applied to $A|_{G^{\prime}(\mathbb{Q}_{p})}$.
Proof.
By applying Corollary 4.4, we obtain a decomposition
$A\simeq\bigoplus_{\phi}A_{\phi}$
of $A$ as a $G(\mathbb{Q}_{p})$-module. We restrict to $G^{\prime}(\mathbb{Q}_{p})$ and apply the localization map $(-)_{\phi^{\prime}}$. This gives an isomorphism
$(A|_{G^{\prime}(\mathbb{Q}_{p})})_{\phi^{\prime}}\simeq\bigoplus_{\phi}(A_{\phi}|_{G^{\prime}(\mathbb{Q}_{p})})_{\phi^{\prime}},$
where we have used that localization commutes with direct sums since it is a left adjoint by definition. Now, using the compatibility of the Fargues-Scholze correspondence with central isogenies [16, Theorem IX.6.1], either $\phi^{\prime}=g\circ\phi$ and $A_{\phi}|_{G^{\prime}(\mathbb{Q}_{p})}\in\mathrm{D}(G^{\prime}(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})_{\phi^{\prime}}$ and, by the idempotence of the localization map, we have that $(A_{\phi}|_{G^{\prime}(\mathbb{Q}_{p})})_{\phi^{\prime}}=A_{\phi^{\prime}}|_{G^{\prime}(\mathbb{Q}_{p})}$ or $(A_{\phi}|_{G^{\prime}(\mathbb{Q}_{p})})_{\phi^{\prime}}$ is 0. The claim follows. ∎
This passage to derived subgroups allows us to pass to connected components of Shimura varieties. To see this, fix a connected component $X^{+}\subset X$. This also fixes a $X^{+}_{1}\subset X_{1}$, and an isomorphism $X^{+}\simeq X^{+}_{1}$, since $\mathbf{G},\mathbf{G}_{1}$ have isomorphic adjoint quotients. For any compact open subgroup $K\subset\mathbf{G}(\mathbb{A}_{f})$, we let $\mathrm{Sh}^{+}(\mathbf{G},X)_{K}$ be the geometrically connected component which is the image of $X^{+}\times 1$. Moreover, we will let
$\mathrm{Sh}^{+}(\mathbf{G},X)_{K_{p}}=\mathop{\rm lim}\nolimits\limits_{{\begin{subarray}{c}\leftarrow\\ K^{p}\end{subarray}}}\mathrm{Sh}^{+}(\mathbf{G},X)_{K_{p}K^{p}}.$
Note that since all the transition morphisms are finite étale and hence affine, $\mathrm{Sh}^{+}(\mathbf{G},X)_{K_{p}}$ is also a qcqs scheme by [61, Lemma 01YX].
Now, observe that the action of $H^{\prime}_{K^{\prime}_{p}}$ on $\mathrm{Sh}(\mathbf{G},X)_{K_{p}}$ via the inclusion $H^{\prime}_{K^{\prime}_{p}}\subset H_{K_{p}}$ preserves the geometric connected component, as it acts trivially on the set of geometric connected components. To see this, recall from [14, §2.1.3] that the connected components are given by the profinite group
$\pi_{0}(\mathrm{Sh}_{K_{p}}(\mathbf{G},X))=\mathbf{G}(\mathbb{Q})^{-}_{+}\backslash\mathbf{G}(\mathbb{A}^{f})/K_{p},$
where as in the notation of [14] we denote $\mathbf{G}(\mathbb{R})_{+}$ the inverse image of $\mathbf{G}^{\mathrm{ad}}(\mathbb{R})$, $\mathbf{G}(\mathbb{Q})_{+}=\mathbf{G}(\mathbb{R})_{+}\cap\mathbf{G}(\mathbb{Q}$ and $\mathbf{G}(\mathbb{Q})^{-}_{+}$ is the closure of $\mathbf{G}(\mathbb{Q})_{+}$ in $\mathbf{G}(\mathbb{A}_{f})$. Since this identification is $\mathbf{G}(\mathbb{A}_{f})$-equivariant, and $\pi_{0}(\mathrm{Sh}_{K_{p}}(\mathbf{G},X))$ is abelian, we thus we see that the action of $\mathbf{G}(\mathbb{A}^{f})$ must factor through the maximal abelian quotient, so the action of $\mathbf{G}^{\mathrm{der}}(\mathbb{A}^{f})$ is trivial.
Now, using the notation of [36, §3.3] we see that there exist groups $\mathscr{A}(\mathbf{G})^{\circ},\mathscr{A}(\mathbf{G}_{1}),\mathscr{A}(\mathbf{G})$, whose definition we do not precisely recall here, such that from [36, Proposition 3.3.10], we know that we have an isomorphism
$\mathrm{Sh}(\mathbf{G}_{1},X_{1})_{K_{1p}}\simeq(\mathrm{Sh}^{+}(\mathbf{G},X)_{K_{p}}\times\mathscr{A}(\mathbf{G}_{1})/\mathscr{A}(\mathbf{G})^{\circ}.$
We further note that since $G_{1}^{\mathrm{der}}=G^{\mathrm{der}}$, we can simplify this to
$\mathrm{Sh}(\mathbf{G}_{1},X_{1})_{K_{1p}}\simeq\mathrm{Sh}(\mathbf{G},X)_{K_{p}}^{+}\times\mathscr{A}(\mathbf{G}_{1})/\mathscr{A}(\mathbf{G})^{\circ},$
since from [36, §3.3.9] we see that the map $\mathscr{A}(\mathbf{G})^{\circ}\simeq\mathscr{A}(\mathbf{G}^{\mathrm{der}})^{\circ}\simeq\mathscr{A}(\mathbf{G}^{\mathrm{der}}_{1})^{\circ}\hookrightarrow\mathscr{A}(\mathbf{G}_{1}))$ is injective, using that $\mathbf{G},\mathbf{G}_{1}$ have isomorphic derived subgroups.
Thus, we have an isomorphism
$R\Gamma(\mathrm{Sh}(\mathbf{G}_{1},X_{1})_{K_{1p}},\overline{\mathbb{F}}_{\ell})_{\mathfrak{m^{\prime}}}\simeq R\Gamma(\mathrm{Sh}(\mathbf{G},X)^{+}_{K_{p}},\overline{\mathbb{F}}_{\ell})_{\mathfrak{m^{\prime}}}\otimes^{\mathbb{L}}_{\overline{\mathbb{F}}_{\ell}}C(\mathscr{A}(\mathbf{G}_{1})/\mathscr{A}(\mathbf{G})^{\circ})$
where $C(\mathscr{A}(\mathbf{G}_{1})/\mathscr{A}(\mathbf{G})^{\circ})$ denotes the space of all continuous $\overline{\mathbb{F}}_{\ell}$-valued functions on the profinite set $\mathscr{A}(\mathbf{G}_{1})/\mathscr{A}(\mathbf{G})^{\circ}$.
Thus, observe that if $R\Gamma(\mathrm{Sh}^{+}(\mathbf{G},X)_{K_{p}},\overline{\mathbb{F}}_{\ell})_{\mathfrak{m^{\prime}}}$ is concentrated in degrees $[d,2d]$, then so is $R\Gamma(\mathrm{Sh}(\mathbf{G}_{1},X_{1})_{K_{1p}},\overline{\mathbb{F}}_{\ell})_{\mathfrak{m^{\prime}}}$, where we note that the above tensor product has no higher derived terms. Summarizing the above discussion, we thus have:
Proposition 5.4.
Suppose that we have two Shimura data $(\mathbf{G},X)$, $(\mathbf{G}_{1},X_{1})$ such that $(\mathbf{G}^{\mathrm{ad}},X^{\mathrm{ad}})\simeq(\mathbf{G}^{\mathrm{ad}}_{1},X^{\mathrm{ad}}_{1})$, $\mathbf{G}^{\mathrm{der}}\simeq\mathbf{G}_{1}^{\mathrm{der}}$, and both $G,G_{1}$ are unramified. Then the action of $H^{\prime}_{K^{\prime}_{p}}\subset H_{K^{\mathrm{hs}}_{p}}$ stabilizes each connected component, and for any maximal ideal $\mathfrak{m}^{\prime}\subset H^{\prime}_{K^{\prime}_{p}}$ if $R\Gamma(\mathrm{Sh}^{+}(\mathbf{G},X)_{K_{p}},\overline{\mathbb{F}}_{\ell})_{\mathfrak{m^{\prime}}}$ is concentrated in degrees $[d,2d]$, then so is $R\Gamma(\mathrm{Sh}(\mathbf{G}_{1},X_{1})_{K_{1p}},\overline{\mathbb{F}}_{\ell})_{\mathfrak{m^{\prime}}}$.
Proposition 5.5.
Suppose that $(\mathbf{G},X)$ is a PEL-type Shimura datum such that Conjecture 1.2 holds for $(\mathbf{G},X)$, and we have another Shimura datum $(\mathbf{G},X)$ such that $(\mathbf{G}^{\mathrm{ad}},X^{\mathrm{ad}})\simeq(\mathbf{G}^{\mathrm{ad}}_{1},X^{\mathrm{ad}}_{1})$, $\mathbf{G}^{\mathrm{der}}\simeq\mathbf{G}_{1}^{\mathrm{der}}$, and both $G,G_{1}$ are unramified. Then Conjecture 1.2 holds for $(\mathbf{G}_{1},X_{1})$.
Proof.
Since we assume that Conjecture 1.2 is true for $(\mathbf{G},X)$, we will first show that a maximal ideal $\mathfrak{m}$ of $H_{K_{p}}$ is generic if and only if the maximal ideal $\mathfrak{m}^{\prime}$ of $H^{\prime}_{K^{\prime}_{p}}$ is generic. To see this, we will reformulate this in terms of $L$-parameters. This is equivalent to showing that an $L$-parameter
$\phi=\phi_{\mathfrak{m}}:W_{\mathbb{Q}_{p}}\rightarrow\phantom{}^{L}T(\overline{\mathbb{F}}_{\ell})$
attached to $\mathfrak{m}$ is generic if and only if the composition $\phi^{\prime}$ with the map $g:\phantom{}^{L}T(\overline{\mathbb{F}}_{\ell})\rightarrow\phantom{}^{L}T^{\prime}(\overline{\mathbb{F}}_{\ell})$ induced by the inclusion of tori $T^{\prime}\hookrightarrow T$ is generic. (Here, $T^{\prime}=G^{\mathrm{der}}\cap T$). This follows from the observation that any coroot $\alpha$ factors through $G^{\mathrm{der}}$, and hence the composition $\alpha\circ\phi$ is equal to $\alpha\circ\phi^{\prime}$.
By Lemma 5.3 applied to $G^{\mathrm{der}}=G^{\prime}\subset G$, we have a natural decomposition
$H^{i}(\mathrm{Sh}(\mathbf{G},X)_{K^{p}},\overline{\mathbb{F}}_{\ell})_{\phi^{\prime}}\simeq\bigoplus_{\begin{subarray}{c}\phi\\ \phi^{\prime}=g\circ\phi\end{subarray}}H^{i}(\mathrm{Sh}(\mathbf{G},X)_{K^{p}},\overline{\mathbb{F}}_{\ell})_{\phi}.$
Hence, we see that $H^{i}(\mathrm{Sh}(\mathbf{G},X)_{K^{p}},\overline{\mathbb{F}}_{\ell})_{\phi^{\prime}}$ vanishes for $i<d$, since all the $\phi$ appearing on the right-hand side are generic, and hence we can apply Theorem 1.8. Applying $R\Gamma(K_{p}^{\prime},-)$ and Lemma 4.3 (3), we see that
$H^{i}(\mathrm{Sh}(\mathbf{G},X)_{K_{p}^{\prime}K^{p}},\overline{\mathbb{F}}_{\ell})_{\mathfrak{m}^{\prime}}$
vanishes for $i<d$. Now, we can further apply the Hochschild-Serre spectral sequence to the $K_{p}/K_{p}^{\prime}$-torsor $\mathrm{Sh}(\mathbf{G},X)_{K_{p}^{\prime}K^{p}}\rightarrow\mathrm{Sh}(\mathbf{G},X)_{K_{p}K^{p}}$ to get that $H^{i}(\mathrm{Sh}(\mathbf{G},X)_{K_{p}K^{p}},\overline{\mathbb{F}}_{\ell})_{\mathfrak{m}^{\prime}}$ vanishes for $i<d$.
Now, we can consider the connected component $\mathrm{Sh}^{+}(\mathbf{G},X)_{K_{p}K^{p}}$, and observe that $\mathrm{Sh}(\mathbf{G},X)_{K_{p}K^{p}}$ is a finite disjoint union of isomorphic copies of $\mathrm{Sh}^{+}(\mathbf{G},X)_{K_{p}K^{p}}$. Moreover, exactly as explained above the action of $H^{\prime}_{K_{p}^{\prime}}$ stabilizes each connected component, so we see that $H^{i}(\mathrm{Sh}^{+}(\mathbf{G},X)_{K_{p}K^{p}},\overline{\mathbb{F}}_{\ell})_{\mathfrak{m}^{\prime}}$ vanishes for $i<d$ as well.
Finally, taking colimits over $K^{p}$, we see that $H^{i}(\mathrm{Sh}^{+}(\mathbf{G},X)_{K_{p}},\overline{\mathbb{F}}_{\ell})_{\mathfrak{m}^{\prime}}$ vanishes for $i<d$ as well. Applying Proposition 5.4, $H^{i}(\mathrm{Sh}(\mathbf{G}_{1},X_{1})_{K_{1p}},\overline{\mathbb{F}}_{\ell})_{\mathfrak{m^{\prime}}}$ vanishes for $i<d$.
Now, we let $\mathbf{Z}\subset\mathbf{G}$ be the center, and we denote by $\mathbf{Z}(\mathbb{Q})^{-}$ the closure of $\mathbf{Z}(\mathbb{Q})$ in $\mathbf{Z}(\mathbb{A}_{f})$. Denote by $\overline{K_{1}^{p}}$ the quotient
$\overline{K_{1}^{p}}=(K_{1p}K^{p}_{1}/K_{1p}K^{p}_{1}\cap\mathbf{Z}(\mathbb{Q})^{-})/(K^{p}_{1}/K^{p}_{1}\cap\mathbf{Z}(\mathbb{Q})^{-}).$
By the Hochschild-Serre spectral sequence (applied to the quotients $\overline{K_{1}^{p}}$), we see that for all sufficiently small $K_{1}^{p}$, $H^{i}(\mathrm{Sh}(\mathbf{G}_{1},X_{1})_{K_{1p}K_{1}^{p}},\overline{\mathbb{F}}_{\ell})_{\mathfrak{m}^{\prime}}$ vanishes for $i<d$.
Now, consider a generic maximal ideal $\mathfrak{m}_{1}$ for the spherical Hecke algebra $H_{K_{1p}}$ of $G_{1}$. This corresponds to a generic maximal ideal $\mathfrak{m}^{\prime}$ of $H^{\prime}_{K^{\prime}_{p}}$. It remains to observe that for any finitely-generated $H_{K_{1p}}$-module $A$, if the localization $A_{\mathfrak{m}^{\prime}}=0$, then there is some element in $r\in H^{\prime}_{K^{\prime}_{p}}\backslash\mathfrak{m}^{\prime}$ such that $rA=0$. Thus we must have $A_{\mathfrak{m}_{1}}=0$ as well since $H^{\prime}_{K^{\prime}_{p}}\backslash\mathfrak{m}^{\prime}\subset H_{K_{1p}}\backslash\mathfrak{m}_{1}$. Taking $A=H^{i}(\mathrm{Sh}(\mathbf{G}_{1},X_{1})_{K_{1p}K_{1}^{p}},\overline{\mathbb{F}}_{\ell})$, we can conclude. ∎
We obtain the following by combining this Proposition with Theorem 1.8.
Corollary 5.6.
Suppose $(\mathbf{G}_{1},X_{1})$ is an abelian-type Shimura datum which has an associated PEL-type datum $(\mathbf{G},X)$ of type AC satisfying Assumption 1.13 and such that $\mathbf{G}^{\mathrm{der}}\simeq\mathbf{G}_{1}^{\mathrm{der}}$, $G_{1}:=\mathbf{G}_{1,\mathbb{Q}_{p}}$ is unramified and $G:=\mathbf{G}_{\mathbb{Q}_{p}}$ is a product of groups as in Table (1) with $p$ and $\ell$ satisfying the corresponding conditions. We assume that $\mathfrak{m}_{1}\subset H_{K_{1p}^{\mathrm{hs}}}$ is a generic maximal ideal. Then, for a level $K_{1}=K_{1}^{p}K_{1p}^{\mathrm{hs}}\subset\mathbf{G}(\mathbb{A}_{f})$, the cohomology of $R\Gamma(\mathrm{Sh}(\mathbf{G}_{1},X_{1})_{K_{1},\overline{E}},\overline{\mathbb{F}}_{\ell})_{\mathfrak{m}}$ (resp. $R\Gamma_{c}(\mathrm{Sh}(\mathbf{G}_{1},X_{1})_{K_{1},\overline{E}},\overline{\mathbb{F}}_{\ell})_{\mathfrak{m}}$) is concentrated in degrees $d\leq i\leq 2d$ (resp. $0\leq i\leq d$).
Proof.
This follows from Theorem 1.8, once we show that the generic condition on $\phi^{T}_{\mathfrak{m}}$ is also satisfied by $\phi^{T_{1}}_{\mathfrak{m}_{1}}$. To see this, we can in fact see that this condition is satisfied if and only if it is satisfied by $G^{\mathrm{der}}$, where we look at the composition
$\phi^{T^{\mathrm{der}}}_{\mathfrak{m^{\prime}}}:W_{\mathbb{Q}_{p}}\rightarrow\phantom{}^{L}T(\overline{\mathbb{F}}_{\ell})\rightarrow\phantom{}^{L}T^{\mathrm{der}}(\overline{\mathbb{F}}_{\ell}).$
In fact, observe that we can descend any coroot $\alpha$ to a coroot $\alpha^{\mathrm{der}}:\phantom{}^{L}T^{\mathrm{der}}(\Lambda)\rightarrow\overline{\mathbb{F}}_{\ell}^{\times}$, since $G,G^{\mathrm{der}}$ have the same (co-)root systems, and we have $\alpha\circ\phi_{\mathfrak{m}}^{T}=\alpha^{\mathrm{der}}\circ\phi^{T^{\mathrm{der}}}_{\mathfrak{m^{\prime}}}$. ∎
In particular, we can strengthen previous results of Caraiani-Tamiozzo [9, Theorem B], who previously showed torsion vanishing for Hilbert modular varieties under the additional assumption that $p$ was split in the totally real field $F$ (though we also note that they showed torsion vanishing under a hypothesis on $\mathfrak{m}$ which is weaker than the genericity considered here, see Remark 1.7).
Corollary 5.7.
Conjecture 1.2 is true for Hilbert modular varieties, quaternionic Shimura varieties assuming $(\ell,2[L:\mathbb{Q}_{p}])=1$, and Hilbert-Siegel Shimura varieties (attached to $\mathrm{Res}_{F/\mathbb{Q}}\mathrm{GSp}_{4}$) assuming that $\ell\nmid[L:\mathbb{Q}_{p}]$, where for both cases $L$ is the completion of $F$ at some unramified prime above $p$.
Proof.
Observe that for Hilbert-Siegel Shimura varieties (attached to $\mathrm{Res}_{F/\mathbb{Q}}\mathrm{GL}_{2}$, $\mathrm{Res}_{F/\mathbb{Q}}\mathrm{GSp}_{4}$), there is a cover by a PEL-type Shimura variety with local group $G$ of the form $G(\mathrm{SL}_{2})$ and $G(\mathrm{Sp}_{4})$ respectively. For the case of quaternionic Shimura varieties, we can relate their geometric connected components to unitary PEL-type Shimura varieties with local group $G=\mathrm{G}(U_{2}\times\dots\times U_{2})$, as described in [64, Corollary 3.11]. This group does not appear in Table 1, but we show below that it satisfies the additional assumptions to apply Theorem 4.23. Therefore, the result follows from Corollary 5.6. ∎
Lemma 5.8.
The group $G=\mathrm{G}(U_{2}\times\dots\times U_{2})$ satisfies Assumption 4.5, and genericity implies regularity and strong $\mu$-regularity for any toral parameter.
Proof.
Firstly observe that $G$ is a subgroup of $\mathrm{Res}_{L/\mathbb{Q}_{p}}\mathrm{GU}_{2}$. Thus, we can apply Proposition 4.9 because we know the result for $\mathrm{Res}_{L/\mathbb{Q}_{p}}\mathrm{GU}_{2}$. The additional assumptions required to invoke Theorem 4.23 can be verified after passing to an extension where the group splits, from which we see that $\mathrm{G}(U_{2}\times\dots\times U_{2})$ is isomorphic to $G^{\prime}=(\mathbb{G}_{m}\times\mathbb{G}_{m}\times\mathrm{GL}_{2}\times\dots\times\mathrm{GL}_{2})/\mathbb{G}_{m}$, where the $\mathbb{G}_{m}$ in the quotient is the diagonal copy of $\mathbb{G}_{m}$. Moreover, we observe that the maximal torus $T^{\prime}$ of $G^{\prime}$ can be further identified with $\mathbb{G}_{m}\times T\times\dots\times T$ by setting the element in the first copy of $\mathbb{G}_{m}$ to 1, and where each $T$ is the diagonal torus on $\mathrm{GL}_{2}$. In particular, we can reduce to the case of a product of $\mathrm{GL}_{2}$’s, from which the additional conditions follow as in the proof of Lemma 4.25. More precisely, we write an element of $T$ as $(t,t_{11},t_{12},\dots,t_{d1},t_{d2})$, and the character $\chi$ as $\nu(t)\chi_{11}(t_{11})\dots\chi_{d2}(t_{d2})$. We get regularity by observing that if any element of $W_{G}$ is non-trivial on some $T$ factor (say the $i$th one), then by setting $t=1$, $t_{j1}=t_{j2}=1$, we get the relation $\chi_{i1}\chi^{-1}_{i2}\simeq\mathbf{1}$, a contradiction. We get $\mu$-regularity for any dominant cocharacter $\mu$, since we see that the strong $\mu$-regularity condition again holds for any minuscule $\mu$, and any dominant cocharacter of $G$ can be written as a sum of such minuscule $\mu$. ∎
Remark 5.9.
As remarked by a referee, there are Shimura data other than the ones listed in Corollary 5.7 to which we can apply Proposition 5.5. For instance, one can consider the Shimura data where the group $\mathbf{G}_{1}$ is an inner form of $\mathrm{Res}_{F/\mathbb{Q}}\mathrm{GSp}_{4}$ for $F$ totally real, and which is split at all real places, and $X_{1}$ is the product of $F:\mathbb{Q}]$ copies of the Siegel upper half space $\mathcal{H}_{2}$. In this case, one can show that there exists a PEL-type datum $(\mathbf{G},X)$ satisfying $(\mathbf{G}^{\mathrm{ad}},X^{\mathrm{ad}})\simeq(\mathbf{G}^{\mathrm{ad}}_{1},X^{\mathrm{ad}})$, and $\mathbf{G}^{\mathrm{der}}\simeq\mathbf{G}_{1}^{\mathrm{der}}$.
6. Conjectures and Concluding Remarks
In this section, we give some further complements to the main results established in the previous section. In particular, in §6.1, we explain how, by combining Theorem 1.17, with some further results of [24] one can obtain a more precise description of the contribution of a basic unramified element to the cohomology of the torsion cohomology of the global Shimura variety, that is compatible with an analogous description provided by Xiao and Zhu in [67]. In §6.2, we discuss possible generalizations of our results, introducing the notion of a Langlands-Shahidi type parameter, which should be the most general condition for which the analogue of all of our main results hold (Conjectures 6.4, 6.6).
6.1. Relationship to Xiao-Zhu
We keep the running notation and assumptions of the previous section. In particular, the pair $(\mathbf{G},X)$ will be a Shimura datum of PEL type AC satisfying Assumption 1.13, and $K=K^{p}K_{p}\subset\mathbf{G}(\mathbb{A}_{f})$ will denote a sufficiently small level. We assume that the local group $G:=\mathbf{G}_{\mathbb{Q}_{p}}$ is an unramified group of the form described in (1) with $p$ and $\ell$ satisfying the corresponding conditions and that $\mathfrak{m}\subset H_{K_{p}^{\mathrm{hs}}}$ is a generic maximal ideal in the spherical Hecke algebra of the form described in Theorem 1.11. We will assume that the basic element $b\in B(G,\mu)_{\mathrm{un}}$ is unramified (See [67, Remark 4.2.11] for a classification). Let us look at the middle degree cohomology $H^{d}(R\Gamma_{c}(\mathcal{S}(\mathbf{G},X)_{K^{p},C},\overline{\mathbb{F}}_{\ell})_{\phi_{\mathfrak{m}}})$. By Theorem 1.17, it has a summand isomorphic to
$H^{d}(R\Gamma_{c}(G,b,\mu)_{\phi_{\mathfrak{m}}}\otimes^{\mathbb{L}}_{\mathcal{H}(J_{b})}R\Gamma_{c-\partial}(\mathrm{Ig}^{b},\overline{\mathbb{F}}_{\ell})).$
To describe this, let $\mathbf{G}^{\prime}$ be the unique $\mathbb{Q}$-inner form of $\mathbf{G}$ such that $\mathbf{G}(\mathbb{A}^{p\infty})\simeq\mathbf{G}^{\prime}(\mathbb{A}^{p\infty})$, $\mathbf{G}^{\prime}(\mathbb{R})$ is compact modulo center, and $\mathbf{G}_{\mathbb{Q}_{p}}\simeq J_{b}$ (See [29, Proposition 3.1] for the existence). We write $C(\mathbf{G}^{\prime}(\mathbb{Q})\backslash\mathbf{G}^{\prime}(\mathbb{A}_{f})/K^{p},\overline{\mathbb{F}}_{\ell})$ for the set of all continuous functions on the profinite (since $\mathbf{G}^{\prime}_{\mathbb{R}}$ is compact modulo center) set $\mathbf{G}^{\prime}(\mathbb{Q})\backslash\mathbf{G}^{\prime}(\mathbb{A}_{f})/K^{p}$. It is easy to show that one has an isomorphism
$C(K^{p}\backslash\mathbf{G}^{\prime}(\mathbb{A}_{f})/\mathbf{G}^{\prime}(\mathbb{Q}),\overline{\mathbb{F}}_{\ell})\simeq R\Gamma_{c-\partial}(\mathrm{Ig}^{b},\overline{\mathbb{F}}_{\ell})$
[…]
(24) ${}R\Gamma_{c}(\mathrm{Sht}(G,b,\mu)_{\infty,C}/\underline{K_{p}^{\mathrm{hs}}},\overline{\mathbb{F}}_{\ell})_{\mathfrak{m}}\otimes_{\mathcal{H}(J_{b})}^{\mathbb{L}}C(K^{p}\backslash\mathbf{G}^{\prime}(\mathbb{A}_{f})/\mathbf{G}^{\prime}(\mathbb{Q}),\overline{\mathbb{F}}_{\ell})$ $\simeq$
$C(K^{p}K_{p}^{\mathrm{hs}}\backslash\mathbf{G}^{\prime}(\mathbb{A}_{f})/\mathbf{G}^{\prime}(\mathbb{Q}),\overline{\mathbb{F}}_{\ell})_{\mathfrak{m}}\otimes V_{\mu}|_{\hat{G}^{\Gamma}}(b_{T})[-d](-\frac{d}{2})$
of $H_{K_{p}^{\mathrm{hs}}}$-representations [^12] One should also be able describe the Weil group action, as in [24, Conjecture 1.20]. up to taking the semi-simplifications of both sides as $H_{K_{p}^{\mathrm{hs}}}$-representations, where we note that $J_{b}\simeq G$ if $b\in B(G,\mu)_{\mathrm{un}}$ since $b$ is basic, and $J_{b}$ must be quasi-split since $b$ is unramified. Note that, if we were working with $\overline{\mathbb{Q}}_{\ell}$-coefficients [^13] For this comparison, it would have been more natural to consider an analogue of Theorem 1.17 with $\overline{\mathbb{Q}}_{\ell}$-coefficients. This is indeed doable assuming that $\phi_{\mathfrak{m}}$ admits a $\overline{\mathbb{Z}}_{\ell}$-lattice as in [24, Theorem 1.13]. This integrality condition is however an artifact of the theory of solid $\overline{\mathbb{Q}}_{\ell}$-sheaves not being properly understood (e.g excision fails) and should be removable with more technology., the space of automorphic forms $C(K^{p}\backslash\mathbf{G}^{\prime}(\mathbb{A}_{f})/\mathbf{G}^{\prime}(\mathbb{Q}),\overline{\mathbb{Q}}_{\ell})$ with rational coefficients will be semisimple as a $G(\mathbb{Q}_{p})$-representation, by virtue of the fact that $\mathbf{G}^{\prime}_{\mathbb{R}}$ will be compact modulo center, and in this case the semi-simplification would be unnecessary.
In particular, by arguing as in [37, Page 6], we know that $R\Gamma_{c}(\mathrm{Sht}(G,b,\mu)_{\infty,C}/\underline{K_{p}^{\mathrm{hs}}},\overline{\mathbb{F}}_{\ell})_{\mathfrak{m}}$ will have cohomology sheaves with irreducible constituents given by the representations of $J_{b}(\mathbb{Q}_{p})$ with Fargues-Scholze parameter equal to $\phi_{\mathfrak{m}}$ as conjugacy classes of parameters. Moreover, using that Assumption 4.5 holds for the groups appearing in Table (1), we know by Proposition 4.6 that they have to be constituents of $i_{B}^{G}(\chi)$, which will also be irreducible under the genericity assumption and the constraints appearing in Table (1) (See the proof of Corollary 4.27). In particular, the only irreducible $J_{b}(\mathbb{Q}_{p})=G(\mathbb{Q}_{p})$-representations that will contribute to the $H_{K_{p}^{\mathrm{hs}}}$-semisimplification of the LHS of (24) will be given by $i_{B}^{G}(\chi)$, and [24, Conjecture 1.20] would imply that $R\Gamma_{c}(G,b,\mu)[i_{B}^{G}(\chi)]\simeq i_{B}^{G}(\chi)\otimes V_{\mu}|_{\hat{G}^{\Gamma}}(b_{T})[-d](\frac{-d}{2})$ as $G(\mathbb{Q}_{p})$-modules. If we assume $\ell$ is banal (i.e coprime to the pro-order of $K_{p}^{\mathrm{hs}}$) then passing to $K_{p}^{\mathrm{hs}}$-invariants, recalling that this is exact under the banal hypothesis, gives us the isomorphism (24) after semi-simplification as $H_{K_{p}^{\mathrm{hs}}}$-modules.
Remark 6.1.
If $B(G,\mu)_{\mathrm{un}}$ consists of only the basic element and the $\mu$-ordinary element and $\phi_{T}$ is strongly $\mu$-regular (Definition 4.16) then [24, Conjecture 1.20] is true. In particular, it follows from [24, Theorem 1.21] that the isomorphism (24) can be made unconditional. More precisely, for $(\mathbf{G},X)$ any PEL type datum of type AC satisfying Assumption 1.13 such that the local group $G$ is of the form described in (1), and $\mathfrak{m}\subset H_{K_{p}^{\mathrm{hs}}}$ a generic maximal ideal of the form described in Theorem 1.11 with associated semisimple toral parameter $\phi_{\mathfrak{m}}^{T}$ such that $\phi_{\mathfrak{m}}^{T}$ is strongly $\mu$-regular in the sense of Definition 4.16, then if $B(G,\mu)_{\mathrm{un}}$ consists of only the basic element and the $\mu$-ordinary element the isomorphism (24) holds after semi-simplifying as a $H_{K_{p}^{\mathrm{hs}}}$-module.
We note that this description of the $K_{p}^{\mathrm{hs}}$-invariants of the middle degree cohomology on the generic fiber of the Shimura variety at hyperspecial level parallels Theorem [67, Theorem 1.14 (1)], describing the middle degree cohomology on the special fiber of the natural integral model.
6.2. A General Torsion Vanishing Conjecture
Consider now a general Shimura datum $(\mathbf{G},X)$. Let $\Lambda\in\{\overline{\mathbb{Q}}_{\ell},\overline{\mathbb{F}}_{\ell}\}$. If $\Lambda=\overline{\mathbb{F}}_{\ell}$ assume that $\ell$ is very decent as in Definition 4.11. We can then look at the $G(\mathbb{Q}_{p})$-representation.
$R\Gamma_{c}(\mathcal{S}(\mathbf{G},X)_{K^{p},C},\Lambda)$
defined by the cohomology at infinite level. By applying Corollary 4.4, we obtain a $G(\mathbb{Q}_{p})$-equivariant decomposition of this
$R\Gamma_{c}(\mathcal{S}(\mathbf{G},X)_{K^{p},C},\Lambda)=\bigoplus_{\phi}R\Gamma_{c}(\mathcal{S}(\mathbf{G},X)_{K^{p},C},\Lambda)_{\phi}$
running over semi-simple $L$-parameters $\phi:W_{\mathbb{Q}_{p}}\rightarrow\phantom{}^{L}G(\Lambda)$. For such a $\phi$, we let $(\phi_{M},M)$ denote a cuspidal support. I.e $M$ is a Levi of $G$ and $\phi_{M}:W_{\mathbb{Q}_{p}}\rightarrow\phantom{}^{L}M(\Lambda)$ is a supercuspidal $L$-parameter such that $\phi$ is induced (up to $\hat{G}-$conjugacy) by composing with the natural (up to $\hat{G}-$conjugacy) embedding $\phantom{}{}^{L}M(\Lambda)\rightarrow\phantom{}^{L}G(\Lambda)$. We want to describe the degrees of cohomology that $R\Gamma_{c}(\mathcal{S}(\mathbf{G},X)_{K^{p},C},\Lambda)_{\phi}$ sits in for suitably nice $\phi$. The case where $\phi$ factors through $M=T$ is covered by Conjecture 1.2. To go beyond this, we give the following definition.
Definition 6.2.
For a semi-simple $L$-parameter $\phi$ with a cuspidal support $(M,\phi_{M})$, we let $P$ be a parabolic with Levi factor $M$ and unipotent radical $N$. We consider the representation $r$ given by looking at the action of $\phantom{}{}^{L}M$ on the Lie algebra of $\phantom{}{}^{L}N$ via the adjoint action. We say $\phi$ is of Langlands-Shahidi type if the Galois cohomology groups
$R\Gamma(W_{\mathbb{Q}_{p}},r\circ\phi_{M})$
and
$R\Gamma(W_{\mathbb{Q}_{p}},r\circ\phi_{M}^{\vee})$
are trivial. Similarly, we say $\phi$ is of weakly Langlands-Shahidi type if
$H^{2}(R\Gamma(W_{\mathbb{Q}_{p}},r\circ\phi_{M}))$
and
$H^{2}(R\Gamma(W_{\mathbb{Q}_{p}},r\circ\phi_{M}^{\vee})$
are trivial.
Remark 6.3.
We note that since we enforced this condition on both $r\circ\phi_{M}$ and $r\circ\phi_{M}^{\vee}$ that this is independent of the choice of parabolic $P$ and the choice of cuspidal support. Moreover, it is easy to check that, if $M=T$, this precisely recovers Definition 1.1.
The terminology of "Langlands-Shahidi type" comes from the fact that the representation $r\circ\phi_{M}$ is precisely the representation which appears in the description of the constant terms of the usual Eisenstein series via the Langlands-Shahidi method ([43], [59], [60]). The motivation for this definition comes from considering the behavior of geometric Eisenstein series over the Fargues-Fontaine curve for general parabolics, by making analogies with the classical theory over function fields, as developed in [5, 44]. In particular, this should be the correct definition that guarantees that the eigensheaves $\mathcal{S}_{\phi}$ on $\mathrm{Bun}_{G}$ with eigenvalue $\phi$ are as simple as possible, and the analysis carried out in [24] generalizes to the non-principal case. This is discussed in more detail in [26, Chapter 3]. In addition, we expect that the consequences derived from the analysis in [24] in the principal case should also generalize. More precisely, we conjecture the following generalization of Proposition 4.18 and Corollary 4.27
Conjecture 6.4.
Let $B(G)_{M}:=\mathrm{Im}(B(M)_{\mathrm{basic}}\rightarrow B(G))$ be the set of $M$-reducible elements, and let $\phi$ be a semi-simple $L$-parameter of Langlands-Shahidi type with cuspidal support $(M,\phi_{M})$. The category $\mathrm{D}_{\mathrm{lis}}^{\mathrm{ULA}}(\mathrm{Bun}_{G},\Lambda)_{\phi}$ of $\phi$-local (as defined in Appendix A) lisse-étale ULA $\Lambda$-sheaves breaks up as direct sum
$\mathrm{D}_{\mathrm{lis}}^{\mathrm{ULA}}(\mathrm{Bun}_{G},\Lambda)_{\phi}\simeq\bigoplus_{b\in B(G)_{M}}\mathrm{D}_{\mathrm{lis}}^{\mathrm{ULA}}(\mathrm{Bun}_{G}^{b},\Lambda)_{\phi}$
via excision [^14] We note that a priori there are no excision triangles for the HN-stratification of $\mathrm{Bun}_{G}$, since there is in general no well-behaved !!-pushforward in the lisse formalsim with non-torsion coefficients. However, one can show that such a pushforward exists for the inclusions of Harder-Narasimhan strata, and satisfies the usual excision triangles (See for example the discussion on [33, Page 11]), and so this is well-defined in the present context., and the !! and * pushhforwards agree for any smooth irreducible representation $\rho$ of $J_{b}(\mathbb{Q}_{p})$ lying in $\mathrm{D}_{\mathrm{lis}}(\mathrm{Bun}_{G}^{b},\Lambda)_{\phi}$ for $b\in B(G)_{M}$.
Given a tilting module $V\in\mathrm{Tilt}_{\Lambda}(\phantom{}^{L}G^{I})$ (where $\mathrm{Tilt}_{\Lambda}(\phantom{}^{L}G^{I})$ is defined as in [24, Section 9]), if $\phi$ is of weakly Langlands-Shahidi type then the map induced by associated the Hecke operator
$T_{V}:\mathrm{D}_{\mathrm{lis}}^{\mathrm{ULA}}(\mathrm{Bun}_{G},\Lambda)_{\phi}\rightarrow\mathrm{D}_{\mathrm{lis}}^{\mathrm{ULA}}(\mathrm{Bun}_{G},\Lambda)^{BW_{\mathbb{Q}_{p}}^{I}}_{\phi}$
on the $\phi$-localized ULA subcategory is perverse $t$-exact, where the fact the Hecke operator preserves this subcategory is proven as in Lemma 4.3 (2).
Remark 6.5.
During the preparation of this manuscript, Hansen formulated similar conjectures with rational coefficients [30]. He refers to Langlands-Shahidi parameters as generous parameters [30, Definition 2.5] and to weakly Langlands-Shahidi parameters as generic semi-simple parameters [30, Section 2.3]. One can show that these two definitions are equivalent. Indeed, note that the Galois cohomology $H^{1}(R\Gamma(W_{\mathbb{Q}_{p}},r\circ\phi_{M}))$ controls the lifts of a semi-simple parameter $\phi_{M}:W_{\mathbb{Q}_{p}}\rightarrow\phantom{}^{L}M(\Lambda)$ to a $\phantom{}{}^{L}P(\Lambda)$-valued parameter and that such lifts correspond to finding parameters whose semi-simplification is equal to $\phi$. Moreover, insisting that $H^{1}(R\Gamma(W_{\mathbb{Q}_{p}},r\circ\phi_{M}))$ is trivial is equivalent to insisting that $R\Gamma(W_{\mathbb{Q}_{p}},r\circ\phi_{M})$ is trivial using local Tate-duality and that the Euler-Poincaré characteristic of this complex is 0. This shows the equivalence of the generous condition with the Langlands-Shahidi type condition, using that the stack of Langlands parameters with rational coefficients is reduced. Lastly, the set of such lifts coming from classes in $H^{0}(R\Gamma(W_{\mathbb{Q}_{p}},r\circ\phi_{M}))$ will give rise to non Frobenius semi-simple L-parameters allowing one to see that weakly Langlands-Shahidi is equivalent to generic semi-simple.
In particular, by combining this with a generalization of Theorem 1.15 to arbitrary Shimura varieties and the analysis carried out in §5, we could deduce the following as a consequence.
Conjecture 6.6.
Let $\phi$ be a semi-simple $L$-parameter of weakly Langlands-Shahidi type with cuspidal support $(M,\phi_{M})$. Then the complex $R\Gamma_{c}(\mathcal{S}(\mathbf{G},X)_{K^{p},C},\Lambda)_{\phi}$ (resp. $R\Gamma(\mathcal{S}(\mathbf{G},X)_{K^{p},C},\Lambda)_{\phi}$) is concentrated in degrees $0\leq i\leq d$ (resp. $d\leq i\leq 2d$).
Remark 6.7.
For $(\mathbf{G},X)$ of PEL type AC, assuming 1.13 and that $\phi$ of Langlands-Shahidi type, we should also obtain a $W_{E_{\mathfrak{p}}}\times G(\mathbb{Q}_{p})$-equivariant direct sum decomposition
$R\Gamma_{c}(\mathcal{S}(\mathbf{G},X)_{K^{p},C},\Lambda)_{\phi}\simeq\bigoplus_{b\in B(G,\mu)_{M}}(R\Gamma_{c}(G,b,\mu)_{\phi}\otimes_{\mathcal{H}(J_{b})}^{\mathbb{L}}V_{b})[2d_{b}],$
where $R\Gamma_{c}(G,b,\mu):=\mathop{\rm colim}\nolimits_{K_{p}\rightarrow\{1\}}R\Gamma_{c}(\mathrm{Sht}(G,b,\mu)_{\infty,C}/\underline{K_{p}},\Lambda(d_{b}))$ and $R\Gamma_{c}(G,b,\mu)_{\phi}$ is the projection applied to the complex viewed as a $G(\mathbb{Q}_{p})$-representation. This should also generalize once one has appropriate general definitions of $\mathrm{Ig}^{b}$ and $\mathrm{Ig}^{b,*}$ so that one can actually define $V_{b}:=R\Gamma_{c-\partial}(\mathrm{Ig}^{b},\Lambda)$. Under possible additional constraints on $\phi$, one should also be able to describe the contribution of $R\Gamma_{c}(G,b,\mu)_{\phi}$ in terms of the decomposition $V_{\mu}|_{Z(\hat{M}^{\Gamma})}=\mathcal{T}_{\mu}|_{Z(\hat{M}^{\Gamma})}$ for $b\in B(G)_{M}$ (along the lines of [24, Conjecture 1.20]), as is explained in the toral case in §6.1. It would be interesting to formulate an optimal conjecture.
Remark 6.8.
We believe that this conjecture should be true under just the weakly Langlands-Shahidi condition. However, we strongly suspect that the splitting of the semi-orthogonal decomposition and in turn the splitting of Mantovan’s filtration discussed in the previous Remark should not hold unless the set $B(G,\mu)_{M}$ is a singleton. In particular, in [30, Section 2.2] Hansen conjectures the existence of perverse sheaves lying $\mathrm{D}_{\mathrm{lis}}(\mathrm{Bun}_{G},\Lambda)_{\phi}$, for which the semi-orthogonal decomposition does not split. Nonetheless, one still expects perverse $t$-exactness of Hecke operators to hold in these cases [30, Conjecture 2.32].
Appendix A Spectral Decomposition of Sheaves on $\mathrm{Bun}_{G}$, by David Hansen
Let $G/\mathbb{Q}_{p}$ be a connected reductive group, $\Lambda/\mathbf{Z}_{\ell}$ an algebraically closed field. If $\mathrm{char}(\Lambda)\neq 0$ we assume $\ell\nmid|\pi_{0}(Z(G))|$, where $Z(G)$ denotes the center of $G$, as in [16, Theorem I.10.1].
[…]
Definition A.1.
Given any $\phi$ as above, $\mathrm{D}(\mathrm{Bun}_{G})_{\phi}\subset\mathrm{D}(\mathrm{Bun}_{G})$ is the full subcategory of sheaves $A\in\mathrm{D}(\mathrm{Bun}_{G})$ such that for every $f\in\mathcal{O}(X_{\hat{G}})\smallsetminus\mathfrak{m}_{\phi}$, $A\overset{\cdot f}{\to}A$ is an isomorphism. Here $\cdot f$ is the endomorphism of $A$ induced by $\Psi_{G}$.
We will call objects of $\mathrm{D}(\mathrm{Bun}_{G})_{\phi}$ $\phi$-local sheaves.
By construction, $\mathrm{D}(\mathrm{Bun}_{G})_{\phi}$ is a full subcategory of $\mathrm{D}(\mathrm{Bun}_{G})$ stable under arbitrary limits and colimits, and the tautological inclusion functor $\iota_{\phi}:\mathrm{D}(\mathrm{Bun}_{G})_{\phi}\hookrightarrow\mathrm{D}(\mathrm{Bun}_{G})$ commutes with limits and colimits. Since $\mathrm{D}(\mathrm{Bun}_{G})$ is presentable, $\mathrm{D}(\mathrm{Bun}_{G})_{\phi}$ is then presentable by [51, Theorem 1.1]. By the $\infty$-categorical adjoint functor Theorem [45, Cor. 5.5.2.9.(2)], the inclusion $\iota_{\phi}$ therefore admits a left adjoint $\mathcal{L}_{\phi}:\mathrm{D}(\mathrm{Bun}_{G})\to\mathrm{D}(\mathrm{Bun}_{G})_{\phi}$. [^15] To see that $\iota_{\phi}$ is accessible, use [45, Prop. 5.4.7.7] together with the fact that $\iota_{\phi}$ admits a right adjoint, which follows from [45, Cor. 5.5.2.9.(1)]. The unit of the adjunction gives a map $A\to\iota_{\phi}\mathcal{L}_{\phi}A=:A_{\phi}$ functorially in $A$. Since $\iota_{\phi}$ is fully faithful, $\mathcal{L}_{\phi}\iota_{\phi}=\mathrm{id}$, so $(A_{\phi})_{\phi}=A_{\phi}$, i.e. the endofunctor $A\rightsquigarrow A_{\phi}$ is idempotent. We remark that $\mathrm{D}(\mathrm{Bun}_{G})_{\phi}$ is a localization of $\mathrm{D}(\mathrm{Bun}_{G})$, and the map $A\to A_{\phi}$ is the initial map from $A$ to a $\phi$-local sheaf.
Proposition A.2.
[…]
Here $\mathrm{supp}(A)\subset|\mathrm{Bun}_{G}|=B(G)$ denotes the set of points $b$ such that $i_{b}^{\ast}A\neq 0$.
Proof.
[…]
Proposition A.3.
[…]
Proof.
Notation as in the previous proof, we have
$\mathrm{Hom}(C,A_{\phi})$ $\cong\mathrm{Hom}(C,\mathrm{colim}_{i\in\mathcal{I}_{\phi}}F(i))$
$\cong\mathrm{colim}_{i\in\mathcal{I}_{\phi}}\mathrm{Hom}(C,F(i))$
$\cong\mathrm{Hom}(C,A)_{\mathfrak{m}_{\phi}}$
where the second isomorphism follows from the compactness of $C$ and the third isomorphism is immediate from the definition of $(-)_{\mathfrak{m}_{\phi}}$. ∎
Proposition A.4.
If $A$ is ULA, then also $A_{\phi}$ is ULA.
Proof.
Recall from [16, Prop. VII.7.9] that $B\in\mathrm{D}(\mathrm{Bun}_{G})$ is ULA iff $R\mathrm{Hom}(C,B)\in\mathrm{Perf}(\Lambda)$ is a perfect complex for all compact objects $C\in\mathrm{D}(\mathrm{Bun}_{G})$. Now, if $C$ is compact, $R\mathrm{Hom}(C,-)$ commutes with filtered colimits, so
$R\mathrm{Hom}(C,A_{\phi})$ $\simeq R\mathrm{Hom}(C,\mathrm{colim}_{i\in\mathcal{I}_{\phi}}F(i))$
$\simeq\mathrm{colim}_{i\in\mathcal{I}_{\phi}}R\mathrm{Hom}(C,F(i))$
with notation as in the proof of Proposition A.2. Since $F(i)\simeq A$ for all $i$, $\mathrm{colim}_{i\in\mathcal{I}_{\phi}}R\mathrm{Hom}(C,F(i))$ is a filtered colimit of perfect complexes $P_{i}$ which vanish outside a finite interval independent of $n$, and with $\mathrm{dim}_{\Lambda}(H^{j}(P_{i}))$ bounded independently of $i$. It then easily follows that $\mathrm{colim}_{i\in\mathcal{I}_{\phi}}R\mathrm{Hom}(C,F(i))$ is perfect, whence the claim. ∎
Proposition A.5.
If $A$ is ULA, the natural maps $A\to\prod_{\phi}A_{\phi}\leftarrow\oplus_{\phi}A_{\phi}$ are isomorphisms, where the direct sum and direct product are taken over all semi-simple $L$-parameters. In particular, $A_{\phi}$ is functorially a direct summand of $A$ for ULA sheaves $A$, and the functor $(-)_{\phi}$ on ULA sheaves is perverse t-exact.
Remark A.6.
The isomorphism $\oplus_{\phi}A_{\phi}\overset{\sim}{\to}\prod_{\phi}A_{\phi}$ may be surprising at first glance. To put this in context, we remind the reader that if $(\pi_{i})_{i\in I}$ is a collection of admissible smooth $\Lambda[G(\mathbb{Q}_{p})]$-modules whose product $\prod_{i}\pi_{i}$ is admissible, then $\oplus_{i}\pi_{i}\overset{\sim}{\to}\prod_{i}\pi_{i}$ automatically, because admissibility of $\prod_{i}\pi_{i}$ implies that for any given compact open subgroup $K\subset G(\mathbb{Q}_{p})$ we have $\pi_{i}^{K}=0$ for all but finitely many $i$. A similar argument occurs in the following proof, which actually shows that if $(A_{i})_{i\in I}$ is any collection of ULA sheaves on $\mathrm{Bun}_{G}$ whose product $\prod_{i}A_{i}$ is ULA, then $\oplus_{i}A_{i}\overset{\sim}{\to}\prod_{i}A_{i}$ automatically.
Proof.
We first show that $A\to\prod_{\phi}A_{\phi}$ is an isomorphism. Let $C$ be any compact object. It suffices to prove that the natural map
$\mathrm{Hom}(C,A)\to\prod_{\phi}\mathrm{Hom}(C,A_{\phi})\cong\mathrm{Hom}(C,\prod_{\phi}A_{\phi})$
[…]
$\mathrm{Hom}(C,A)$ $=\oplus_{\phi\in S}\mathrm{Hom}(C,A)_{\mathfrak{m}_{\phi}}$
$=\oplus_{\phi\in S}\mathrm{Hom}(C,A_{\phi})$
$=\prod_{\phi}\mathrm{Hom}(C,A_{\phi})$
where the first equality follows from general nonsense about finite length modules over commutative rings, the second equality follows from Proposition A.3, and the third equality follows from the vanishing of $\mathrm{Hom}(C,A_{\phi})$ for all but finitely many $\phi$. This also shows that $\mathrm{Hom}(C,\oplus_{\phi}A_{\phi})\cong\oplus_{\phi}\mathrm{Hom}(C,A_{\phi})\to\prod_{\phi}\mathrm{Hom}(C,A_{\phi})$ is an isomorphism (here again the first isomorphism follows from compactness of $C$), which implies that $\oplus_{\phi}A_{\phi}\overset{\sim}{\to}\prod_{\phi}A_{\phi}$ is an isomorphism. ∎
Next, recall the Verdier duality functor $\mathbb{D}_{\mathrm{Bun}_{G}}$ on $\mathrm{D}(\mathrm{Bun}_{G})$, which induces an involutive anti-equivalence on the subcategory of ULA sheaves. Recall also that, for any $A$, the diagram
${\mathcal{O}(X_{\hat{G}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Psi_{G}}$$\scriptstyle{f\mapsto f^{\vee}}$${\mathrm{End}(A)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{O}(X_{\hat{G}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Psi_{G}}$${\mathrm{End}(\mathbb{D}_{\mathrm{Bun}_{G}}(A))}$
[…]
Corollary A.7.
If $A$ is ULA, then $\mathbb{D}_{\mathrm{Bun}_{G}}(A_{\phi})\cong(\mathbb{D}_{\mathrm{Bun}_{G}}(A))_{\phi^{\vee}}$.
Proof.
By Proposition A.5 and the remarks preceding its proof, the decomposition $A=\oplus_{\psi}A_{\psi}$ dualizes to a decomposition
$\mathbb{D}_{\mathrm{Bun}_{G}}(A)=\prod_{\psi}\mathbb{D}_{\mathrm{Bun}_{G}}(A_{\psi})\cong\oplus_{\psi}\mathbb{D}_{\mathrm{Bun}_{G}}(A_{\psi})$
where the second isomorphism follows from the discussion in Remark A.6. On the other hand, applying Proposition A.5 directly to $\mathbb{D}_{\mathrm{Bun}_{G}}(A)$ gives a decomposition
$\mathbb{D}_{\mathrm{Bun}_{G}}(A)\cong\oplus_{\psi^{\prime}}(\mathbb{D}_{\mathrm{Bun}_{G}}(A))_{\psi^{\prime}},$
so comparing these we get a natural isomorphism
$\oplus_{\psi}\mathbb{D}_{\mathrm{Bun}_{G}}(A_{\psi})\cong\oplus_{\psi^{\prime}}(\mathbb{D}_{\mathrm{Bun}_{G}}(A))_{\psi^{\prime}}.$
Applying $(-)_{\phi^{\vee}}$ to both sides, we get $\mathbb{D}_{\mathrm{Bun}_{G}}(A_{\phi})$ on the left side (using that $\mathbb{D}_{\mathrm{Bun}_{G}}(A_{\phi})$ is $\phi^{\vee}$-local), and $(\mathbb{D}_{\mathrm{Bun}_{G}}(A))_{\phi^{\vee}}$ on the right side. This gives the claim. ∎